A Complete Bibliography of Publications in Information Processing Letters: 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

20 October 2023 Version 2.16

Title word cross-reference

#3 [Kut07]. #3-SAT [Kut07]. #P [Liv09]. #P-completeness [Liv09].

(0, j) [CF09]. (1, r) [LL02a]. (1, X_{p}) [SW09]. (2, 1) [WC09]. (a, b) [JLN02]. (D − J) [ZB07]. (Δ + 4) [TJH07]. (f, n) [ZS09]. (γ, κ) [Efr08]. (λ, ω) [vDKST06]. (s, t) [Nag05a]. (SIR_{∞}) [CM07]. (t, n) [LHC^{+}01]. 0 [Bar01a]. 0, 1 [ELS06]. 1 [Bar01a, CHC02, DMN09, DBLS06, Hal03, HC01, KH03, Riz03]. 1, 2 [SK08a]. 1.235 [Kur04]. 1.5 [HL08]. 12 [HM09]. 13 [HSL^{+}02]. 1_{p} [KH03]. 2 [ABM09, BG08, Che00a, CKL^{+}03, CY02, Dal09, DLS08, EO07, HP00b, LCL08, LMMM04, MB08, OP08, Rom09, Tan06b, TZ08, ŠŽ04].

2/3 − ϵ [PS04b]. 28 [Sch06]. 2^{28} [Bih02]. 2^{n} [Pre00]. 2^{O(k)}poly(n) [Raz07]. 3 [BCP09, Cai09, DLP04, HM09, KKSŠ08, LR08, LCW07, MRWW08, NY09, Sea04, SW07a, She07, Sku02a, WLC08, WWW08, Wol06, ZS08]. 3/2 [CL03, Riz03]. 4 [EG03, HM09, HKKL09, Sku02a, WWW08, Xia07, Yun08]. 47 [Pec07]. 5 [WWW08]. 5/3 [GGK03, ZNI03]. 5/4 [ŠŽ04]. 78 [HR02]. 8/(7 + 1k − 1) [FK00b]. ^{+} [Lan08]. ^{∘} [KV08a]. ^{NP} [GP01]. _{5} [CZ02, Fou02, LM09]. A^{∗} × N^{m} [CG06]. aa → bc [HW06]. α [BLL02]. AM_{exp} ⊈ (NP ∩ coNP)/poly [Vin04]. b [PT00]. bb → ac [HW06]. c [KR04]. C_{7} [VŽ02]. C_{E} [Fer05a]. C_{k} [KOH09]. cc → ab [HW06]. co − NP [log^{2} n] [KS03a]. D [NB05, ABPR06, CGK08, Jun05]. Δ [Sku06, CIL^{+}03]. ℓ_{∞} [CM07]. ϵ [LMST04, Lif03]. ∃∀^{2} [Tal00]. f [CQ09]. f < n [FMR04]. FewP [RRW94]. g^{x^{2}} [Shp02]. G_{n,p} [COK06]. GF ((2^{n})^{m}) [CKKC04]. GF (2^{m}) [SJ08]. GF(2n) [HT00]. H_{1} [GL05]. i [OF03, RR00]. j [Kou06]. K [CR09, AGM03, AIM^{+}03, Ara07, BCNR02, BN01, CC01a, Cha05, CPX06, CKY06, DGN05, EA07, Eis09, FGR03, FKW09, Gør08, GDN07, GL09a, HHP08, JVA06, Jia05, KLP04, KL04a, Kri02, LM08a, LC03b, LMZ08, LEP07, Maf09, MNRR09, Nag09, NON09, OT03, PCW09, PC05b, QPV05, Smo08, SG03, TAI04, TZ08, TLZ05, TM09, VGG^{+}00, WL02, Wes08, Wil09, XUT00, XUT01, YAM04, YSK04]. K_{2,n} [BCLO03]. K_{3,3} [CZ02]. K_{4} [WY06]. K_{n,n} [CF09]. κ [HZ09]. k ≤ ϵ log n [Seg05]. l [QPV05, Vid03]. L(2, 1) [KV05]. L(p, q) [WY06]. L^{2} [Ram06]. L^{k} [Ram06]. L_{1} [LP08b, SS00]. L_{∞} [LP08a, XDZ05]. Λ [JN06, eF02]. λ^{μ} [eF02]. λ′ [BCD^{+}05, WL08]. λ_{c} [TM09]. ≤ [DG02]. log_{2} 14 [Via08]. log log N [Geo08]. m [GZN06]. n [GZN06, KPC07, Kur04, PCW09, Lav08]. n + C + o(n) [GS03]. o(log m) [TZ05]. O(log n) [Tse05]. O(log n/ log log n) [Li09]. O(m) [HJS01]. O(mn) [RR00]. O(n(log n)^{2}/ log log n) [SW09]. O(n) [WYL06, HR06a]. O(N^{2}) [Kun00]. O(n^{3}/ log n) [Han08]. O(n^{3} log log n/ log n) [Tak05]. O(n⌈m/w⌉) [GF08]. O(n log n) [DM08b, PT06]. O(|V |^{2}) [Khu99, Khu00]. O^{∗}(2^{k}) [Wil09]. ω [KDTS01, Ser03, Tao06]. Ω(n log n) [MRST09]. ⊕ [BDHW07]. p [DLPS08]. P_{3} [KMŻ06]. P_{5} [BLL02, Maf09]. P_{n}^{2} [NON09]. Π [VV01]. Π_{1}^{0} [Tal00]. r [JKŠ05, Jha03, Yus06]. R^{3} [BD09]. R^{n} [BMCK04, KR05]. Res(k) [Seg05]. Res(k + 1) [Seg05]. s [FP04, GHS02]. t [AOS^{+}09, CP05a, GY08, ZLM09]. τ [BMX05]. Θ(log n) [GS09a]. Θ(N^{2}) [MM05].

T L[EF] [Wu07]. V [Kou06]. W [1] [Ces02]. x [Sma01]. Z_{n} [SK08b].

-approximable [HL08]. -approximation [CL03, GGK03, Riz03, Tan06b, ZNI03, GS09a, HR06a, HR02]. -ary [HHP08, JVA06, Jun05, PCW09, XUT00, XUT01, NB05]. -automata [Tao06]. -Background [LR03a]. -calculus [eF02]. -Canadian [Wes08]. -center [Gør08, Hal03]. -chains [DG02]. -choosability [WLC08, Xia07, ZS08]. -choosable [HM09, SW07a, WWW08]. -CNF [AIM^{+}03]. -color [MRWW08]. -colorability [Kri02]. -colorable [LCW07, Maf09, CR09]. -coloring [HKKL09]. -competitive [ŠŽ04]. -competitiveness [Geo08]. -complementing [AOS^{+}09]. -complete [Ces02]. -constraint [ELS06]. -coordinate [Sma01]. -corner [DMN09, DBLS06]. -Covered [ABM09]. -critical [ZS09]. -cubes [PCW09]. -customer [Nag09]. -cuts [Nag05a]. -cycles [EG03, WWW08]. -D [BCP09]. -deap* [Jun05]. -decompositions [NY09, vDKST06]. -degenerate [LCL08]. -dimensional [CGK08]. -distance [FGR03]. -domination [Jha03]. -edge [CHC02, KH03, TZ08]. -edge-coloring [KKSŠ08, Sku02a, TJH07]. -exclusion [TAI04, Vid03]. -factors [MNRR09]. -fair [HC01]. -fault [CQ09]. -fold [EA07, Eis09]. -form [Efr08]. -free [Lif03, LM09, BLL02, CZ02, KOH09, Maf09]. -from [Mad03]. -gon [CPX06]. -grammars [BN01]. -Hamiltonian [KH03]. -hard [Tal00]. -Helly [DLPS08]. -Hitting [Cai09]. -hop [LM08a]. -inductive [Smo08]. -institutions [VV01]. -keyword [KLP04]. -label [Tse05]. -labeling [KV05]. -labelling [WY06]. -languages [FP04]. -List [Sku06]. -local [ŠŽ04]. -matching [CIL^{+}03, PT00]. -median [CKY06, VGG^{+}00]. -medianoid [SW09]. -minor [WY06]. -monotonic [ZB07]. -MST [PC05b]. -mutual [CC01a].

-NAF [BMX05]. -nearest [GDN07]. -NN [TLZ05]. -OBDDs [BDHW07]. -optimal [WL08]. -optimality [BCD^{+}05]. -Optimally [TM09]. -outerplanar [EO07]. -packing [KMŻ06]. -partite [ZLM09]. -partitions [GHS02]. -perfect [JKŠ05]. -point [DGN05]. -polytope [Kou06]. -prisms [CF09]. -rankings [NON09]. -rationality [Lav08]. -redundant [BLL02]. -regular [Ser03]. -relaxed [CP05a]. -Restricted [LEP07]. -round [CKL^{+}03]. -searchers [YSK04]. -secting [KPC07]. -semidistributive [JN06]. -server [BCNR02, FKW09]. -simplex [Kou06]. -states [Yun08]. -Steiner [WL02, Riz03]. -subcolorability [LL02a]. -subgraph [LMZ08]. -termination [Fer05a]. -time [PT06]. -Total [HSWC09, WC09]. -trajectories [KDTS01]. -transitions [LMST04]. -tree [SG03, YAM04]. -trees [JLN02, OP08]. -triangulated [RR00]. -triangulation [KV08a]. -tuple [Ara07, KL04a, LC03b]. -uniform [Yus06]. -vertex [Kur04, TZ08]. -vertex-ranking [KR04]. -wise [AGM03, GY08].

05 [ZC09a]. 0CFA [LYP02].

1-maximal [SGH04]. 1-variable [BNvO09]. 1.5-Approximation [BG09]. 104 [LPWZ14]. 128 [JJ02]. 180 [RRW00].

2- [Loz05]. 2-approximation [FD04]. 2-bit [HL05]. 2-constraint [Koi06]. 2-dipath [MW06]. 2-facial [MR06]. 2-factor [Pan05]. 2-layered [NY04a]. 2-query [SL06a]. 2-splittable [Kol05]. 2-variable [BNvO09]. 2003 [Ton06]. 2004 [ZC04]. 2PM [RSGK08].

3-approximation [PS09]. 3-choosability [Mon06]. 3-CNF [BB09]. 3-colorability [Loz05]. 3-leaf [BL06]. 30 [KV08a].

5-cycles [CHZL09]. 52 [Cha02b, RRW00].

7-round [Pha04]. 72 [Khu00].

81 [AS02a]. 82 [Kwo03]. 87 [GP03a].

Absolute [TZ05]. Absorbant [SCK07]. Abstract [BBD02, ADM04, Bla05, DM08a]. Abstracting [HK00]. abstraction [Dil02]. accelerate [Kel02]. Accelerated [KM02]. acceptance [BHW06, NSW02a]. accepting [MM07a]. access [BEL08, Bol08b, CW04, IY06, LHS01, MF07, PS02, DFM04, Sha09a, hY09a]. accumulating [BA03b, ZHW01]. accurate [KLN^{+}03]. achieve [AK08, Alb02, Geo08]. ACISP [Ton06]. Acknowledged [FP08]. acknowledgement [NSW02b]. ACO [Gut02]. active [BCP05, Fuj03]. Acyclic [FGR03, FR08, JMV06, Sku04, Šte00, XKS06, AF08, BFK09, BLME02, FGR02, FHN08, IN04, KdV01, MP04, Sub06, ZM03]. acyclically [HM09]. ad [FP08, GL02, WW09, ZC09a]. ad-hoc [ZC09a]. Adams [CWTHR01]. adaptation [Nau01, TC02]. Adaptive [PG06, BHW03, GL02, Hie06, IM06, JUZ05, SM02, ZH07, BHW03]. Addendum [Khu00]. Adding [BHM08]. Addition [PP04, HT00, TV02]. additional [IAIH03]. additive [BD00]. address [Lop00, LT03]. adequate [CK07]. Adjacency [Kow07]. Adjacent [CG09]. adjustable [FKM06]. adjusting [CC02]. adjustment [RK09]. admissible [Göb00]. admission [YC09]. adOPTed [LC03c]. Advanced [Pha04]. Adversarial [ÀBD^{+}04]. advice [NY04b]. AES [Pha04]. after [CT03a, Deo05]. against [XH05]. agent [LCLP09]. agents [JX09]. aggregate [Mel07a]. aggregation [SZM09]. aggressive [Dai09]. Agrawal [Sze08]. Agreement [Cha05, AFG07, CWT08, Gue01, LHC^{+}05, PT06, Tsa06, ZC04, AFG07]. ahead [FKV04]. Aho [DL06]. al [LHC^{+}01]. alarms [YCKK07]. algebra [BP01, CCVP01, Gam00, NNP07, PU01]. Algebraic [KDTS01, CKN06, HL05, Hon02, Kat04, Win03, Yun08]. algebraically [RBF08]. algebras [CWTHR01, Dia09]. Algorithm [ZM03, ADMP01, Ala05, AS07b, AZ08, Alo03, AYTH08, Ars08, AF08, ARdAdS07, BCNR02, Bas08, BHY09, BG01, BCP08, BFK09, Ber00, BG09, BDL09, BMX05, Boj07, BS08b, BdFdS06, BS09, BM01, Bre04, BUŽ02, CKK03a, CCGP05, dSCLP08, CK03, CKKC04, CC01b, Che03, Che04a, CC06a, Che06a, CLZ09, CNK06, CW07, CW02, CW03, CDF^{+}04, CS00, CKY06, Chu08, CY01b, COK06, CT05, CT07, CIPR01, DDR04, Dai09, DR06, DMN09, DFG^{+}03, Deo09, Det09, DH03, DM08b, DOS08, Dur03, ER03, Eng04, FKM06, FG06, FB04, FD04, FA05, GS05a, GGK03, God02, GH05, GD04, HR06a, Hal03, Han06, Han08, HS00, HR02, HR06b, HLY09, HZ03, HR05, HR03, Hes03, Hes07, Heu08, HT00, Hu07, HTT03, Iba09, IAIH03, IS08, IMS00, IP09, IM06, JCL04, JN07, JCSX09]. algorithm [Jia06b, JLY07, Joh04, JXL04, KS05, KS06, KR04, KZ06, KY03a, KBEG07, Khu99, Khu00, Kim03, KLP04, KS00, KMS^{+}02, KKL03, Kol03, KCYL06, KM07, KMŻ06, KOH09, Kou05, KPS02, KW01a, KW01b, Kun00, KHLL09, Kwo02b, Kwo03, LPW00, LKL^{+}00, LKY02, LHK02, LHC^{+}05, LTH05, LLP09a, LTT00, LH04, LSC06, LSL08, LMZ08, LZ09, Lin02, LH05, Low02, Low06, LWC01, Luc04a, LC03c, Mad06, MTF02, Mar02, McC01, MNRR09, Mis03, MC02b, Nag09, iNM07, NP07, NN05, NP00, NSW02b, OK04, PD03, PD09, Pan05, Pec07, PT06, PS09, Qua02, Raz07, RS07c, Ric02, RRSY07, RGDN04, Sak06, Sas02, Sas08, Sea04, SKA05, Sed07, Seg08, SGH04, SLL^{+}05, ST06, Smo08, SW09, SLT08, Tak00, TAI04, Tak05, TH02, Tan06b, TYW01, TC02, Tre03, TLH02, Tsi07, TH09, TJH07, TJH09]. algorithm [UPGM09, Vah07, Vid03, VGG^{+}00, Vog08, Wan01, WL02, Weg02, Wes04, WS07, Wot01, WS01a, WS01b, WW03, WS06, WYL06, Wu09a, XX05a, YHC05, Yan08, Yij04, ZSZY08, ZLZ05, ZH07, K^{+}02, ŠŽ04, AD06, Mis01]. Algorithmic [SAOKM01, Goo09, Hon07]. Algorithms [AMT09, ADLP04, FMSS01, Gav02, GHZ02, KK02, MUK04, Ver02, AKL07, Alb02, ALA06, BFL00, BL00, BHK^{+}00, BEP09, BLdR04, BA03b, BdW03, BZ02, CP02, CWC00, COW05, CC05a, CLC09, CLT07, CDL^{+}04, CHLN05, DF03, DLPS08, DH04, DD06, Eid02, FFFdP07, FL02, FH06, FGK^{+}09a, GIS04, GL09a, GM07, GNU08, Gur08, Gut02, Gut08, HR01, HYT^{+}07, HYT^{+}08, IR08b, KM02, Kel02, KL04a, Koi06, KR08, KC02, Lec07, LPC02, LCKL06, LCLP09, LM06, LJ03, Lin07, LC08, LL09, LSY07, MC06, NR00, RS07a, Reg02, Sep05, SJ05, SLL^{+}07, SG03, Šte00, Tsu08, Tur07, UC00, Vas09, Ver06, WCH08, Wu00, WHZ04, Wu08, ZC09b, Zie08, dFdF03]. aligned [Jun05]. alignment [CLT07, Man05, LSZU03]. all-farthest-segments [DM08b, MCL06]. all-optical [CQ09]. All-pairs [MPSS02, Sib04, Tak05]. all-to-all [SZQ08]. all different [AMM05]. allocation [CCY^{+}07, FSW03, HG06, KLN^{+}03, LCKL06]. allow [RRB00]. Almost [AGM03, SKA01, GH05, HIM09, LS07a, dFdF09, LH05]. almost-linear [GH05]. almost-sure [LS07a]. almost-unit [dFdF09]. alpha [HTT03]. alpha-connected [HTT03]. alphabet [JS07]. alphabets [DL06, Fre03b, Gag06b]. Alternating [HT03, Göb00, JS07, Sze06, jXyL09]. alternative [Nun04]. Alternators [KBOR05, HC01]. always [Kle08]. ambients [GYY01]. AMO [ST06]. among [HI03, Vig03, Wan08]. amplification [Mon08]. analysed [Mad06]. analyses [Cha02a]. Analysis [Heu08, Hon07, KV04, Sub06, Tan06a, AS07b, AZ08, ACN03, AF08, BG01, BLMM06, BDL09, CS00, DM03, DOS08, Dur03, Eng04, FA05, GM08b, JLO01, JLY07, KY03a, LHS01, LW09, LJL07, MB08, NSZ00, RK09, Ros02, ST06, SYL04, TZ05, Tre03, WNR00, XH05, XZX07, ZH07, dFdF03]. analytical [GRH09, Wu09b, YWW09]. analyzer [YCKK07]. Analyzing [Bre01, ST09]. Ancestor [RP03]. angle [AW04, KPC07]. angle-restricted [AW04]. Annealing [Mee07]. annotated [Dam06]. Annotation [FJL01]. anomalous [LMVV05]. anonymity [KMST07, LV07]. Anonymous [BBMR08, CC03, DB06, HHT05, Mia03, Sae02, Sha03a, SGH04, SWCC04, TJH07, ZC09a]. answering [Mad05]. answers [CV03]. Ant [CA08, AF08]. antennas [ABM09]. antisymmetry [BKS01]. Antwerpen [Min05]. any [Göb00, DFM04]. API [Zea00]. applicability [CDP02]. Application [PRSS01, Cha03, CZ02, DM08a, HKL01, Jha03, YSK03, GNAJT09, Zea00]. Applications [BK05, ACN03, BHM08, CI08, HLSW02, Jia05, LCL06, MB08, MCS^{+}09, MC02b, SKP^{+}02]. applied [Ohl00]. Applying [AK08, Hie06, JUZ05, vDKST06]. approach [BdATdC04, BEF^{+}00, CTGN05, FJ08a, FMCW02, GNAJT09, GL07, HŽR08, LR05, OB03, QH01, Ves06]. approaches [LMMM04]. Approximability [Suo07, EMP06, Fuj01a, Fuj01b, KZ09, Man05, Pen02]. approximable [HL08]. Approximate [ALP02, BBS06a, KKS04, LP08a, AGS07, BM01, CN00, DD08, GHL02, Hyy08, KR08, Liu04b, MN05, SSZZ00, SW07b]. approximated [NEK03]. Approximately [KPC07, Kuj09]. approximates [WS07]. Approximating [AHS01, BS00a, CR09, Cha05, FMT09, FKN02, HV06b, HN08, Jia06a, LM08a, NO07, NR08, Wan08, Zie09, CM07, COK06, DS00, DS04a, FKL01, FR06, KRS01, OB03, Shi00, SB06, TZ05, XX05a]. Approximation [AH08, ABG05, BG09, BHK^{+}00, BEP09, CDL^{+}04, CHLN05, Eid02, FL02, HR01, KC02, LCLP09, LSY07, Yan08, Zie08, AFT08, BCG09, BHY09, Bre04, BDHW07, CKK03a, COW05, CL03, CKR06, CKW09, DFMR08, DH03, DH04, EL08, FL07, FFFdP07, FD04, GGK03, GR03, GS09a, Gut08, HR06a, HIM09, HR02, HR06b, HG01, IS08, JQQ^{+}03, Jia06b, KZ06, KL04a, KKSŠ08, KMŻ06, LLP09a, LW06, LMZ08, LL09, LR08, Low06, MR09a, Mon02, Nag09, NN05, PS04b, PS09, PHG08, Riz03, Sep05, Tan06b, UC00, UPGM09, WL02, WCH08, ZNI03]. approximations [GL04a, HR00]. approximative [Jon03b]. APX [DP06]. APX-hardness [DP06]. arank [IJN09]. arbitrary [DGN05, Dia04, FJ08a, Man05, MVM07, Xir06]. arborescence [BFK09]. arc [BFK09, CQ09, CLM09, Dam06, GHM07, Gup08, IN04, LM08c, Mel07a, NBCS07, PS06, Vol07]. arc-annotated [Dam06]. arc-connectivity [Vol07]. arc-disjoint [BFK09]. arc-forwarding [CQ09]. architecture [CWCL02, KLY00]. architectures [CHU06, DIPS00, HC03b]. arcs [Dam06]. area [GG02, MKGB08, PC00]. areas [FJ08a]. arising [AFK08, Gus02, SK08b]. arithmetic [Duq07, HT00, To09, LM04a]. arithmetics [DLX09, Kuj09]. arm [Pal03b]. Arom [CT03a]. arrangement [FL07, HR01]. arrangements [Liu04a]. array [CWCL02, Dil02, LKP01]. arrays [Kas04a, Maa07, MNR04, Rao02]. arrivals [CB07]. Art [ACHP09]. Articles [Ano01g]. artificial [Tam00]. ary [HHP08, JVA06, Jun05, NB05, PCW09, XUT00, XUT01]. Asian [ST06]. ASL [Bor00]. Aspect [Asa07]. Aspect-ratio [Asa07]. aspects [DPS06]. Assembling [Kuj09]. assembly [MAL08]. Assignment [CKR06, AHS01, AAER03, Aud01, BG09, BA06, CT05, CT07, Hwa00, KKS04, Low02, DFM04, TLH02, hYScH05, hY08, hY09a, ZB07]. assignments [LT07]. association [HC02]. associative [Kno05]. associativity [BF01, Pal09]. assumptions [BW01]. assurance [Amb01]. Asymmetric [Gør08, Gag08a, LH02, Mel07b, Miy08, Rez05]. Asymptotic [Dur03, EL08]. Asynchronous [FMR05, FMR04, Hoe01, IM06, MRT00]. Atkin [KCYL06]. ATM [NSZ00]. atomic [DGP04, Tal00]. atomicity [IR09, KBOR05]. Attack [Sti04, ZC04, GM09, HL05, JJ02, Pad05, XH05, hYScH05]. attacks [DK08, GS04a, Shi08]. attractor [BBS06b]. attractor-property [BBS06b]. Attribute [MP09, Kas04a]. attributed [FM00]. auction [BDQ08]. auctions [PT00]. audio [LH04]. augmentation [PN01a, PN01b, RL02]. augmented [BS08b, Che04b, CC05b, MLX08, MTXL09, WMX07, XX07, HLT07]. augmenting [HLS03]. authenticated [HWW02, YPKL08, ZC04]. authentication [CWJT01, GS09b, LC03a, PCC03, Zha06]. Author [Ano01a, Ano01b, Ano01c, Ano02d, Ano02a, Ano02b, Ano02c, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano06a]. automata [AS07a, ANR05, BD00, BEF^{+}00, BBM06, CY04, CS03b, DPS02, Doy07, DSV08, GTM03, GdPL08, GNR08, GING09, GRT00, Héa08, JS07, Lit03a, Löd01, MVM01, MP02, MSS06, NSW02a, Sku02b, SHH07, Tao06, To09, Tri06, UN07, Yun08, ZLZ05, tBK05, LMST04]. automatic [KMT04]. automation [MC02b]. automaton [Bol09b, DL06]. automorphism [Pae03]. availability [Bre04, CL03, TH02, TK02, WC01, WL03]. Available [HM01, Pre00]. avalanche [Mai02, Sun01]. Average [LJL07, AAK09, BF09, CC05a, CLC09, DW07, EHRS09, Rez05, TZ05].

Average-case [LJL07].

average-constrained [CLC09]. avoidance [LLK08]. avoiding [Bar09, CD09]. axiomatic [PSL^{+}01]. Axiomatizing [HHLS03]. axis [Ary02, BS00b]. axis-parallel [Ary02, BS00b].

Background [LR03a]. backup [RK03]. Baeza [Hyy08]. Baeza-Yates [Hyy08]. balanced [BJ08, CCKL02, Chu04, FY03, HLT07, LLP09a, Low06, Mai02, Sun01]. balancedness [FSGM03]. Balancing [LTT06, AKE00, KM02, LLP09b, MCS^{+}09, QH01]. ball [dFdF09]. Bandwidth [LCKL06, MC02a, SS03a]. bar [JCP08]. barycenter [LS02]. based [AGGZP09, BHMR01, BdATdC04, Bij01, Cha02a, CTN03, CC09, Chu08, CL06, CDP02, DEPZ09, FMR04, Gab00, GNAJT09, GT09, HS00, HC08a, HC03b, IKKY00, IM06, JCSX09, KLN^{+}03, KHL09, KR03b, KY06, LHS01, LM06, LPM05, MHK07, Mel07b, MK05, Min07b, RG09, Sae02, ST00, Sha03a, Sha03b, SLL^{+}05, SLL^{+}07, SB02, Tat06, VS01, XL00, YWW09, Yan00, hY08, YC09, YPKL08, Zea00, ZC04, ZC09a, ZLZ05, ZH07]. Bases [BRSV05, ELR07, Gal03, Göb00, Gol04, Ham07]. basic [ACIL08]. basis [BMP^{+}00, JN06, LR05, YZ08]. batch [CHK^{+}09, MO06, PYZ09, TCNY09]. batch-dependent [MO06]. batching [HLY09]. Bayesian [Zie08]. BC [TYP08]. BCH [LWC01]. BDD [WNR00]. BDDs [GD00]. be [BM02b, BdW03, MM07a, NEK03, SW07a]. becoming [SWCC04]. before [GMNP09]. behavior [BCP05, CWC00]. behaviors [Lit03a, Nun04]. Bendix [SL06b]. Bernoulli [Giu04]. best [AD06]. best-so-far [AD06]. beta [DN05]. beta-redexes [DN05]. Beth [Stu09]. Better [HR00, DL08, Lev06]. Between [Loz05, ASW05, ADM04, Aku06, AL02, ARdFPS02, BFB09, BP06, BB01, Bro05, CW04, Cap06, CB04, Cle02, DM08a, FJ08b, FB04, HYT^{+}08, KP06, Lan08, Las09, LSL08, LEP07, Pal03b, Par08, RR01, Seg05, Tan00, Wan01]. Betweenness [DEPZ09]. beyond [Rah08]. BFS [BDLP08]. bias [KN04a]. biclustering [PHG08]. bicompatible [PD09]. biconnected [Gab00]. biconnectivity [RL02, Tsi07]. bicriteria [ABG05]. bidirectional [Car02]. bidual [EIM02]. big [DM08a]. big-step [DM08a]. bimatrix [CŠ05]. bin [CP05b, CDL^{+}04, EL08]. Binary [Ary02, CV03, MR09b, MA06, Wan07, BHW06, Cle02, Dur04, EQ04, FR04, FSGM03, Gag05, HI03, Kuj09, Lev06, LCC06, LWC01, LEP07, LP09, Pal00, Pal03a, Pal03b, PT06, Pio02, ZC09b, ZC09c, LJL07, MRT00]. BioHash [TJC^{+}06]. biology [CT05, CT07]. biometrics [CTGN05]. Bipanconnectivity [LTTH03]. bipancyclicity [Che09b, KA06, Lai09, LTTH03, STH07, TJ07, XDX05]. Bipartite [GL04b, Alo03, CDS09, CZ02, Daw03, FKM06, GHNP02, GHM07, Gup08, KS00, LLP09a, LS02, LT06, Lie08, Low06, Loz02, MTF02, Mon05, Sas08, TJH09, XKS06]. Bisection [FKL01]. bisimilarity [Frö05]. bisplit [Van03]. Bit [GF08, CIPR01, EM08, HL05, Hyy08, SJ08]. Bit-parallel [GF08, Hyy08]. bit-vector [CIPR01]. bits [Shp02, Via08]. bivalency [WTC05]. black [Goo09, Shi00]. black-box [Shi00]. black-peg [Goo09]. block [BF09, Cha02b, CCKL02, Chu04, GO05, HL08, PC05a, SHH07, XH05, YC94]. blocking [BA06, CKS09, DGP04, RK03]. Bloom [ABPH07, BGK^{+}08]. Board [Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano04x, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Cur04, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano03w, Ano03x, Ano03y, Ano03z, Ano03-27, Ano04b, Ano04c, Ano04d, Ano04e, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l]. Board [Ano06m, Ano06n, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano07n, Ano07o, Ano07p, Ano07q, Ano07r, Ano07s, Ano07t, Ano07u, Ano07v, Ano07w, Ano07x, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano08t, Ano08u, Ano08v, Ano08w, Ano08x, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q, Ano09r, Ano09s, Ano09t, Ano09u, Ano09v]. bondage [ZLM09]. Boneh [Hes04]. bookmark [BFL00]. books [FL05a]. boolean [Bol05, BRSV05, FMSS01, FA05, GKL09, HK00, KSS00, KS03a, Mai02, QPV05, Sar00, Shi00, SKA01, Sun01, Wol06]. Boosting [IR09]. border [Frö05]. both [Hoe01]. Bottleneck [PZ09, EK00, PC05b, TZ08, WL02]. Bottom [SPKL07, AZ08, FKV04, SV05]. Bottom-up [SPKL07, AZ08, FKV04, SV05]. bound [AM05, Aro02, BB05, BCLO03, Bol05, Cha03, CS03a, CZ02, CM07, CG04a, DS00, ELS06, FK00b, HIM09, HH07, KR03a, Kur04, Kut07, Li09, Lif03, LP08b, Liu04b, Lon03, MRST09, PCW08, Pal00, Riz03, Sep05, SL06a, SK02, SM02, Tse05, VŽ02, WTC05, WCH08, Wu09b, XDZ05, XL00, hY08, ZCF^{+}06, dWH03]. Bounded [CP05a, GT09, ZC09b, AFT08, Ala05, BYSP02, BCG09, CR02, Dar05, DKV09, ELO08, FMR05, GK00, Hås00a, HP05, IP09, MZ00, Mos08a, NR08, PP04, PP05a, SW05b, ST04a, YAM04]. bounded-degree [ELO08]. bounded-depth [ST04a]. Boundedness [DV06]. Bounding [CT03b, HRWZ07, BE05]. Bounds [Bac07, LY07, Pet02, Shp07, AAK09, ACI^{+}06, Asa07, BME02, BP06, BZLN07, BW01, BV00b, CKK02a, CLL04, Cho08, ELR07, FS09, FKW09, FM05, Fun08, GING09, GHL02, HV06a, LH02, LLC09, Las09, LS02, LM04b, MP08a, NB05, NDG02, PS02, PCW09, Par11, Pud00, Rau05, San01b, Saw07, Shi00, Shi02, ST05, WLLS08, WCK04]. box [LW05, Shi00]. boxes [BK00, BS00b, GS05b]. Boyer [DR06]. BPP [CW04]. brain [Kno05]. branching [BW03, Bol03, Bol08b, Bro05, CDC04, NEK03]. branchings [BBS06a, Zie09]. breadth [Hes03]. Breaking [Küh08]. bridge [BB01, Tsi07, Wan01]. bridge-connectivity [Tsi07]. bridges [Tan00]. broadcast [CKK03a, DN08, FK00a, Fer00, HGYZ09, KM03a, Mor06, RG09, SS03a, ZC09c, BHY09]. Broadcasting [PP05a, Che04a, DP01, FP08, KM03a, MSS09]. Brother [KH03]. browsing [CB07, PHK01]. Bruijn [Ara07, KFS01, KS03b, Mor05, SCK07, SDC09]. bubble [KA06]. bubble-sort [KA06]. BubbleSearch [LM06]. Büchi [LW09, Sku02b, UN07]. bucket [UPGM09]. Bucking [Bra01]. budget [Lev06, Lev07]. buffer [KV04, KLN^{+}03]. bug [YCKK07]. bug-finding [YCKK07]. build [dSCLP08]. build-up [dSCLP08]. built [Par05]. Bundle [FMCW02]. bundles [BŽ06]. Burrows [Deo05, MRS03]. butterflies [CKK02a, CLL04, WMG00]. bypasses [Ala05]. Bypassing [WNR00]. Byzantine [ACKM07, Att02, PP05a, TK02].

C [Bor00, KV03, YCKK07]. Cache [dBT08, HC08a, Jun05, KLN^{+}03, KY03b, SSYM04, SB04, Vog08]. cache-aligned [Jun05]. Cache-oblivious [dBT08]. caching [GIS04]. cactuses [Sid07]. Calabrese [CWTHR01]. Calculating [Nau01]. calculational [MM01]. calculi [Gol04]. calculus [eF02, Gol00, MS06, PRSS01]. Calinescu [FK00b]. Calinescu-Karloff-Rabani [FK00b]. call [ADM04]. call-by-need [ADM04]. calls [GTM03]. Camellia [XH05]. Can [Hoe01, MSSW08, Alb02, BM02b, LSS07, MM07a, PSL^{+}01, ZXZ08]. Canadian [Wes08]. cancelable [CTGN05]. cancellation [ZCF^{+}06]. cancer [GNAJT09]. cannot [BdW03, NEK03]. Canonical [IK07, Lec04, ZH05]. capacitated [JR04, ZLJ06, AD06]. capacities [CKP00, Cho08, Mel07a]. capacity [PZ09, VŽ02]. capping [JN07]. Capturing [AMS08]. cardinality [BHM08, CFM09, KK04]. care [CIL^{+}03]. cares [LS09b]. Cartesian [BŽ06, CLM09, DPP05, JHfS08, LM08c, LCM09, SZQ08, XXH05]. case [ACRS04, BG01, CFvG08, DW07, GF08, GPS01, Irv07, KY03a, LJL07, TZ05, VSSV05, Weg00]. cases [CVEOV08, Hie06, JUZ05]. CAT [ER03]. catastrophic [MRN02, MNR04, NRS00]. categorical [Dia08, OK04, dH01]. Category [CT00b, HP05]. Cauchy [Sim04]. Cayley [AD09, LM08b, XP06]. CCA2 [RG09]. CCS [ACIL08, KJ04]. cellular [AS07a, AD06, CCY^{+}07, SHH07, Yun08]. center [AAK09, Gør08, Hal03, HTT03, KRD08, LPM05, MU05]. central [GRH09, GLS08]. Centrality [DEPZ09]. centroid [WS07]. certain [CH06, CMO05, Lav08, Sun01, ZS08]. certification [DS04b]. certified [Sae03, Sha09b, Wan04]. certifying [Chu08]. Chain [FMT09, Geo08, Tan07].

Chain-splay [Geo08]. chaining [PCC03]. chains [BBS06b, DG02, DHS03, HLS03, PP04]. chair [BLV03, BLdR04]. chair- [BLV03]. chair-free [BLdR04]. Chang [CWJT01]. change [Fer05a]. changes [Her04]. channel [GS04a, Sch02a, YWW09, ZXC05]. channels [EM08]. Characterising [Dav04]. characteristic [HKL01, HI03, Sma01]. characteristics [Mai02, Sun01]. Characterization [BG08, MNR04, Pre02, AA05, Bou02, CMS03, Cap06, Che04b, DW07, Eng07, ET03, FP04, FKV04, KS02b, May02, Mia03, NNR04, Pap09, QPV05, Sar00, Sku02b, Win03, Wu07]. characterizations [FR04]. Characterizing [JKŠ05, CB01, RH02]. chasing [FIY08]. check [KLK06]. checkers [FJL01]. checking [Lib06, MS06, MMV08, Ste08, ZLZ05]. checkpointing [BHMR01]. Chen [CW07]. chess [vdBW08]. Chinese [DLX09]. chip [ACRS04]. chips [MB08]. choice [FHK06]. Cholesky [BVG02]. choosability [CHZL09, Mon06, WLC08, Xia07, ZS08]. choosable [HM09, MR09b, SW07a, WWW08]. chop [LS06]. chord [OPR07]. chordal [CR09, CLZ09, DN06, LLC09, LMZ08, Par08]. chosen [CHH^{+}09]. chosen-ciphertext [CHH^{+}09]. Chromatic [GDN07, CG09, COK06, DLS08, DS06, FRR03, GK00, MW06, RW09, Sid07, Sop02, ST04b, dWH03]. cipher [SHH07, XH05]. ciphers [DK08]. ciphertext [CHH^{+}09]. Circle [DP07, DP06, Gav08, Jia06b, KRD08, SK02]. circles [JQQ^{+}03]. circuit [FM05, HS01, Kwo02a, Poo03]. circuit-switched [Kwo02a]. circuits [BW06, FS09, KJ04, Pud00, ST04a, Wol06]. circulant [AS02a, AS02b, OPR07, TTH04, YEM05]. circulants [Ara03]. circular [AAN^{+}07, Deo09, HP00b, NBCS07]. circumscription [DH08]. CKY [RP08]. clairvoyant [KC03a]. Clarifying [CVEOV08]. class [ELO08, Gus02, Hon05, HL07, Lav08, Li00, LL01, PC04, PR05, Sep05, SCHT08, XZX07]. classes [Bar09, BM02b, BLdR04, Cap06, CHL^{+}08, CT00b, EHHR00, Eis09, HRV00, IT03, KV03, Mos08a, Pal03a, RRW94, RRW00, YC09, Živ09]. classical [Bla05]. classification [IN04, Nik00, YCKK07]. classifier [CTN03, HLSW02]. clauses [Mad06]. claw [BM02b, HLS03]. claw-free [HLS03]. Clean [GMNP09]. clickstream [AMT09]. clients [LY07]. Clique [GS00, AH08, BV09, KHN05, LC09a, LS07b, Vas09]. clique-covering [BV09]. clique-Helly [LS07b]. clique-transversal [LC09a]. Cliques [CCF04, Gav00, Yus06]. clock [BHMR01, BD00, BW01]. clock-based [BHMR01]. clones [BRSV05]. close [AV07]. Closed [BW01, Hon05, LMC^{+}09, RSGK08]. closed-2PM [RSGK08]. closed-form [LMC^{+}09]. closer [RS06]. closest [Ric02, WCH08]. Closure [ANR05, HRV00, KK02, Kun00, VL04, Gou01]. Clouds [VS01]. cluster [HR06b]. clustered [EQ04, KLK04, TLZ05]. clustering [AMT09, BDKW07, GM07, Gus02, HS00, HR06b, LC01, OK04, SKA05, Tan08]. clusterings [BDKW07]. clusters [Tan01b]. CNF [AIM^{+}03, BB09]. CNFs [HI03]. CNN [IY04]. co [BRSV05, BM02b, Juk05, Wir03, BLV03]. co-claw [BM02b]. co-clones [BRSV05]. co-design [Wir03]. co-NP [Juk05]. co-P- [BLV03]. Coarse [GS05a]. Coarse-Grained [GS05a]. cocolorings [FKN02]. Code [dH01, Abd09, KS02b, Lyo02, Pec06, Ces02]. codes [Bar09, CKK02b, Gro06, HP00a, HL04, HL07, JKŠ05, Juk09, LS09a, LWC01, MA06, NB05, Neb07, Pes08, Sav04, SL06a, Suo07, Vaj08, XUT00]. Coding [LM02, Gag07, Pol04]. coefficients [Shp07]. cographs [Nik00]. coherence [Fre05]. coin [BHW03, KN04a]. coincides [Dar05]. collapse [Lib03, ZŽ09]. collapsing [CW04]. collecting [Gut08, ABL08, FFFdP07]. collection [KdV01]. collective [JX09]. collision [BBMR08, Her04]. Colony [AF08, CA08]. color [MRWW08]. colorability [Kri02, Loz05, Rot03]. colorable [CR09, LCW07, Maf09]. coloration [HHT05]. colored [GDN07]. coloring [Alo03, BV08, BEP09, CCF04, CP05a, CRW07, FGR02, FGR03, FR08, FKŻ08, GRW06, HK09, HKKL09, Jia06a, KKSŠ08, MR06, NGI08, NDG02, NP00, Sku02a, Sku06, Smo08, TJH07, XL00, Yan08, EMP06]. colorings [CF02, EO07, HJS03, JMV06, LHL08, Och04, OP08, PS06, Sku04]. colors [FR08]. colouring [DSSV00, DOS08, Sid07, Sub06]. colourings [FHN08]. column [BS00a, CG04b]. column-restricted [BS00a]. Combinatorial [DF03, CdL02, Cha03, DFG^{+}03, ET03, KZ06, Mia03]. combinators [Dil02]. combined [BH08b, Ohl00]. commands [ST00]. comment [Woe09]. Comments [GZN06, Par11, WW09]. commerce [BHM06]. commit [DGP04, RK03]. Common [CIPR01, HS05, LSZU03, Stu09, AFT08, AGS07, AYTH08, FB04, GM07, GH05, HYT^{+}07, KV08b, NKS09, Ric00, Sak06, Tsa03, YAM04, YHC05]. communicating [MVM01]. communication [BHMR01, COP00, DW07, Gag08a, HSZZ06, Juk05, LH02, Las09, LPM05, SY08, YWW09]. communication-free [Las09]. communication-induced [BHMR01]. communications [HWW02]. commutative [Lit03b]. Commutativity [YCFD07, MRG05]. commute [AL02]. commuting [SL06b]. Compact [Dur04, HRC08, DGG03, Lai07].

Comparative [SKP^{+}02]. compare [QW06]. Comparison [KS02a, Lyo02, SHC08, ZLR^{+}05]. comparisons [GS03]. competitive [Bun04, DD08, GIS04, GM08b, LM04b, ŠŽ04]. competitiveness [Geo08]. complementary [Bra07]. Complementing [MSS06, AOS^{+}09]. complements [Chu08]. complete [And09, Ces02, CK07, Fu08a, Gri01, GHM07, Ham07, HSTH09, MV08, MS04, PW07, Pol03, TW06, TJH09, ZLM09]. complete/star [CY00]. completed [Fec04]. Completeness [NG05, Agr02, LL02a, Liv09, Sze08, TA08]. completion [ADMP01, BJK02, HLY09, LCLP09, LLPZ07, NCY02, NCBJ02, SL06b, BdFdS06, FMT09]. completions [RST08]. complex [GL04b, WS06]. Complexities [RK05a]. Complexity [BS02, DPS06, DOS08, JCP08, KKSŠ08, KV03, MK03, NSW02a, PC00, ÀBD^{+}04, Amb00, Aro02, Asa07, BGHP09, BHY09, BMTV09, Bol05, Bol08b, BR03, CGK08, CT00b, CŠ05, CS03b, DH06, DW07, DT09, DS04b, DH08, Efr08, EMP06, EHK04, FY03, FSW03, FM03, FHK06, GP03a, GP03b, GVFS00, Goo09, Gro06, GCC00, HSZZ06, IS09, Juk05, Kat04, KMRR06, KM07, Kou06, Lee09, Lib06, Lon03, Mad05, Mar02, Mar01, MM07b, Mas08, MMV08, May02, MW03, Miy08, Mon08, NP07, Oku09, Pro08, RS07b, Ram06, RB02, Rot03, San01b, Sch00a, Sch02a, SS02, SJ08, Shi00, Shi02, Shp00, SW05b, Sta05, Tat06, Ver07, Wol06]. component [HRWZ07, Min07b, Par05]. component-based [Min07b]. components [ACKM07, CC06b, Gab00, Weg02]. composability [FVC03]. composed [BdW03]. Composition [LH08, BP01]. compressed [Car02, CY01b, CY02, IS09, MS04, Neb07, Rao02]. compressible [DKV09]. Compressing [Gag06a, Kas04a]. compression [FG06, KMCC04, LL02b, LD04, Vu05, WYL01a, WYL01b]. Compror [LL02b]. Computability [Lut04, Eng07, HKL01]. computable [FL05b]. computably [Dav04]. Computation [CT03a, AB09, DD08, JNK^{+}02, KRD08, LJ03, LY07, Mar01, MPSS02, Mis01, NPW01, NY04b, Ric00, Tan01a, Tan01b, WLP07, PBD^{+}02]. Computational [GCC00, Tat06, ARdFPS02, BA03b, CL06, CŠ05, CT05, CT07, GD00, NP07]. computations [BM02a, BH08a, RSV07]. compute [DDR04, LR05]. computer [Min07a, XUT01]. Computing [AD06, BFB09, BMCK04, BF03, Blä03, CI08, EK00, EKF00, Elm09, Lec04, Maa07, MOS04, iO09, SW05a, AM05, AYTH08, BB01, BS09, BZ02, CHL^{+}08, CT05, CT07, DD06, The02, ES02, GM02, GRW06, KY03b, KL01, KCYL06, KM07, Kou06, Kun00, LSL08, LY07, MNRR09, MRST09, Nag05a, Rug03, Sak06, Sch00b, Tak00, UPGM09, Weg02, Wol06, YHC05]. concast [Sti04]. concatenation [AK09b]. concentration [Pes00]. concept [Eis09, Yan00]. concept-based [Yan00]. Concepts [Jon03a]. conceptual [Lin08a]. concerning [AK09b, Ric02]. concurrency [KM03a, KJ04]. concurrent [CA08, Fec04, LHS01, LMST04]. condition [BV09, DPR00, GLS08, IM06, Li06, MHK07, RC06, SW07a, ZLZ09]. condition-based [IM06]. Conditional [Fu08a, HSTH09, BCD^{+}05, BP01, CFM09, CWTHR01, LMM05, MM07b, Oku09, STH07, SCHT08]. conditions [BCD^{+}05, FVC03, Koi02, LJ09, Min07a, NSW02a, Ser03]. Condorcet [Pro08]. cones [BE05, vdZ01]. conference [CCKL02, Chu04]. configurations [WCK04]. conflict [LHS01]. confluence [Jac03b, Sim04]. Confusion [Mos08b]. congestion [BS00a, LH03]. congruence [EG07]. conjecture [AK09b, SK08a, Val04a]. conjunctive [CK03, Okh03]. connected [Fuj01a, Fuj01b, FD04, HTT03, HRWZ07, LM08c, LCM09, MP04, iNM07, TTH04, TH09, Weg02, ZNI03]. connectedness [Khu99, Khu00]. connection [FJ08b, GHZ02]. connections [YSK03]. connectives [CWTHR01]. Connectivity [Sha09a, CLM09, Fan02, HS00, HRV04, HV06a, HSTH09, MLX08, MTXL09, NII01, SCHT08, TZ08, Tsi07, Vol07, XLMH05, XXZ05, CL00a]. conscious [SSYM04]. Consecutive [HL08, HP00b, HG02]. consecutive- [HP00b]. consensus [Att02, BDKW07, FMR04, IR09, IM06, KR03a, WTC05, ZC09b, ZC09c, MRT00]. consequences [CS03a]. conservative [GH00b, Jai09a]. consideration [BLME02]. considered [Heh01]. considering [TMSO08]. consistency [CKS09, CDFG01, FVC03, SG05]. conspiratorial [RCA07]. Constant [KL04b, MR09a, Bas08, BA03a, HY09b, Kim05, LMZ08, MM07a, PG09, PP05b]. constant-factor [BA03a]. Constrained [PT00, BME02, BD08, CLC09, CDF^{+}04, CLT07, DLP04, HM08, IR08b, Kwo02a, LCKL06, PT06, Rem04, SLT08, Tsa03]. constraint [Bre04, Cha02a, CC05a, Dal09, ELS06, Hås00a, Koi06, LMG07, Lev07, Lib03, WC01, WL03, DPR00]. constraint-based [Cha02a]. constraints [BD00, BHM08, CFM09, CL03, Eps00, EN07, Fis02, HK00, HC08b, KG04, LPM05, NTT00, RS06, Sey01, Tal00, TH02, Wan07, ZŽ09]. constructing [BDKW07, KTK02, Pio02, TH09, Wan01]. Construction [DL06, GS05b, AS07b, BDLP08, CK07, CPX06, CZ09, GdPL08, Gol04, HP00b, JLN02, KK07, LR06, MC02a, NKSK09, Pec06, Pes08, RRB00, SAOKM01, UN07, WNR00, ZSN02]. constructions [CFW01, LCC06, vDKST06]. Constructive [Sta05, Hit03, May02, SM03, Ves06, Völ04, DPR00]. constructor [Luc04b, MOY06]. consumption [CKK03a]. containing [GH00a, HP00a, Ram07]. content [BdATdC04, LZS04]. content-based [BdATdC04]. contention [JLO01]. contention-less [JLO01]. Context [Deo05, BEF^{+}00, DPS02, EN07, FM03, Fuj05, LS06, MK02, NTT00, Ram07]. context-free [BEF^{+}00, DPS02, Fuj05, LS06]. context-freeness [Ram07]. continuations [BDcS05]. continued [Eis01]. continuous [JXL04, LKP01, Lon08]. contrast [CDFM05]. control [BEL08, CC02, IK07, KM03a, PR05, DFM04, Tri04, YC09]. controllability [UW03]. conventions [Pen02]. Convergence [ZH07, Gou01, Gut02, JLY07, JX09, RK09, Tre03]. convergence-related [RK09]. convergent [Sim04]. converging [Sol08]. convertible [Sae00]. convex [ASW05, AAK09, BB01, CW09, CW02, CW03, CSW03, FL05b, Hu07, KSY00, LR08, MKGB08, Tan00, Wan08, AMS08, Raz07]. Convexifying [CKM^{+}01]. Cook [CN00]. cooperating [BH08b]. coordinate [Sma01]. coordinates [HT00]. coordinator [VB08]. Corasick [DL06]. core [JR07, SG03]. corner [DMN09, DBLS06]. corona [LC04]. correct [JN07, LPW00, Sin03]. correcting [Gro06, LWC01, Pes08]. correctness [EHRS09, LYP02, LC03c, Weg02]. correlation [GM09, JJ02, Sar00, Tan08, WWGF05]. Correspondence [HHHK08, ADM04, Rah08]. corresponding [Lif03]. corridor [DMN09, DBLS06]. Corrigendum [Cha02b, CT07, LPWZ14, RRW00]. cosine [CY02]. cost [ABBG06, BS08a, Che06b, IMS00, Kol05, LTT06, NII01, NK07, SWCC04, WYL01a, WYL01b, ZC09b]. cost-sharing [ABBG06]. costs [Dam02, GN00, SY08, SB06]. count [Lin07, SM03]. countability [Tao06]. counterexample [AIM08, AK09b]. counterfeit [BHW03]. Counting [LC09b, NY04a, DH08, FMSS01, KKM00, LC08, Lin02, Lin08b, Rah08, Yus06]. cover [BDL09, CG05, CG04b, CKW09, DL08, DS04a, FD04, GS09a, LH08, Mes06, NO07, PS09, Mie08]. coverability [KK07]. Coverage [CK08, FJ08a, Gør08, SW09]. coverage-evaluating [FJ08a]. Covered [ABM09]. Covering [BS00b, GNU08, SD09, AFK08, BV09, DGHS07, KTK02, LCL06, LWŻ07, LPWZ14, PN01a, PN01b, TZ05]. Coverings [BV00a, Blä03]. covers [CL00b, CHLN07, Lec04, Lin07, LC08, LC09b]. cow [JŁ09]. cpo [FMCW02]. CPS [DN05, IN06]. CPS-translations [IN06]. criterion [QPV05]. critical [ZS09]. cross [Daw03]. cross-perfect [Daw03]. crossed [DYZ08, FLJ05, XML06, YLTH03]. crossing [Ber00]. crossings [AMS08, NY04a]. Crowds [SWCC04]. Cryptanalysis [CWJT01, GS09b, Hwa00, Joy03, LC03a, MS03, Sha03a, SHH07, ZC09a, CKN06, CKL^{+}03, Pha04]. cryptanalytic [Gol01]. Cryptographic [DFM04, QPV05, RBF08]. cryptography [BDN00, CDFM05]. cryptosystem [Pae03, VS01, hY08]. cryptosystems [GH08]. CSP [DFG^{+}03]. CT [ZC09a]. CT-RSA [ZC09a]. CTL [Lan08]. cube [AG03, DYZ08, YCML06]. cubes [Fan02, FLJ05, HLT07, KMCC04, LT08b, LT08a, LTHS01, MLX08, MTXL09, PCW09, WMG00, WMX07, XX05b, XXZ05, XML06, XX07, YLTH03, Yan09]. cubic [FH06, LL02a]. cubicity [CMO05, CDS09]. Cuckoo [DM03]. curse [CN03, Pes00]. curve [CS09b, GS04a, HT00, Sma01]. curves [Duq07, HRC08, LMVV05]. customer [Nag09]. cut [FK00b, HR06a]. cuts [ACN03, Cha04b, FY03, Nag05a]. Cycle [CC05b, Niv04, ELR07, Fan02, Gal03, Ham07, Hsi08, Irv07, KFS01, LR05, Tam00, WC08, jXyL09, YLTH03, LH03]. cycle-embedding [Fan02, jXyL09, YLTH03]. cycle-rooted [KFS01]. Cycles [Che07, ML09, Tsa07, ČKRR09, CHZL09, CRW07, Che09a, EG03, EKF00, Fu08b, HM09, HKC09, Iba09, JKŠ05, Jha03, KMU05, KV05, LT08b, LL09, LHL08, LCW07, Mar04, MHK07, MPSS02, RK05b, RR09, WWW08, WCFC08, ZS08]. Cyclic [Lin08b, WMG00, Lin02]. Cyclic-cubes [WMG00]. cyclicity [GK00].

D [BCP09, Che00a, CY02, DLP04, LMMM04, LR08, MB08, Rom09]. dag [ANR05]. dangerous [PK02]. Darts [Cur04]. Data [HL03, Mad05, AMT09, AFK08, BLMM06, Bou02, BA03b, CWC00, Cho02, DDD08, EM08, GRH09, GT09, GL04a, KZ06, LA01, LC01, LHK02, LK02, LL02b, LH04, OK04, PC00, PHK01, SKA05, Seg08, Wan07, WWGF05, DK08]. data-accumulating [BA03b]. Database [Cho02, BEL08, CLW02, GCC00, Van01, VL04]. databases [GR03, KPL04, LKL^{+}00, SL01, WC03]. datasets [TLZ05]. datatypes [MG01]. dates [LC02]. DAWGs [AL02]. Day [CL00a]. Day-Tripathi [CL00a]. DBM [ZLZ05]. DBM-based [ZLZ05]. Deadline [SB02, Fun08, KC03b]. Deadline-based [SB02]. deadlines [HGYZ09]. deadlock [CB01, Min07b, Šte00]. deadlock-detection [Min07b]. deadlock-free [CB01, Šte00]. deap [Jun05]. deap* [Jun05]. decentralised [DOS08]. decentralized [Tri04]. Decidability [CDFG01, Frö05, HK04]. decidable [BHK04, Cap06, CG06, EM06, PR05, Rah08, ŢM05]. Deciding [GL05, Kri02, ÀBD^{+}04]. decipherable [Sav04]. Decision [FL09, RRV07, ANR05, BHW06, HI03, HTT03, MOY06, PK02, Shi02]. Declustering [IKKY00]. decodable [SL06a]. decoding [CKK02b, Gro06, LWC01]. decomposition [Cha08, GHL02, Tsi04]. decompositions [BD08, NY09, vDKST06]. decreasing [JC07, LY07, WX05]. decrypt [Bih02]. decryption [GH08, Mil03, Shp04]. deduction [Del06, HR03]. default [Lib06]. deficiency [BZ02]. Definability [CWTHR01]. definable [CKL01]. definition [Kur08, LHS01, SG05]. degenerate [LCL08]. degree [BCG09, BHLV09, CR02, CS09b, ELO08, FR08, GS05b, HY09b, HSWC09, IP09, Kim05, Lib03, Liu05, MHK07, QPV05, Sea04, Shi00, Sku02a, UN07, Val04a, WC09]. degree-bounded [IP09]. degrees [DG08, GMM^{+}07, KY00]. Delaunay [ZSN02, ZJ05]. delay [KJ04, NSZ00]. delay-insensitive [KJ04]. delayed [BN07, CV03]. deletion [DC07, OB03]. delimited [BDcS05]. delivery [Jia05, LC02]. demand [Che04a, GN06]. deniable [Zha06]. denoising [JCSX09]. dense [Elm09]. densest [LMZ08]. density [HC08b, IK07, Kim03, MB08, Sch06, SLT08, Wu09a]. dependable [LLK08]. dependencies [CFM09, GR03, GY08, HHLS03, Lon08]. dependent [FL02, KY07, MO06, SB06]. deployment [DEPZ09]. Depth [JVA06, ZM03, AM05, Gab00, HR05, Mon01, Rez05, She07, ST04a, Tsi02, Wol06, RRV07]. depth- [Wol06]. Depth-First [ZM03, Gab00, HR05, Tsi02]. depths [Has00b]. derangements [KL04b]. Derek [HPW08]. Derivation [HS01, Mis03]. DES-encrypted [Bih02]. Descendants [Vág06]. descent [BtN05]. Description [AZ08, CDFG01, HZ03]. Descriptional [MM07b, Mas08, FM03, Oku09]. design [CCKL02, Chu04, Cur04, DM00c, JR04, Li09, MY09, MC02b, NY09, Sin03, Wir03]. designated [LV07, Shi08]. designs [CHC02, GO05, KTK02]. desirability [MM01]. Desmedt [CHH^{+}09]. detect [BHW03, CK03]. Detecting [EG03, ST00, HP00a, Pes08]. detecting/correcting [Pes08]. detection [GS05a, Her04, KG05, Min07b, Niv04, ST00]. detector [FMR04]. detector-based [FMR04]. detectors [FMR05, JAF06, RRT08]. deteriorating [BJK02, NCBJ02, WL03]. deterioration [WX05]. determinant [DT09, LR06, Pud00]. Determination [Gol03]. Determining [RC06]. determinism [UN07, VM01]. Deterministic [COP00, GTM03, GM08a, HM04, YHT04, COW05, CC07, EM06, FKV04, GL02, GRT00, Löd01, MSS06, NN05, WS07, XP06]. Deterministically [BHLV09]. determinization [LW09, Tri06]. developed [Mai03]. development [HŽR08]. devices [Pet07]. DFA [AZ08]. Diagnosability [ZLR^{+}05, SHC08, yZyLZ09]. diagnosable [YSK03]. diagnosis [GNAJT09, ZLR^{+}05]. diagonal [BP06, DIPS00, Jha03]. diagonal-flip [BP06]. diagonal/toroidal [Jha03]. diagram [Aro02, Asa07, BMO08, CPX06, JMM09, KS02a, Kol04, Tan01b, AHS^{+}03]. diagrams [ADK06, BHW06, CFS09, CKCB02, HI03, WCK04]. diameter [BCD^{+}05, BŽ06, BS06, CW09, XXH05, XY07]. diamonds [CCGP05]. Dicing [Hor07]. dictionaries [RR01]. dictionary [BV00b, MN05]. difference [CL00b, Fis02]. different [CKP00, SW05b]. Differential [CKL^{+}03, Mon02, CTN03, Nag09, Pha04]. difficult [Jac03a]. Diffie [Tsa06]. DiffServ [AKAE04]. diffusion [KM02]. Digit [LKY02]. Digit-serial-in-serial-out [LKY02]. Digital [Kwo02b, Kwo03, WHLH03]. digraph [BR03, DV09, TKS07]. digraphs [Ara07, BLME02, CLM09, HV06a, KFS01, KS03b, LM08c, MP04, NII01, Nag05a, SCK07, SDC09, SK08b, Vol07, WL08, ZLM09]. Dijkstra [Mis01]. dilation [Tse05]. dimension [AKL07, CKL01, DG08, EA07, ERS05, Hit03, HP05, KMF04, May02, Sta05, WWGF05]. dimensional [AB09, AK08, BK00, CGK08, DYZ08, Eis09, Gia09, JŁ09, LHK02, LK02, MNR04, NKSK09, SSYM04]. dimensionality [CN03, Pes00]. dimensioning [LZS04]. dimensions [ABPR06, AH08, BE05, EK00, Vah07]. dipath [MW06]. Direct [Kur08, Bol08b, BS02, CZ09, Ham07, JKŠ05, Luc04a]. directed [AF08, BFK09, Ber00, DF03, EG03, Gol03, HR06a, HJR02, LW06, LR05, RS07b, VGG^{+}00, ZM03]. direction [BV00a, DLP04, SB06]. direction-dependent [SB06]. disc [AH08]. discipline [DM00c]. discount [Dam02]. discounting [DSV08]. discovering [CA08]. Discovery [ZM03, KMT04]. discrepancy [MB08, SSZZ00]. discrete [BMCK04, Chu00, CY02, Sch01]. discretization [SB06]. Discriminative [LSS07]. discussion [AV07]. disguised [EIM02]. Disjoint [ČKRR09, BFK09, CW02, CW03, CSW03, DYZ08, Has00b, HT03, LH08, Tan01b]. disjunction [MST02]. Disjunctive [HR03, NG05]. disk [AKL07, IAIH03, LKP01, NN05, NN08]. disks [KSY00]. disorder [BYSP02]. Dispersed [CCZT08]. Distance [AA05, AAK09, Aku06, ALM02, BP06, dSCLP08, CWCL02, Cle02, CT03b, CS09a, FGR03, Fre05, GF08, Gus02, HSZZ06, JHfS08, KS06, LSL08, LCC06, LHWL07, Luc04a, LEP07, Pal00, Pal03b, PHK01, RS07c, Tou03, Tsu08]. distance-increasing [LCC06]. distances [BFB09]. distant [DLS08]. distant- [DLS08]. Distinctness [Pet02]. distinguishing [CG09]. Distortion [MP08a]. Distributed [BCG09, CK03, LKL^{+}00, SG03, UW03, WW09, AFG07, BM02a, BH08b, BDLP08, CC01a, CWC00, CHU06, DM00c, Gho07, IM06, JS01, LK00, NC05, QH01, Sas02, Tsi02, Tur07]. distribution [CCKL02, Chu04, DHS09, GDN07, LZS04, LT07, Lon03, MCS^{+}09, Sch00a, Shp04, Sol08, Sri08]. distributions [Gag06a]. distributive [LM02]. distributively [Tsi04]. divergence [ACIL08]. Diversity [LLK08]. divisible [Det09]. divisor [SK08b]. DL [Sch01]. DL-keys [Sch01]. DNA [MAL08]. do [Šte00]. document [CTN03]. documents [MAC03]. does [Der09, LPW00, Stu09]. Domains [FM00, HRC08, Liu05]. domatic [RRSY07]. Dominance [EN07, Sax09]. dominating [DFMR08, Fuj01a, Fuj01b, HN08, Lie08, Suo07, Tur07, TH09]. domination [Ara07, CLZ09, DP06, FGK^{+}09a, Jha03, KS03b, KL04a, KMRR06, Lee09, LC03b, OPR07, Sea04, SDC09, Wes04, XKS06]. Domingo [CKN06]. don’t [CIL^{+}03, LS09b]. don’t-cares [LS09b]. dotted [Jia06a, Yan08]. Double [Bur05, GH08, Yos09]. down [AS07b, FKV04]. downloading [HJ05]. DP [HLY09]. DPDAs [GTM03]. draw [GMNP09, Kur04]. Drawing [BCLO03, LS02, NY04a, Pen02, WCY00]. drawings [DLP04, LE02]. dream [Dij01]. drinking [BBF01]. driven [Bol03]. DTDs [CK02]. dual [FFFdP07, LT08b, ZH06]. dual-cubes [LT08b]. dualization [EIM02, KS03a]. dummy [BLME02]. duplicate [AKE00]. duplicated [Low02]. duplicating [EM08]. duration [Bac07]. durations [AAER03]. during [RB07, SV05]. Dwork [Zha06]. Dynamic [BMO08, Gag07, Gag08a, HYT^{+}07, TZ08, TC04, AB09, BDcS05, Cho08, DC07, DGG03, GM02, IKKY00, KM02, Kow07, LS06, Mel07a, NC05, SG03, SM02, WS06]. dynamical [Sin03]. dynamics [BBF01].

ear [Tsi04]. easier [MY09]. Easy [Del06, KS00]. ECB [Sar08]. ECB-Mix-ECB [Sar08]. economic [FSW03]. economy [CL06]. Edge [Ara03, CHZL09, GK00, KA06, STH07, XDX05, XX05b, Yan09, Alo03, AC02, Bas08, Ber00, CG09, CHC02, FLJ05, FHN08, Fuj01a, Fuj01b, Has00b, HYCH07, IP09, KH03, KS02b, KKSŠ08, LE02, LTTH03, LHL08, LCM09, MZ00, Mar04, MR09a, MR09b, NY04a, iNM07, NPW01, Pre02, RL02, RS07a, Sku02a, Sub06, TZ08, TJH07, Via08, XDZ05, Zel07, ZLJ06, LM02]. edge-biconnectivity [RL02]. Edge-bipancyclicity [KA06, STH07, XDX05]. Edge-choosability [CHZL09]. Edge-chromatic [GK00]. edge-coloring [Alo03]. edge-crossing [Ber00]. edge-disjoint [Has00b]. Edge-fault-tolerant [XDX05, Yan09, KA06, LTTH03]. edge-firing [Pre02]. edge-length [MZ00]. edge-node [LE02]. Edge-pancyclicity [Ara03, XX05b, FLJ05, HYCH07]. edge-weighted [IP09]. edges [AG03, BS08b, CZ02, Che07, DLP04, HKL^{+}04, Nik00, WC08]. Edit [ALM02, LHWL07, Aku06, KS06, Tou03, Tsu08]. editing [Ars08]. Editorial [Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano07n, Ano07o, Ano07p, Ano07q, Ano07r, Ano07s, Ano07t, Ano07u, Ano07v, Ano07w, Ano07x, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano08t, Ano08u, Ano08v, Ano08w, Ano08x, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j]. Editorial [Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q, Ano09r, Ano09s, Ano09t, Ano09u, Ano09v, Tar04]. Editors [AG01, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano03w, Ano03x, Ano03y, Ano03z, Ano03-27, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano04x, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q]. EDZL [CLL08]. effect [JR07, KY06]. effective [DG08, GD04, KHN05, LC01]. effectively [FL05b]. Effects [KM03a, KY07]. efficiency [BS08a, LLH09]. Efficient [Abd09, BP06, BLdR04, CW02, CW03, CKK03b, CLT07, Dil02, DD06, FM06, HYT^{+}08, KPL04, KLK04, KR08, KMT04, Lei05, LL01, LW08, LL09, Löd01, MAC03, iN02, OPR07, RRB00, RY08, Vas09, Wu08, XUT00, ZSN02, ARdAdS07, BCP08, BEF^{+}00, BDLP08, CC03, Che03, CKK02b, CY01b, CKR06, Gam00, GdPL08, GM02, HZ03, IR08b, JCL04, KA02a, KHLL09, Kur03, LHK02, LK02, LTH05, LTT00, Lin02, Low02, MK05, Pal00, PHK01, PCC03, Pet05, Ser01a, Ser01b, SSYM04, TYW01, WS01a, WS01b, YZ08, ZJ05]. Efficiently [ST06, BD08, KKM00, Kuj09]. Ehrenfeucht [Zho08]. Eigen [LT07]. Eigen-distribution [LT07]. Elastic [YC09, Qua02]. election [MR08]. elections [MSZ02, RRB00]. electrical [BCNR02]. electronic [BHM06, OF03, RRB00]. Element [Pet02]. Elementary [Gut08, AD07, MM05, Yos09]. Elements [Gou01, Fu08b, Hsi08, KP09, MRST09, Pec07]. elephants [VLK06]. eliminating [Amt08]. elimination [PD09, SS03b]. elliptic [CS09b, Duq07, GS04a, HT00, LMVV05, Sma01]. embedded [SKP^{+}02]. Embedding [DYZ08, FL05a, LT08a, LH03, LTH08, AGGZP09, CH02, CS09b, Fan02, Hsi08, HHLH03, LT08b, LW06, ML09, Tsa07, jXyL09, YLTH03, ZH06]. embeddings [BKS01, HS03, MP08a]. emitting [BtN05]. empirical [YCKK07]. employing [CCKL02, Chu04]. emptiness [JS07]. empty [DMN09, DBLS06]. ENA [KKL03]. encapsulation [CHH^{+}09]. Enclosing [dFdF09, DGN05, KRD08, NN05, NN08]. enclosures [MP08b]. encode [BM02a]. encoded [AYTH08, ALM02, FB04, LHWL07]. encoding [Dia09, SV08, Via08]. encountering [UW03]. encrypted [Bih02, Hes04]. Encryption [Pha04, HWW02, KHL09, Küh08, LV07, RG09, Shp04]. end [EL08, YWW09]. end-to-end [YWW09]. endomorphisms [Ric03, SL06b]. energy [CKK03a, Mar01]. Enforcing [hY09a]. engineering [Par03]. enhance [KKL03]. Enhanced [Ver06, Che04a]. enough [FR08]. Ensuring [RSV07, BHMR01, Nun04]. entailment [VFV02]. enter [CH03b]. Entity [HHLS03]. Entity-Relationship [HHLS03]. Entropy [Shi02]. enumerable [Dav04, Pol04]. Enumerating [BS00c, Hag03, Ver07]. Enumeration [SS00, God02, MRN02, RT00, SV09]. enumerators [KL01]. environment [Cho02, LLPZ07, Lop00]. environments [KY03b, KM03a, Lin08a]. episodes [CGM06]. EQP [GP01]. equal [GM08b, KP09, LH08]. equal-length [GM08b]. Equality [Dar05, HHL05, NNS01, Hon02, Kat04, Wan07]. equally [SJ08]. equally-spaced [SJ08]. equals [Sta05]. equational [ACIL08, BB05]. equations [BNvO09, HM02, NR08]. equilibria [CŠ05, DH06, FSW03, FHK06, Ves06]. equilibrium [BDQ08, DD08]. Equivalence [PU01, BF01, BCP09, CDP02, DM08a, EM06, Hag07, Nag05b, NRW09]. equivalences [ARdFPS02, Fec04]. Equivalency [Li00]. equivalent [FR04, ZLJ06]. ergodic [GRT00]. Erlang [HM02]. Errata [JLW09]. Erratum [AS02a, CW03, Chu04, GP03a, Kwo03]. Error [HP00a, NEK03, BV00b, BdW03, Gro06, LK00, LWC01, MN05, Pes08, TZ05]. error-correcting [Gro06]. Error-detecting [HP00a]. errors [KM03a, WYL06]. escalation [CLW02]. escape [TA08]. estate [Lei01]. Estimating [AD00, CPC03, HM03, PC00, RB07]. estimation [MP09, WWGF05]. estimator [Giu04]. eta [IN03]. eta-expansions [IN03]. Euclidean [ASW05, BZLN07, BD04, EK00, Hes07, KPL04, LM08a, WL02]. Euler [Aku06]. European [ST06]. European-Asian [ST06]. evaluating [FJ08a]. evaluation [AHV03, DR06, SKP^{+}02, Yan00, Zea00]. evaluators [ADM04]. evasion [KR05]. even [Bih02, MRST09]. even-ranked [MRST09]. Event [BCP05, CWTHR01, FMCW02]. Every [HM09, KV08a]. evolution [GPS01, LR03a]. Evolutionary [DSSV00, ES02, JXL04, PK02, RL02, Alb02, CP02, JTG04, MM07a, SM02, ZH07]. evolving [GYY01, SSSS02]. Exact [CT00a, Rot03, TLZ05, WHZ04, FH06, Lec07, Lut04, Pap09, RRSY07, FGK^{+}09a, Mad06, Rot03]. Exact-Four-Colorability [Rot03]. exactly [Mai03, iO09]. examples [Weg00]. exchange [BC07, CWJT01, DCS07, Irv07, Mad05, MS03, YPKL08, BD05]. exchanged [ML09]. exclusion [Ala05, AIM^{+}03, CC01a, DL00, LPW00, Pet05, TAI04, TMSO08, Vid03]. Execution [AS07b, DGG03]. executions [RCA07]. exercise [EG07]. Exhaustive [LMMM04]. exhumation [Deo05]. exist [Sop02]. existence [JLN02, MSSW08, RH02]. existential [Bar01a]. exit [AFI03]. expansion [KAS08]. expansions [IN03]. expected [Ars08, Bac07, Kri02, Mar02, RRV07]. experimental [AK02]. explicit [MW03]. Exploiting [Fre05]. exploration [BBMR08]. explosion [LS07a]. exponent [GD05, Shp04]. Exponential [CKW09, Seg05, Bol03, BEP09, Lan08, ST02]. Exponential-time [CKW09]. exponentials [BS09]. exponentiation [CHK^{+}09, Kle08, KR08, LY00]. expressibility [AI08, tC09]. expression [CLT07, Jon03a, KSS00, Lif03, SKA05]. expressions [CY04, RZ00, tC09]. expressive [LTV04]. expressiveness [AFI03]. EXPSPACE [Las09]. exptime [Cap06]. extendable [YZ03]. extended [Gol04, JNK^{+}02, MN08, QPV05]. Extending [dSCLP08]. extension [CKKC04, Lib06, ST04a, ŢM01]. extensions [HLS03]. extent [BDcS05]. extents [KLK04]. External [HR05, Sib04, IAIH03, LTT00]. External-memory [HR05]. Extracting [CK02]. extraction [MAC03, Rom09]. extractors [TS02]. extremal [BD09]. extreme [JLN02]. extremely [AK08].

f [Mon01]. f-depth [Mon01]. facial [MR06]. facilitate [Lon08]. facility [AB09, XX05a]. factor [BA03a, CdL02, JQQ^{+}03, KW05, LL02b, NEK03, Pan05, CI08]. factor-3 [JQQ^{+}03]. Factoring [Dob04, MCS^{+}09, Sha03b]. Factorization [CF09, KFS01, BVG02, TKS07]. factors [MNRR09, Sed07, Sri08]. fail [PSL^{+}01]. fails [Bar01a, LSS07]. failure [FMR04, FMR05, Gue01, JAF06, RRT08]. failure-sensitive [Gue01]. failures [KS09, ST09, Tsa07]. Fair [ABBG06, VLK06, HC01, MS03, ZC09c]. fair-lossy [ZC09c]. false [BGK^{+}08]. false-positive [BGK^{+}08]. families [CP05a, CIL^{+}03, Elm09, Mon06, SSZZ00]. family [DYZ08, KH03, LT08a]. far [AD06]. Farey [MGLA06]. Farmer [Vea00]. Farthest [ADK06, BD09, MCL06, DM08b]. Fast [AHM08, BN04, CY02, DLX09, DGP04, KRD08, Lec07, Lin07, Seg08, Tan01a, Tsa06, WWGF05, AB09, AK08, AYTH08, BS09, CIPR01, FG06, Her04, HT00, HTT03, IMS00, Joh04, JJ02, Jun05, KLP04, LY00, NN05, Ric00, RBDN09, Tak00, YHC05, ALO^{+}09, CC01b, VB08]. Faster [DP01, GKL09, IR08a, JLW09, LS07b, Tsu08, EG03, KC03b, Kou05, NPW01, NSW02b, SW05a]. fat [CDKK02]. fatness [dBvdS02]. Fault [BŽ06, CC01b, Fu08b, GH00a, Shi03a, TTLH02, WMX07, XXH05, XY07, XZX07, YLTH03, BMP^{+}00, CQ09, CHC02, Fu08a, HP00a, HSTH09, HZ09, HKC09, KA06, Lat07, LTTH03, LLK08, MRN02, NRS00, Shi03b, WC08, XDX05, jXyL09, Yan09]. Fault-containing [GH00a, HP00a]. Fault-diameter [BŽ06]. Fault-free [Fu08b, HKC09, WC08]. Fault-tolerant [CC01b, Shi03a, TTLH02, WMX07, XZX07, YLTH03, BMP^{+}00, CHC02, Shi03b, jXyL09]. faults [MNR04, PP05a, RB07, Shi03b, SCHT08]. faulty [Che07, CH02, Fu08b, Hsi08, HKC09, HHLH03, LT08b, STH07, WC08]. FCFS [Pet05]. FDs [Lec04]. feasibility [KLK06]. feasible [YZ08]. feature [GNAJT09, Hes07]. Feedback [CCGP05, FLP00, GHM07, Gup08, WWC04, CKK02a, CLL04, DF03, FGR02, Gav08, IN04, KAS08, KR03b, KHLL09, Sas08]. Fermat [AAK09, LP09]. Ferrer [CKN06]. few [KMU05, Mar04, SW08]. FewP [RRW00]. FFS [ÀBD^{+}04]. FFT [NP07]. Fibonacci [Joh04, Tak00]. fields [CG08, KCYL06, LC06]. Fighting [CN03]. filament [ES07]. filaments [Gav00]. file [ALO^{+}09]. files [ALO^{+}09, Che02, IKKY00]. filling [HRC08]. filtering [BdATdC04, MK05]. filterings [GVFS00]. filters [BGK^{+}08, ABPH07]. find [BFK09, Iba09, MSSW08, WS07]. Finding [DB06, FK00a, FGLW04, HLS03, HC08b, IK05, Kim05, KKM00, KN04b, Lie08, NKS09, Pol03, TH03, Wil09, WCFC08, YAM04, Yus06, BGHP09, Chu00, DBLS06, FY03, FHK06, Gal03, HLP03, HIM09, HYT^{+}08, IP09, Kim03, KOH09, LSC06, Pan05, RB02, SW08, SG03, SLT08, Tsi04, VGG^{+}00, Win04, YCKK07]. finite [BB05, CG08, DJ07, DPS02, DV04, Fis02, GTM03, Göb00, GRT00, Hie04, JS07, Kir02, KCYL06, Las09, Lev07, LC06, Lit03a, MVM01, To09, Ver07, Zho08]. finite-state [Las09]. finitely [AHM08, Ric03]. finiteness [Dia08]. Firing [LM02, Pre02]. First [ZM03, CDC04, Gab00, HR05, Hes03, NP00, Tal00, TLH02, Tsi02, Smo08]. first-fit [NP00, TLH02, Smo08]. first-order [CDC04, Tal00]. fit [NP00, TLH02, Smo08]. Five [NP07, FR08]. Five-step [NP07]. fix [BH08a]. fix-point [BH08a]. fixed [AKL07, Aud01, BZ02, CS09a, ER03, GN06, GNU08, Hag07, HM02, LTH08, MR09a, Mis01, NR00, RS07a, Shp04, Ste08, Wes04]. fixed-parameter [CS09a, GN06, GNU08, Hag07, MR09a, Ste08, Wes04]. fixed-parameter-tractable [NR00]. fixed-point [HM02, Mis01]. fixpoint [LS06, LM04a]. flashlights [LPC02]. flat [Jac03b]. Flaw [OF03, BHM06]. Flaws [DHS09]. flexible [JCL04, LLPZ07]. flight [Her04]. flip [BP06]. flipping [KN04a]. floodlight [DHS08]. floor [Kur03]. floor-planning [Kur03]. Flow [ZLJ06, BBD02, BLMM06, Bre04, Bun04, GN06, IMS00, KC03a, Kol05, Kol03, KW01a, KW01b, MC06, WYL01a, WYL01b, XL00]. Flows [Mel07a, PZ09]. flowshop [CL03, WC01, ZSZY08]. FLP [Völ04]. Flying [ZZ08]. foci [vdZ01]. fold [EA07, Eis09]. folded [CQ09, Che09c, Fu08b, Hsi08, HKC09]. Folk [Tri06]. Fooling [Sze08]. forbidden [PĆB04]. force [Ber00]. force-directed [Ber00]. Ford [ARdAdS07, Pec07]. forest [PS04a, Cha05]. Foreword [FMT03]. forgery [Wan04]. Fork [CKS09]. form [BW01, Efr08, HG06, LMC^{+}09, Yos09]. Formal [LZ04, Lav08, Rao01, vDR09]. Formality [Hoo01]. formally [Heh01]. format [CY01b, CY02, ME04, MRG05]. formation [DP07]. forms [Hag07, Wan07]. formula [Fei01]. formulas [AIM^{+}03, Amb01, BZ02]. formulation [GK04]. Forward [BDQ08]. forwarding [CQ09, XX07]. foundation [TJC^{+}06]. foundedness [CK07]. Four [Rot03, GRW06, LCW07]. four-coloring [GRW06]. Fourier [CC01b, CY01b, Høy00, Shp07]. FPT [Sas08]. FPTAS [ESZ02, JC07, JC08]. Fractal [KMF04, WWGF05]. Fraction [MGLA06]. fractional [ZS09]. fractions [Eis01]. fragment [MAL08, Tal00, BG01]. fragments [KV03]. Fraïssé [Zho08]. Frames [CVEOV08]. framework [ARdFPS02, HC03b, KR05, Ohl00, ZH02]. Free [GO05, IT03, ACKM07, Att02, BEF^{+}00, BLL02, BLV03, BLdR04, CZ02, CB01, DPS02, Fu08b, Fuj05, HHK08, HLS03, HKC09, KOH09, Kur08, LS06, Las09, Lif03, LH05, LM09, Maf09, MK02, NNP07, Och04, Ric03, Šte00, WC08, WY06, ŠŽ04, tC09]. freedom [Mil03]. freeness [Ram07]. Frequency [EHRS09, CCY^{+}07, Che04a, TLH02]. fugitive [Der09]. Fully [RG09, EL08, LCL08]. fun [Tur03]. function [Bol03, Bol05, Bol08b, DIPS00, LPM05, Saw07, Shp00, Shp07]. functional [ADM04, AHV03, CFM09, HHLS03]. Functions [BW03, BW06, Bol09a, Bra09, CFW01, EIM02, FMSS01, FJ08b, FM05, FA05, GK04, Hon05, HK00, IT03, LC09a, LC06, Mai02, OT03, QPV05, Sar00, SKA01, SM03, Sun01, Wol06]. fundamental [AFG07, Gal03, Saw07]. further [EMP06, HKC09, DM03]. fusion [DM08a]. futures [VM01]. fuzzy [Yan00].

GA [SLL^{+}05]. gains [KLN^{+}03]. Gales [Hit03]. galleries [EHP06]. Gallery [ACHP09]. Galois [AS07a, BKS01]. Game [RW09, ABBG06, Bac07, CV00, Goo09, HJR02, HPW08, JXL04, LT07, Sid07]. games [CH08, CŠ05, KR05, Li09, Pre02, Ser03, Sin03, Zho08, AR06, LM02]. Ganzinger [Ano04y]. gap [Bro05, GHL02, Lan08, ST02]. gaps [Tou03]. garbage [KdV01]. gate [ST02]. gathering [Seg08]. gem [BLV03]. gem-free [BLV03]. gene [SKA05]. general [CN00, Cha02c, FG06, IY06, KS00, KC02, LPSR08, Mor05, NR00, PG06, Sar08, UC00, KZ02]. generalization [HWW02, LTHS01, RS05, Yos09]. generalizations [FS09]. Generalized [Bor00, CH03a, Par11, WLLS08, YELM04, BBF01, Bol09a, CC01a, DM00b, DPS06, FGK^{+}09a, KS03b, KLP04, KCYL06, Par08, SZM09, SCK07, SDC09, SJ05, SLL^{+}07, YCML06, CKR06, CK08, JP09]. Generalizing [AFG07]. generate [YZ08]. generated [AHM08, GVFS00, Ric03, RBF08, SJ08]. Generating [KSS00, LMVV05, MP04, Pal03a, RCA07, ER03, PD09, Sol08, XUT01]. generation [Abd09, DDD08, KAS08, KL04b, iN02, Neb07, Orl09, XUT00]. generative [LSS07]. generator [Gol01, GM09, HL05]. generators [FSGM03, Tsa06]. generic [HG01, Sch01, yZyLZ09]. genetic [ALA06, TC02, AD06]. Genome [JN07]. genomic [SL01]. Gentry [Hes04]. Geodesic [HLT07]. Geometric [AK09a, ABE^{+}09, AK08, ABF^{+}08, dSCLP08, CW09, DS00, DLS08, LW08]. geometry [dSCLP08, Pes00]. Getting [JR07]. GF [KLY00]. Gilbert [SYL04].

Gilbert-model [SYL04]. Girault [Sae03]. girth [Cha04b]. given [DV09, HR06b, JC04, KLP04]. Glivenko [Pal03a]. Global [CLW02, Doe04, Che03, KBEG07, Mai02, Min07b, Sun01]. Go [Kol04, FS08]. Gödelization [Gol00]. goes [Mai03]. Gold [LZ04]. Gold-style [LZ04]. Goldreich [ACI^{+}06]. Golomb [MY09]. gon [CPX06]. good [Mai02, Sun01]. Goodman [CWTHR01]. Goodman-Nguyen-Walker [CWTHR01]. gossip [Fer00]. Gossiping [KS09, GL02]. grade [JC08, Woe09]. gradient [BtN05]. Graham [MU05]. Grained [GS05a]. grammar [BH08b]. grammars [BN01, BEF^{+}00, DPS02, FM03, Fuj05, MM07b, Mas08, MK02, Oku09, RP08]. granularity [Cha02a]. Graph [LD04, ADMP01, BBMR08, BŽ06, Bas08, BFK09, Bol03, Bol08a, BM02b, BLdR04, ČKRR09, CS03a, CZ02, CY00, CDP02, DDR04, DV09, DSSV00, DOS08, FM09, Fu08a, GMNP09, GHJ^{+}08, HKL01, HS00, HSTH09, HM09, yH02, HHP08, KY03a, KV08a, KN04b, Kun00, Lat07, LLP09b, LS02, LR05, Lin07, LC08, LC09b, Ohl00, Sha09a, SW07a, SK08b, TK09, Tsi04, WWW08, Win04, YAM04, ZH06, yZyLZ09]. graph-based [CDP02]. graph-driven [Bol03]. graph-rewriting [TK09]. graphic [PN01a, PN01b]. Graphical [WHO09]. graphics [AK08]. graphs [AKL07, AOS^{+}09, AH08, AA05, AD09, AS02a, AS02b, ABF^{+}08, AC02, AF08, BCD^{+}05, BCG09, Bar01a, BK04, BS06, BJ08, BV08, BM02b, BLL02, BLV03, BV09, BB09, BS02, CHZL09, CH06, CR02, CR09, Cha04a, CMO05, CDS09, Cha02b, CW09, CRW07, CLZ09, CL00a, CF02, Che04b, CC05b, Chu08, CGŻ09, CHLN05, DP06, Dar05, Daw03, DF03, DN06, DLS08, DGHS07, DS06, ES07, EO07, ELO08, FS08, FKL01, FR08, FHN08, FH06, Fou02, Gav00, Gav08, GK00, Göl08, GHNP02, GRW06, Gus02, Ham07, HRV04, HV06a, HR05, HLS03, HJR02, HZ09, HL07, HH09, HHLH03, Iba09, Jia06a, JHfS08, KH03, KA06, KS00, KL04a, KKSŠ08, KK07, KMŻ06, KOH09, Kow07, KHLL09, Kur04, LE02, LC04, LCL08, LL02a, Lee09, LLP09a, LC09a, Li06, LT06, LM08b, LC03b, LMZ08]. graphs [Lie08, LTT06, LS07b, LW05, LL09, LHL08, LCM09, Low06, Loz02, LR03b, LM09, LCW07, Maf09, MTF02, MV08, Mar04, MW06, MR08, Mon05, Mon06, MR06, Mor05, NSI04, NK07, iNM07, NBCS07, NP00, Nik00, NNR04, OPR07, Och04, OHNT02, PC05a, PD03, PD09, Pan05, PS06, RS07b, RR09, RR00, Saw07, Sea04, Sku02a, Sku04, Smo08, Sop02, Sza03, TYP08, TM09, TTH04, TJH07, UC00, Van03, Via08, WWC04, WLC08, WY06, XP06, Xia07, XXH05, XLMH05, XKS06, XY07, jXyL09, YC94, Yan08, Zem08, ZH05, ZLJ06, ZS08, ZLZ09, ZLR^{+}05, ZM03, ZS09, ZLT02, dlVT09, ŠŽ04]. Gray [LS09a, Abd09, Bar09, KS02b, Vaj08, XUT00]. Greedy [CCY^{+}07, IJN09, CKY06, FA05, KS05, KS06, KY03a, Kol03, LM06, LR05, Sea04, KOH09]. Green [SKP^{+}02]. Greibach [Yos09]. grey [BDN00]. grid [CC01a, CL06, FGR03, Fuj03, HR05, WA01, ZH06]. grid-packing [Fuj03]. grids [FRR03, KSRI03, RW09, ST04b, ZH07]. group [CWJT01, Göb00, He02, HWW02, LHC^{+}00, LHC^{+}01, LPM05, Miy08, Pet05, Sae00, SL06b, TAI04, Vid03, Wan04, jXyL09, YSK04, ZC04, BD05, DEPZ09]. group-intersection [Miy08]. group-oriented [CWJT01, LHC^{+}01]. Groupies [dlVT09]. groups [LM08b, Pae03]. growth [Bol09b]. guarantee [DM00c, LPW00]. guaranteed [Gut02]. guarantees [BC00]. Guarding [EHP06, LHT00, Eid02]. guards [Tan07]. guesses [JP09]. guided [Tsu08]. guillotine [ABPR06, HKKL09]. Gusfield [Heu08].

Hadwiger [NBCS07]. half [TM09]. halfspaces [Lon03]. Halin [DS06, MW06]. Hamilton [ČKRR09, Fan02]. Hamilton-connectivity [Fan02]. Hamiltonian [ADMP01, Che09a, FS09, HSTH09, HHP08, Iba09, KH03, LTH08, MHK07, MU05, Par05, PC04, RK05b, Tam00, TTLH02, WC08, YELM04]. Hamiltonians [SAOKM01]. Hamiltonicity [BV09, DPP05, Fu08a, Li06, Par05]. Hamming [AA05, GF08, HSZZ06]. handling [GGK03]. handshakes [Zem08]. hanger [dFdFC04]. Hanoi [BF01, BS06, WW03]. Harald [Ano04y]. Hard [TMSO08, DD08, FJ05, Irv07, Li00, Min07b, Tal00, WYL01a, WYL01b]. hardcore [Sch01]. hardest [FM05, Okh03]. Hardness [DCS07, KL04a, CN00, DP06, DS04a, FR06, Gue01, NCY02, RRV07]. Hardware [Wir03, AK08, WS06]. Hardware/software [Wir03, WS06]. harmonic [ABF^{+}08, BCNR02, CS00]. hash [Bra09, ZHW01]. hashing [Bur05, DM03, HC02, KR06]. Hashiwokakero [And09]. Hausdorff [LSL08, LR08, May02]. Havel [MRWW08]. having [BK04, IN04]. HCPO [ZJ05]. HDT0L [Hon05]. healing [DHS09]. heap [dFdF03]. heap-ordered [dFdF03]. heaps [Bra07, HM01, MTT03, Pio02]. heaviest [BD08]. height [AFT08, Ary02, KK04, LJL07]. Hellman [MH09, Tsa06]. Helly [BUŽ02, DPS05, DPS06, DLPS08, LS07b]. help [Der09, ZXZ08]. Herbrand [Dia04]. hereditary [DLPS08, Frö05, LS07b]. Herman [MM05]. heterogeneous [LKL^{+}00, QH01]. heuristic [CV00, DL08, Kel02, KW01a, KW01b, LM06, LS02, MZ00, Riz03, RBDN09, Sub06]. Heuristics [WC01, CP05b]. hexagonal [HL04, ŠŽ04]. hidden [EHK04]. Hiding [PSL^{+}01]. Hierarchical [Fer00, Ohl02, OK04, SKA05, WCFC08, hY08, hY09a]. hierarchically [CDL^{+}04]. hierarchies [AK09b, BA03a]. hierarchy [AFI03, Fec04, Hwa00, RC06]. high [AH08, BK00, Gam00, Gia09, GS05b, LHK02, LK02]. high-dimensional [BK00, Gia09, LHK02, LK02]. Higher [CFS09, EK00, HHLS03]. Higher-order [CFS09, HHLS03]. Highly [Mai02, CFW01]. highway [AHS^{+}03]. Hilbert [HRC08]. history [DGG03, Frö05, Wu09b]. Hitting [ERS05, Cha03, LJ03, Wot01, Cai09]. hitting-set [Wot01]. hoc [FP08, GL02, WW09, ZC09a]. hold [IT03]. hole [CZ09]. holes [CCF04]. homeomorphic [AGGZP09]. Homogeneous [MK02, AKAE04, HLP03, HS01, LY07, MF07, MA06, NC05, PS02, TYW01, BdFdS06, dFdS05]. homomorphism [CKN06]. homomorphisms [AD07, BS02, Del06]. homophonic [Sav04]. Honeycomb [AD09, CH02, CH03a, YELM04]. hoopla [Cur04]. hop [LM08a, SZQ08, ZXC05]. hops [DM09]. horizon [Lev07]. horizontal [JCP08]. Horn [EIM02, HZ03]. hot [KMT04]. hotlink [KKS04]. house [FSW03]. hub [GHJ^{+}08]. Huffman [CKK02b]. hull [FL05b]. hulls [AMS08]. Hwang [hYScH05]. hybrid [HC08a, LPM05, MM07a, SLL^{+}05]. hyper [NG05]. hyper-resolution [NG05]. hypercube [Che07, FL05a, FS09, GG02, HH07, HYCH07, KS02b, NDG02, PCW08, PC04, Sep05, Shi03a, SCHT08, WC08, WCFC08]. hypercube-like [PC04, SCHT08]. hypercubes [CLHH04, CC01b, CG09, CQ09, Che09a, Che09b, Che09c, FLP00, Fu08b, GD08, Hsi08, HKC09, Lai09, LTH08, LTTH03, ML09, STH07, TTLH02, Tsa07, TJ07, XDX05, YZ07]. Hypergraph [LH03, KBEG07, LW06, NGI08]. hypergraphs [DPS06, DLPS08, Yus06]. hyperplanes [Liu04a]. hyperspherical [LHK02]. hypersurfaces [BGHP09].

i.i.d [MY09]. I/O [KV04, tBK05]. ID [KHL09, ZC04, ZC09a]. ID-based [KHL09, ZC04, ZC09a]. Ideal [CDFM05]. ideals [Abd09]. identical [Eps00, Kas04b, TH02]. identification [LZ04, ZH02, ZC09a]. identifying [HL04, HL07, Suo07]. identities [BHLV09, Kwo02b, Kwo03]. identity [IN03, RG09, Sae02, Sha03a]. identity-based [Sae02, Sha03a]. if [Boj07]. II [Mai03, Ver02]. illumination [DHS08]. image [CY01b, CY02, JCSX09]. images [BDN00]. immediate [RRT08]. immune [Sar00]. impact [Min07a]. imperfect [Fou02]. implement [WW03]. implementation [DIPS00, GS04a, KR08, ST06]. implementations [GM07, Hie06]. Implementing [JAF06, ZC09c]. implication [BMTV09, Wan07]. implicational [JN06]. Impossibility [BHMR01, Bro05, RRT08]. Impossible [Pha04, VM01]. imprecise [WLP07]. Improved [AAK09, AFT08, AGS07, BBM06, BV00b, COW05, CHK^{+}09, DMN09, DLPS08, GING09, GL09a, Han06, KCYL06, LH02, LLC09, LSC06, LR06, LLH09, RS07a, Sas08, VŽ02, Yij04, vDKST06, Bre04, CM07, CV00, COK06, ESZ02, FL07, FB04, GS05b, HIM09, HL05, HR06b, HH07, Kou06, LKP01, LHC^{+}05, McC01, MHK07, NRS00, PCW08, RRSY07, SLL^{+}05, SLT08, Tan06a, Tsi07, WCH08, ZSZY08, dFdF03]. Improvement [WHLH03, GD04, Pec06, SM02, ZC09a]. Improvements [HWW02]. Improving [Duq07, Hyy08, KP09, NKSK09, Sun01, ALO^{+}09, LM06]. in-arborescence [BFK09]. In-memory [ZHW01]. in-network [GRH09]. in-place [Che06a, Vah07]. in-situ [Kel02]. Inapproximability [BC07, AH08, Kou06, Tan08]. Inclusion [AIM^{+}03, CC06a, CW07, CC05b, Kat04]. incomplete [Cho02, CCKL02, Chu04, FL05a]. incompressibility [DKV09, Gag06b, LJL07, Sch09]. inconsistency [WYL06]. Incorporating [CFM09]. increasing [AAN^{+}07, BS00c, Deo09, LCC06, Sak06, YHC05]. Incremental [BHY09, DEPZ09, GL04a, KK07, CWT08, LY07, ZM03, ZJ05]. indefinite [Yan00]. independence [AGM03, GY08, Win03, BCP05]. independent [Cha04a, Fuj01a, Fuj01b, Gav00, HH09, HN08, Jia06a, KOH09, Lai07, LW05, LM09, SSZZ00, SGH04, Tur07, Win04, Yus06]. independent/connected [Fuj01a, Fuj01b]. Index [Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01h, Ano01l, Ano01i, Ano01j, Ano01k, Ano02d, Ano02a, Ano02b, Ano02c, Ano02o, Ano02l, Ano02m, Ano02n, Ano03a, Ano03b, Ano03c, Ano03d, Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano04a, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano06a, Ano06y, Ano06z, Ano06-27, IR08a, JLW09, SSYM04]. Indexed [Ano01g]. indexes [Fre05]. indexing [ALU02, CK02, MN05, RY08]. indices [CQ09, HRC08, XX07]. induced [BHMR01, CHZL09, ELO08, Gav02, Hon07, KKM00, Loz02, LR03b, PK02, RS07b, YEM05]. induced-universal [ELO08]. inductive [Jai09b, Smo08]. industrial [DD06]. industrial-strength [DD06]. inequalities [BNvO09, Mad05, Rav07]. Inequality [Lit03a, BC00, BHK^{+}00]. inert [Der09]. Inf [LPWZ14]. inference [BEL08, FJL01, Jai09b]. inferring [FA05, Mäk00]. infinite [ALS09, BBS06b, Bra09, DJ07, KMST07, Kur08, NRW09, LM04a]. infinite-state [BBS06b, KMST07]. infinite-valued [NRW09]. Infinity [Tao06]. influence [FHK06, Shi00]. Inform [AS02a, Chu04, CT07, Kwo03]. Information [Cha02b, GP03a, Khu00, RRW00, BM02a, BBD02, BdATdC04, CK02, HM03, KdW04, Kno05, NN08, PS02, RC06, TS02]. inherent [DGP04]. initial [DK08]. innermost [Fer05b]. input [HM08, NSZ00]. input-constrained [HM08]. inputs [JUZ05]. insecurity [Gho07]. insensitive [KJ04]. Insertion [LJL07, ZJ05]. instability [Gia09]. instance [Mee07]. instances [Rah08, ZC09b]. institutions [AD07, Dia04, VV01]. insufficient [EM08]. Integer [BNvO09, CM07, Asa08, BCG09, BHM08, Bol08a, CKSY07, DL06, FL05b, GS04b, SW05b, XL00]. integers [QW06, Rug03, Sri08]. integral [CKSY07, WWGF05]. integrality [FK00b]. Integrating [CCG08]. integration [GT09]. integrity [AFK08, KSRI03]. intensity [Wu08]. intensity-modulated [Wu08]. interactive [KHL09]. interchanges [BF09, HKL^{+}04]. interconnect [HS01]. interconnection [ÁAC^{+}08, KR03b]. interesting [RK05a]. Interface [Zea00, Gam00]. interleaving [HYT^{+}08]. internal [BK04, OR05]. Internet [LZS04]. interpolation [Bor00, MGLA06, VFV02]. interpretation [BBD02, Bla05, eF02]. interpretations [MVM07]. intersected [Juk05]. intersecting [MKGB08, WWW08, Xia07]. intersection [AC02, CH06, Cha04a, Gav00, Miy08, SJ05]. intersections [AMS08, KMU05, LE02, Vig03]. interval [AGS07, DDR04, GH00b, Iba09, Jia06a, KZ06, Lin07, LC08, Orl09, PD03, PD09, RST08, RY08, Sha09a, Tse05, Yan08]. intervals [HS05]. Intractability [KG05]. intrinsic [Amb00, BGHP09]. intruder [Del06]. intrusion [ST00]. invariant [Deo06, Ete00, TD09]. invariants [Göb00, MOS04]. Inverse [LH05]. inversions [ER03]. inverted [HC02]. investor [Lev07]. Invisible [CG08]. involutional [SHH07]. IP [AKAE04, HC08a]. Irrelevant [VL04, Amt08]. Isolated [VSSV05]. Isometric [PC05a]. Isometric-path [PC05a]. Isomorphic [TKS07]. isomorphism [Dar05, RZ00]. isomorphisms [Par08]. issue [AG01]. issues [CLL08]. item [Det09]. itemsets [JCL04, KMT04, Tat06]. iterated [RRT08, BG01, Riz03]. iterates [GK04]. iteration [AFI03, HM02, Joh04]. iterative [IK05].

Jackson [CVEOV08]. Jacobi [Duq07]. Jan [LHC^{+}00]. Java [Gam00, SKP^{+}02, Zea00]. Java-based [Zea00]. Jebelean [Sed07]. job [ABG05, CCG08, LC02, ZCF^{+}06]. jobs [BJK02, GM08b, JC07, NS02, NCY02, NCBJ02, WX05, WL03, YZ07]. Johnson [ARdAdS07, Eng04, Pec07]. join [KA02b, PC00]. joining [AG03]. joins [CC09, KLK04, MK05]. Joint [LZS04]. joy [Fei01]. jump [McC01]. juntas [CG04a].

K-set [GM08a]. Kautz [Ara07, KS03b]. KEM [LLH09]. kernelizations [Cai09]. Key [BD05, CWJT01, CHH^{+}09, CCKL02, Chu04, DHS09, GMLS02, GH08, Hwa00, LPM05, MSS09, DFM04, Shi08, Tan06a, Tsa06, VS01, WW09, XH05, hYScH05, hY08, hY09a, YPKL08, ZC04]. keyed [Küh08]. keys [Sch01, Wan04]. keyword [BDQ08, KLP04]. kidney [BC07, Irv07]. Kiltz07 [LLH09]. Kiltz07-KEM [LLH09]. Kinetic [Her04, dFdF03, dFdFC04]. kings [HC03a, WS01a, WS01b]. Klee [Vah07]. Kleene [DL04, Eng07, NNS01]. knapsack [CKP00, Det09]. knowledge [Stu09]. Known [CKN06, Pad05, Weg02]. Known-plaintext [CKN06]. Knuth [SL06b]. Kolmogorov [DV04, CF02, May02, Sta05, Ver07].

Koyama [Pad05]. KP [OF03]. Kronecker [Jha03, MV08, TKS07]. Kurosawa [CHH^{+}09].

label [DS04a, Tse05, WCX02]. label-cover [DS04a]. labeled [Mon05]. Labeling [JC04, JQQ^{+}03, Jia06b, KV05, RSGK08, RGDN04, RBDN09, SK02]. labelings [CGŻ09]. labelled [ABF^{+}08]. labelling [BS07, HSWC09, WY06]. laceability [CLHH04, TTLH02]. ladders [CHC02]. lambda [Gol00, Gol04]. language [ALA06, BN04, Boj07, LZ04, Mor05, Nag05b, Okh03, Vág06]. languages [ALS09, ARdFPS02, Bou02, CDFG01, Dar05, FJ05, FP04, GdPL08, HHL05, IMP06, LS09a, Mäk00, Tri04]. LANs [ZXC05]. Large [Gag06b, BMCK04, Bol05, JCL04, NC05, Tak00]. largest [AFT08, HH09, Kou06, OT03]. latency [SS03a, Wu00, WHZ04]. lattice [CN00, FS08, SS00]. lattices [Eis01, JN06, LM02, Pal03a]. law [Bar01a]. layer [BG08]. layer- [BG08]. layered [NY04a]. layering [BLME02]. Layland [DG00]. layouts [GG02]. lazy [ADM04]. LBFS [Chu08]. LBFS-based [Chu08]. LC [Rez05]. LC-tries [Rez05]. LCS [IR08b, LSZU03]. lead [Šte00]. leading [Juk04]. leaf [BL06, LT06]. learnable [SKA01]. learned [Car02]. Learning [BK00, HLSW02, Amb00, Amb01, CG00, GL04a, Jai09a, JR07, KY06, KY07, LZ04, Lon03, LSS07]. least [Sop02]. leaves [SW08]. left [HR03, SV05]. Legacy [Hoa01]. Leibniz [Dij01]. lemma [Gri01, HG01, Sch09]. length [AYTH08, ALM02, BD08, ELR07, FB04, GM08b, HC08b, Hu07, LHWL07, MZ00, SV08, SLT08, Wil09]. length-constrained [BD08, SLT08]. lengths [HRC08, HM09, LCW07, MR09b, OPR07]. less [JLO01, Pec07]. Lett [AS02a, Chu04, CT07, Kwo03, LPWZ14]. letter [JS07]. Letters [Cha02b, GP03a, Khu00, RRW00]. level [BDN00, LA01, SSSS02, RP03]. Level-Ancestor [RP03]. Levin [ACI^{+}06]. lexicographic [Hal03]. lexicographical [Neb07]. lifetime [FMR05]. lightweight [DM08a]. like [BN07, DC07, HSL^{+}02, PC04, SCHT08]. likely [CG04b]. LILI [JJ02]. LILI-128 [JJ02]. limit [Amb01, BH08a, NNR04]. Limited [BFL00, BLN07, DPS05]. Limiting [ST04a]. limits [BA03b, Ser01a, Ser01b]. line [ADMP01, AAER03, Ary02, Asa08, ADK06, BF03, BS07, CP05a, Car02, Eps00, Fuj03, GN00, GDN07, GD04, HGYZ09, HRV04, HT03, KRD08, Kol04, LL02b, LSL08, MP08a, MKGB08, NKSK09, NS02, NSW02b, Sas02, TKS07, TLH02, XLMH05, YZ03, ZH06, ADLP04]. Linear [AKL07, Cai09, HJS03, Kim03, LY00, LC06, LC08, MR08, Shp00, Tur07, ADMP01, Alb02, AN09, Ars08, BFK09, BF03, Bla05, Bol09b, BM02b, BL06, CCGP05, Cha08, CDC04, CLZ09, CWCL02, Chu08, CKL^{+}03, CI08, DDR04, DN08, DL06, Efr08, FKM06, FL07, Gam00, GVFS00, GH05, Hal03, HR01, Hes07, HS03, Hu07, IP09, JC07, Juk09, KM07, LL01, Mäk00, MW03, MR01a, MR01b, NKS09, NR08, OT03, Okh03, PD03, PS04b, Pud00, Ric00, Sak06, SS03b, Sch00a, SL06a, SW05b, Sku02a, Sku06, Vág06, WX05, WS07, WL03, Yun08, ZNI03, Zie09]. Linear-time [AKL07, Kim03, BFK09, CLZ09, DDR04, Hes07, KM07]. Linearity [Fuj05]. linearly [GD00]. lines [CC06b, Jon02, LT03, WCK04]. link [AMP03, Cho08, EM08, LCKL06]. links [KKL03, Maa07, MNR04, ZC09c]. lintime [Cap06]. Linux [SKP^{+}02]. Liquid [VSSV05]. List [Ano01g, LHL08, Mar04, AAN^{+}07, DL08, PYZ09, Sku06]. lists [DCS07]. literals [MST02]. Little [Min05]. Liu [DG00]. live [Che04b]. liveness [LJ09]. LL [BN01]. load [AKE00, KM02, LLP09a, Low06, MCS^{+}09, QH01]. load-balanced [LLP09a, Low06]. Local [Min07b, yZyLZ09, GY08, HM03, KHN05, KA02a, Mai02, MSZ02, PS09, RL02, RB02, Sch09, Sun01, VV01, VFV02, ŠŽ04, Ver06]. local-spin [KA02a]. localization [WA01]. locally [Boj07, MA06, PP05a, SL06a]. locating [CC05a, Suo07]. location [AB09, Liu04a, NII01, NWZ02, SW09, XX05a, vDR09]. lock [CLW02, KA02a]. Löf [Dav04]. Log [Val04a]. Log-Rank [Val04a]. logarithm [Sch01]. logarithmic [CKK03a, NEK03]. Logic [DPR00, NNS01, AHV03, Bar01a, Bla05, CDP02, LS06, Lib06, NG05, NRW09, SSSS02, WHO09, PBD^{+}02]. logical [Bou02, GL07]. logics [CDC04, tC09]. LogP [ZLT02]. LogP-model [ZLT02]. logspace [HRV00]. long [LR03b]. Longest [BG01, FB04, AAN^{+}07, AYTH08, BS00c, CC05a, Deo09, The02, GH05, HYT^{+}07, Kim05, NKS09, RS07c, Ric00, Sak06, Tsa03, YHC05, CIPR01, CI08]. look [BV00b, FKV04, RS06]. look-ahead [FKV04]. look-up [BV00b]. lookahead [BN01, ZXZ08]. looking [BDQ08]. lookup [MN05]. loopless [XUT00]. loops [Amt08]. loss [AKAE04, SYL04]. lossless [LL02b]. lossy [Sch02a, ZC09c]. lot [KPS02]. lot-size [KPS02]. Loveland [DV04]. Low [SSZZ00, SJ08, WS06, BS00a, Eis09, Val04a]. Low-complex [WS06]. low-congestion [BS00a]. low-dimensional [Eis09]. Lower [BME02, BZLN07, Fun08, GHL02, HV06a, PS02, Rau05, San01b, Saw07, Shi00, ACI^{+}06, AM05, Aro02, BP06, BCLO03, Bol05, Cha03, CS03a, CM07, CG04a, DS00, FKW09, FK00b, HIM09, KR03a, Kur04, LLC09, Las09, Lif03, LP08b, LM04b, Liu04b, MN08, MP08a, MRST09, Pud00, Shi02, SL06a, SM02, Tse05, VŽ02, WTC05, WCH08, XDZ05, XL00, ZCF^{+}06]. LP [Mel07b, XX05a]. LP-based [Mel07b]. LR [BN04, BN07]. LR-like [BN07]. LSI [CTN03]. LTL [Ete00]. Lucky [CGŻ09]. lumping [DHS03]. Lunch [GO05, IT03]. Lyndon [DP07, Vaj08, VS01]. Lynn [Hes04].

M [Sze08, KLY00]. machine

[ABG05, Bre04, CL03, CNK06, GM08b, HLY09, Hie04, JC08, KY06, KY07, LCL06, MO06, Mul09, NCY02, Pap09, TH02, WYL01a, WYL01b, WC01, Woe09, WL03, ZXZ08, VSSV05]. machine-model [Pap09]. machines [ADM04, DM08a, Eps00, FL09, GL09b, KC03b, LLPZ07, Pet02, Rem04, San01b, Sch08, SJ05, Sze06, Sze08, TH02, TCNY09, YZ03, LM04a, Pol04, RP03]. Macro [PS04a]. MAD [DDR04]. made [Sax09]. Magnus [HPW08]. mail [Dam02]. mail-order [Dam02]. Mailman [LZ09]. maintainability [VL04, dH01]. maintaining [SG03]. makespan [BS08a, CNK06, HLY09, JC07, JC08, KY06, Reg02, TCNY09, Woe09, WL03]. malicious [Tsa06]. man [DCS07]. man-exchange [DCS07]. Managed [AFG07]. management [CLW02, HC08a, KV04, KY03b, LPM05, WW09]. Managing [Cha02a, Lon08, CWC00]. manipulation [Fei01]. many [KP09, PP05b]. many-to-one [PP05b]. map [SK02, Wu08]. mappings [LCC06]. maps [Pol03]. marginal [KLN^{+}03]. marked [Che04b, CC05b]. market [DH06]. markets [FSW03]. Markoff [GLS08]. Markov [BBS06b, DHS03, ST04a]. marriage [DCS07, DS04b]. Martin [Dav04]. Martin-Löf [Dav04]. martingales [Lut04]. masquerades [ST00]. Master [Ano01d, Ano01e, Ano04z, Ano04-27, MCS^{+}09]. Mastermind [Goo09, JP09]. Matching [LS09b, AK08, AK09a, ALP02, AN09, BCG09, BK04, BEL04, CGK08, CGM06, Cha02c, CC07, CIL^{+}03, CS03b, Dai09, DR06, DS00, Deo06, DH03, FKM06, Fre03b, FM06, GF08, HJS01, HR03, Hyy08, IR08a, JLW09, KPL04, Lec07, LLP09a, LP08a, LP08b, Low06, MTF02, MS04, Mis03, Mon05, PT00, Pet07, PS04b, UC00, LH08]. matchings [EK00, HIM09, HV06b, KS00, Loz02]. mathematical [TJC^{+}06]. matrices [AHM08, ELS06, GP03a, GP03b, ME04, NKSK09, RS07c, dBT08]. matrix [BD04, FR04, GKL09, IS09, LZ09, MN08, Sib04, Wes04]. matting [LMC^{+}09]. max [HR00, Hu07, MTT03, MK03, DFG^{+}03, COW05, FKL01, FR06, HR02]. Max-Bisection [FKL01]. max-min [Hu07]. Max-Satisfy [FR06]. MAXCUT [RS07a]. MAXDAG [RS07a]. Maximal [HJS01, Gal03, HH09, KSS00, KS00, NSI04, Raf01, SGH04, Win04]. Maximizing [YZ07]. Maximum [AG03, CK08, Der09, Gav00, LM09, YEM05, AH08, AHS01, BBS06a, BCG09, BG09, BHM08, CR09, Cha04a, Che03, CLC09, DH06, FKM06, FR08, Gav02, HIM09, HR01, HR06b, HC08b, HSWC09, Jia06a, KHN05, Kim03, KMŻ06, LH03, LHC^{+}05, LCKL06, LT06, LC08, LW05, Loz02, MC02a, Mon01, NR08, PT06, PS04b, RB02, Sea04, Sku02a, SW09, SLT08, TK02, TJH09, UC00, WC09, Wu09a, YAM04, BM02b, BLdR04, Cha05]. maximum-density [HC08b, Kim03, SLT08, Wu09a]. maximum-sum [CLC09]. Mean [BDL09, CH08, KW01a, KW01b]. mean-payoff [CH08]. meaningful [Hoe01]. means [JCSX09]. measure [CT00b, Dav04, IKKY00, LW08, Mos08a, Pes00, Vah07]. measures [FM03, GR03]. measuring [LTH05, KSRI03]. mechanism [GH08]. mechanizing [MM01]. media [LKP01]. median [BF03, CKY06, PG09, VGG^{+}00]. medianoid [SW09]. medians [Chu00]. meet [KG04]. membership [JAF06, SKA01]. memorization [CG05]. Memory [DK08, CB04, HL05, HR05, HC03b, Kno05, Mos08b, ZHW01]. Menger [SCHT08]. merging [Che06a]. mesh [Che00a, HL04, JNK^{+}02, Jha03, PP05b, TV02]. meshes [CKK02a, DYZ08, LT08a]. Message [NSR06, ÁAC^{+}08, MSS09, MRT00]. message-passing [MRT00]. messages [Bih02, Car02, EM08, Kwo02a]. metablocks [RT00]. metering [BCM02]. Method [Nun04, Gol01, GM02, KMCC04, KTK02, KA02b, LYP02, MN05, MK05, NR00, Sch09, XDZ05, YZ08]. methods [ABBG06, CP02, Mel07b, Rao01, SB06, SM02, YCKK07]. metric [CN03, CKY06, Fre05, HR02, HR06b, KPL04, LTT06, LP08a, OB05]. metrical [BLN07]. metrics [KRS01, LM08a, Man05]. Metropolis [Mee07]. mice [VLK06]. migration [Che00a, dH01]. min [ACN03, BEP09, Hu07, Kun00, MTT03, MK03, SSZZ00, Van07, Vog08, Hes03]. min-cuts [ACN03]. min-max [MTT03]. min-transitive [Kun00]. Min-tree [Hes03]. min-wise [SSZZ00]. Minimal [BS07, KSRI03, NON09, PĆB04, RST08, BM02a, BZ02, CT00a, CB01, CHLN07, Fou02, Hie04, LC08, TH09]. minimization [AZ08, HL08, Löd01, Reg02, Tri06]. minimize [CNK06, JC07, JC08, LCLP09, NCBJ02, TCNY09, WYL01a, WYL01b, Woe09, WL03]. Minimizing [BJK02, JUZ05, KY06, LLPZ07, LPM05, OT03, HLY09, KK04, KW01a, KW01b, LH03, LSY07]. Minimum [AMP03, BZLN07, Cha04b, DV09, FGR02, Gav08, Ham07, JN06, Kol05, KR03b, Liu05, MP08b, NII01, NK07, YCML06, ABM09, ABBG06, CKK02a, CKK03a, CCGP05, Che03, CG04b, DS00, FL07, FKN02, GMM^{+}07, Gør08, GHL02, HN08, IN04, Jia06a, JR04, Juk04, KHLL09, LM08a, LSL08, LR05, LTT06, LC08, MZ00, Mar02, MKGB08, Nag05a, iNM07, OHNT02, Poo03, Tan07, WCX02, WCY00, Wu00, WHZ04, WYL06, Wu09b, XL00, ZNI03, ZSN02, FMT09]. Minimum-cost [Kol05]. Minimum-link [AMP03]. Minimum-perimeter [MP08b]. minimum-weight [WCY00]. minimum/maximum [LC08]. Mining [HC02, JCL04, RC06]. minislots [JLO01]. Minkowski [PS03, dBvdS02]. minmax [KZ06, KZ09, NO07]. minor [Par03, WY06]. minus [Lee09, LC09a]. mismatches [LS09b]. missing [CV03]. Mix [Sar08]. mixed [BJ08, FKŻ08, LLC09]. mixing [Sar08]. Miyazaki [WHLH03]. mobile [BBMR08, Bol09b, DGR05, KY03b, LC03a, WW09, HC08a]. Möbius [Fan02, XX05b, XML06]. MOD [HRV00]. modal [tC09]. mode [PG09, Sar08]. Model [Bij01, ÀBD^{+}04, CH03b, Duq07, GHZ02, Hoe01, KLN^{+}03, Lan08, Lin08a, MMV08, Mul09, Pap09, RRT08, Reg02, Sae03, SHC08, Ste08, SYL04, YWW09, Zel07, ZLR^{+}05, ZLT02, vDR09, HHLS03]. Model-based [Bij01]. model-checking [Ste08]. model-theoretic [Lan08]. Modeling [NC05, WC03, CB07, Kno05]. Modelling [KJ04]. models [ARdFPS02, BFMZ01a, BFMZ01b, CA08, CB04, DC07, HLSW02, HI03, KSS00, Ros02, WHO09, XP06]. moderately [BEP09]. modern [Amt08]. modes [BH08b, BHW06]. modification [Heu08, OF03]. Modified [CP05b]. modify [Tho05]. Modifying [hYScH05]. modular [FJL01, KLY00, Kle08]. Modularity [tBK05, VV01, VFV02]. modularization [DM00b]. modularized [LYP02]. modulated [Wu08]. module [Lon08]. modules [PK02]. moduli [Pre00, SW05b]. modulus [GS00]. molecular [dSCLP08, PRSS01]. moment [JCSX09]. moment-based [JCSX09]. monadic [Bar01a, DJ07, Fuj05, MOY06, Vág06, Van01]. monitors [DEPZ09]. monoids [Ric03]. monopolist [Bac07]. Monotone [BW06, HY09b, KS03a, DDD08, FMSS01, GP03a, GP03b, Hag07, LWŻ07, LPWZ14, SKA01]. monotonic [Jai09a, Par11, WLLS08, ZB07]. monotonicity [AA05, Fer05b]. Montgomery [LH05, LKY02]. moody [vdBW08]. Moore [DR06]. morphisms [CIL^{+}03, HHL05]. mosaic [HYT^{+}07]. most [NPW01, Shp02]. motif [Rom09]. moves [KS06, WW03]. moving [CPC03, Kol04, WCK04]. MPLS [AKAE04]. MREP [FGLW04]. mRNA [Gur08]. MSO [EM06]. MST [MZ00, PC05b]. Mu [MS06]. Mu-calculus [MS06]. much [QW06, ZXZ08]. Multi [AB09, Che00b, LV07, SZQ08, AFI03, CY00, Dil02, DYZ08, GMLS02, GKL09, GR03, GHZ02, JNK^{+}02, JX09, LCLP09, LY07, Mas08, Mes06, NCY02, PC00, PT00, SSYM04, Shi08, ZXC05, ZH07]. multi-agent [LCLP09]. multi-agents [JX09]. multi-channel [ZXC05]. multi-clients [LY07]. multi-complete [CY00]. multi-complete/star [CY00]. Multi-designated [LV07, Shi08]. Multi-dimensional [AB09, DYZ08, SSYM04]. multi-exit [AFI03]. Multi-hop [SZQ08]. multi-mesh [JNK^{+}02]. multi-object [PT00]. multi-objective [ZH07]. multi-operation [NCY02]. multi-parallel [Mas08]. multi-path [GHZ02]. multi-radius [Mes06]. Multi-subsequence [Che00b]. multi-user [GMLS02]. multi-valued [GR03]. multi-variate [Dil02]. multi-way [PC00]. multi-witnesses [GKL09]. multicast [AKAE04, CC03, GHZ02, Kwo02a, MD00, Seg08]. multicasting [JLO01]. multicolored [KMU05]. multicoloring [Mar04, NPZ01, ŠŽ04]. multicommodity [GN06]. Multicomputer [GS05a]. multicomputers [Che00a]. multicut [GL07, MR09a]. Multidimensional [Bra07, LC01]. multigraphs [Alo03]. multikey [KP09]. multimedia [AKE00, KA02b, LKL^{+}00]. multinomial [KM07]. multiobjective [GNAJT09]. Multiple [CGM06, ALO^{+}09, BDKW07, BtN05, Dai09, Det09, LTT06, Man05, MVM07, Rem04, TC04, TV02, YC09, CKP00, Kub06]. multiple-file [ALO^{+}09]. multiple-source [LTT06]. multiplexer [GM09]. multiplication [AHM08, Bol08a, GKL09, GD05, IS09, KLY00, LZ09, Sib04]. multiplicative [SK08a]. multiplicities [Alh06]. multiplier [LY00, LKY02, SJ08]. multiplier/squarer [LY00]. multiprocessor [CLL08]. multiprocessors [BVG02, Bar01b, SB02]. multisets [Gag08b, Hag03]. multisignature [He02]. multisystem [Cho02]. multivalued [ZC09b, MRT00]. multivariate [Lit03b]. multiway [FK00b]. Murg [Vág08]. music [LW08, McI01]. mutation [JTG04]. mutual [Ala05, CC01a, DL00, LPW00, Pet05, Vid03, YSK03].

NA-trees [CC09]. NAF [BMX05]. Name [Lai07, NNP07, PRSS01]. Name-independent [Lai07]. name-passing [PRSS01]. names [Lei01]. Naor [Gho07, Mar08, Shp00, Zha06]. narrowing [GH00b, RSV07]. Nash [BDQ08, CŠ05, FHK06, Ves06]. Natural [ALA06, LMC^{+}09]. Navarro [Hyy08]. NC [FD04, Win04]. Near [BVG02, CD09, Che04a, SJ05]. near-optimal [SJ05]. nearest [AW04, Gia09, GDN07, HKL^{+}04, KS02a, KA02b, LK02, Liu04b, SPKL07, Wan08]. nearest-point [KS02a]. Nearly [FS09]. necessarily [BW03]. necessary [DPR00, RC06]. necklaces [Vaj08]. need [ADM04]. needed [BF09, Kur04, QW06, RSV07]. negation [ST02, DPR00]. negations [Juk04, ST04a]. Negative [Mie08, MPSS02]. negotiating [Lin08a]. Nei [XDZ05]. neighbor [AW04, Gia09, HKL^{+}04, KA02b, LK02, Liu04b, SPKL07, Wan08]. neighborhood [YCML06]. neighborhoods [YSK03]. neighboring [AHS^{+}03]. neighbors [GDN07]. neighbourhood [FHK06, KV08a]. nervous [Hon07]. nested [ASW05, Dam06, MG01]. net [Las09, Win03]. nets [Dar05, DV06, Jia08, LL01, LJ09, MSSW08, OR05, ŢM05]. network [Bol09b, DEPZ09, FK00a, GRH09, Gol03, HM03, IK05, JNK^{+}02, JR04, KZ09, LC03a, Li09, Min07a, MC06, NY09, Sas02, SG03, Tam00, XL00, Zie08, LH08]. network-flow-based [XL00]. networking [Zea00]. networks [AKAE04, ACN03, BCNR02, BK05, CCY^{+}07, Cho08, COP00, DN08, DP01, DC07, DB06, FJ08a, Fer00, FP08, GRH09, GL02, GL04b, HHT05, HN08, HRWZ07, KR03b, Kwo02a, LLC09, LTH05, MM07a, Mel07a, Mor06, NC05, Par05, Par08, PC04, Pio02, PZ09, Ros02, SY08, Seg08, SHC08, Shi03a, Shi03b, SCHT08, SZQ08, TJH09, WCFC08, WW09, XP06, XZX07, YC09, Nut09]. Neural [Kno05, BK05, IK05, OR05, Tam00]. neurons [BtN05]. next [KN04b, LSC06]. next-to-shortest [KN04b, LSC06]. NFA [Hyy08, Lif03]. Nguyen [CWTHR01]. nine [AD06, FR08]. NML [Giu04]. NN [TLZ05]. No [Fou02, ZXC05, Agr02, CL03, Dal09, Göb00, IAIH03, MY09, Sch02b, Sze08, WC01, GO05, IT03]. No-wait [ZXC05, CL03, WC01]. Node [FLJ05, HYCH07, LE02, LZS04, MK05, OB03, Tsa07, Yan09, ZLJ06]. node-edge-capacitated [ZLJ06]. Node-pancyclicity [FLJ05, HYCH07, Yan09]. nodes [BLME02, JVA06]. noise [Hon07]. noise-induced [Hon07]. noisy [FK00a]. Non [HKL^{+}04, KC03a, Man05, AN09, BHW03, CN00, CHZL09, DGP04, GRT00, GRW06, JLN02, KLK06, KHL09, MK02, NC05, Ric03, Ros02, RCA07, Sol08, UN07, hY09a]. non-adaptive [BHW03]. Non-approximability [Man05]. non-blocking [DGP04]. Non-clairvoyant [KC03a]. non-conspiratorial [RCA07]. non-context-free [MK02]. non-converging [Sol08]. non-determinism [UN07]. non-extreme [JLN02]. non-hierarchical [hY09a]. non-homogeneous [NC05]. non-induced [CHZL09]. non-interactive [KHL09]. non-linear [AN09]. non-NP-hardness [CN00]. non-preemptive [KLK06]. non-probabilistic [Ros02]. non-self-reducible [GRW06]. Non-shared [HKL^{+}04]. non-uniform [GRT00]. nonconvex [Shi03b]. nondeletion [Fuj05]. Nondeterministic [BHW06, Bol03, CCVP01, GING09, Hie06, Juk09]. nonlinear [CFW01, FSGM03, GVFS00, Mai02, MW03]. nonlinearity [GS05b, Sun01]. nonlocal [JCSX09]. nonlocal-means [JCSX09]. Nonnegative [CKSY07, Kim05]. nonprimitive [Sch02a]. Nonstationary [CB07]. nonsuperstring [NKS09]. norm [CM07, SS00, Sza03]. norm-graphs [Sza03]. normal [Hag07, Wan07, Yos09]. normalization [IN06]. normalized [BEL08]. Note [Ano02k, HRV04, LWŻ07, LPWZ14, Ser03, dFdS05, AFI03, AOS^{+}09, ABF^{+}08, AR06, BBS06b, BCP09, BMX05, BCM02, Bol08a, BH08b, CN00, CT04, Cha04a, CHL^{+}08, CT00b, CHLN07, DLP04, DHS08, Eps00, Ete00, FKL01, Fuc03, FKŻ08, GP03a, GP03b, GR03, GLS08, Gro06, HLP03, HG02, Han08, HGYZ09, Héa08, HRV00, Hsi08, IR09, IS09, JS07, Jia08, Kol03, KOH09, Lai09, LLP09a, LLP09b, LZ09, LW05, Lit03b, Liu04a, Liv09, LT03, MTXL09, MC02a, MZ00, MF07, Mon06, MR06, MP02, Mul09, NBCS07, Nut09, Oku09, Pap09, PN01a, PN01b, Pes08, Pio02, Pro08, Pud00, Ric02, Ros02, Sae03, Sar00, SZM09, Sma01, Smo08, Sri08, ST04b, Sze06, Sze08, TYP08, Tan08, TD09, ŢM01, Tse05, WCX02, WLC08, Wes08, Wu07, ZŽ09, Zha06, tC09, PCW08]. notion [Cha02c, Daw03]. novel [CTGN05, SLL^{+}05]. NP [And09, CN00, Fou02, GHM07, Juk05, LL02a, Liv09, MM07a, Min07b, NCY02, Pol03, TW06, WYL01a, WYL01b]. NP-complete [And09, GHM07, Pol03, TW06]. NP-completeness [LL02a]. NP-hard [Min07b, WYL01a, WYL01b]. NP-hardness [NCY02]. NP-problems [MM07a]. NP-witnessing [Liv09]. Number [YSK03, AG03, ABPR06, ADMP01, AC02, AD00, BLN07, BF09, BLV03, CZ02, CY01a, CY00, COK06, DLS08, DS06, ER03, ELS06, FS08, FS09, FRR03, Fre03a, GHJ^{+}08, Han06, Hie04, HH09, JP09, JUZ05, Juk04, Kur04, Lin07, LC08, LC09b, LS09b, Mad06, MZ00, McC01, MK02, MW06, MUK04, Mor06, NBCS07, NNR04, NON09, Orl09, PP04, RW09, RRSY07, SDC09, Sid07, SK08b, Sop02, Sri08, ST04b, Tan07, WC09, Wu09b, YAM04, YZ03, ZLM09, dWH03]. numbers [CLL04, CG09, IJN09, KS03b, MW03, PC05a, Tak00]. numeric [HK00]. numerical [HKL01]. numerically [Sav04]. NV [Luc04b]. NV-sequentiality [Luc04b].

O [KV04, tBK05]. OBDD

[CZ09, GPS01, Saw07]. OBDDs

[Bol08a, Bol09a, BDHW07, Weg00]. object [AAK09, LZS04, PT00]. objection [RRB00]. objective [ZH07]. objects [Alh06, CPC03, JS01, Vig03]. oblivious [GZN06, Gho07, dBT08]. observability [CHU06, UW03]. observation [PR05, Tri04]. observations [BN07]. obstacles [ASW05, CDKK02]. obtain [Sch09]. obtaining [DR06, iNM07]. Occam [LTV03]. occupation [Der09]. occurrence [Cha02c, Hås00a]. odd [Fou02]. oddity [CT03a]. off [DGP04, MH09, NSW02b, Nor09]. off-line [NSW02b]. offs [DS01, Rao02]. offsets [KLP04]. Oja [AM05]. omega [Löd01]. omega-automata [Löd01]. omission [KS09]. On-line [AAER03, Fuj03, GN00, LL02b, YZ03, BS07, CP05a, Eps00, HGYZ09, NKSK09, NS02, TLH02, ADLP04]. once [BW03, Bol03, Bol08b, GRT00]. one [BV00b, Dav04, Gol04, GM08b, HS03, JS07, Kle08, LPM05, MN05, MSS09, Pet02, PP05b, RH02, SS02, SV08, ST02, Sze06]. one-letter [JS07]. one-page [HS03]. one-pebble [Sze06]. one-point [Gol04]. one-round [MSS09]. one-shot [SS02]. one-tape [Pet02]. one-time [SV08]. one-way [LPM05, RH02]. Ones [HG02]. oneway [FY03]. Online [ALU02, Asa08, FIY08, KC03b, NGI08, Sch08, TCNY09, AGGZP09, ABL08, BFL00, BL00, BDL09, CCY^{+}07, CNK06, Du04, FKW09, Fun08, GL09b, JCL04, PYZ09, Smo08, TH02, ZXZ08]. online-list [PYZ09]. only [CF02]. open [DM00c, EL08, Jai09b, Jai09a, MSSW08, RRB00]. open-end [EL08]. operation [Hoe01, NCY02, Sar08]. Operational [Jon03a, LMM05, BBD02]. operations [Weg00]. operators [MAL08]. optical [CQ09, SZQ08]. Optimal [BYSP02, BE05, BA06, CC05a, CLC09, CHC02, CIR08, DHS03, DM05, Du04, GM09, GM07, HG06, HS03, HC01, KdV01, Koi06, KCH03, KM03b, PYZ09, PP05b, Rom09, RSGK08, RGDN04, SJ05, Tan01b, TH02, Zel07, ZB07, BVG02, BDKW07, BB01, BCM02, Bol08b, BDLP08, DHS08, EQ04, FGLW04, Gut02, HZ09, IP09, JN07, KH03, KR04, KS00, KMS^{+}02, Kuj09, Kun00, MY09, ME04, NB05, PN01a, PN01b, Pre00, Sas02, Sch02b, SKA05, SL06a, Sin03, Šte00, SB06, Tan00, Wan01, WL08, WW03, Wu09a, jXyL09, K^{+}02]. optimal- [HZ09]. optimality [BCD^{+}05, Vog08]. optimally [Ala05, Bun04, TM09]. Optimization [GO05, CCZT08, CA08, CHLN05, JLY07, JXL04, KZ06, KZ09, LLC09, MY09, MC02b, SLL^{+}07, TZ08, TC04, Tre03, Van03, YC09, ZSZY08, AF08]. optimization-based [SLL^{+}07]. Optimized [Orl09, Dur03, LC03a]. Optimizing [Dam02, HŽR08]. optimum [Zie09]. options [LPSR08, ST06]. options-ultimate [ST06]. oracle [CW04, LL02b, Živ09]. oracles [KW05, RG09]. Order [CP02, Abd09, Bar01a, CFS09, CDC04, CA08, Dam02, Elm09, Göb00, MMV08, QPV05, Tal00, YHT04, ZJ05, HHLS03]. ordered [Aku06, BHW06, Héa08, HI03, dFdF03]. ordering [Kas04a, SS03b, ZH05]. orderings [PD09]. orders [CK07, LLPZ07, UPGM09]. orientability [LCL08]. orientation [CL00a, DGN05]. orientations [BJ08, DGHS07, Šte00]. Oriented [EO07, Och04, OP08, PS06, CWJT01, DS06, FRR03, LHC^{+}01, Lin08a, NWZ02, Sae02, Sha03a, Sop02, ST04b]. orienteering [FL02]. orthogonal [BVG02, DLP04, IY04, LWŻ07, LPWZ14, Woe01]. orthographic [CKL01]. oscillations [Nik06]. other [GM07]. OTIS [Par05]. out-arborescence [BFK09]. outerplanar [EO07, iNM07, PS06]. Output [Pol04, Sol08]. overall [ACN03]. Overcoming [CHU06]. overlap [CHL^{+}08, ES07]. overlaps [Ram07]. overload [Min07a]. ownership [GS09a].

P

[Alh06, BLV03, GP01, GRW06, Juk05, LM09]. PAC [CG00, Lon03]. PAC-learning [Lon03]. package [RB07]. Packet [AKAE04, Ros02, Šte00, SYL04]. Packing [CR02, Kas04b, BS00a, CP05b, CDL^{+}04, DG08, EL08, FGK09b, Fuj03, JCP08, KMŻ06, Kou05, RR09]. packings [BR03, LMMM04]. PACS [CG00]. page [Cho02, HS03]. paging [BFL00]. pair [BFK09, Cha08, KTK02, Ric02, BdFdS06]. pair-wise [KTK02]. paired [CLZ09]. paired-domination [CLZ09]. pairings [KP06]. pairs [BMCK04, Fou02, Han08, Jia06b, MPSS02, Sib04, SK02, Tak05, Yij04]. pairwise [Dam06]. PalmHashing [CTGN05]. pancake [HHLH03, MR08]. pancyclic [ZLZ09]. Pancyclicity [AS02a, AS02b, EHHR00, Ara03, FLJ05, HLT07, HYCH07, WMX07, XX05b, Yan09]. Parallel [Alb02, CWCL02, JNK^{+}02, TC02, UC00, Ary02, BME02, BN04, BS00b, BM01, BB09, FKM06, Gam00, GF08, GH00b, HV06b, Hyy08, JC08, KV04, KR04, KBEG07, KS00, KC02, MVM01, Mas08, Mis03, MKGB08, NS02, PD09, SD09, San01a, SJ08, SV09, TCNY09, Woe09, WW03, YZ07]. parallel-batch [TCNY09]. parallel-machine [JC08, Woe09]. parallel-serial [BB09]. parallelism [CCG08, FB04]. parallelizing [HC03b]. parameter [CS09a, GN06, GNU08, Hag07, JLY07, JXL04, Mad06, MR09a, NR00, RS07a, RK09, Ste08, TC02, Tre03, Wes04]. parameterization [RK09]. Parameterized [AN09, DFMR08, KMRR06, Kou06, RS07b, BC00, CGK08, FM06, Kou05, Mon08, Pap09, Raz07, Ste08]. parameters [MV08]. Parametric [ACN03, Doy07, Lev06]. Parametric-search [Lev06]. Parikh [FR04, Hon02]. parity [CH08]. parse [SV05]. parser [RP08]. parsing [BN07, SV05]. part [KV08a, Par03]. Partial [Mie08, Pad05, AHV03, BFB09, Blä03, CA08, DR06, Dia09, Elm09, FJ08b, FM00, HHK08, HR03, Kub06, MMV08, OP08, PR05, RH02, YAM04]. partially [Héa08]. participant [SWCC04]. Particle [SLL^{+}07, CCZT08, JLY07, Tre03, ZSZY08]. partite [ZLM09]. Partition [Gus02, Cha02b, YC94]. Partition-distance [Gus02]. Partitioned [KLY00]. Partitioning [KK04, BS08b, Jia05, MC02b, WS06]. partitions [ABPR06, Ary02, GHS02, GD08, Sav04]. partner [MSSW08]. pass [Bas08]. passbits [Bur05]. passing [ÁAC^{+}08, Che07, Che09a, MRT00, PRSS01, WC08]. password [YPKL08]. past [KY07]. past-sequence-dependent [KY07]. Patchworks [RT00]. Path [Gab00, GD08, TJ07, AIM07, BD08, Cha02b, CW09, Che09a, Che09b, DM09, The02, GGK03, GT09, GHZ02, HC08b, Kim05, LH08, Lai09, MS06, MK05, Mis01, NPW01, NPZ01, PC05a, RS07c, San01a, Sib04, Tak05, Wu09a, YC94, YZ08, ESZ02]. Path-based [Gab00, MK05]. path-partition [Cha02b, YC94]. Paths [GNU08, XML06, ABM09, AF08, Che09a, FGLW04, Gav02, GL09a, Han08, HT03, KG04, KN04b, LTH08, LSC06, LCL06, LR03b, MC02a, MPSS02, RK05b, SB06, Wil09, YZ08, Yij04]. Pathwidth [FH06]. patience [BS00c]. Pattern [CY04, AK08, AK09a, Bar09, CGK08, Cha02c, LP08b, MVM07]. patterns [CS03b, MRN02, Nag05b, NRS00, Rom09, YHT04]. Paxos [VB08]. pay [Joy03]. pay-as-you-watch [Joy03]. payload [SWCC04]. payment [OF03]. payoff [CH08]. pebble [Sze06]. peer [HJ05, SY08]. peer-to-peer [HJ05, SY08]. peg [Goo09]. Peled [Ete00]. penalties [Fun08, LC02, XX05a]. per-edge [Zel07]. Perfect [Jha03, BK04, BCP08, Chu08, Daw03, EKF00, Gus02, JKŠ05, LMMM04, Mia03, Mon05, Sch00a, Ces02]. perfectly [PG06, NSR06]. Performance [BC00, FA05, JLO01, SM02, Zea00, ALO^{+}09, Alb02, Gam00, KKL03, NP00, SKP^{+}02]. perimeter [MP08b]. period [Mar08, Par11, SW07b, WLLS08]. Periodic [IMP06, Bar01b, KLK06, PG06, Pio02, SB02]. periodically [Par08]. permanent [LR06]. permanents [SW05a]. permutation [BF09, Bro05, EKF00, Gol01, Kel02, NP00, SSZZ00, XKS06]. permutations [AOS^{+}09, Bar09, ER03, LCC06, MR01a, MR01b, RH02, SV09]. permuted [BD04, BEL04]. persistent [Dar05, LJ09]. personal [Mad03]. Personalized [KY03b]. perspective [Hon07, Mad03, PSL^{+}01, dH01]. pessimistic [JP09]. Petersen [Par08]. Petri [Dar05, Jia08, Las09, LL01, Win03]. PH [CW04]. phase [Hon07, RK03, RB07]. Phased [LKP01]. philosophers [BBF01, DGR05]. phylogenetic [XDZ05]. pickup [Jia05]. pickup-delivery [Jia05]. pictorial [GCC00]. picture [CDFG01, RP08]. piecewise [JC07]. pigeon [CZ09]. Pinkas’ [Gho07]. pipelined [GIS04]. Piterman [LW09]. PKC [ZC04]. place [Che06a, Vah07]. placement [IS08, LZS04]. Placing [KSY00]. plain [KP06]. plaintext [CKN06, Pad05]. plan [Her04]. Planar [AIM08, FGK09b, AKL07, AIM07, BMCK04, CHZL09, CRW07, DM08b, Eis01, FHN08, GDN07, GRW06, HM08, HM09, HHT05, KRS01, Kur04, LHL08, LCW07, Mon06, MCL06, NSI04, Och04, SW07a, Sop02, TJH07, WLC08, WWW08, ZLJ06, ZH06, ZS08]. planarization [TJH09]. plane [AK09a, AMS08, Hal03, Jon02, KMU05, MR06, iN02, SD09, Via08, WL02, WCK04, ZH05]. planning [Kur03]. plapackJava [Gam00]. players [vdBW08]. PLCN [YWW09]. PLTL [TD09]. plumb [GD04]. plumb-line [GD04]. plurality [ST05]. plus [WYL01a, WYL01b]. point [AK09a, BGHP09, BH08a, CGK08, DGN05, Gol04, GDN07, HM02, HM08, JTG04, KS02a, KMU05, LSL08, Liu04a, MB08, Mis01, RB02, Ric02, RSGK08, RGDN04]. point-labeling [RGDN04]. Pointer [RP03, KdV01, Mul09, SJ05]. points [AB09, Asa08, BCP09, BS00b, BD09, CW09, CC06b, Chu00, DM08b, JQQ^{+}03, Jia06b, JC04, Kol04, Kur04, MZ00, MCL06, NWZ02, RBDN09, SD09, SS00, Sha09a, Sma01, WCK04, dFdF09]. Poisson [CB07]. polar [BMO08]. policies [hY09a]. policy [KCH03, LCKL06, DFM04, Sch02b]. Polychromatic [HKKL09]. polygon [GHL02, Hu07, KG04, KSY00, LPC02, LHT00, MKGB08, PS03]. polygon-meet [KG04]. polygonal [YSK04]. polygons [BB01, CKM^{+}01, CW02, CW03, CSW03, FIY08, KG04, LWŻ07, LPWZ14, Tan01a, Tan01c, TH03, Tan07, Wan01]. polyhedra [ELS06]. polyhedral [FL05b, ZZ08]. polyhedron [AD00]. polymatroids [PN01a, PN01b]. Polynomial [BZ02, Gur08, NY04b, Wu00, BHK04, BHLV09, BUŽ02, CZ09, Det09, DD06, EL08, EHRS09, EHK04, Göb00, HG06, Juk04, KPS02, Kri02, LW06, MM07a, MC02b, MOS04, iNM07, WYL06]. polynomial-size [CZ09]. polynomial-time [MC02b]. polynomially [SKA01, YAM04]. polynomials [GS00, GS09b, HY09b, LD04, Shi00, Val04a]. polyominoes [Woe01]. polytope [CCF04, Kou06]. polytopes [LR08]. pop [SV09]. populations [Wu09b]. poset [Abd09, CdL02]. position [CW09, LTH08, Wan08]. positions [FS08, Ser03]. positive [BGK^{+}08, DG08, GL04a, Mie08]. Post [HHHK08, Rah08]. postdomination [BLMM06]. Power [GS04a, ABM09, BS09, BA03b, CDC04, GD00, Hon02, Kir02, KMRR06, LTV04, Sch00b, UN07, XH05, BHY09]. power-set [UN07]. Powering [She07]. powers [BL06]. PPA [Gri01]. PPAD [DD08]. PPAD-hard [DD08]. practical [CIPR01, KLN^{+}03, MC02a, QH01]. practice [LR03a]. PRAM [Hoe01]. prebisimilarity [ACIL08]. precedence [BME02, Eps00, TMSO08]. precedence-constrained [BME02]. precision [QW06]. precomputation [CKKC04]. preconditions [Lei05, YCFD07]. predecessor [Mul09]. predicate [KG05]. predicates [AMM05, CK03]. prediction [EQ04, PK02, SYL04]. preemption [ABG05, Fun08, LC02]. Preemptive [GL09b, Du04, KLK06]. Preface [AF01, Ano01f]. preference [DCS07]. Preferential [DC07]. prefetching [GIS04]. prefix [CIR08, Göl08, HHL05, JNK^{+}02, KMCC04, KV08b, Pes08, TV02]. prefix-recognizable [Göl08]. prefix-sum [KMCC04]. prefixes [CK07]. preorder [CFvG08, Las09]. preprocessing [CKK03b]. prescribed [Che07, Che09a, WC08]. presence [BA06, MPSS02, VLK06]. preserve [BV00a]. preserves [Ber00]. preserving [Frö05]. Previous [CI08]. price [Li09]. pricing [ST06]. Prim [Mar02]. Primal [FFFdP07]. Primal-dual [FFFdP07]. prime [Sri08]. primes [DM05]. primitive [DD06]. principle [RS05]. Priority [Reg02, AI08, AMR01, Aud01, BBF01, BA06, GH00a, LM06, YC09, ZB07]. priority-based [LM06]. priority-scheduled [AMR01]. prisms [CF09]. privacy [CKN06, Tan06a, vDR09]. private [KdW04]. prize [Gut08, ABL08, FFFdP07]. Prize-Collecting [ABL08, FFFdP07]. Probabilistic [BHK04, CN03, GS04b, CCVP01, CTN03, Gol01, LS07a, Mos08a, MP02, Ros02]. probability [AKAE04, Gag06a, MH09, Mon08, Sol08]. problem [ACI^{+}06, AW04, ACRS04, ABG05, AHS01, AFK08, BF01, BFL00, BL00, BMCK04, BD08, BDL09, BC07, CKK03a, dSCLP08, Cha02b, CZ09, CLZ09, CLC09, CY01a, CKY06, CDL^{+}04, CT05, CT07, CDFG01, CG04b, CIPR01, Dam06, Deo09, DCS07, Det09, DH03, DH04, DM08b, EM06, EHK04, FMT09, FM09, FL07, Fuc03, GS09a, GH05, Gup08, Gus02, HG02, Hal03, Han06, HR06b, Hon02, HTT03, HSZZ06, HYT^{+}07, Irv07, IY04, JR07, JCP08, JŁ09, Jia05, JR04, JHfS08, KHN05, KV08b, KBEG07, KZ02, KLP04, Kir02, Kol03, KY06, Kut07, LH08, LHC^{+}05, LLP09a, Lev07, LT06, LMZ08, LX02, Low06, MS04, MR09a, McC01, Mes06, Mie08, MAL08, Min05, MUK04, Miy08, Mon01, MRWW08, MC02b, MCL06, NII01, Nag09, NS02, NCY02, NDG02, NSW02b, PT06, Pro08, Rah08, RL02]. problem [RS07b, RP03, Raz07, RRSY07, RRV07, RGDN04, Sas08, SW09, TAI04, Tak05, TA08, Tam00, TYW01, TC04, Tsa03, TJH09, Vah07, WL02, WCH08, WHZ04, WYL06, XX05a, XL00, YC94, ZSZY08, ZNI03, tC09, vdBW08, ADLP04, ABL08, BdFdS06, BLdR04, CVEOV08, CKP00, CKR06, CK08, FFFdP07, HHHK08, Jon02, Jon03b, Pet02, Tan06b, Wes08, dFdS05]. problems [AKL07, ANR05, AFG07, BS00a, BEF^{+}00, CN00, Cai09, CHU06, CDF^{+}04, CHLN05, Dal09, DP06, Del06, DF03, DFMR08, FKW09, FL09, Fuj01a, Fuj01b, FIY08, GL09a, GL07, Gue01, Gut08, IR08b, Jai09b, Jai09a, KZ06, KZ09, LLP09b, LMG07, Li00, LW05, MM07a, MVM01, MK03, MOY06, Mon02, NCBJ02, OB03, OB05, PR05, RK05a, RS07a, Sch09, Ste08, Tao06, Tri04, TLH02, Tur07, UW03, Van03, Vas09, Ver02, Wu00]. Process [AS02a, BP01, Chu04, CT07, Kwo03, LPWZ14, CCVP01, CFvG08, NNP07, PU01, Vu05]. processes [CCVP01, PRSS01, RS06, Win03]. Processing [Cha02b, GP03a, Khu00, KA02b, RRW00, Bas08, GRH09, JC07, KPL04, LHK02, Zel07]. processor [HC03b]. processor-in-memory [HC03b]. processors [BME02, Du04, KS09, MM07a, WLP07]. Producing [SV05]. product [CLM09, CC06b, DPP05, Jha03, JHfS08, Koi06, LM08c, LCM09, SZQ08, TKS07, XXH05, XY07]. productions [MK02]. products [BS02, Ham07, JMV06, JKŠ05, KV05, MV08]. profile [LC04, MW03, Sch00a]. program [Amt08, Bol08b, GM02, MOS04, dH01].

Programming [Zea00, BNvO09, Efr08, HYT^{+}07, JTG04, LL01, Par03, TC04, WS06, DPR00]. programs [AHV03, BHM08, BW03, Bol03, Bro05, CB01, CDP02, HG06, LS06, NG05, NEK03, NRW09]. progressions [To09]. projections [CKM^{+}01, Woe01]. projective [HT00, Sza03]. Prominence [JX09]. prone [KS09]. Proof [LC03c, AMR01, CZ09, Høy00, KR03a, Lan08, LYP02, MM05, Rav07, Vog08, Völ04, WTC05, Weg02, Win04, Yos09, vDR09, vdZ01]. proofs [IN06, MM01]. propagation [LK00, QPV05]. proper [Iba09, PD03, PD09, RST08]. Properties [FJ05, ANR05, Ber00, Che09c, DS04b, GK04, Giu04, HRV00, IN03, KDTS01, PC04, Sch00a, TYP08, Živ09]. property [BBS06b, BHMR01, BUŽ02, CdL02, CT03a, CC05b, CA08, CB01, DPS05, IR08a, JLW09, Stu09, TTH04]. Propositional [LS06, BMTV09, DH08, Kra07]. Protection [Lop00, Gou01]. protein [Gur08]. protocol [BHM06, CC03, FMR04, JLO01, LC03a, MS03, MSS09, Tsa06, YPKL08]. protocols [BHMR01, Bla05, CWJT01, LH02, LM04b, MD00, OF03]. Provable [HSL^{+}02]. prove [Geo08]. proving [Fer05a, KMST07, Nor09, Pud00]. provision [JC08, Woe09]. proximity [CN03, DLS08, KLP04]. Proxy [HC08a, Sha03b, KHL09]. Proxy-based [HC08a]. pruning [HC02, NEK03]. Pseudo [KAS08, AAHK06, Ber04, Bol09b, Shp00]. pseudo-random [Bol09b, Shp00]. Pseudo-relevance [KAS08]. pseudo-triangulations [AAHK06, Ber04]. pseudorandom [MW03]. PSO [RK09, SLL^{+}05]. PSO-GA-based [SLL^{+}05]. PSPACE [PW07, TA08]. PSPACE-complete [PW07]. PSPACE-completeness [TA08]. PTAS [CKP00]. PTIME [MS04]. PTIME-complete [MS04]. Public [GMLS02, GH08, VS01, Wan04, hY08]. Publisher [Ano02k]. pure [Dal09, FHK06, RP03]. purely [Lan08]. pursuit [KR05]. push [CH03b]. push-enter [CH03b]. pushdown [Ser03]. pyramid [yZyLZ09]. pyramidal [LH04]. pyramids [ACHP09, SAOKM01].

QBDs [YWW09]. QoS [CL06, YC09].

quadratic

[AHS01, BW03, CT04, PG09, ZLZ05]. quadratic-time [ZLZ05]. quadratically [KL01]. quadtree [CY01b, CY02]. quality [CC02]. quantification [CDC04]. Quantified [PBD^{+}02]. quantization [KMF04]. Quantum [ACI^{+}06, BdW03, KdW04, KL01, SW07b, BMP^{+}00, DT09, EHK04, NY04b, Shi00, Shi02, YCFD07]. quartet [WYL06]. quasi [GdPL08, GHNP02, MP09, RSV07]. quasi-bipartite [GHNP02]. quasi-reversible [GdPL08]. quasi-symmetry [MP09]. quasi-termination [RSV07]. queries [CKK03b, CIR08, DS01, Fre05, GT09, Kow07, LTV04, Lib03, Mad05, PHK01, PG09, SKA01, Tat06, TLZ05, Van01]. query [AB09, Bol05, CG00, CPC03, DT09, EHK04, KRD08, KAS08, LZ04, LHK02, Pro08, RB02, SL06a, TC04, Yan00]. question [Ete00, Juk05]. Queueing [ÀBD^{+}04]. queuenumber [HH07, PCW08, PCW09]. queues [KCH03, PP05b]. Quick [Via08]. Quickselect [Kub06]. quicksort [Dur03, LJL07, KP09]. quorum [CC01a, TK02]. Quorums [CL00b]. Quota [ADLP04].

R [AS07b, Lit03a, SPKL07]. R-tree [AS07b]. R-trees [SPKL07]. Rabani [FK00b]. radiation [LSY07, Wu08]. radio [DN08, DP01, FP08, GL02, TLH02]. radius [DV09, Mes06, XDZ05]. RAID [Tho05]. Rajagopalan [Riz03]. ramp [IY06].

Ramsey [Sch09]. Ramsey-type [Sch09]. Random [AKE00, BCNR02, Bol09b, CMS03, CF02, Dav04, DDD08, DLS08, ES06, JVA06, LT07, Low02, MP04, Orl09, RG09, Sha09a, Shp00, TZ05, Zem08, dlVT09]. Randomized [AK02, DL00, GS03, MSZ02, BFL00, KV04, Pan05, UPGM09]. Range [PG09, ABM09, Bra07, CDP02, Fre05, LHK02]. rank [GP03a, GP03b, NON09, iO09, SZM09, Val04a]. rank-width [iO09]. ranked [MRST09]. Ranking [MR01a, MR01b, DS01, yH02, KR04, iNM07]. rankings [BFB09, DN06, IJN09, NON09]. rate [BGK^{+}08, LCKL06, PS02, Par11, SYL04, WLLS08]. ratio [Asa07, BHY09, FL07, FK00b, LM04b, MR09a, PHG08, Sep05]. rational [CG06, Hon05, Lit03a, Lit03b]. rationality [Lav08]. rationals [KM03b]. ratios [Par11, WLLS08]. raytracing [QH01]. razor [LTV03]. RDT [BHMR01]. re [KHL09]. re-encryption [KHL09]. Reachability [Göl08, Jac03b, Doy07, LS07a, MOY06, TK09]. reachable [Dar05]. read [BW03, Bol03, Bol08b, GRT00, KBOR05, Tho05]. read-modify [Tho05]. read-once [BW03, Bol03, Bol08b, GRT00]. read/write [KBOR05]. Ready [CFvG08, Kol04]. Real [Lei01, Lev06, BGHP09, CLL08, CCG08, KdV01, KLK06, MP02, TMSO08]. real-time [CLL08, CCG08, KdV01, KLK06, MP02, TMSO08]. reals [BK00, DG08]. rearrangements [JN07]. reasoning [FJ08b, Li00]. receive [ZLT02]. receive-graphs [ZLT02]. receiver [Car02]. reception [CC03]. reciprocals [Sch00b]. Recognition [EIM02, BN04, BL06, LS07b, NSI04, PD03, VSSV05]. recognizable [CG06, Göl08, Vág06]. Recognizing [Nik00, RR00, Chu08, DLPS08, San01b]. Recoloring [Raz07]. recombination [MAL08]. recombinations [Wu09b]. reconfigurable [MNR04, TV02]. Reconstruct [Tho05]. reconstructing [Heu08, RS07c]. reconstruction [SW07b, Woe01, XDZ05]. Recontamination [Der09]. Recovering [JS01]. recovery [Cho02]. rectangle [Cha04a, DGN05, IS08]. rectangles [Cha03, JC04, KMS^{+}02, SD09, K^{+}02]. rectangular [CH02, LMMM04, Liu05]. rectify [RB07]. rectilinear [Hal03, MC02b, Wan01]. recurrence [CT00a]. recursion [Xir06]. recursive [AS02a, AS02b, Ara03, Che04a, GTM03, Heu08, JVA06, PU01, Sch02a, TTH04, YEM05]. redefinition [Nun04]. redexes [DN05]. reduce [BLMM06]. reduced [MK02, NP07, NEK03]. reducible [GRW06]. Reducing [BEL08, RK03, SY08, UN07]. Reduction [CH08, BN07, CA08, FM03, Lec04, SS03a]. reductions [CN00, CS09b, Ver02]. Redundancy [MD00, Wan07]. redundant [BLL02]. Reference [Ano01g, Lin02, Lin08b]. Refined [CG05]. refinement [LA01]. reflectable [LS09a]. Reflected [MTT03]. reflective [Jai09b]. regarding [Ete00]. region [YSK04]. regions [SB06, Tan00]. register [HG06]. registers [PP04]. regression [BF03]. regret [KZ06, KZ09]. regular [ACKM07, ARdFPS02, BH08a, Boj07, CLT07, CDFG01, FKL01, FL05b, FIY08, FKV04, GT09, HHL05, Jia08, Lif03, MTF02, Mia03, Par08, Ser03, SHC08, Tri04, tC09]. regularity [BM01, TM09]. Reingold [Mar08, Shp00]. Reiter [Wot01]. related [CY00, CIL^{+}03, LW05, MOY06, Mon02, OB05, RRW94, RRW00, RK09, Sch08]. relation [Aku06, CM07]. relational [Dal09, GR03]. relations [CG06, Dob04, Liv09, TMSO08]. Relationship [HHLS03, HYT^{+}08, KP06]. Relationships [CB04]. relative [KV08a, Par11, WLLS08]. Relativized [CW04, GP01]. relaxation [FK00b, MRWW08]. Relaxations [LMG07]. relaxed [CP05a]. Relaxing [Fer05b]. release [GGK03, LC02]. relevance [KAS08]. reliability [Min07a, NC05, WNR00]. Reliable [NSR06, AKAE04, ZC09c]. rely [DM00c]. rely-guarantee [DM00c]. remain [Jac03a]. remaindering [DLX09]. remark [Dam06]. Remarks [MVM07, TJC^{+}06, SK08b, Tsi02]. remotely [Küh08]. removal [MD00, RSGK08, ST02]. Removing [LE02]. rental [LPSR08]. Rényi [CV00]. repeat [KW05]. repeated [BF03, BHW06, Kle08]. repeats [CD09, Raf01]. repetition [Ric03]. repetition-free [Ric03]. repetitions [GS05a]. replacement [GL09a, Vog08]. replicated [CWC00]. replication [HL03, LLK08]. Reporting [Vig03, KMS^{+}02, SJ05, K^{+}02]. representation [BZ02, Dil02, PRSS01, XUT01, ZH05]. representations [CKSY07, CDP02, DJ07]. required [Fre03a, LTH08]. requirements [BLMM06, NII01]. requires [Bol03, She07]. reservation [KCH03]. reservoir [ES06]. resets [Hie04]. residualizing [AHV03]. residuals [Sim04]. residuosity [CT04]. resilient [GS05b]. resist [hYScH05]. resistant [GS04a]. resolution [ET03, GL05, NG05, Nor09]. Resolving [LK00]. Resource [HP05, Mos08a, Rem04, CL06]. Resource-bounded [HP05, Mos08a]. resources [PSL^{+}01]. respect [Göb00, Hon05, Lon03]. Response [CWC00]. Ressel [LC03c]. restarts [GM08b]. Restricted [Cle02, Vol07, AW04, AAER03, BS00a, BLME02, Cai09, CLM09, CY01b, CY02, CT03b, DGHS07, Fis02, Gur08, KY00, LCM09, Luc04a, Pal09, Zie08, ESZ02, LEP07]. restriction [CT05, CT07, Qua02]. restrictions [LLP09b]. restrictive [Bar09, Vaj08]. Restructuring [Gag05]. result [BF09, CPC03, GN06, HKC09, Lib03, PC00, jXyL09]. results [AH08, AK02, ABG05, BHY09, BS08a, BBM06, Che09c, DCS07, EMP06, Gia09, Goo09, IT03, JCP08, KG05, KL04a, LR03b, Mon02, Nor09, Sch09, ZŽ09, AIM08]. retrieval [GCC00, KdW04, Kno05, LKP01, Low02, SL01]. Retrieving [CC06b]. return [Hoe01]. reusable [CC06b]. Reversals [LM08b, Ars08, BCP08]. reverse [CKY06]. reversible [Che04b, GdPL08]. reversing [CDFM05]. Review [LJ03]. Reviewing [FM05]. revised [FMCW02]. revisited [BLV03, DG00, Gag05, HKL^{+}04, HPW08, Kir02, Ohl02, Vea00]. revocation [DHS09]. reward [YC09]. rewrite [NTT00, Vág06, Vág08]. Rewriting [Van01, Fer05a, Jac03b, LMM05, Ohl00, Sim04, TK09, Ver02]. RFID [vDR09]. RGSS [LKP01]. Rhee [Küh08]. rhythmic [CT03a]. rich [Amb00]. Right [Pal03b, HR03]. Right-arm [Pal03b]. right-to-left [HR03]. Rigidity [MN08]. Ring [CH02, HHLH03, LLC09, LTH05, LW06, SK08b, MM05]. Rings [Göb00, AS07a, BG09, Par08]. RNS [DIPS00]. Robot [WA01]. robots [BBMR08, DP07]. Robust [Doy07, BLdR04, SY08]. Robustness [CPX06, PW07]. Rogue [Shi08]. Rogue-key [Shi08]. role [Rao01]. roommates [ABE^{+}09, Irv07]. root [GS00]. rooted [KFS01, NO07, Nut09]. rooted-tree [NO07]. roots [DD06, IK05, KCYL06, QW06]. Rotation [CS09a, RBF08, Cle02, CT03b, HHP08, Luc04a, LEP07, Pal00, Pal03b, Pal09, SM03]. rotator [HS01, KHLL09]. round [CKL^{+}03, HSL^{+}02, MSS09, Pha04]. rounding [XX05a]. roundings [Doe04]. routed [Che00a]. routes [DB06, Tan01a, Tan01c, TH03]. Routing [NPZ01, BS00a, BG09, FKW09, FGLW04, GHZ02, Lai07, LTT06, Nag09, PP05b, Sch02b, Shi03a, Shi03b, Šte00, Tse05]. routings [SZQ08]. RS [PHK01]. RS-tree [PHK01]. RSA [Shp04, hY08, YPKL08, ZC09a]. RSA-based [YPKL08]. rule [BtN05, Cha02a, VB08]. ruler [MY09]. rules [CFM09, HC02, Nau01, Vág06]. run [AYTH08, ALM02, FB04, LHWL07, SV08]. run-length [AYTH08, ALM02, SV08]. run-length-encoded [FB04, LHWL07]. runners [CG08]. running [Ars08, AF08]. Runtime [KLK06].

S [GS05b]. S-boxes [GS05b]. safari [TH03]. safe [Xir06]. safely [AV07]. safety [BH08a, CA08]. SAGBI [Göb00]. Saitou [XDZ05]. Salembier [Hes03]. Salesman [ADLP04, ABL08, Jon02, Mon02]. sample [Lon03]. Sampling [Høy00, ES06]. Sandwich [BdFdS06, dFdS05, TYW01]. sandwiches [HLP03]. SAT [Kut07, Li00, Pap09, San01b]. satellite [KKL03]. satisfaction [Dal09, Hås00a, Koi06, LMG07, YC09]. satisfiability [BHM08, Mad06]. satisfiable [NR08]. Satisfy [FR06]. satisfying [CT03a, CC05a, QPV05]. savings [Juk04]. Scaffold [CT07, CT05]. scalability [Lyo02]. scalable [BD05, SB04, ABPH07]. scalar [BHMR01]. Scaled [BEL04]. Scaling [TV02, IMS00, TMSO08]. scattered [FM03]. Schay [CWTHR01]. Schay-Adams-Calabrese [CWTHR01]. Schedulability [CLL08, DG00, Par11, WLLS08]. schedule [XH05]. scheduled [AMR01]. scheduler [Tur07]. Scheduling [Bar01b, Kwo02a, LC02, WLP07, WX05, WL03, ABG05, Aud01, BME02, Bre04, CLL08, CL03, CNK06, CL06, CCG08, Du04, Eps00, Fun08, GN00, GGK03, GH00a, GM08b, GL09b, HGYZ09, JC07, JC08, KC03a, KC03b, KPS02, KY06, KY07, KC02, LKP01, LCLP09, LLPZ07, Lev06, MO06, NS02, NCY02, NCBJ02, PG06, PYZ09, QH01, Rem04, ST09, Sch08, SB02, TH02, TMSO08, TCNY09, WYL01a, WYL01b, WC01, Woe09, YZ03, ZSZY08, ZCF^{+}06, ZXZ08, ZXC05, ZLT02]. schemas [BEL08, MAC03, Van01]. scheme [ALO^{+}09, CC02, CKN06, GH00b, GS09b, Hes04, Hwa00, KHL09, LW06, NRS00, Pad05, PCC03, Sae00, Sae02, Sha03a, Shi03b, Shi08, SB04, Sti04, Tan06a, Ton06, Wan04, WHLH03, hYScH05, hY08, ZC04, ZC09a]. schemes [BCM02, CDFM05, DHS09, DM00a, EL08, He02, IY06, Küh08, LHC^{+}00, LHC^{+}01, MF07, Mia03, Mor06, PS02, RK09, DFM04, Sha03b, hY09a, vDKST06]. scopes [Heh01]. Search [LH04, CN03, CV03, EQ04, Fre03a, Gab00, Gag05, Gia09, Gur08, HR05, Hes03, JŁ09, KS02a, KHN05, KW05, KA02b, Kuj09, KM03b, LKL^{+}00, LK02, Lev06, Pes00, RR01, RL02, SPKL07, Tsi02, FGK^{+}09a, Ver06]. searchable [Dur04]. searchers [YSK04]. Searching [YSK04, Bra07, Che00b, Che06b, DS01, FG06, LPC02, Liu04b, Mul09]. second [Bar01a, YHT04]. second-order [Bar01a, YHT04]. secret [DM00a, IY06, MF07, PS02, Sch01, vDKST06]. secrets [GD05]. secting [KPC07]. Secure [GD05, BBD02, BD05, CHH^{+}09, IY06, MSS09, RG09, hY08]. securing [Kwo02b, Kwo03]. Security [Bla05, KHL09, LHC^{+}00, LHC^{+}01, Sha09b, Shp02, GH08, Gou01, Hes04, HSL^{+}02, LC03a, Pae03, Rao01, Sae00]. Seed [MY09]. Seeding [MAL08]. segment [AMS08, ADK06, CLC09, GDN07, KRD08, Kim03, PS03]. segmentation [Sep05]. segments [Ary02, BD09, CC05a, DM08b, HT03, LSY07, MKGB08, MCL06]. selection [CP02, GNAJT09, GS03, JLY07, Low02, Tre03, dBT08]. Self [HHT05, CC02, DHS09, DL00, GH00a, God02, GRW06, HJS03, HP00a, LTH05, NSZ00, Sae03, Sha09b, SGH04, SM02, Tsi07, Tur07, TH09, TJH07, TJH09, VL04, Wan04, YSK03, ZH07]. self-adaptive [SM02, ZH07]. self-adjusting [CC02]. self-certified [Sae03, Sha09b, Wan04]. self-diagnosable [YSK03]. self-healing [DHS09]. self-maintainability [VL04]. self-similar [NSZ00]. self-stabilization [GH00a, HP00a]. Self-stabilizing [HHT05, DL00, God02, HJS03, LTH05, SGH04, Tsi07, Tur07, TH09, TJH07, TJH09]. semantics [AHV03, BBD02, CVEOV08, CFvG08, CKCB02, Jon03a, MG01, NG05, NRW09]. semi [Du04, LLP09a, Low06, MM07b, MOY06, Oku09, Rug03, Sch08, Zel07]. semi-conditional [MM07b, Oku09]. semi-constructor [MOY06]. semi-matching [LLP09a, Low06]. semi-online [Du04]. semi-related [Sch08]. semi-streaming [Zel07]. semi-sum [Rug03]. semidistributive [JN06]. semijoin [LTV04]. semirandom [BV08]. semiring [AHM08, LMG07]. send [ZLT02]. send-graphs [ZLT02]. Sending [Car02]. sense [BV00a]. sensing [FJ08a, IS09]. sensitive [Gue01]. Sensitivity [GNAJT09]. sensor [DB06, FJ08a, GRH09, KSRI03]. Separability [CG06, SS02]. separated [Cha08]. Separating [BFMZ01a, BFMZ01b]. separation [GP01, RRW94, RRW00, Seg05, WHO09]. separations [HY09b]. sequel [Mai03]. Sequence [SL01, CDF^{+}04, CKCB02, CLT07, HC03a, Kim03, KY07, Kur08, Man05, Mar08, OK04, WS01a, WS01b]. sequences [CMS03, Dam06, Dav04, DKV09, Doe04, FSGM03, GVFS00, Hon05, HYT^{+}08, LC01, LY07, MA06, Mor05, RY08, Sch00a, SW05b, SW07b, Wu09b]. Sequential [HM02, HP00b, CKS09, IS08, SG05, Ves06]. sequentiality [Luc04b]. serial [BB09, CGM06, LKY02]. series [BS09, HLY09, Hon02, KPL04, Lav08, Lit03b, Sch00b]. series-batching [HLY09]. server [BCNR02, BL00, FKW09]. servers [AKE00, CS00, Low02]. service [CC02, JC08, Lin08a, VLK06, Woe09]. service-oriented [Lin08a]. services [AHS^{+}03, BB05, Che04a]. servicing [YC09]. session [CB07]. Set [BdFdS06, BLdR04, Kol04, LHS01, Mie08, Sey01, dFdS05, ABM09, BS00b, CGK08, CKK02a, Cha03, CHL^{+}08, CKW09, DM00a, DFMR08, DM08b, FGR02, FL05b, Fis02, FLP00, Fuj01a, Fuj01b, GM08a, GS09a, GDN07, GHM07, Gup08, HLP03, IN04, Jia05, Jia06a, Kat04, Kou05, Lie08, LW05, MB08, MCL06, MKGB08, Pre00, Ram07, SD09, Sas08, SGH04, Tal00, TYW01, TZ05, Tur07, UN07, Win04, Wot01, Zho08, BHM06, BM02b, Cai09, CCGP05]. Set-based [LHS01]. sets [AK09a, Cha04a, CKSY07, CKL01, Dav04, DDD08, Elm09, ERS05, Gav00, Gav08, GRW06, HHL05, HM08, Hes07, HH09, HN08, KMU05, Kas04b, KOH09, KR03b, KHLL09, LSL08, LJ03, LM09, Maf09, PW07, SSZZ00, TH09, Vág06, Ver07, WWC04, Yus06, dWH03]. setting [GMLS02]. setup [GN00, KY07, MO06]. setups [KW01a, KW01b]. several [FM03]. Shacham [Hes04]. shading [CY01b, CY02]. Shannon [Gag07, VŽ02]. shaped [LWŻ07]. Shapley [Li09]. shared [GD05, HKL^{+}04]. sharing [ABBG06, BS08a, Cho02, IY06, LCKL06, MF07, MSS09, PS02, SB04, vDKST06]. sharp [Nau01]. sharpened [BHK^{+}00]. Sharpening [LTV03]. Shellsort [Bre01]. Shift [Fre03b]. Shift-or [Fre03b]. Shin [Küh08, Küh08]. shipping [Dam02]. shop [Bre04]. Short [Eis01, CRW07, DCS07, HHHK08, LHL08, Tse05, Vog08, Sch01]. Shortest [ESZ02, KG04, Tan01c, AF08, CC05a, DM09, Gal03, GL09a, Han08, KS05, KN04b, LSC06, LL09, MPSS02, Mis01, NPW01, RR09, San01a, Sib04, Tak05, Tan01a, TH03, Yij04, CM07]. shortest-paths [MPSS02]. shot [SS02]. Should [Kle08]. shuffle [EKF00, KDTS01, KR03b, LTHS01, XXZ05]. shuffle-based [KR03b]. shuffle-cubes [LTHS01, XXZ05]. sibling [BS08b]. side [GS04a, HRC08]. side-channel [GS04a]. sign [Koi02]. signals [DL04]. Signature [Che02, Che06b, Hes04, HWW02, IKKY00, Kwo02b, Kwo03, LHC^{+}00, LHC^{+}01, Sae00, Sae02, Sha03a, Sha03b, Shi08, SV08, Sti04, Wan04, WHLH03]. signatures [CKK03b, GMLS02, LV07, Sha09b]. signcryption [Tan06a]. Signed [LC09a, KL01, Lee09]. signers [LHC^{+}01, Sae02, Sha03a]. significant [Shp02]. Simha [CT03a]. similar [BM02b, NSZ00]. similarities [MY09, Par08]. similarity [AK09a, GCC00, Gur08, LKL^{+}00, LW08, Pes00]. Simple [CC07, Kur03, LPC02, Ric00, Alo03, AMR01, AYTH08, BCP09, BW01, Bol03, Bol05, BS09, CKM^{+}01, CG00, Che06a, CDF^{+}04, CS00, Chu08, DM08a, DH03, FJ08a, GL09a, GD04, Hon02, Iba09, JQQ^{+}03, Joh04, Jun05, KV04, KR03a, KG04, Lec04, LM06, LCC06, Lin07, LHT00, MN08, MTF02, MNRR09, NSI04, Pan05, Pet07, PS09, Pre02, PS03, SS03b, Sax09, Tan01a, Tan01c, TH03, Tan07, Vid03, Wan01, WCK04, Win04, XH05]. simple-equivalence [BCP09]. simple-PAC [CG00]. simpler [GM08b, PS04b]. simplex [Kou06]. simplification [MC06]. Simplified [Eng04, Høy00, Nor09, Weg02]. Simulated [Mee07]. Simulating [DPS02]. Simulation [Xir06, BHK04, Las09, PRSS01]. simulations [LA01]. simultaneous [FM03, Xir06, ZH06]. Single [EM08, KY07, MO06, WYL01a, WYL01b, ABG05, Bas08, Fuj03, JTG04, Khu99, Khu00, Kol05, KY06, Lop00, Mäk00, NCY02, SW09, WL03, ZXZ08, ZXC05]. single-address-space [Lop00]. Single-bit [EM08]. single-hop [ZXC05]. single-point [JTG04]. single-source [Kol05]. single-tree [Mäk00]. singular [BGHP09]. sites [RK03]. situ [Kel02]. situated [JX09]. sixteen [Sop02]. Size [Ary02, BN01, BW03, Bol03, Bol08a, Bol09a, CKK02a, Cha03, CZ09, CPC03, CHLN07, DV09, DPS05, Fer05a, FM05, GPS01, HŽR08, HM03, KPS02, LTH05, Lif03, LC08, MM07a, OHNT02, Pud00, Saw07]. size-change [Fer05a]. Size-height [Ary02]. Size/lookahead [BN01]. sizes [Det09, HR06b, HRWZ07, PC00]. skewed [CPC03, PC00]. Ski [LPSR08]. Skipjack [HSL^{+}02]. Skipjack-like [HSL^{+}02]. SkipNet [HM04]. slices [GM02]. Slicing [Amt08, DGG03]. SLP [MS04]. SLP-compressed [MS04]. Small [Sch01, ALO^{+}09, BCD^{+}05, Blä03, COP00, DM08a, ERS05, Has00b, KN04a, KKM00, Lib03, LS09b, WS07, XP06, YZ03]. small-step [DM08a]. small-world [COP00, XP06]. Smaller [Sch06, WWW08]. Smallest [DGN05, GS00, GRW06, KRD08, NN05, NN08]. smooth [Pol03]. snapshot [Che03, RRT08]. snippet [KAS08]. snow [LMC^{+}09]. SNP [Wu09b]. social [DH06]. society [Sae02, Sha03a]. Software [LR03a, CC06b, HŽR08, Jac03a, KTK02, Mai03, Par03, PK02, RB07, Wir03, WS06]. solid [HR05]. solution [CT00a, Gut02, Jon03b, LMC^{+}09, Mel07b, TC04, Yun08]. solutions [AD06, BHW03]. solve [BD08, IR09, Li00]. solved [BM02b, MM07a]. Solving [Fis02, Tam00, Deo09, MAL08, MUK04, TYW01]. Some [BN07, Che09c, FR04, LR03b, MVM01, Ric03, SK08b, Tsi02, Van03, Bar09, BFMZ01a, BFMZ01b, BEF^{+}00, Che07, DHS09, DS01, Giu04, He02, Jai09b, Jai09a, Mon06, RK05a, TYP08, Wu00, ZŽ09]. Sort [FGK^{+}09a, BF09, KA06]. sortable [SV09]. sorted [dBT08]. Sorting [Gag08b, HC03a, BCP08, BS00c, GS04b, IAIH03, KP09, Kub06, LTT00, RS05, Sas02, WS01a, WS01b, ZM03]. SOS [MRG05]. soundness [ŢM05]. source [Kol05, LTT06, NII01]. Space [BDLP08, DS01, Pet05, Agr02, AK08, Ary02, ARdAdS07, BLMM06, CTN03, DHS03, ET03, GH05, IAIH03, Juk09, KA02a, LHK02, LTH05, Lop00, PG09, Rao02, Ric00, Sak06, Sze06, Sze08, Val04b, WCK04, HRC08]. space- [KA02a]. space-based [CTN03]. Space-efficient [Pet05, LTH05]. Space-filling [HRC08]. Space-time [DS01, Juk09]. spaced [SJ08]. spaces [CN03, LK02]. spanned [BD09]. spanners [Bas08]. Spanning [BZLN07, GMM^{+}07, AAHK06, ABBG06, BFK09, CY00, Has00b, JR04, KY00, KMU05, LM08a, LT06, LTT06, Mar02, Mon01, NNR04, RK05a, SW08, WCX02, YAM04, ZNI03, ZSN02]. Sparse [LSZU03, BHLV09, HSWC09, Kow07, ME04, Pan05, SW05a]. sparsest [HR06a]. Spatial [CC09, PC00]. special [AG01, CIL^{+}03, RGDN04]. specification [BB05, Bij01]. specifications [Dia08, MP02, PU01, SSSS02]. specificity [GNAJT09, RC06]. specified [He02]. Spectral [QPV05, CS03a, Cha04b, Sar00]. spectrum [Sza03]. speed [AL02, Fre05, NR00]. speed-up [AL02]. Speeding [Deo06]. Speedup [TAI04]. spell [Dij01]. Sperner [Gri01]. spider [FGK09b]. spikes [BtN05]. Spiking [OR05, BK05, BtN05, IK05]. spin [KA02a]. spine [BD08]. splay [AK02, Geo08]. splaying [Geo08]. splicing [BFMZ01a, BFMZ01b]. splitable [DL04]. splittable [Kol05]. splitting [Che04a, Wu08]. SPNs [CKL^{+}03]. SPT [Bun04]. spurious [Sed07]. Square [HHK08, DGN05, KCYL06, Kur08, NKSK09, QW06]. Square-free [HHK08, Kur08]. squared [KP06]. squarer [LY00]. squaring [Kle08]. SSA [HG06]. SSA-form [HG06]. Stability [Cho08, ÀBD^{+}04, AC02, BLV03, Li09]. stabilization [GH00a, HP00a]. stabilizes [HJS01]. Stabilizing [DGR05, DL00, God02, HJS03, HM01, HHT05, LTH05, SGH04, Tsi07, Tur07, TH09, TJH07, TJH09]. Stable [BM02b, BLdR04, Maf09, ABE^{+}09, CF02, DCS07, HIM09, Poo03, dWH03, DS04b]. stack [Niv04]. stacks [SV09]. stages [DHS08]. Standard [Pha04, JLY07, KS02b]. standards [Tsa06]. Stanley [BF09]. star [ČKRR09, CL00a, CY00, HHL05, Lat07, LWŻ07, LP09, Nag05b, WWC04, ZLR^{+}05, yZyLZ09, tC09, KZ02]. star-free [tC09]. star-patterns [Nag05b]. star-pyramid [yZyLZ09]. star-shaped [LWŻ07]. starlike [yH02]. start [Kol04]. State [LS07a, LA01, VSSV05, BBS06b, BB05, Dar05, DHS03, Hie04, Hoe01, KMST07, Las09, OR05, Ram06]. State-level [LA01]. state-space [DHS03]. statecharts [MST02]. statement [HC03b]. states [BLN07, CDC04, Pre00, Yun08]. static [Dur04, HL03, YCKK07]. stationary [YSK04]. Statistical [NSZ00, KAS08]. statistics [CP02]. Steiner [FFFdP07, KZ02, DH04, Fuc03, GHNP02, Gut08, HRWZ07, Kas04b, LX02, MZ00, NWZ02, Riz03, WL02]. step [DM08a, Mai03, NP07]. step-by-step [Mai03]. steps [Bih02]. still [EG03, Pec07]. Stochastic [DM09, JLY07, CH08, Dob04, GL09b, KM07, PRSS01]. stochasticity [DV04]. storage [ACKM07, AKE00, Bol08b, LPM05, Low02, ME04]. Storing [TS02]. straight [ZH06]. straight-line [ZH06]. Strategical [ALS09]. Strategies [HJ05, AKE00, MAL08]. strategy [CC01a, Sar08]. Stratified [AD07, EN07]. stream [DK08, PCC03, WWGF05]. streamed [Gag08b]. Streaming [Bas08, GNR08, Zel07]. streams [McI01]. Streett [Hor07]. strength [DD06]. strict [HK09]. String [Pet07, Ars08, BEL04, Dai09, DR06, Deo06, Fre03b, FM06, GF08, HR03, Hyy08, Lec07, LHWL07, Mis03]. string-matching [DR06, HR03]. stringent [CW04]. strings [Aku06, AYTH08, ALM02, DKV09, DV04, FB04, Giu04, KP09, Raf01]. Strong [IN06, Luc04b, NCY02, NRW09, SHC08, SCHT08, BUŽ02, CK03, DM05, Gab00, HK09, HP05, HC03a, JHfS08, KV05, Liu04b, SS03b]. Strongly [GIS04, IY06, CLZ09, Weg02, ZNI03]. strongly-connected [ZNI03]. Structural [ŢM05, Živ09, EG07, IN03, KLK04, MK05]. Structure [BLV03, BL06, Amb00, CH06, GM08a, GL04b, HSL^{+}02, MF07, PS02, PHK01, SSYM04, Vu05, Xir06, ZH02]. Structured [Par03, AFK08, CKK03b, CDL^{+}04]. structures [AN09, Amt08, FMCW02, Gur08, HL03, IY06, KdV01, LC06, Lib03, Zie08]. study [Dia08, GRH09, Lat07, VSSV05, YCKK07]. Sturmian [IMP06, MRS03]. stutter [Ete00, TD09]. stutter-invariant [Ete00, TD09]. style [LZ04]. sub [PG09]. sub-quadratic [PG09]. subclasses [LM09]. subcolorability [LL02a]. subcubic [KMŻ06, Sku04]. subdivisions [HKKL09]. subfamilies [DPS05]. subgraph [CKK03a, LMZ08, RS07b, YAM04, YEM05, ZNI03]. subgraphs [ABF^{+}08, CR09, KKM00]. subgroup [EHK04]. Subject [Ano01e, Ano01h, Ano01l, Ano01i, Ano01j, Ano01k, Ano02o, Ano02l, Ano02m, Ano02n, Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano06y, Ano06z, Ano06-27]. sublogarithmic [Agr02, Sze06, Sze08]. Submatrix [HG02]. submodular [IMS00]. submultisets [Hag03]. subpartitions [NK07]. subsequence [AAN^{+}07, AYTH08, Che00b, Dam06, Deo09, FB04, GH05, HYT^{+}07, KPL04, MK03, Sak06, Tsa03, YHC05, CIPR01]. subsequences [BS00c, Ric00]. Subset [CKP00, CP05b, CKSY07, Elm09, Reg02, WS07]. subsets [Sch01]. substitute [Bih02]. substitution [RBF08]. Substitutions [Kra07, Ars08]. Substring [KW05, LSZU03, Ars08, WCH08]. subsystems [NR08]. Subtraction [LH05]. Subtraction-free [LH05]. Subtree [ES07, AFT08, IP09, LHC^{+}05, PT06, SLT08]. subword [AL02]. subwords [PĆB04]. succeed [LSS07]. Success [MH09]. successful [HLSW02]. successor [ZLZ05]. succinctness [Lan08]. suffice [Hit03]. Sufficient [BCD^{+}05, BV09, Li06, SW07a, ZLZ09]. suffix [CIR08, Maa07, NKSK09, Rao02]. suffixes [BN04]. suite [Nik06]. Sum [CKP00, CP05b, CC05a, CLC09, GK00, KMCC04, Koi06, MRST09, OT03, PS03, Rug03, Shp07]. summation [HL05]. sums [BMCK04, QW06, TV02, dBvdS02]. sums-of-pairs [BMCK04]. Super [LM08c, LCM09, XLMH05, Alb02, CLHH04, Fre03b, HZ09, Juk04, MLX08, MTXL09, TTH04, XXZ05]. super- [HZ09]. super-alphabets [Fre03b]. super-arc-connected [LM08c]. Super-connected [LM08c, TTH04]. super-linear [Alb02]. super-polynomial [Juk04]. superconcentrators [Sch06]. superlogarithmic [CHLN05]. superstrings [KS05]. supports [FG06]. sure [LS07a]. surfaces [CFS09]. Survivable [JR04, Nut09]. SVD [TLZ05]. swapped [ALP02, Par05]. swarm [CCZT08, JLY07, SLL^{+}07, Tre03, ZSZY08]. Sweeping [Tan07]. switch [NSZ00]. switched [Kwo02a]. switching [EHHR00, Ros02]. symbol [Alh06]. symbol-objects [Alh06]. symbolic [BdATdC04, CDFG01, OR05]. symbols [CIL^{+}03]. symmetric [AW04, CCKL02, Chu04, LC06, LM08b, SM03, Wol06, WW09]. symmetrically [KdW04]. symmetrically-private [KdW04]. symmetry [CY01a, Han06, MP09, MUK04, RBF08]. synchronization [BW01, Hon07, JX09, Yun08]. synchronized [DV06, ŢM01]. synchronizing [GING09]. synchronous [KR03a]. synonym [CFM09]. syntactic [MRG05]. system [AMM05, BD05, CH03b, CCKL02, Chu04, Fis02, HP00b, Jia05, Joy03, Las09, RB07, TK09, ALO^{+}09]. systems [Alh06, ÁAC^{+}08, AMR01, BLN07, BH08b, CC01a, CLW02, CLL08, DM00c, HC08a, HLSW02, Hon07, IM06, Jac03b, KMST07, LLK08, LK00, Luc04b, LMM05, MVM01, Min07b, MRT00, MP02, Ohl00, QH01, Sch02a, Sin03, SB02, ŢM01, TK02, TJH07, Vág08, VFV02, YSK03, ZC09c, vdZ01]. systolic [KLY00, LY00, LKY02, MNR04].

table [Fre03a]. tables [RBF08, ZHW01]. tagging [ALA06]. tags [RRB00]. Takaragi [WHLH03]. taken [RB07]. tally [RRB00]. Tamari [Pal03a]. tandem [KCH03]. tape [Pet02]. Task [Che00a, MCS^{+}09, BLN07, KC02, QH01, SB02, TMSO08]. tasks [AAER03, Bar01b, KLK06, KPS02, WLP07]. tautologies [Kra07]. TCP [KKL03, NSW02b]. teachability [Ser01a, Ser01b]. team [HŽR08]. teams [tBK05]. technique [AMR01, CKK02b, Low02, vdZ01]. techniques [BH08a]. temporal [CDC04, SSSS02]. temporary [AAER03]. term [Jac03b, LMM05, Ohl00, RC06, Vág06, Vág08]. terminal [DH04, Fuc03, Kas04b, LX02]. Termination [HW06, AGGZP09, Fer05a, Fer05b, GK04, LMM05, Ohl02, RSV07]. terms [RC06]. ternary [Kur08]. terrain [Eid02, ZZ08]. terrains [EHP06]. Test [Nik06, CHU06, DG00, Hag07, Hie06, JUZ05]. Test&Set [IR09]. testable [Boj07, BW03]. testing [BZLN07, BHLV09, Boj07, CG04a, Hie04, KTK02, Lon08, MP09, NRS00, RB07, UW03, YZ08]. tests [BHW06, Fre03a, Lyo02]. Tetravex [TW06]. Tetris [HK04]. text [ALU02, ZHW01]. texts [CIR08]. their [Chu08, Jon03a, LTHS01, Nik00, Sav04, Woe01]. Theorem [ACHP09, KMST07, DM00b, DL04, GRT00, Pol04, ST04a, Høy00, Vea00]. Theorem-proving [KMST07]. theorems [BK04, Dia04, GO05, Koi02, Tri06]. theoretic [Lan08]. theoretical [BH08a, AK02]. theories [SL06b]. theory [ACIL08, Amt08, CCG08, DJ07, LR03a, SK08b, Tal00, Zho08]. therapy [LSY07, Wu08]. There [Dal09, Sch02b, Sop02]. thermal [MB08]. thick [Vig03]. thickness [KY03a]. third [HH09]. Thread [SKP^{+}02]. threads [SKP^{+}02]. Three [CRW07, NCBJ02, AK08, BE05, JKŠ05, MF07, WC09]. Three-coloring [CRW07]. threshold [BW06, Boj07, Bol09a, Cha04b, HWW02, LHC^{+}01, Mia03, Nik00, She07, WHLH03, YC09]. threshold-based [YC09]. throughput [YZ07]. Tight [CZ02, ST05, Eng04, FS09, WCK04, ZCF^{+}06, dFdF03]. Tighter [BA03a, LW09]. Time [DKV09, LM04a, Rao02, RK09, AKL07, AS07b, Ars08, AF08, BJK02, BHK04, Bas08, BHY09, BA03a, BFK09, BF03, BDLP08, BM02b, BL06, BZ02, CCGP05, Cha08, CLL08, CDC04, CWC00, CLZ09, Chu08, CCG08, CS03b, CI08, CKW09, DDR04, Der09, DL06, DD06, DS01, EHRS09, FKM06, FL02, Gol01, GF08, GH05, HG06, Hal03, Han08, HLY09, HJS01, HJS03, Hes07, HS03, Hu07, IP09, Juk09, KdV01, KA02a, Kim03, KC03a, KPL04, KLK06, KM07, KL04b, KW01a, KW01b, Kri02, KY06, Kwo02a, LCLP09, LLPZ07, Lev06, LL01, LW06, LC08, MM07a, MP02, MC02b, MR01a, MR01b, NKS09, iNM07, NCY02, NCBJ02, NY04b, OT03, PD03, PT06, PG09, PS04b, Qua02, RY08, RR00, RB07, Sas02, SS03b, Sku02a, Sku06, SV08, Tak05, TMSO08, Val04b]. time [Wil09, WS07, Wu00, WYL06, YZ03, hY08, ZNI03, ZLZ05, Zie09, DK08]. time-based [KY06]. time-bound [hY08]. Time-bounded [DKV09]. time-constrained [Kwo02a]. time-dependent [FL02]. time-efficient [KA02a]. Time-Memory-Data [DK08]. time-optimal [Sas02]. time-series [KPL04]. Time-space [Rao02]. time-variant [Gol01]. Time-varying [RK09]. Timed [BD00, LMST04, BBM06, Doy07, Tri06, ZLZ05, Zha06, vdZ01]. times [GGK03, JC07, KY07, LC02, MO06, Zel07]. timestamped [ALU02]. timestamps [BM02a]. tolerance [HZ09, Lat07]. tolerant [BMP^{+}00, CC01b, CQ09, CHC02, KA06, LTTH03, Shi03a, Shi03b, TTLH02, WMX07, XDX05, XZX07, jXyL09, YLTH03, Yan09]. top [AS07b, FKV04]. top-down [AS07b, FKV04]. Topological [GK04, CMS03, Che09c, ZM03]. topologies [BG08]. topology [Gol03, Sey01]. toroidal [AD09, Jha03, RW09, Xia07]. torus [CH02, CH03a, Sch02b, Shi03b, YELM04]. Total [HSWC09, WC09, BJK02, CG09, Dia09, HLY09, LCLP09, Lee09, LLPZ07, LHL08, NCY02, NCBJ02, WYL01a, WYL01b]. tournament [HC03a, IN04, Sas08, WS01a, WS01b]. tournaments [GHM07, Gup08, IN04]. tours [AMP03]. Towers [Min05, WW03, BF01]. trace [MMV08, BCP05]. traceable [LHC^{+}01]. tracing [Ton06]. Trackless [BL00]. tractability [BV08, GN06, Hag07, Ste08]. tractable [CS09a, NR00, RS07a, Wes04]. trade [DS01, DGP04, MH09, Nor09, Rao02]. trade-off [DGP04, MH09, Nor09]. trade-offs [DS01, Rao02]. tradeoff [BN01, Juk09, RR01, SS03a, DK08]. tradeoffs [Ary02]. Trading [Val04b]. traffic [NSZ00]. traitor [Ton06]. trajectories [KDTS01]. transducers [EM06, FM00, FKV04, PS04a]. Transfer [Koi02, Gho07, GZN06]. transfinite [Sim04]. transform [CY01b, CY02, Deo05, Hes07, MRS03, CC01b]. transformation [AK08, AIM07, AIM08, BB09, Cha02a, DN05]. transformations [AK09a]. transformed [GD00]. Transforming [AAHK06, Ber04]. transforms [CWCL02]. transient [ST09]. transition [vdZ01]. transitions [LMST04, MST02]. transitive [DGHS07, KK02, Kun00, Lec04, TM09, VL04]. Translation [HI03, LSL08, LCL06]. translational [SS02]. translations [IN06]. transmission [ABM09, LM04b, NSR06]. transportation [OB05]. transposition [Deo06]. transposition-invariant [Deo06]. Transversal [CSW03, CW02, CW03, KBEG07, LC09a]. trapezoid [LC09b]. Traveling [ADLP04, ABL08, Mon02, Jon02]. Traveller [Wes08]. Treatment [DK08]. Tree [Cha02c, FFFdP07, PBD^{+}02, Tou03, ADMP01, AIM08, AS07b, ABBG06, AFK08, BDLP08, CC06a, CW07, DDR04, DH04, The02, EQ04, EM06, Fuc03, FD04, Fuj05, FM00, FKV04, FKŻ08, GNR08, Gut08, HL03, Héa08, Hes03, HS03, HC08b, IP09, Kim05, KRS01, Kwo02a, LT06, LX02, Mäk00, Mar02, MSS09, Mon01, MGLA06, MSS06, NO07, NSW02a, PCC03, Pec06, Pen02, Shi02, Sku02b, SG03, SLT08, Tsu08, Vág06, VGG^{+}00, WCX02, WL02, Wu09a, YAM04, ZH05, ZSN02, KZ02, LJL07, PHK01]. tree-covering [AFK08]. Tree-Star [KZ02]. tree-structured [AFK08]. tree-walking [MSS06]. treelike [ET03]. trees [AAHK06, Aku06, AFT08, AK02, AL02, BD08, BS08b, Bra09, Cha05, Che02, Che06b, CC09, CY01a, CY00, Cle02, DN06, Dur04, FGK09b, Gag05, Geo08, GMM^{+}07, GK00, GHNP02, GN06, HK09, Han06, Has00b, HS05, Heu08, HSWC09, HHP08, JLN02, JMV06, JVA06, KS02a, KY00, KMU05, KH03, KR04, Kas04b, KFS01, KV08b, KK04, Kuj09, Lai07, LM08a, LTT06, LCL06, LT07, LEP07, LP09, Maa07, MZ00, MR09b, MUK04, NKSK09, iN02, NWZ02, NNR04, OP08, Pal00, Pal03a, Pal03b, PT06, PK02, RK05a, RS07c, SW08, SGH04, SPKL07, SW09, WC09, XUT00, XUT01, YAM04, ZLJ06, dFdF03, BZLN07, GNU08]. treespan [Rau05]. treewidth [CS03a, CHLN05]. trends [Bra01]. triangle [BC00, BHK^{+}00, Och04, ŠŽ04]. triangle-free [Och04, ŠŽ04]. triangles [BD09, CR02, Wan08, WWW08, Xia07]. triangular [MN08]. triangulated [CFS09, RR00]. Triangulating [HM08]. triangulation [Hu07, KV08a, Liu05, ZSN02, ZJ05]. Triangulations [WCY00, AAHK06, BP06, Ber04, MU05]. tries [Rez05]. triggers [MST02]. trinomials [SJ08]. Tripathi [CL00a]. triple [LWC01]. triple-error-correcting [LWC01]. triples [Yan00]. trivial [CF02]. trivially [Chu08]. TRSs [MOY06]. true [Fec04]. truncated [GS09b]. Tseng [LHC^{+}00]. TSP [ASW05, BC00, BHK^{+}00, BD04, COW05, HR00, HR02, Mee07, Mel07b, SLL^{+}07]. Tumble [Joh04]. tuple [Ara07, KL04a, LC03b]. Turing [LM04a, DG08, FL09, Pet02, San01b, Sze06, Sze08]. turned [Hes03]. turnpike [FM09]. Tutte [BK04]. twin [SDC09]. twisted [LT08a, Yan09]. Two [GNU08, MC06, ASW05, AK08, AFT08, AFG07, AMM05, AC02, BMCK04, BS00b, BB01, Bre04, CL03, CS00, Du04, FJ08b, GL09b, HTT03, HRWZ07, JKŠ05, JŁ09, Jha03, KSY00, KPS02, KCH03, LC04, LH08, LSL08, LPSR08, Lyo02, MNR04, NKSK09, iNM07, OPR07, Pen02, QW06, RS07a, RK03, Rug03, SD09, Saw07, SSSS02, Sma01, Tan00, TH02, UN07, Vah07, WC01, Wan01]. two-center [HTT03]. two-connected [HRWZ07, iNM07]. two-dimensional [JŁ09, MNR04, NKSK09]. two-equal-disjoint [LH08]. two-level [SSSS02]. two-machine [Bre04, CL03, WC01]. two-phase [RK03]. two-watchtower [BMCK04]. Type [AGGZP09, BK04, CH03b, KS02b, KPS02, Sch09]. Type-based [AGGZP09]. Typing [GYY01]. TYT [CKKC04].

Ulam [CV00]. ultimate [ST06]. ultrametric [Heu08]. Unary [To09]. unbeaten [Pec07]. unbiased [CFW01]. unbounded [BHLV09, TCNY09]. uncapacitated [XX05a]. Uncertain [JMM09]. uncertainty [BW01]. uncompressed [LHWL07]. undecidability [BBM06, DV06, MSSW08]. Undecidable [Tri04, Jac03b, MVM01]. undefined [Sol08]. undirected [ACN03, Bar01a, Li09, LL09, Tsi04, ZLJ06]. unequal [HRC08]. unfair [Tur07]. unfolding [CK07]. unidirectional [MNR04]. unification [EN07, NTT00]. unified [ZH02]. Uniform [BLN07, Bar01b, CNK06, Du04, DL00, GRT00, GL09b, KR03a, LLPZ07, Lon03, Ohl00, SK02, TJH07, WLP07, WTC05, Yus06, ZC09c]. uniform-machine [CNK06]. uniformity [Asa08, Shp04]. uniformly [FJ05, GHNP02, MP04, NWZ02]. uniformly-oriented [NWZ02]. uninitialized [Val04b]. union [EA07]. unions [BK00, Eis09]. uniprocessor [Bun04]. unique [BG08, GS00, Lec04, AR06]. unit [KW01a, KW01b, PZ09, dFdF09]. Universal [Bra09, Wan04, BMP^{+}00, ELO08, KR06, Sol08]. universality [LR06]. universe [Bra09]. unknown [AAER03, DP01, JAF06]. unlimited [DM09]. unordered [AFT08, Fre03a]. unpopular [ST00]. unranking [MR01a, MR01b]. unreliable [JAF06]. unsatisfiable [BZ02]. unscrambling [LT03]. unsplittable [Kol03]. update [BF03, Hoe01, RR01]. updates [VL04]. Updating [DG02]. UPGMA [GM07]. uplink [SYL04]. Upper [ELS06, PCW09, Sep05, BP06, BB05, CLL04, CZ02, GING09, HH07, Kut07, LLC09, Li09, LM04b, Lon03, PCW08, Pal00, Wu09b]. UPS [Jia05]. UPS- [Jia05]. Upward [RRW94, RRW00]. use [CVEOV08, Kle08, Pud00]. used [CK07]. user [GMLS02]. Using [BLMM06, BD08, Hie04, Sch09, dWH03, CKKC04, Cho02, CKK03b, DLX09, Dav04, Dil02, EQ04, GH00a, GS09b, HT00, HC02, KJ04, KW05, KS02b, Kuj09, LL01, LJL07, MCS^{+}09, Niv04, Pae03, PCC03, Par11, RC06, SD09, SKA01, SV08, SHH07, Tur07, Van01, Wan04, WLLS08, ZC09b]. utilization [Pre00].

value [CVEOV08, DK08, Hoe01, LA01]. value-level [LA01]. valued [GR03, NRW09, ZŽ09]. values [Kas04a, Poo03]. Varadhrajan [MS03]. variable [BNvO09]. Variables [Heh01, Fis02, Shi00]. variant [ARdAdS07, Gol01, Wot01]. variants [Bre01, EL08, HYCH07, VV01]. variate [Dil02]. variation [CC06b]. Variationally [KR06]. Variations [LT06, Sch00b]. various [BHW06, KS02a]. varying [RK09]. Vazirani [Riz03]. VC [EA07, ERS05]. VC-dimension [ERS05]. vector [CIPR01, KMF04, LZ09]. vectors [BE05, Eis01, FL05b, LCC06]. Veen [BD04]. vehicle [Nag09]. vehicles [GGK03]. velocity [CPC03]. verifiably [Hes04]. verifiably-encrypted [Hes04]. verification [BH08a, KJ04, LL01, MP02]. verifiers [He02, LV07, Shi08]. Verifying [Poo03, Sch02a]. version [SK08a, TK09]. versions [Cho02]. versus [AGM03, BHW03, EHRS09, HP05, Juk05, Lut04, Mee07, Tho05]. Vertex [CCGP05, DN06, HZ09, LCL06, MV08, BDL09, CCF04, CKK02a, CG05, CLL04, CG09, DL08, Der09, FGR02, FLP00, FD04, Gav08, HV06a, yH02, KR04, KR03b, KHLL09, Kur04, LTH08, Lin07, LC08, LC09b, NII01, PS06, PS09, Sas08, Sku06, TZ08, TM09, WWC04, GNU08]. vertex-connectivity [HV06a, NII01]. vertex-distinguishing [CG09]. vertices [AG03, AD00, BLL02, ELS06, HSWC09]. very [Bol03]. via [Eis01, Koi02, Koi06, Lec04, NG05, NY09, Tam00, ZH05]. video [Che04a, LMC^{+}09, Low02]. video-on-demand [Che04a]. Vidyasankar [TAI04]. view [GT09]. view-based [GT09]. viewed [Mis01]. views [Van01, VL04]. violations [CA08]. virtual [WC03]. Visibility [ZH05, OHNT02]. visible [KMS^{+}02, K^{+}02]. Visual [BDN00, CDFM05]. visualize [ARdFPS02]. vital [NPW01]. VLSI [MB08, MC02b]. vocabularies [ZHW01]. voltage [TMSO08]. Volume [Ano01a, Ano01b, Ano01c, Ano01h, Ano01i, Ano01j, Ano01k, Ano02d, Ano02a, Ano02b, Ano02c, Ano02o, Ano02l, Ano02m, Ano02n, Ano03a, Ano03b, Ano03c, Ano03d, Ano03-28, Ano03-29, Ano03-30, Ano03-31, Ano04a, Ano04-28, Ano04-29, Ano04-30, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano06a, Ano06y, Ano06z, Ano06-27]. Volumes [Ano01d, Ano01e, Ano01l, Ano04z, Ano04-27]. Voronoi [AHS^{+}03, Aro02, Asa07, ADK06, CFS09, CPX06, JMM09, KS02a, Kol04, Tan01b, WCK04]. vote [RRB00]. vote-tags [RRB00]. voting [CWC00]. VRP [AD06]. vs [LZ04, SKP^{+}02, To09]. vulnerability [KSRI03, KHL09, MV08].

Wait

[ACKM07, Att02, CL03, WC01, ZXC05]. Wait-free [ACKM07, Att02]. Waiting [CS03b]. walk [Mis01]. Walker [CWTHR01]. Walking [CDKK02, MGLA06, MSS06]. walks [BCNR02]. Wang [LHC^{+}01]. was [Tur03]. watch [Joy03]. watchman [AMP03, Tan01a]. watchtower [BMCK04]. wavelength [BG09]. wavelet [LH04]. way [LPM05, Nor09, PC00, RH02]. ways [FJ08b]. Weak [KN04a, Pal09, BHK04, CFvG08, DP07, Löd01, NSW02a, Van03]. weak-bisplit [Van03]. weakest [FMR04, Lei05, YCFD07]. weakly [IM06, LJ09, TH09]. weakly-adaptive [IM06]. Weakness [YPKL08]. Weaknesses [He02]. web [DC07, CB07, KAS08, SB04]. web-like [DC07]. Weber [AAK09, Sed07]. Weight [BM02b, BLdR04, BBS06a, BCG09, CR09, DS00, FKM06, Gav00, Gav02, Gav08, HC08b, KL01, Lit03a, PS04b, WCY00]. Weighted [DN08, DSV08, ES06, EMP06, BJK02, BW06, BG09, BBM06, Chu00, CKW09, DN06, DH03, GMM^{+}07, Hal03, HV06b, IP09, KC03a, Kun00, LCKL06, LCLP09, LLP09a, LLPZ07, Low06, Man05, Shp07, UC00, WYL01a, WYL01b, XL00, dFdF09, CCGP05, LH03]. welfare [DH06]. Well [BJ08, Cha08, CK07, Weg02]. Well-balanced [BJ08]. well-foundedness [CK07]. well-known [Weg02]. Well-separated [Cha08]. Wheeler [Deo05, MRS03]. where [LSS07]. which [IT03]. while [JUZ05]. widest [DMN09, DBLS06]. Width [BLME02, Bro05, Dal09, NR08, iO09, Pap09]. width-4 [Bro05]. width-5 [Bro05]. width-parameterized [Pap09]. Width-restricted [BLME02]. wildcard [CC07]. Wilke [Ete00]. will [Jac03a]. window [BMX05]. winner [Pro08]. winning [Ser03]. Wireless [NY09, FJ08a, GRH09, HN08, KM03a, Seg08, YC09]. wise [AGM03, KTK02, SSZZ00, GY08]. with/without [KS02a]. within [NEK03]. without [Alh06, ABG05, BBMR08, BM02b, CHZL09, CRW07, HM09, KS02a, LV07, LHL08, LR03b, LCW07, RG09, Sed07, UW03, WCY00, WWW08, Xia07, ZS08, ZSN02]. witnesses [GKL09]. witnessing [Liv09]. WMT [Mad03]. Word [EQ04, CdL02, CD09, VSSV05]. words [ALS09, CT03a, DJ07, DP07, FR04, GLS08, HHHK08, HHK08, MRS03, MS04, Ram07, Vaj08, VS01]. Workbench [KJ04]. Worker [MCS^{+}09]. workflow [LHS01, ŢM05]. working [BH08b, YZ03]. works [CT03a, Hoo01]. world [COP00, XP06]. worlds [WC03]. wormhole [Che00a, Shi03a, Shi03b]. wormhole-routed [Che00a]. Worst [BG01, KY03a, Weg00, ACRS04, GF08, GPS01]. Worst-case [BG01, ACRS04, GPS01]. wrap [WMG00]. wrap-around [WMG00]. Wright [Rav07]. write [ALO^{+}09, KBOR05]. writes [Tho05]. writing [Jac03a]. Wu [CWJT01, MS03].

X [dBT08]. XML

[CK02, CKK03b, MAC03].

Yates [Hyy08]. YCN [Hwa00, hYScH05]. yield [SSZZ00].

zero  [BdW03,  SK08b].

zero-divisor [SK08b]. zero-error [BdW03]. zeros [LD04]. zigzag [CFW01]. zookeeper [Tan01c, Jon03b, Tan06b].

References

Aider:2005:DMN

[AA05] Méziane Aïder and Mustapha Aouchiche. Distance monotonicity and a new characterization of Hamming graphs. Information Processing Letters, 96(6): 207–213, December 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alvarez:2008:IMP

[ÁAC^{+}08] A. Álvarez, S. Arévalo, V. Cholvi, A. Fernández, and E. Jiménez. On the interconnection of message passing systems. Information Processing Letters, 105(6): 249–254, March 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Armon:2003:LRA

[AAER03] Amitai Armon, Yossi Azar, Leah Epstein, and Oded Regev. Online restricted assignment of temporary tasks with unknown durations. Information Processing Letters, 85(2):67–72, January 31, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aichholzer:2006:TST

[AAHK06] Oswin Aichholzer, Franz Aurenhammer, Clemens Huemer, and Hannes Krasser. Transforming spanning trees and pseudotriangulations. Information Processing Letters, 97(1):19–22, January 16, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Abu-Affash:2009:IBA

[AAK09] A. Karim Abu-Affash and Matthew J. Katz. Improved bounds on the average distance to the Fermat–Weber center of a convex object. Information Processing Letters, 109(6):329–333, February 28, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Albert:2007:LIS

[AAN^{+}07] M. H. Albert, M. D. Atkinson, Doron Nussbaum, Jörg-Rüdiger Sack, and Nicola Santoro. On the longest increasing subsequence of a circular list. Information Processing Letters, 101(2):55–59, January 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Abravaya:2009:MDD

[AB09] S. Abravaya and D. Berend. Multi-dimensional dynamic facility location and fast computation at query points. Information Processing Letters, 109(8): 386–390, March 31, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Angel:2006:FCS

[ABBG06] Eric Angel, Evripidis Bampis, Lélia Blin, and Laurent Gourvès. Fair cost-sharing methods for the minimum spanning tree game. Information Processing Letters, 100(1):29–35, October 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alvarez:2004:CDS

[ÀBD^{+}04] C. Àlvarez, M. Blesa, J. Díaz, A. Fernández, and M. Serna. The complexity of deciding stability under FFS in the Adversarial Queueing model. Information Processing Letters, 90(5): 261–266, June 15, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Abdo:2009:EGI

[Abd09] Mohamed Abdo. Efficient generation of the ideals of a poset in Gray code order. Information Processing Letters, 109(13):687– 689, June 15, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arkin:2009:GSR

[ABE^{+}09] Esther M. Arkin, Sang Won Bae, Alon Efrat, Kazuya Okamoto, Joseph S. B. Mitchell, and Valentin Polishchuk. Geometric stable roommates. Information Processing Letters, 109(4): 219–224, January 31, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Araujo:2008:NHS

[ABF^{+}08] G. Araujo, J. Balogh, R. Fabila, G. Salazar, and J. Urrutia. A note on harmonic subgraphs in labelled geometric graphs. Information Processing Letters, 105 (3):98–102, January 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Angel:2005:ARB

[ABG05] Eric Angel, Evripidis Bampis, and Laurent Gourvès. Approximation results for a bicriteria job scheduling problem on a single machine without preemption. Information Processing Letters, 94(1):19–27, April 15, 2005. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Ausiello:2008:OPC

[ABL08] Giorgio Ausiello, Vincenzo Bonifaci, and Luigi Laura. The online Prize-Collecting Traveling Salesman Problem. Information Processing Letters, 107(6): 199–204, August 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Abellanas:2009:CPS

[ABM09] M. Abellanas, A. L. Bajuelos, and I. Matos. 2-covered paths by a set of antennas with minimum power transmission range. Information Processing Letters, 109(14):768–773, June 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Almeida:2007:SBF

[ABPH07] Paulo Sérgio Almeida, Carlos Baquero, Nuno Preguiça, and David Hutchison. Scalable Bloom Filters. Information Processing Letters, 101(6):255–261, March 31, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ackerman:2006:NGP

[ABPR06] Eyal Ackerman, Gill Barequet, Ron Y. Pinter, and Dan Romik. The number of guillotine partitions in d dimensions. Information Processing Letters, 98(4): 162–167, May 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arbib:2002:SNE

[AC02] Claudio Arbib and Alberto Caprara. On the stability number of the edge intersection of two graphs. Information Processing Letters, 83(3):173–174, August 16, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Abellanas:2009:AGT

[ACHP09] M. Abellanas, S. Canales, and G. Hernández-Peñalver. An “Art Gallery Theorem” for pyramids. Information Processing Letters, 109(13):719–721, June 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Adcock:2006:QLB

[ACI^{+}06] Mark Adcock, Richard Cleve, Kazuo Iwama, Raymond Putra, and Shigeru Yamashita. Quantum lower bounds for the Goldreich–Levin problem. Information Processing Letters, 97 (5):208–211, March 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aceto:2008:ETP

[ACIL08] Luca Aceto, Silvio Capobianco, Anna Ingolfsdottir, and Bas Luttik. The equational theory of prebisimilarity over basic CCS with divergence. Information Processing Letters, 108(5):284– 289, November 15, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Abraham:2007:WFR

[ACKM07] Ittai Abraham, Gregory Chockler, Idit Keidar, and Dahlia Malkhi. Wait-free regular storage from Byzantine components. Information Processing Letters, 101(2):60–65, January 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aneja:2003:PAO

[ACN03] Y. P. Aneja, R. Chandrasekaran, and K. P. K. Nair. Parametric analysis of overall min-cuts and applications in undirected networks. Information Processing Letters, 85(2):105–109, January 31, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alonso:2004:WCC

[ACRS04] Laurent Alonso, Philippe Chassaing, Edward M. Reingold, and René Schott. The worst-case chip problem. Information Processing Letters, 89(6):303–308, March 31, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Avis:2000:ENV

[AD00] David Avis and Luc Devroye. Estimating the number of vertices of a polyhedron. Information Processing Letters, 73(3–4):137–143, February 29, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.

nl/gej-ng/10/23/20/62/27/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/62/27/34/article.
pdf.

Alba:2006:CNN

[AD06] Enrique Alba and Bernabé Dorronsoro. Computing nine new best-so-far solutions for Capacitated VRP with a cellular Genetic Algorithm. Information Processing Letters, 98(6):225– 230, June 30, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aiguier:2007:SIE

[AD07] Marc Aiguier and Răzvan Diaconescu. Stratified institutions and elementary homomorphisms. Information Processing Letters, 103(1):5–13, June 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alspach:2009:HTG

[AD09] Brian Alspach and Matthew Dean. Honeycomb toroidal graphs are Cayley graphs. Information Processing Letters, 109 (13):705–708, June 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aurenhammer:2006:FLS

[ADK06] F. Aurenhammer, R. L. S. Drysdale, and H. Krasser. Farthest line segment Voronoi diagrams. Information Processing Letters, 100(6):220–225, December 31, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ausiello:2004:ALQ

[ADLP04] G. Ausiello, M. Demange, L. Laura, and V. Paschos. Algorithms for the On-Line Quota Traveling Salesman Problem. Information Processing Letters, 92(2):89–94, October 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ager:2004:FCB

[ADM04] Mads Sig Ager, Olivier Danvy, and Jan Midtgaard. A functional correspondence between call-by-need evaluators and lazy abstract machines. Information Processing Letters, 90(5): 223–232, June 15, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Agnetis:2001:LAH

[ADMP01] A. Agnetis, P. Detti, C. Meloni, and D. Pacciarelli. A linear algorithm for the Hamiltonian completion number of the line graph of a tree. Information Processing Letters, 79(1):17–24, May 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.elsevier.nl/gej%2Dng/
10/23/20/80/25/28/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/23/20/80/25/
28/article.pdf.

Aceto:2001:P

[AF01] Luca Aceto and Wan Fokkink.

Preface. Information Processing Letters, 80(1):1, October 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/84/27/27/abstract.html.

Attiratanasunthron:2008:RTA

[AF08] Nattapat Attiratanasunthron and Jittat Fakcharoenphol. A running time analysis of an Ant Colony Optimization algorithm for shortest paths in directed acyclic graphs. Information Processing Letters, 105(3):88–92, January 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anceaume:2007:MAG

[AFG07] Emmanuelle Anceaume, Roy Friedman, and Maria Gradinariu. Managed Agreement: Generalizing two fundamental distributed agreement problems. Information Processing Letters, 101(5):190–198, March 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aceto:2003:NEH

[AFI03] Luca Aceto, Wan Fokkink, and Anna Ingólfsdóttir. A note on an expressiveness hierarchy for multi-exit iteration. Information Processing Letters, 87(1):17–23, July 16, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Atallah:2008:TCP

[AFK08] Mikhail J. Atallah, Greg N. Frederickson, and Ashish Kundu. A tree-covering problem arising in integrity of tree-structured data. Information Processing Letters, 109(1):79–82, December 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Akutsu:2008:IAL

[AFT08] Tatsuya Akutsu, Daiji Fukagawa, and Atsuhiro Takasu. Improved approximation of the largest common subtree of two unordered trees of bounded height. Information Processing Letters, 109(2):165–170, December 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Almstrum:2001:ESI

[AG01] Vicki L. Almstrum and David Gries. From the Editors of this special issue. Information Processing Letters, 77(2–4):49– 51, February 28, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/68/29/
                                        24/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/68/29/24/article.
                                        pdf.

Abdel-Ghaffar:2003:MNE

[AG03] Khaled A. S. Abdel-Ghaffar. Maximum number of edges joining vertices on a cube. Information Processing Letters, 87 (2):95–99, July 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Albert:2009:TBH

[AGGZP09] Elvira Albert, John Gallagher, Miguel Gómez-Zamalloa, and Germán Puebla. Type-based homeomorphic embedding for online termination. Information Processing Letters, 109(15):879– 886, July 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alon:2003:AWI

[AGM03] Noga Alon, Oded Goldreich, and Yishay Mansour. Almost k-wise independence versus k-wise independence. Information Processing Letters, 88(3):107–110, November 15, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Agrawal:2002:CSS

[Agr02] Manindra Agrawal. For completeness, sublogarithmic space is no space. Information Processing Letters, 82(6):321–325, June 30, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See note [Sze08].

Amir:2007:IAC

[AGS07] Amihood Amir, Leszek Gasieniec, and Riva Shalom. Improved approximate common interval. Information Processing Letters, 103(4):142–149, August 16, 2007. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Afshani:2008:AIR

[AH08] Peyman Afshani and Hamed Hatami. Approximation and inapproximability results for maximum clique of disc graphs in high dimensions. Information Processing Letters, 105(3):83– 87, January 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Andren:2008:FMM

[AHM08] Daniel Andrén, Lars Hellström, and Klas Markström. Fast multiplication of matrices over a finitely generated semiring. Information Processing Letters, 107(6):230–234, August 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arkin:2001:AMQ

[AHS01] Esther M. Arkin, Refael Hassin, and Maxim Sviridenko. Approximating the maximum quadratic assignment problem. Information Processing Letters, 77(1): 13–16, January 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/24/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/24/26/article.
pdf.

Abellanas:2003:VDS

[AHS^{+}03] M. Abellanas, F. Hurtado, V. Sacristán, C. Icking, L. Ma, R. Klein, E. Langetepe, and B. Palop. Voronoi Diagram for services neighboring a highway. Information Processing Letters, 86(5):283–288, June 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Albert:2003:RSP

[AHV03] Elvira Albert, Michael Hanus, and Germán Vidal. A residualizing semantics for the partial evaluation of functional logic programs. Information Processing Letters, 85(1):19–25, January 16, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aceto:2008:EP

[AI08] Luca Aceto and Anna Ingolfsdottir. On the expressibility of priority. Information Processing Letters, 109(1):83–85, December 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Amano:2003:IEC

[AIM^{+}03] Kazuyuki Amano, Kazuo Iwama, Akira Maruoka, Kenshi Matsuo, and Akihiro Matsuura. Inclusion–exclusion for k-CNF formulas. Information Processing Letters, 87(2):111–117, July 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Akl:2007:PPT

[AIM07] Selim G. Akl, Md. Kamrul Islam, and Henk Meijer. On planar path transformation. Information Processing Letters, 104 (2):59–64, October 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Akl:2008:PTT

[AIM08] Selim G. Akl, Kamrul Islam, and Henk Meijer. Planar tree transformation: Results and counterexample. Information Processing Letters, 109(1):61– 67, December 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Albers:2002:RST

[AK02] Susanne Albers and Marek Karpinski. Randomized splay trees: Theoretical and experimental results. Information Processing Letters, 81(4):213– 221, February 28, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/85/35/
                                        34/abstract.html.

Aiger:2008:AGH

[AK08] Dror Aiger and Klara Kedem. Applying graphics hardware to achieve extremely fast geometric pattern matching in two and three dimensional transformation space. Information Processing Letters, 105(6):224–230, March 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aiger:2009:GPM

[AK09a] Dror Aiger and Klara Kedem.

Geometric pattern matching for point sets in the plane under similarity transformations. Information Processing Letters, 109(16):935–940, July 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Almeida:2009:CCC

[AK09b] Jorge Almeida and Ondřej Klíma. A counterexample to a conjecture concerning concatenation hierarchies. Information Processing Letters, 110(1):4–7, December 1, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

AlWehaibi:2004:PLP

[AKAE04] Abdullah AlWehaibi, Michael Kadoch, Anjali Agarwal, and Ahmed ElHakeem. Packet loss probability for DiffServ over IP and MPLS reliable homogeneous multicast networks. Information Processing Letters, 90(2):73–80, April 30, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aerts:2000:RDS

[AKE00] Joep Aerts, Jan Korst, and Sebastian Egner. Random duplicate storage strategies for load balancing in multimedia servers. Information Processing Letters, 76(1–2):51–59, November 20, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/67/24/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/67/24/32/article.
                                        pdf.

Abu-Khzam:2007:LTA

[AKL07] Faisal N. Abu-Khzam and Michael A. Langston. Lineartime algorithms for problems on planar graphs with fixed disk dimension. Information Processing Letters, 101(1):36–40, January 16, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Akutsu:2006:RBE

[Aku06] Tatsuya Akutsu. A relation between edit distance for ordered trees and edit distance for Euler strings. Information Processing Letters, 100(3):105–109, November 15, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Apostolico:2002:SCB

[AL02] Alberto Apostolico and Stefano Lonardi. A speed-up for the commute between subword trees and DAWGs. Information Processing Letters, 83(3):159–161, August 16, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alagarsamy:2005:MEA

[Ala05] K. Alagarsamy. A mutual exclusion algorithm with optimally bounded bypasses. Information Processing Letters, 96(1):36–40, October 16, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alba:2006:NLT

[ALA06] Enrique Alba, Gabriel Luque, and Lourdes Araujo. Natural language tagging with genetic algorithms. Information Processing Letters, 100(5):173–182, December 16, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alba:2002:PEA

[Alb02] Enrique Alba. Parallel evolutionary algorithms can achieve super-linear performance. Information Processing Letters, 82 (1):7–13, April 15, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/86/27/
28/abstract.html.

Alhazov:2006:PSM

[Alh06] Artiom Alhazov. P systems without multiplicities of symbolobjects. Information Processing Letters, 100(3):124–129, November 15, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arbell:2002:EDR

[ALM02] Ora Arbell, Gad M. Landau, and Joseph S. B. Mitchell. Edit distance of run-length encoded strings. Information Processing Letters, 83(6):307–314, September 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alon:2003:SAE

[Alo03] Noga Alon. A simple algorithm for edge-coloring bipartite multigraphs. Information Processing Letters, 85(6):301–302, March 31, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ahn:2009:MFW

[ALO^{+}09] Woo Hyun Ahn, Kyungjae Lee, Jaewon Oh, Kyungsub Min, and Joon Sung Hong. A multiplefile write scheme for improving write performance of small files in Fast File System. Information Processing Letters, 109(18): 1021–1026, August 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Amir:2002:ASM

[ALP02] Amihood Amir, Moshe Lewenstein, and Ely Porat. Approximate swapped matching. Information Processing Letters, 83 (1):33–39, July 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arfi:2009:SLI

[ALS09] M. Arfi, B. Ould M. Lemine, and C. Selmi. Strategical languages of infinite words. Information Processing Letters, 109(13):749– 753, June 15, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Amir:2002:OTT

[ALU02] Amihood Amir, Gad M. Landau, and Esko Ukkonen. Online timestamped text indexing. Information Processing Letters, 82 (5):253–259, June 15, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aloupis:2005:LBC

[AM05] Greg Aloupis and Erin McLeish.

A lower bound for computing Oja depth. Information Processing Letters, 96(4):151–153, November 30, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ambainis:2000:HRS

[Amb00] Andris Ambainis. How rich is the structure of the intrinsic complexity of learning. Information Processing Letters, 75(3): 109–112, August 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/28/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/28/27/article.
pdf.

Ambainis:2001:LFL

[Amb01] Andris Ambainis. On learning formulas in the limit and with assurance. Information Processing Letters, 77(1):9–11, January 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/24/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/24/25/article.
pdf.

Appa:2005:STA

[AMM05] G. Appa, D. Magos, and I. Mourtos. On the system of two all different predicates. Information Processing Letters, 94 (3):99–105, May 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arkin:2003:MLW

[AMP03] Esther M. Arkin, Joseph S. B. Mitchell, and Christine D. Piatko. Minimum-link watchman tours. Information Processing Letters, 86(4):203–207, May 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anderson:2001:SPT

[AMR01] James H. Anderson, Mark Moir, and Srikanth Ramamurthy. A simple proof technique for priority-scheduled systems. Information Processing Letters, 77(2–4):63–70, February 28, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/

10/23/20/68/29/26/article.
pdf.

Arkin:2008:CCC

[AMS08] Esther M. Arkin, Joseph S. B.

Mitchell, and Jack Snoeyink. Capturing crossings: Convex hulls of segment and plane intersections. Information Processing Letters, 107(5):194–197, August 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Amtoft:2008:SMP

[Amt08] Torben Amtoft. Slicing for modern program structures: a theory for eliminating irrelevant loops. Information Processing Letters, 106(2):45–51, April 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Antonellis:2009:ACC

[AMT09] Panagiotis Antonellis, Christos Makris, and Nikos Tsirakis. Algorithms for clustering clickstream data. Information Processing Letters, 109(8):381–385, March 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Amir:2009:PMN

[AN09] Amihood Amir and Gonzalo Navarro. Parameterized matching on non-linear structures. Information Processing Letters, 109(15):864–867, July 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Andersson:2009:HNC

[And09]  Daniel Andersson.  Hashi-

wokakero is NP-complete. Information Processing Letters, 109(19):1145–1146, September 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2000:Ia

[Ano00a] Anonymous. Index. Information Processing Letters, 73(5–6): 227–228, March 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/62/28/
                                        37/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/62/28/37/article.
                                        pdf.

Anonymous:2000:Ib

[Ano00b] Anonymous. Index. Information Processing Letters, 73(5–6): 229–230, March 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/62/28/
                                        38/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/62/28/38/article.
                                        pdf.

Anonymous:2000:Ic

[Ano00c] Anonymous. Index. Information Processing Letters, 74(5–6): 265–266, June 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.

nl/gej-ng/10/23/20/63/28/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/28/38/article.
pdf.

Anonymous:2000:Id

[Ano00d] Anonymous. Index. Information Processing Letters, 74(5–6): 267–268, June 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/28/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/28/39/article.
pdf.

Anonymous:2000:Ie

[Ano00e] Anonymous. Index. Information Processing Letters, 75(6):283– 284, November 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/31/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/31/31/article.
pdf.

Anonymous:2000:If

[Ano00f] Anonymous. Index. Information Processing Letters, 75(6):285– 286, November 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/31/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/23/20/64/31/32/article.
                                        pdf.

Anonymous:2000:Ig

[Ano00g] Anonymous. Index. Information Processing Letters, 76(4–6):183– 184, December 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/67/28/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/67/28/29/article.
                                        pdf.

Anonymous:2000:Ih

[Ano00h] Anonymous. Index. Information Processing Letters, 76(4–6):185– 186, December 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/67/28/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/67/28/30/article.
                                        pdf.

Anonymous:2000:Ii

[Ano00i] Anonymous. Index. Information Processing Letters, 76(4–6):187– 290, December 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/67/28/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/67/28/31/article.
                                        pdf.

Anonymous:2001:AIVa

[Ano01a] Anonymous. Author index — volume 77 (2001). Information Processing Letters, 77(5–6): 293–294, March 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/32/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/32/35/article.
pdf.

Anonymous:2001:AIVb

[Ano01b] Anonymous. Author index — volume 79. Information Processing Letters, 79(6):303–304, September 30, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/80/37/
36/abstract.html.

Anonymous:2001:AIVc

[Ano01c] Anonymous. Author index — volume 80 (2001). Information Processing Letters, 80(6):319– 321, December 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/84/37/
35/abstract.html.

Anonymous:2001:MIV

[Ano01d] Anonymous. Master index— volumes 71–80. Information Processing Letters, 80(6):333– 342, December 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/84/37/
                                        37/abstract.html.

Anonymous:2001:MSI

[Ano01e] Anonymous. Master subject index — volumes 71–80. Information Processing Letters, 80 (6):323–331, December 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/gej-ng/10/23/
                                        20/84/37/36/abstract.html.

Anonymous:2001:P

[Ano01f] Anonymous. Preface. Information Processing Letters, 78(1–6):1–3, April 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.elsevier.nl/gej%2Dng/
                                        10/23/20/79/25/25/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej%2Dng/10/23/20/79/25/
                                        25/article.pdf.

Anonymous:2001:RLI

[Ano01g]  Anonymous.  Reference list

of indexed articles. Information Processing Letters, 78 (1–6):337–448, April 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.elsevier.nl/gej%2Dng/
                                        10/23/20/79/25/27/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej%2Dng/10/23/20/79/25/
                                        27/article.pdf.

Anonymous:2001:SIVa

[Ano01h] Anonymous. Subject index — volume 77 (2001). Information Processing Letters, 77(5–6): 291–292, March 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/32/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/32/34/article.
pdf.

Anonymous:2001:SIVc

[Ano01i]  Anonymous.  Subject index

— volume 79. Information Processing Letters, 79(6):301– 302, September 30, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/80/37/
35/abstract.html.

Anonymous:2001:SIVd

[Ano01j]  Anonymous.  Subject index

— volume 79. Information Processing Letters, 79(6):301– 302, September 30, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/80/37/
35/abstract.html.

Anonymous:2001:SIVe

[Ano01k] Anonymous. Subject index — volume 80 (2001). Information Processing Letters, 80(6):317– 318, December 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/84/37/
                                        34/abstract.html.

Anonymous:2001:SIVb

[Ano01l]  Anonymous.  Subject index

to volumes 1–75. Information Processing Letters, 78 (1–6):5–336, April 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.elsevier.nl/gej%2Dng/
                                        10/23/20/79/25/26/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej%2Dng/10/23/20/79/25/
                                        26/article.pdf.

Anonymous:2002:AIVb

[Ano02a] Anonymous. Author index — - volume 82 (2002). Information Processing Letters, 82(6): 333–334, June 30, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2002:AIVc

[Ano02b] Anonymous. Author index — - volume 83 (2002). Information Processing Letters, 83(6): 355–356, September 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2002:AIVd

[Ano02c] Anonymous. Author index — - volume 84 (2002). Information Processing Letters, 84(6): 347–349, December 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2002:AIVa

[Ano02d] Anonymous. Author index — volume 81 (2002). Information Processing Letters, 81(6): 347–349, March 31, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/37/
37/abstract.html.

Anonymous:2002:BEa

[Ano02e] Anonymous. Board of Editors.

Information Processing Letters, 84(1):CO2, October 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2002:BEb

[Ano02f] Anonymous. Board of Editors.

Information Processing Letters, 84(2):CO2, October 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2002:BEc

[Ano02g] Anonymous. Board of Editors.

Information Processing Letters, 84(3):CO2, November 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2002:BEd

[Ano02h] Anonymous. Board of Editors.

Information Processing Letters, 84(4):CO2, November 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2002:BEe

[Ano02i] Anonymous. Board of Editors. Information Processing Letters, 84(5):CO2, December 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2002:BEf

[Ano02j] Anonymous. Board of Editors. Information Processing Letters, 84(6):CO2, December 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2002:NP

[Ano02k] Anonymous. Note from the Publisher. Information Processing Letters, 83(4):181, August 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2002:SIVb

[Ano02l] Anonymous. Subject index — - volume 82 (2002). Information Processing Letters, 82(6): 331–332, June 30, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2002:SIVc

[Ano02m] Anonymous. Subject index — - volume 83 (2002). Information Processing Letters, 83(6): 353–354, September 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2002:SIVd

[Ano02n] Anonymous. Subject index — - volume 84 (2002). Information Processing Letters, 84(6): 345–346, December 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2002:SIVa

[Ano02o] Anonymous. Subject index — volume 81 (2002). Information Processing Letters, 81(6): 345–346, March 31, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/37/
36/abstract.html.

Anonymous:2003:AIVa

[Ano03a] Anonymous. Author index — - volume 85 (2003). Information Processing Letters, 85(6): 345–347, March 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:AIVb

[Ano03b] Anonymous. Author index — - volume 86 (2003). Information Processing Letters, 86(6): 345–347, June 30, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:AIVc

[Ano03c] Anonymous. Author index — - volume 87 (2003). Information Processing Letters, 87(6): 337–339, September 30, 2003.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:AIVd

[Ano03d] Anonymous. Author index — - volume 88 (2003). Information Processing Letters, 88(6): 313–314, December 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEa

[Ano03e] Anonymous. Board of Editors. Information Processing Letters, 85(1):CO2, January 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEb

[Ano03f] Anonymous. Board of Editors. Information Processing Letters, 85(2):CO2, January 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEc

[Ano03g] Anonymous. Board of Editors. Information Processing Letters, 85(3):CO2, February 14, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEd

[Ano03h] Anonymous. Board of Editors. Information Processing Letters, 85(4):CO2, February 28, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEe

[Ano03i] Anonymous. Board of Editors.

Information Processing Letters, 85(5):CO2, March 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEf

[Ano03j] Anonymous. Board of Editors.

Information Processing Letters, 85(6):CO2, March 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEg

[Ano03k] Anonymous. Board of Editors.

Information Processing Letters, 86(1):CO2, April 15, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEh

[Ano03l] Anonymous. Board of Editors.

Information Processing Letters, 86(2):CO2, April 30, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEi

[Ano03m] Anonymous. Board of Editors.

Information Processing Letters, 86(3):CO2, May 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEj

[Ano03n] Anonymous. Board of Editors.

Information Processing Letters, 86(4):CO2, May 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEk

[Ano03o] Anonymous. Board of Editors. Information Processing Letters, 86(5):CO2, June 15, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEl

[Ano03p] Anonymous. Board of Editors. Information Processing Letters, 86(6):CO2, June 30, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEm

[Ano03q] Anonymous. Board of Editors. Information Processing Letters, 87(1):CO2, July 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEn

[Ano03r] Anonymous. Board of Editors. Information Processing Letters, 87(2):CO2, July 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEo

[Ano03s] Anonymous. Board of Editors. Information Processing Letters, 87(3):CO2, August 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEp

[Ano03t] Anonymous. Board of Editors. Information Processing Letters, 87(4):CO2, August 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEq

[Ano03u] Anonymous. Board of Editors.

Information Processing Letters, 87(5):CO2, September 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEr

[Ano03v] Anonymous. Board of Editors.

Information Processing Letters, 87(6):CO2, September 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEs

[Ano03w] Anonymous. Board of Editors.

Information Processing Letters, 88(1–2):CO2, October 17, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEt

[Ano03x] Anonymous. Board of Editors.

Information Processing Letters, 88(3):CO2, November 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEu

[Ano03y] Anonymous. Board of Editors.

Information Processing Letters, 88(4):CO2, November 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEv

[Ano03z] Anonymous. Board of Editors. Information Processing Letters, 88(5):CO2, December 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:BEw

[Ano03-27] Anonymous. Board of Editors. Information Processing Letters, 88(6):CO2, December 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:SIVa

[Ano03-28] Anonymous. Subject index — - volume 85 (2003). Information Processing Letters, 85(6): 343–344, March 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:SIVb

[Ano03-29] Anonymous. Subject index — - volume 86 (2003). Information Processing Letters, 86(6): 343–344, June 30, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:SIVc

[Ano03-30] Anonymous. Subject index — - volume 87 (2003). Information Processing Letters, 87(6): 335–336, September 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2003:SIVd

[Ano03-31] Anonymous. Subject index — - volume 88 (2003). Information Processing Letters, 88(6): 311–312, December 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:AIV

[Ano04a] Anonymous. Author index — volume 90 (2004). Information Processing Letters, 90(6): 315–316, June 30, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEa

[Ano04b] Anonymous. Board of Editors.

Information Processing Letters, 89(1):CO2, January 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEb

[Ano04c] Anonymous. Board of Editors.

Information Processing Letters, 89(2):CO2, January 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEc

[Ano04d] Anonymous. Board of Editors.

Information Processing Letters, 89(3):CO2, February 14, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEd

[Ano04e] Anonymous. Board of Editors. Information Processing Letters, 89(4):CO2, February 28, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEe

[Ano04f] Anonymous. Board of Editors. Information Processing Letters, 89(5):??, March 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEf

[Ano04g] Anonymous. Board of Editors. Information Processing Letters, 89(6):??, March 31, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEg

[Ano04h] Anonymous. Board of Editors. Information Processing Letters, 90(1):??, April 15, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEh

[Ano04i] Anonymous. Board of Editors. Information Processing Letters, 90(2):??, April 30, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEi

[Ano04j] Anonymous. Board of Editors. Information Processing Letters, 90(3):??, May 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEj

[Ano04k] Anonymous. Board of Editors.

Information Processing Letters, 90(4):??, May 31, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEk

[Ano04l] Anonymous. Board of Editors.

Information Processing Letters, 90(5):??, June 15, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEl

[Ano04m] Anonymous. Board of Editors.

Information Processing Letters, 91(1):??, July 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEm

[Ano04n] Anonymous. Board of Editors.

Information Processing Letters, 91(2):??, July 31, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEn

[Ano04o] Anonymous. Board of Editors.

Information Processing Letters, 91(3):??, August 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEo

[Ano04p] Anonymous. Board of Editors.

Information Processing Letters, 91(4):??, August 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEp

[Ano04q] Anonymous. Board of Editors. Information Processing Letters, 91(5):??, September 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEq

[Ano04r] Anonymous. Board of Editors. Information Processing Letters, 91(6):??, September 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEr

[Ano04s] Anonymous. Board of Editors. Information Processing Letters, 92(1):??, October 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEs

[Ano04t] Anonymous. Board of Editors. Information Processing Letters, 92(2):??, October 31, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEt

[Ano04u] Anonymous. Board of Editors. Information Processing Letters, 92(3):??, November 15, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEu

[Ano04v] Anonymous. Board of Editors. Information Processing Letters, 92(4):??, November 30, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEv

[Ano04w] Anonymous. Board of Editors.

Information Processing Letters, 92(5):??, December 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:BEw

[Ano04x] Anonymous. Board of Editors.

Information Processing Letters, 92(6):??, December 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:HG

[Ano04y] Anonymous. Harald Ganzinger.

Information Processing Letters, 91(4):155, August 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:MIV

[Ano04z] Anonymous. Master index — volumes 81–90. Information Processing Letters, 90(6):319– 330, June 30, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:MSI

[Ano04-27] Anonymous. Master subject index — volumes 81–90. Information Processing Letters, 90(6): 317–318, June 30, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:SIVa

[Ano04-28] Anonymous. Subject index — volume 89 (2004). Information Processing Letters, 89(6): 315–316, March 31, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:SIVb

[Ano04-29] Anonymous. Subject index — volume 91 (2004). Information Processing Letters, 91(6): 301–302, September 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2004:SIVc

[Ano04-30] Anonymous. Subject index — volume 92 (2004). Information Processing Letters, 92(6):317– 318, December 31, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:BEa

[Ano05a] Anonymous. Board of Editors. Information Processing Letters, 93(1):??, January 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:BEb

[Ano05b] Anonymous. Board of Editors. Information Processing Letters, 93(2):??, January 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:BEc

[Ano05c] Anonymous. Board of Editors. Information Processing Letters, 93(3):??, February 14, 2005.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:BEd

[Ano05d] Anonymous. Board of Editors.

Information Processing Letters, 93(4):??, February 28, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:BEe

[Ano05e] Anonymous. Board of Editors.

Information Processing Letters, 93(5):??, March 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:BEf

[Ano05f] Anonymous. Board of Editors.

Information Processing Letters, 93(6):??, March 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:BEg

[Ano05g] Anonymous. Board of Editors.

Information Processing Letters, 94(1):??, April 15, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:BEh

[Ano05h] Anonymous. Board of Editors.

Information Processing Letters, 94(2):??, April 30, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:BEi

[Ano05i] Anonymous. Board of Editors.

Information Processing Letters, 94(3):??, May 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:BEj

[Ano05j] Anonymous. Board of Editors. Information Processing Letters, 94(4):??, May 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:BEk

[Ano05k] Anonymous. Board of Editors. Information Processing Letters, 94(5):??, June 15, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:BEl

[Ano05l] Anonymous. Board of Editors. Information Processing Letters, 94(6):??, June 30, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:BEm

[Ano05m] Anonymous. Board of Editors. Information Processing Letters, 95(1):??, July 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:BEn

[Ano05n] Anonymous. Board of Editors. Information Processing Letters, 95(2):??, July 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:BEo

[Ano05o] Anonymous. Board of Editors. Information Processing Letters, 95(3):??, August 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:BEp

[Ano05p] Anonymous. Board of Editors.

Information Processing Letters, 95(4):??, August 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:BEq

[Ano05q] Anonymous. Board of Editors.

Information Processing Letters, 95(5):??, September 15, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:EBa

[Ano05r] Anonymous. Editorial Board.

Information Processing Letters, 95(6):??, September 30, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:EBb

[Ano05s] Anonymous. Editorial Board.

Information Processing Letters, 96(1):??, October 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:EBc

[Ano05t] Anonymous. Editorial Board.

Information Processing Letters, 96(2):??, October 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:EBd

[Ano05u] Anonymous. Editorial Board. Information Processing Letters, 96(3):??, November 15, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:EBe

[Ano05v] Anonymous. Editorial Board. Information Processing Letters, 96(4):??, November 30, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:EBf

[Ano05w] Anonymous. Editorial Board. Information Processing Letters, 96(5):??, December 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:EBg

[Ano05x] Anonymous. Editorial Board. Information Processing Letters, 96(6):??, December 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:SIVa

[Ano05y] Anonymous. Subject index — volume 93 (2005). Information Processing Letters, 93(6): 315–316, March 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:SIVb

[Ano05z] Anonymous. Subject index — volume 94 (2005). Information Processing Letters, 94(6): 283–284, June 30, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:SIVc

[Ano05-27] Anonymous. Subject index — volume 95 (2005). Information Processing Letters, 95(6): 559–560, September 30, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2005:SIVd

[Ano05-28] Anonymous. Subject index — volume 96 (2005). Information Processing Letters, 96(6):230– 231, December 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:AIV

[Ano06a] Anonymous. Author index — volume 100 (2006). Information Processing Letters, 100(6):246– 247, December 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBa

[Ano06b] Anonymous. Editorial Board.

Information Processing Letters, 97(1):??, January 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBb

[Ano06c] Anonymous. Editorial Board. Information Processing Letters, 97(2):??, January 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBc

[Ano06d] Anonymous. Editorial Board. Information Processing Letters, 97(3):??, February 14, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBd

[Ano06e] Anonymous. Editorial Board. Information Processing Letters, 97(4):??, February 28, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBe

[Ano06f] Anonymous. Editorial Board. Information Processing Letters, 97(5):??, March 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBf

[Ano06g] Anonymous. Editorial Board. Information Processing Letters, 97(6):??, March 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBg

[Ano06h] Anonymous. Editorial Board. Information Processing Letters, 98(1):??, April 15, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBh

[Ano06i] Anonymous. Editorial Board.

Information Processing Letters, 98(2):??, April 30, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBi

[Ano06j] Anonymous. Editorial Board.

Information Processing Letters, 98(3):??, May 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBj

[Ano06k] Anonymous. Editorial Board.

Information Processing Letters, 98(4):??, May 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBk

[Ano06l] Anonymous. Editorial Board.

Information Processing Letters, 98(5):??, June 15, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBl

[Ano06m] Anonymous. Editorial Board.

Information Processing Letters, 98(6):??, June 30, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBm

[Ano06n] Anonymous. Editorial Board.

Information Processing Letters, 99(1):??, July 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBn

[Ano06o] Anonymous. Editorial Board. Information Processing Letters, 99(2):??, July 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBo

[Ano06p] Anonymous. Editorial Board. Information Processing Letters, 99(3):??, August 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBp

[Ano06q] Anonymous. Editorial Board. Information Processing Letters, 99(4):??, August 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBq

[Ano06r] Anonymous. Editorial Board. Information Processing Letters, 99(5):??, September 15, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBr

[Ano06s] Anonymous. Editorial Board. Information Processing Letters, 99(6):??, September 30, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBs

[Ano06t] Anonymous. Editorial Board.

Information Processing Letters, 100(1):??, October 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBt

[Ano06u] Anonymous. Editorial Board.

Information Processing Letters, 100(2):??, October 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBu

[Ano06v] Anonymous. Editorial Board.

Information Processing Letters, 100(3):??, November 15, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBv

[Ano06w] Anonymous. Editorial Board.

Information Processing Letters, 100(4):??, November 30, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:EBw

[Ano06x] Anonymous. Editorial Board.

Information Processing Letters, 100(5):??, December 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:SIVa

[Ano06y] Anonymous. Subject index — volume 97 (2006). Information Processing Letters, 97(6): 251–252, March 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:SIVb

[Ano06z] Anonymous. Subject index — volume 98 (2006). Information Processing Letters, 98(6): 263–264, June 30, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2006:SIVc

[Ano06-27] Anonymous. Subject index — volume 99 (2006). Information Processing Letters, 99(6): 252–253, September 30, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBa

[Ano07a] Anonymous. Editorial Board. Information Processing Letters, 101(1):??, January 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBb

[Ano07b] Anonymous. Editorial Board. Information Processing Letters, 101(2):??, January 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBc

[Ano07c] Anonymous. Editorial Board.

Information Processing Letters, 101(3):??, February 14, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBd

[Ano07d] Anonymous. Editorial Board.

Information Processing Letters, 101(4):??, February 28, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBe

[Ano07e] Anonymous. Editorial Board.

Information Processing Letters, 101(5):??, March 16, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBf

[Ano07f] Anonymous. Editorial Board.

Information Processing Letters, 101(6):??, March 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBg

[Ano07g] Anonymous. Editorial Board.

Information Processing Letters, 102(1):??, April 15, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBh

[Ano07h] Anonymous. Editorial Board.

Information Processing Letters, 102(2–3):??, April 30, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBi

[Ano07i] Anonymous. Editorial Board. Information Processing Letters, 102(4):??, May 16, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBj

[Ano07j] Anonymous. Editorial Board. Information Processing Letters, 102(5):??, May 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBk

[Ano07k] Anonymous. Editorial Board. Information Processing Letters, 102(6):??, June 15, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBl

[Ano07l] Anonymous. Editorial Board. Information Processing Letters, 103(1):??, June 30, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBm

[Ano07m] Anonymous. Editorial Board. Information Processing Letters, 103(2):??, July 16, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBn

[Ano07n] Anonymous. Editorial Board. Information Processing Letters, 103(3):??, July 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBo

[Ano07o] Anonymous. Editorial Board.

Information Processing Letters, 103(4):??, August 16, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBp

[Ano07p] Anonymous. Editorial Board.

Information Processing Letters, 103(5):??, August 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBq

[Ano07q] Anonymous. Editorial Board.

Information Processing Letters, 103(6):??, September 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBr

[Ano07r] Anonymous. Editorial Board.

Information Processing Letters, 104(1):??, September 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBs

[Ano07s] Anonymous. Editorial Board.

Information Processing Letters, 104(2):??, October 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBt

[Ano07t] Anonymous. Editorial Board. Information Processing Letters, 104(3):??, October 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBu

[Ano07u] Anonymous. Editorial Board. Information Processing Letters, 104(4):??, November 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBv

[Ano07v] Anonymous. Editorial Board. Information Processing Letters, 104(5):??, November 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBw

[Ano07w] Anonymous. Editorial Board. Information Processing Letters, 104(6):??, December 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2007:EBx

[Ano07x] Anonymous. Editorial Board. Information Processing Letters, 105(1):??, December 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBa

[Ano08a] Anonymous. Editorial Board.

Information Processing Letters, 105(2):??, January 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBb

[Ano08b] Anonymous. Editorial Board.

Information Processing Letters, 105(3):??, January 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBc

[Ano08c] Anonymous. Editorial Board.

Information Processing Letters, 105(4):??, February 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBd

[Ano08d] Anonymous. Editorial Board.

Information Processing Letters, 105(5):??, February 29, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBe

[Ano08e] Anonymous. Editorial Board.

Information Processing Letters, 105(6):??, March 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBf

[Ano08f] Anonymous. Editorial Board.

Information Processing Letters, 106(1):??, March 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBg

[Ano08g] Anonymous. Editorial Board. Information Processing Letters, 106(2):??, April 15, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBh

[Ano08h] Anonymous. Editorial Board. Information Processing Letters, 106(3):??, April 30, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBi

[Ano08i] Anonymous. Editorial Board. Information Processing Letters, 106(4):??, May 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBj

[Ano08j] Anonymous. Editorial Board. Information Processing Letters, 106(5):??, May 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBk

[Ano08k] Anonymous. Editorial Board. Information Processing Letters, 106(6):??, June 15, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBl

[Ano08l] Anonymous. Editorial Board. Information Processing Letters, 107(1):??, June 30, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBm

[Ano08m] Anonymous. Editorial Board.

Information Processing Letters, 107(2):??, July 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBn

[Ano08n] Anonymous. Editorial Board.

Information Processing Letters, 107(3–4):??, July 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBo

[Ano08o] Anonymous. Editorial Board.

Information Processing Letters, 107(5):??, August 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBp

[Ano08p] Anonymous. Editorial Board.

Information Processing Letters, 107(6):??, August 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBq

[Ano08q] Anonymous. Editorial Board.

Information Processing Letters, 108(1):??, September 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBr

[Ano08r] Anonymous. Editorial Board.

Information Processing Letters, 108(2):??, September 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBs

[Ano08s] Anonymous. Editorial Board. Information Processing Letters, 108(3):??, October 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBt

[Ano08t] Anonymous. Editorial Board. Information Processing Letters, 108(4):??, October 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBu

[Ano08u] Anonymous. Editorial Board. Information Processing Letters, 108(5):??, November 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBv

[Ano08v] Anonymous. Editorial Board. Information Processing Letters, 108(6):??, November 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBw

[Ano08w] Anonymous. Editorial Board. Information Processing Letters, 109(1):??, December 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2008:EBx

[Ano08x] Anonymous. Editorial Board.

Information Processing Letters, 109(2):??, December 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBa

[Ano09a] Anonymous. Editorial Board.

Information Processing Letters, 109(3):??, January 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBb

[Ano09b] Anonymous. Editorial Board.

Information Processing Letters, 109(4):??, January 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBc

[Ano09c] Anonymous. Editorial Board.

Information Processing Letters, 109(5):??, February 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBd

[Ano09d] Anonymous. Editorial Board.

Information Processing Letters, 109(6):??, February 28, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBe

[Ano09e] Anonymous. Editorial Board. Information Processing Letters, 109(7):??, March 16, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBf

[Ano09f] Anonymous. Editorial Board. Information Processing Letters, 109(8):??, March 31, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBg

[Ano09g] Anonymous. Editorial Board. Information Processing Letters, 109(9):??, April 15, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBh

[Ano09h] Anonymous. Editorial Board. Information Processing Letters, 109(10):??, April 30, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBi

[Ano09i] Anonymous. Editorial Board. Information Processing Letters, 109(11):??, May 16, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBj

[Ano09j] Anonymous. Editorial Board. Information Processing Letters, 109(12):??, May 31, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBk

[Ano09k] Anonymous. Editorial Board.

Information Processing Letters, 109(13):??, June 15, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBl

[Ano09l] Anonymous. Editorial Board.

Information Processing Letters, 109(14):??, June 30, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBm

[Ano09m] Anonymous. Editorial Board.

Information Processing Letters, 109(15):??, July 16, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBn

[Ano09n] Anonymous. Editorial Board.

Information Processing Letters, 109(16):??, July 31, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBo

[Ano09o] Anonymous. Editorial Board.

Information Processing Letters, 109(17):??, August 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBp

[Ano09p] Anonymous. Editorial Board.

Information Processing Letters, 109(18):??, August 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBq

[Ano09q] Anonymous. Editorial Board. Information Processing Letters, 109(19):??, September 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBr

[Ano09r] Anonymous. Editorial Board. Information Processing Letters, 109(20):??, September 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBs

[Ano09s] Anonymous. Editorial Board. Information Processing Letters, 109(21–22):??, October 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBt

[Ano09t] Anonymous. Editorial Board. Information Processing Letters, 109(23–24):??, November 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBu

[Ano09u] Anonymous. Editorial Board. Information Processing Letters, 110(1):??, December 1, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anonymous:2009:EBv

[Ano09v] Anonymous. Editorial Board.

Information Processing Letters, 110(2):??, December 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anantharaman:2005:CPD

[ANR05] Siva Anantharaman, Paliath Narendran, and Michael Rusinowitch. Closure properties and decision problems of dag automata. Information Processing Letters, 94(5):231–240, June 15, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Adamus:2009:NCP

[AOS^{+}09] Lech Adamus, Beata Orchel, Artur Szymański, A. Paweł Wojda, and Małgorzata Zwonek. A note on t-complementing permutations for graphs. Information Processing Letters, 110(2): 44–45, December 16, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Avidor:2006:NUG

[AR06] Adi Avidor and Ricky Rosen.

A note on Unique Games. Information Processing Letters, 99 (3):87–91, August 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Araki:2003:EPR

[Ara03] Toru Araki. Edge-pancyclicity of recursive circulants. Information Processing Letters, 88 (6):287–292, December 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Araki:2007:TDB

[Ara07] Toru Araki. On the k-tuple domination of de Bruijn and Kautz digraphs. Information Processing Letters, 104(3):86– 90, October 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ayala-Rincon:2007:VFJ

[ARdAdS07] Mauricio Ayala-Rincón, Bruno T. de Abreu, and José de Siqueira. A variant of the Ford–Johnson algorithm that is more space efficient. Information Processing Letters, 102(5):201–207, May 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ayala-Rincon:2002:FVE

[ARdFPS02] Mauricio Ayala-Rincón, Alexsandro F. da Fonseca, Haydée Werneck Poubel, and José Siqueira. A framework to visualize equivalences between computational models of regular languages. Information Processing Letters, 84 (1):5–16, October 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aronov:2002:LBV

[Aro02] Boris Aronov. A lower bound on Voronoi diagram complexity. Information Processing Letters, 83(4):183–185, August 31, 2002.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arslan:2008:ALE

[Ars08] Abdullah N. Arslan. An algorithm with linear expected running time for string editing with substitutions and substring reversals. Information Processing Letters, 106(5):213– 218, May 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arya:2002:BSP

[Ary02] Sunil Arya. Binary space partitions for axis-parallel line segments: Size-height tradeoffs. Information Processing Letters, 84 (4):201–206, November 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Araki:2002:EPR

[AS02a] Tom Araki and Yukio Shibata.

Erratum to: “Pancyclicity of recursive circulant graphs”: [Inform. Process. Lett. 81 (2002) 187–190]. Information Processing Letters, 84(3):173, November 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [AS02b].

Araki:2002:PRC

[AS02b] Toru Araki and Yukio Shibata. Pancyclicity of recursive circulant graphs. Information Processing Letters, 81(4):187– 190, February 28, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/85/35/
                                        29/abstract.html. See erra-

tum [AS02a].

Akin:2007:CAG

[AS07a] Hasan Akın and Irfan Siap. On cellular automata over Galois rings. Information Processing Letters, 103(1):24–27, June 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alborzi:2007:ETA

[AS07b] Houman Alborzi and Hanan Samet. Execution time analysis of a top-down R-tree construction algorithm. Information Processing Letters, 101(1): 6–12, January 16, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Asano:2007:ARV

[Asa07] Tetsuo Asano. Aspect-ratio Voronoi diagram and its complexity bounds. Information Processing Letters, 105(1):26– 31, December 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Asano:2008:OUI

[Asa08] Tetsuo Asano. Online uniformity of integer points on a line. Information Processing Letters, 109(1):57–60, December 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Abrahamson:2005:ETB

[ASW05] Jeff Abrahamson, Ali Shokoufandeh, and Pawel Winter. Euclidean TSP between two nested convex obstacles. Information Processing Letters, 95(2):370– 375, July 31, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Attie:2002:WFB

[Att02] Paul Attie. Wait-free Byzantine consensus. Information Processing Letters, 83(4):221–227, August 31, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Audsley:2001:PAF

[Aud01] N. C. Audsley. On priority assignment in fixed priority scheduling. Information Processing Letters, 79 (1):39–44, May 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.elsevier.nl/gej%2Dng/
10/23/20/80/25/31/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/23/20/80/25/
31/article.pdf.

Avoine:2007:HSC

[AV07] Gildas Avoine and Serge Vaudenay. How to safely close a discussion. Information Processing Letters, 102(4):138–142, May 16, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aggarwal:2004:SAR

[AW04] Alok Aggarwal and Youngcheul Wee. On the symmetric angle-restricted nearest neighbor problem. Information Processing Letters, 92(3):121–126, November 15, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ann:2008:FSA

[AYTH08] Hsing-Yen Ann, Chang-Biau Yang, Chiou-Ting Tseng, and Chiou-Yi Hor. A fast and simple algorithm for computing the longest common subsequence of run-length encoded strings. Information Processing Letters, 108(6):360–364, November 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Almeida:2008:DAB

[AZ08] Jorge Almeida and Marc Zeitoun. Description and analysis of a bottom-up DFA minimization algorithm. Information Processing Letters, 107(2):52–59, July 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ben-Amram:2003:TCF

[BA03a] Amir M. Ben-Amram. Tighter constant-factor time hierarchies. Information Processing Letters, 87(1):39–44, July 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bruda:2003:LCP

[BA03b] Stefan D. Bruda and Selim G. Akl. On limits on the computational power of dataaccumulating algorithms. Information Processing Letters, 86 (4):221–227, May 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bletsas:2006:OPA

[BA06] Konstantinos Bletsas and Neil Audsley. Optimal priority assignment in the presence of blocking. Information Processing Letters, 99(3):83–86, August 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bach:2007:BED

[Bac07] Eric Bach. Bounds for the expected duration of the monopolist game. Information Processing Letters, 101(2):86–92, January 31, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bars:2001:LFM

[Bar01a] Jean-Marie Le Bars. The 0-1 law fails for monadic existential second-order logic on undirected graphs. Information Processing Letters, 77(1): 43–48, January 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/24/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/23/20/68/24/31/article.
                                        pdf.

Baruah:2001:SPT

[Bar01b] Sanjoy Baruah. Scheduling periodic tasks on uniform multiprocessors. Information Processing Letters, 80(2):97–104, October 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/80/37/
                                        31/abstract.html.

Baril:2009:MRG

[Bar09] Jean-Luc Baril. More restrictive Gray codes for some classes of pattern avoiding permutations. Information Processing Letters, 109(14):799–804, June 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baswana:2008:SAG

[Bas08] Surender Baswana. Streaming algorithm for graph spanners— single pass and constant processing time per edge. Information Processing Letters, 106(3):110– 114, April 30, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bhattacharya:2001:COB

[BB01] Binay Bhattacharya and Robert Benkoczi. On computing the optimal bridge between two convex polygons. Information Processing Letters, 79(5):215– 221, September 15, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/80/34/
27/abstract.html.

Bergstra:2005:UBE

[BB05] J. A. Bergstra and I. Bethke.

An upper bound for the equational specification of finite state services. Information Processing Letters, 94(6):267–269, June 30, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bubeck:2009:NCT

[BB09] Uwe Bubeck and Hans Kleine Büning. A new 3-CNF transformation by parallel-serial graphs. Information Processing Letters, 109(7):376–379, March 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barbuti:2002:AIO

[BBD02] R. Barbuti, C. Bernardeschi, and N. De Francesco. Abstract interpretation of operational semantics for secure information flow. Information Processing Letters, 83(2):101–108, July 31, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barbosa:2001:PDG

[BBF01] Valmir C. Barbosa, Mario R. F. Benevides, and Ayru L. Oliveira Filho. A priority dynamics for generalized drinking philosophers. Information Processing Letters, 79(4):189– 195, August 15, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/33/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/80/33/30/article.
pdf.

Bouyer:2006:IUR

[BBM06] Patricia Bouyer, Thomas Brihaye, and Nicolas Markey. Improved undecidability results on weighted timed automata. Information Processing Letters, 98 (5):188–194, June 15, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baldoni:2008:AGE

[BBMR08] Roberto Baldoni, François Bonnet, Alessia Milani, and Michel Raynal. Anonymous graph exploration without collision by mobile robots. Information Processing Letters, 109(2):98–103, December 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bagchi:2006:AMW

[BBS06a] Amitabha Bagchi, Ankur Bhargava, and Torsten Suel. Approximate maximum weight branchings. Information Processing Letters, 99(2):54–58, July 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baier:2006:NAP

[BBS06b] Christel Baier, Nathalie Bertrand, and Philippe Schnoebelen. A note on the attractor-property of infinite-state Markov chains. Information Processing Letters, 97(2):58–63, January 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bender:2000:PGT

[BC00] Michael A. Bender and Chandra Chekuri. Performance guarantees for the TSP with a parameterized triangle inequality. Information Processing Letters, 73(1–2):17–21, January 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/62/25/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/62/25/28/article.
pdf.

Biro:2007:IKE

[BC07] Péter Biró and Katarína Cechlárová.

Inapproximability of the kidney exchange problem. Information Processing Letters, 101(5): 199–202, March 16, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Balbuena:2005:SCO

[BCD^{+}05] C. Balbuena, M. Cera, A. Diánez, P. García-Vázquez, and X. Marcote. Sufficient conditions for λ′- optimality of graphs with small conditional diameter. Information Processing Letters, 95(4): 429–434, August 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Banerjee:2009:DAM

[BCG09] Satyajit Banerjee, Atish Datta Chowdhury, and Subhas Kumar Ghosh. Distributed approximation for maximum weight matching on bounded degree bounded integer weight graphs. Information Processing Letters, 109(14): 790–794, June 30, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Biedl:2003:DLB

[BCLO03] Therese Biedl, Timothy M. Chan, and Alejandro López-Ortiz. Drawing K_{2,n}: A lower bound. Information Processing Letters, 85(6):303–305, March 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blundo:2002:NOM

[BCM02] Carlo Blundo, Stelvio Cimato, and Barbara Masucci. A note on optimal metering schemes. Information Processing Letters, 84 (6):319–326, December 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bartal:2002:MRW

[BCNR02] Yair Bartal, Marek Chrobak, John Noga, and Prabhakar Raghavan. More on random walks, electrical networks, and the harmonic k-server algorithm. Information Processing Letters, 84(5):271–276, December 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bonifati:2005:ETI

[BCP05] Angela Bonifati, Stefano Ceri, and Stefano Paraboschi. Event Trace Independence of active behavior. Information Processing Letters, 94(2):71–77, April 30, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Berard:2008:MEA

[BCP08] Sèverine Bérard, Cedric Chauve, and Christophe Paul. A more efficient algorithm for perfect sorting by reversals. Information Processing Letters, 106(3): 90–95, April 30, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertrand:2009:NDS

[BCP09] Gilles Bertrand, Michel Couprie, and Nicolas Passat. A note on 3-D simple points and simple-equivalence. Information Processing Letters, 109(13):700– 704, June 15, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Berard:2000:TAA

[BD00] Béatrice Bérard and Catherine Dufourd. Timed automata and additive clock constraints. Information Processing Letters, 75 (1–2):1–7, July 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/25/25/article.
pdf.

Burkard:2004:ETP

[BD04] Rainer E. Burkard and Vladimir G. Deineko. On the Euclidean TSP with a permuted Van der Veen matrix. Information Processing Letters, 91(6):259–262, September 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Burmester:2005:SSG

[BD05] Mike Burmester and Yvo Desmedt. A secure and scalable Group Key Exchange system. Information Processing Letters, 94(3):137–143, May 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bhattacharyya:2008:USD

[BD08] Bishnu Bhattacharyya and Frank Dehne. Using spine decompositions to efficiently solve the length-constrained heaviest path problem for trees. Information Processing Letters, 108 (5):293–297, November 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bitner:2009:FSE

[BD09] Steven Bitner and Ovidiu Daescu. Farthest segments and extremal triangles spanned by points in R^{3}. Information Processing Letters, 109(20):1167– 1171, September 30, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bezerra:2004:SAC

[BdATdC04] Byron L. D. Bezerra and Francisco de A. T. de Carvalho. A symbolic approach for contentbased information filtering. Information Processing Letters, 92 (1):45–52, October 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Biernacki:2005:DED

[BDcS05] Dariusz Biernacki, Olivier Danvy, and Chung chieh Shan. On the dynamic extent of delimited continuations. Information Processing Letters, 96(1):7–17, October 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bornstein:2006:PCA

[BdFdS06] Claudson Bornstein, Celina M. H. de Figueiredo, and Vinícius G. P. de Sá. The Pair Completion algorithm for the Homogeneous Set Sandwich Problem. Information Processing Letters, 98(3):87–91, May 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brosenne:2007:AO

[BDHW07] Henrik Brosenne, Carsten Damm, Matthias Homeister, and Stephan Waack. On approximation by ⊕- OBDDs. Information Processing Letters, 102(1):17–21, April 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Berman:2007:COC

[BDKW07] Piotr Berman, Bhaskar Das-Gupta, Ming-Yang Kao, and Jie Wang. On constructing an optimal consensus clustering from multiple clusterings. Information Processing Letters, 104(4): 137–145, November 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Birmele:2009:MAO

[BDL09] Etienne Birmelé, François Delbot, and Christian Laforest. Mean analysis of an online algorithm for the vertex cover problem. Information Processing Letters, 109(9):436–439, April 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boulinier:2008:SET

[BDLP08] Christian Boulinier, Ajoy K. Datta, Lawrence L. Larmore, and Franck Petit. Space efficient and time optimal distributed BFS tree construction. Information Processing Letters, 108 (5):273–278, November 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blundo:2000:VCG

[BDN00] Carlo Blundo, Alfredo De Santis, and Moni Naor. Visual cryptography for grey level images. Information Processing Letters, 75(6):255–259, November 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/31/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/31/27/article.
pdf.

Bu:2008:FLN

[BDQ08] Tian-Ming Bu, Xiaotie Deng, and Qi Qi. Forward looking Nash equilibrium for keyword auction. Information Processing Letters, 105(2):41–46, January 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Buhrman:2003:QZE

[BdW03] Harry Buhrman and Ronald de Wolf. Quantum zero-error algorithms cannot be composed. Information Processing Letters, 87(2):79–84, July 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barequet:2005:OBC

[BE05] Gill Barequet and Gershon Elber. Optimal bounding cones of vectors in three dimensions. Information Processing Letters, 93(2):83–89, January 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bouajjani:2000:EAA

[BEF^{+}00] Ahmed Bouajjani, Javier Esparza, Alain Finkel, Oded Maler, Peter Rossmanith, Bernard Willems, and Pierre Wolper. An efficient automata approach to some problems on contextfree grammars. Information Processing Letters, 74(5–6):221– 227, June 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/28/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/28/31/article.
pdf.

Butman:2004:SPS

[BEL04] Ayelet Butman, Revital Eres, and Gad M. Landau. Scaled and permuted string matching. Information Processing Letters, 92 (6):293–297, December 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Biskup:2008:RIC

[BEL08] Joachim Biskup, David W. Embley, and Jan-Hendrik Lochner. Reducing inference control to access control for normalized database schemas. Information Processing Letters, 106(1):8–12, March 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bourgeois:2009:AMC

[BEP09] Nicolas Bourgeois, Bruno Escoffier, and Vangelis Th. Paschos. Approximation of min coloring by moderately exponential algorithms. Information Processing Letters, 109(16):950–954, July 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertault:2000:FDA

[Ber00] François Bertault. A forcedirected algorithm that preserves edge-crossing properties. Information Processing Letters, 74(1–2):7–13, April 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/25/26/article.
pdf.

Bereg:2004:TPT

[Ber04] Sergey Bereg. Transforming pseudo-triangulations. Information Processing Letters, 90(3): 141–145, May 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Backhouse:2001:AET

[BF01] Roland Backhouse and Maarten Fokkinga. The associativity of equivalence and the Towers of Hanoi problem. Information Processing Letters, 77(2–4):71– 76, February 28, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/29/27/article.
pdf.

Bernholt:2003:CUR

[BF03] Thorsten Bernholt and Roland Fried. Computing the update of the repeated median regression line in linear time. Information Processing Letters, 88 (3):111–117, November 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bona:2009:ANB

[BF09] Miklós Bóna and Ryan Flynn. The average number of block interchanges needed to sort a permutation and a recent result of Stanley. Information Processing Letters, 109(16):927–931, July 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bansal:2009:CDB

[BFB09] Mukul S. Bansal and David Fernández-Baca. Computing distances between partial rankings. Information Processing Letters, 109(4):238–241, January 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Berczi:2009:LTA

[BFK09] Kristóf Bérczi, Satoru Fujishige, and Naoyuki Kamiyama. A linear-time algorithm to find a pair of arc-disjoint spanning in-arborescence and outarborescence in a directed acyclic graph. Information Processing Letters, 109(23–24): 1227–1231, November 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bein:2000:LBR

[BFL00] Wolfgang W. Bein, Rudolf Fleischer, and Lawrence L. Larmore. Limited bookmark randomized online algorithms for the paging problem. Information Processing Letters, 76(4–6):155– 162, December 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/28/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/28/25/article.
pdf.

Bonizzoni:2001:SSSa

[BFMZ01a] P. Bonizzoni, C. Ferretti, G. Mauri, and R. Zizza. Separating some splicing models. Information Processing Letters, 79(6):255–259, September 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/80/37/27/abstract.html.

Bonizzoni:2001:SSSb

[BFMZ01b] P. Bonizzoni, C. Ferretti, G. Mauri, and R. Zizza. Separating some splicing models. Information Processing Letters, 79(6):255–259, September 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/23/
                                        20/80/37/27/abstract.html.

Bekesi:2001:WCA

[BG01] József Békési and Gábor Galambos. Worst-case analysis of the Iterated Longest Fragment algorithm. Information Processing Letters, 79(3):147– 153, July 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/80/32/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/80/32/32/article.
                                        pdf.

Breitbart:2008:CLU

[BG08] Yuri Breitbart and Hassan Gobjuka. Characterization of layer-2 unique topologies. Information Processing Letters, 105(2): 52–57, January 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bian:2009:AAW

[BG09] Zhengbing Bian and Qian-Ping Gu. 1.5-approximation algorithm for weighted maximum routing and wavelength assignment on rings. Information Processing Letters, 109(8):400–404, March 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bank:2009:ICP

[BGHP09] Bernd Bank, Marc Giusti, Joos Heintz, and Luis Miguel Pardo. On the intrinsic complexity of point finding in real singular hypersurfaces. Information Processing Letters, 109(19):1141– 1144, September 15, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bose:2008:FPR

[BGK^{+}08] Prosenjit Bose, Hua Guo, Evangelos Kranakis, Anil Maheshwari, Pat Morin, Jason Morrison, Michiel Smid, and Yihui Tang. On the false-positive rate of Bloom filters. Information Processing Letters, 108(4):210– 213, October 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boichut:2008:TLS

[BH08a] Y. Boichut and P.-C. Héam. A theoretical limit for safety verification techniques with regular fix-point computations. Information Processing Letters, 108 (1):1–2, September 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bordihn:2008:NCD

[BH08b] Henning Bordihn and Markus Holzer. A note on cooperating distributed grammar systems working in combined modes. Information Processing Letters, 108(1):10–14, September 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bockenhauer:2000:AAT

[BHK^{+}00] Hans-Joachim Böckenhauer, Juraj Hromkovič, Ralf Klasing, Sebastian Seibert, and Walter Unger. Approximation algorithms for the TSP with sharpened triangle inequality. Information Processing Letters, 75(3):133–138, August 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/28/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/28/31/article.
pdf.

Baier:2004:PWS

[BHK04] Christel Baier, Holger Hermanns, and Joost-Pieter Katoen. Probabilistic weak simulation is decidable in polynomial time. Information Processing Letters, 89(3):123–130, February 14, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blaser:2009:DTS

[BHLV09] Markus Bläser, Moritz Hardt, Richard J. Lipton, and Nisheeth K. Vishnoi. Deterministically testing sparse polynomial identities of unbounded degree. Information Processing Letters, 109(3): 187–192, January 16, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brlek:2006:FEC

[BHM06] S. Brlek, S. Hamadou, and J. Mullins. A flaw in the electronic commerce protocol SET. Information Processing Letters, 97(3):104–108, February 14, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blaser:2008:ACC

[BHM08] Markus Bläser, Thomas Heynen, and Bodo Manthey. Adding cardinality constraints to integer programs with applications to maximum satisfiability. Information Processing Letters, 105 (5):194–198, February 29, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baldoni:2001:ISC

[BHMR01] Roberto Baldoni, Jean-Michel Hélary, Achour Mostéfaoui, and Michel Raynal. Impossibility of scalar clock-based communication-induced checkpointing protocols ensuring the RDT property. Information Processing Letters, 80(2):105– 111, October 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/80/37/
32/abstract.html.

Born:2003:HDC

[BHW03] Axel Born, Cor A. J. Hurkens, and Gerhard J. Woeginger. How to detect a counterfeit coin: Adaptive versus non-adaptive solutions. Information Processing Letters, 86(3):137–141, May 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brosenne:2006:NOB

[BHW06] Henrik Brosenne, Matthias Homeister, and Stephan Waack. Nondeterministic ordered binary decision diagrams with repeated tests and various modes of acceptance. Information Processing Letters, 98(1):6–10, April 15, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bauer:2009:NRT

[BHY09] Joanna Bauer, Dag Haugland, and Di Yuan. New results on the time complexity and approximation ratio of the Broadcast Incremental Power algorithm. Information Processing Letters, 109(12):615–619, May 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Biham:2002:HDE

[Bih02] Eli Biham. How to decrypt or even substitute DES-encrypted messages in 2^{28} steps. Information Processing Letters, 84 (3):117–124, November 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bijlsma:2001:MBS

[Bij01] Lex Bijlsma. Model-based specification. Information Processing Letters, 77(2–4):77–84, February 28, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/29/28/article.
pdf.

Bernath:2008:WBO

[BJ08] Attila Bernáth and Gwenaël Joret. Well-balanced orientations of mixed graphs. Information Processing Letters, 106(4): 149–151, May 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bachman:2002:MTW

[BJK02] Aleksander Bachman, Adam Janiak, and Mikhail Y. Kovalyov. Minimizing the total weighted completion time of deteriorating jobs. Information Processing Letters, 81(2): 81–84, January 31, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/33/
31/abstract.html.

Beimel:2000:LUH

[BK00] Amos Beimel and Eyal Kushilevitz. Learning unions of highdimensional boxes over the reals. Information Processing Letters, 73(5–6):213–220, March 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/62/28/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/62/28/35/article.
pdf.

Bartha:2004:TTT

[BK04] Miklós Bartha and Miklós Krész. Tutte type theorems for graphs having a perfect internal matching. Information Processing Letters, 91(6):277–284, September 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bohte:2005:ASN

[BK05] Sander M. Bohte and Joost N. Kok. Applications of spiking neural networks. Information Processing Letters, 95(6):519– 520, September 30, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Burghardt:2001:AGE

[BKS01] J. Burghardt, F. Kammüller, and J. W. Sanders. On the antisymmetry of Galois embeddings. Information Processing Letters, 79(2):57–63, June 30, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/31/
26/abstract.html;   http:

//www.elsevier.nl/gej-ng/
10/23/20/80/31/26/article.
pdf.

Bein:2000:TOA

[BL00] Wolfgang W. Bein and Lawrence L.

Larmore. Trackless online algorithms for the server problem. Information Processing Letters, 74(1–2):73–79, April 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/25/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/25/35/article.
pdf.

Brandstadt:2006:SLT

[BL06] Andreas Brandstädt and Van Bang Le. Structure and linear time recognition of 3-leaf powers. Information Processing Letters, 98 (4):133–138, May 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blaser:2003:CSP

[Blä03] Markus Bläser. Computing small partial coverings. Information Processing Letters, 85(6): 327–331, March 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blanchet:2005:SPL

[Bla05] Bruno Blanchet. Security protocols: from linear to classical logic by abstract interpretation. Information Processing Letters, 95(5):473–479, September 15, 2005. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Brandstadt:2004:ERA

[BLdR04] Andreas Brandstädt, Van Bang Le, and H. N. de Ridder. Efficient robust algorithms for the Maximum Weight Stable Set Problem in chair-free graph classes. Information Processing Letters, 89(4):165–173, February 28, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brandstadt:2002:RVF

[BLL02] Andreas Brandstädt, Hoàng-Oanh Le, and Van Bang Le. On α-redundant vertices in P_{5}-free graphs. Information Processing Letters, 82(3):119–122, May 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/86/34/27/abstract.html.

Branke:2002:WRL

[BLME02] Jürgen Branke, Stefan Leppert, Martin Middendorf, and Peter Eades. Width-restricted layering of acyclic digraphs with consideration of dummy nodes. Information Processing Letters, 81(2):59–63, January 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/85/3327/abstract.html.

Bernardeschi:2006:UPR

[BLMM06] C. Bernardeschi, G. Lettieri, L. Martini, and P. Masci. Using postdomination to reduce space requirements of data flow analysis. Information Processing Letters, 98(1):11–18, April 15, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bein:2007:UMT

[BLN07] Wolfgang Bein, Lawrence L.

Larmore, and John Noga. Uniform metrical task systems with a limited number of states. Information Processing Letters, 104(4):123–128, November 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brandstadt:2003:SSN

[BLV03] Andreas Brandstädt, Hoàng-Oanh Le, and Jean-Marie Vanherpe. Structure and stability number of chair-, co-P- and gem-free graphs revisited. Information Processing Letters, 86 (3):161–167, May 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boxer:2001:PAA

[BM01] Laurence Boxer and Russ Miller.

A parallel algorithm for approximate regularity. Information Processing Letters, 80(6):311– 316, December 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/84/37/
33/abstract.html.

Baldoni:2002:MIE

[BM02a] Roberto Baldoni and Giovanna Melideo. On the minimal information to encode timestamps in distributed computations. Information Processing Letters, 84 (3):159–166, November 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brandstadt:2002:MWS

[BM02b] Andreas Brandstädt and Suhail Mahfud. Maximum Weight Stable Set on graphs without claw and co-claw (and similar graph classes) can be solved in linear time. Information Processing Letters, 84(5):251–259, December 16, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ben-Moshe:2004:CAL

[BMCK04] Boaz Ben-Moshe, Paz Carmi, and Matthew J. Katz. Computing all large sums-of-pairs in R^{n} and the discrete planar two-watchtower problem. Information Processing Letters, 89 (3):137–139, February 14, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baev:2002:LBP

[BME02] Ivan D. Baev, Waleed M. Meleis, and Alexandre Eichenberger. Lower bounds on precedenceconstrained scheduling for parallel processors. Information Processing Letters, 83(1):27–32, July 16, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bigham:2008:DPD

[BMO08] Bahram Sadeghi Bigham, Ali Mohades, and Lidia Ortega. Dynamic polar diagram. Information Processing Letters, 109 (2):142–146, December 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boykin:2000:NUF

[BMP^{+}00] P. Oscar Boykin, Tal Mor, Matthew Pulver, Vwani Roychowdhury, and Farrokh Vatan. A new universal and faulttolerant quantum basis. Information Processing Letters, 75 (3):101–107, August 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/23/
20/64/28/26/abstract.html;
 http://www.elsevier.nl/gej-
ng/10/23/20/64/28/26/article.
pdf.

Beyersdorff:2009:CPI

[BMTV09] Olaf Beyersdorff, Arne Meier, Michael Thomas, and Heribert Vollmer. The complexity of propositional implication. Information Processing Letters, 109(18):1071–1077, August 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blake:2005:NWN

[BMX05] Ian F. Blake, V. Kumar Murty, and Guangwu Xu. A note on window τ-NAF algorithm. Information Processing Letters, 95(5):496–502, September 15, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertsch:2001:SLT

[BN01] E. Bertsch and M.-J. Nederhof. Size/lookahead tradeoff for LL(k)-grammars. Information Processing Letters, 80(3):125– 129, November 15, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/84/35/
                                        28/abstract.html.

Bertsch:2004:FPR

[BN04] E. Bertsch and M.-J. Nederhof. Fast parallel recognition of LR language suffixes. Information Processing Letters, 92 (5):225–229, December 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertsch:2007:SOL

[BN07] E. Bertsch and M.-J. Nederhof. Some observations on LRlike parsing with delayed reduction. Information Processing Letters, 104(6):195–199, December 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bodirsky:2009:IPV

[BNvO09] Manuel Bodirsky, Gustav Nordh, and Timo von Oertzen. Integer programming with 2-variable equations and 1-variable inequalities. Information Processing Letters, 109(11):572–575, May 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bojanczyk:2007:NAT

[Boj07] Mikołaj Bojańczyk. A new algorithm for testing if a regular language is locally threshold testable. Information Processing Letters, 104(3):91–94, October 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bollig:2003:VSF

[Bol03] Beate Bollig. A very simple function that requires exponential size nondeterministic graphdriven read-once branching programs. Information Processing Letters, 86(3):143–148, May 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bollig:2005:LLB

[Bol05] Beate Bollig. A large lower bound on the query complexity of a simple boolean function. Information Processing Letters, 95(4):423–428, August 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bollig:2008:NSO

[Bol08a] Beate Bollig. A note on the size of OBDDs for the graph of integer multiplication. Information Processing Letters, 109(1): 41–43, December 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bollig:2008:ORO

[Bol08b] Beate Bollig. The optimal readonce branching program complexity for the direct storage access function. Information Processing Letters, 106(4):171–174, May 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bollig:2009:SGO

[Bol09a] Beate Bollig. On the size of (generalized) OBDDs for threshold functions. Information Processing Letters, 109(10):499– 503, April 30, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bolognesi:2009:PRN

[Bol09b] Tommaso Bolognesi. A pseudorandom network mobile automaton with linear growth. Information Processing Letters, 109(13): 668–674, June 15, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Borzyszkowski:2000:GIC

[Bor00] Tomasz Borzyszkowski. Generalized interpolation in C ASL. Information Processing Letters, 76(1–2):19–24, November 20, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/24/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/24/27/article.
pdf.

Bouyer:2002:LCD

[Bou02] Patricia Bouyer. A logical characterization of data languages. Information Processing Letters, 84(2):75–85, October 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bergstra:2001:PAC

[BP01] Jan A. Bergstra and Alban Ponse. Process algebra and conditional composition. Information Processing Letters, 80(1): 41–49, October 15, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/84/27/
32/abstract.html.

Baril:2006:ELU

[BP06] Jean-Luc Baril and Jean-Marcel Pallo. Efficient lower and upper bounds of the diagonalflip distance between triangulations. Information Processing Letters, 100(4):131–136, November 30, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brewster:2003:CDP

[BR03] Richard C. Brewster and Romeo Rizzi. On the complexity of digraph packings. Information Processing Letters, 86(2):101– 106, April 30, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Braben:2001:BT

[Bra01]  D. W. Braben.  Bucking

the trends. Information Processing Letters, 77(2–4):85–87, February 28, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/68/29/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/68/29/29/article.
                                        pdf.

Brass:2007:MHC

[Bra07] Peter Brass. Multidimensional heaps and complementary range searching. Information Processing Letters, 102(4):152–155, May 16, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brass:2009:UHF

[Bra09] Peter Brass. Universal hash functions for an infinite universe and hash trees. Information Processing Letters, 109(10):461– 462, April 30, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brejova:2001:AVS

[Bre01] Bronislava Brejová. Analyzing variants of Shellsort. Information Processing Letters, 79(5): 223–227, September 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/80/34/28/abstract.html.

Breit:2004:IAA

[Bre04] J. Breit. An improved approximation algorithm for twomachine flow shop scheduling with an availability constraint. Information Processing Letters, 90(6):273–278, June 30, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brodsky:2005:IGB

[Bro05] Alex Brodsky. An impossibility gap between width-4 and width-5 permutation branching programs. Information Processing Letters, 94(4):159–164, May 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bohler:2005:BBC

[BRSV05] Elmar Böhler, Steffen Reith, Henning Schnoor, and Heribert Vollmer. Bases for Boolean co-clones. Information Processing Letters, 96(2):59–66, October 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baveja:2000:ALC

[BS00a] Alok Baveja and Aravind Srinivasan. Approximating lowcongestion routing and columnrestricted packing problems. Information Processing Letters, 74 (1–2):19–25, April 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/25/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/25/28/article.
pdf.

Bespamyatnikh:2000:CSP

[BS00b] Sergei Bespamyatnikh and Michael Segal. Covering a set of points by two axisparallel boxes. Information Processing Letters, 75(3):95– 100, August 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/28/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/28/25/article.
pdf.

Bespamyatnikh:2000:ELI

[BS00c] Sergei Bespamyatnikh and Michael Segal. Enumerating longest increasing subsequences and patience sorting. Information Processing Letters, 76(1–2): 7–11, November 20, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/24/

25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/24/25/article.
pdf.

Buki:2002:CHD

[BS02] Judit Büki and Csaba Szabó.

Complexity of homomorphisms to direct products of graphs. Information Processing Letters, 81 (4):175–178, February 28, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/85/35/27/abstract.html.

Berend:2006:DHG

[BS06] Daniel Berend and Amir Sapir.

The diameter of Hanoi graphs. Information Processing Letters, 98(2):79–85, April 30, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bird:2007:MLL

[BS07] Richard S. Bird and Stefan Sadnicki. Minimal on-line labelling. Information Processing Letters, 101(1):41–45, January 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bleischwitz:2008:NER

[BS08a] Yvonne Bleischwitz and Florian Schoppmann. New efficiency results for makespan cost sharing. Information Processing Letters, 107(2):64–70, July 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bordawekar:2008:APT

[BS08b] Rajesh Bordawekar and Oded Shmueli. An algorithm for partitioning trees augmented with sibling edges. Information Processing Letters, 108(3):136–142, October 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bostan:2009:SFA

[BS09] Alin Bostan and Éric Schost. A simple and fast algorithm for computing exponentials of power series. Information Processing Letters, 109(13):754– 756, June 15, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Booij:2005:GDR

[BtN05] Olaf Booij and Hieu tat Nguyen. A gradient descent rule for spiking neurons emitting multiple spikes. Information Processing Letters, 95(6):552–558, September 30, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bunde:2004:SOC

[Bun04] David P. Bunde. SPT is optimally competitive for uniprocessor flow. Information Processing Letters, 90(5):233–238, June 15, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Burkhard:2005:DHP

[Bur05] Walter A. Burkhard. Double hashing with passbits. Information Processing Letters, 96 (5):162–166, December 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bretto:2002:PAS

[BUŽ02] Alain Bretto, Stéphane Ubéda, and Janez Žerovnik. A polynomial algorithm for the strong Helly property. Information Processing Letters, 81(1):55– 57, January 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/27/
35/abstract.html.

Boldi:2000:CPS

[BV00a] Paolo Boldi and Sebastiano Vigna. Coverings that preserve sense of direction. Information Processing Letters, 75(4):175– 180, September 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/29/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/29/30/article.
pdf.

Brodal:2000:IBD

[BV00b] Gerth Stølting Brodal and S. Venkatesh. Improved bounds for dictionary look-up with one error. Information Processing Letters, 75(1–2):57– 59, July 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/64/25/
                                        34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/64/25/34/article.
                                        pdf.

Bottcher:2008:TCS

[BV08] Julia Böttcher and Dan Vilenchik. On the tractability of coloring semirandom graphs. Information Processing Letters, 108(3):143–149, October 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bretto:2009:CCS

[BV09] Alain Bretto and Thierry Vallée. A clique-covering sufficient condition for Hamiltonicity of graphs. Information Processing Letters, 109(20):1156–1160, September 30, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bansal:2002:NOC

[BVG02] S. S. Bansal, B. Vishal, and P. Gupta. Near optimal Cholesky factorization on orthogonal multiprocessors. Information Processing Letters, 84 (1):23–30, October 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Biaz:2001:CFB

[BW01] Saâd Biaz and Jennifer L. Welch. Closed form bounds for clock synchronization under simple uncertainty assumptions.

Information Processing Letters, 80(3):151–157, November 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/84/35/32/abstract.html.

Bollig:2003:FRO

[BW03] Beate Bollig and Ingo Wegener.

Functions that have read-once branching programs of quadratic size are not necessarily testable. Information Processing Letters, 87(1):25–29, July 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Beimel:2006:MCM

[BW06] Amos Beimel and Enav Weinreb. Monotone circuits for monotone weighted threshold functions. Information Processing Letters, 97(1):12–18, January 16, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baeza-Yates:2002:OBD

[BYSP02] Ricardo A. Baeza-Yates and Héctor Soza-Pollman. Optimal bounded disorder. Information Processing Letters, 83(3): 151–157, August 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Buning:2002:PTA

[BZ02] Hans Kleine Büning and Xishun Zhao. Polynomial time algorithms for computing a representation for minimal unsatisfiable formulas with fixed deficiency. Information Processing Letters, 84(3):147–151, November 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Banic:2006:FDC

[BŽ06] Iztok Banič and Janez Žerovnik. Fault-diameter of Cartesian graph bundles. Information Processing Letters, 100(2):47–51, October 31, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ben-Zwi:2007:LBT

[BZLN07] Oren Ben-Zwi, Oded Lachish, and Ilan Newman. Lower bounds for testing Euclidean Minimum Spanning Trees. Information Processing Letters, 102(6):219–225, June 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chicano:2008:ACO

[CA08] Francisco Chicano and Enrique Alba. Ant colony optimization with partial order reduction for discovering safety property violations in concurrent models. Information Processing Letters, 106(6):221–231, June 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cai:2009:LKR

[Cai09] Xuan Cai. Linear kernelizations for restricted 3-Hitting Set problems. Information Processing Letters, 109(13):730–738, June 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Caporaso:2006:DCC

[Cap06] Salvatore Caporaso. A decidable characterization of the classes between lintime and exptime. Information Processing Letters, 97(1):36–40, January 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carpentieri:2002:SCM

[Car02]  Bruno Carpentieri.  Send-

ing compressed messages to a learned receiver on a bidirectional line. Information Processing Letters, 83(2):63–70, July 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cholvi:2001:MPC

[CB01] Vicent Cholvi and Pablo Boronat.

A minimal property for characterizing deadlock-free programs. Information Processing Letters, 77(5–6):283–290, March 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/32/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/32/33/article.
pdf.

Cholvi:2004:RBM

[CB04] Vicent Cholvi and Josep Bernabéu.

Relationships between memory models. Information Processing Letters, 90(2):53–58, April 30, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chlebus:2007:NPM

[CB07] Edward Chlebus and Jordy Brazier. Nonstationary Poisson modeling of Web browsing session arrivals. Information Processing Letters, 102(5):187–190, May 31, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:2001:GGQ

[CC01a] Ye-In Chang and Bor-Hsu Chen. A generalized grid quorum strategy for k-mutual exclusion in distributed systems. Information Processing Letters, 80(4): 205–212, November 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/84/34/31/abstract.html.

Chen:2001:FTA

[CC01b] Yu-Wei Chen and Kuo-Liang Chung. Fault-tolerant algorithm for Fast Fourier Transform on hypercubes. Information Processing Letters, 79(1):11–16, May 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.elsevier.nl/gej%2Dng/
10/23/20/80/25/27/abstract.
html;   http://www.elsevier.

nl/gej%2Dng/10/23/20/80/25/
27/article.pdf.

Cheng:2002:SAQ

[CC02] Sheng-Tzong Cheng and Ing-Ray Chen. A self-adjusting quality of service control scheme. Information Processing Letters, 83(6):337–344, September 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:2003:EPA

[CC03] Chin-Chen Chang and Chi-Yien Chung. An efficient protocol for anonymous multicast and reception. Information Processing Letters, 85(2):99–103, January 31, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2005:OAL

[CC05a] Kuan-Yu Chen and Kun-Mao Chao. Optimal algorithms for locating the longest and shortest segments satisfying a sum or an average constraint. Information Processing Letters, 96 (6):197–201, December 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheung:2005:CIP

[CC05b] K. S. Cheung and K. O. Chow.

Cycle inclusion property of augmented marked graphs. Information Processing Letters, 94 (6):271–276, June 30, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2006:NTI

[CC06a] Yangjun Chen and Yibin Chen. A new tree inclusion algorithm. Information Processing Letters, 98(6):253–262, June 30, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chou:2006:RRC

[CC06b] Shih-Chien Chou and Yuan-Chien Chen. Retrieving reusable components with variation points from software product lines. Information Processing Letters, 99(3):106–110, August 16, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Clifford:2007:SDW

[CC07] Peter Clifford and Raphaël Clifford. Simple deterministic wildcard matching. Information Processing Letters, 101(2):53– 54, January 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2009:SJB

[CC09] Hue-Ling Chen and Ye-In Chang. Spatial joins based on NA-trees. Information Processing Letters, 109(13):713–718, June 15, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Campelo:2004:CHV

[CCF04] Manoel Campêlo, Ricardo Corrêa, and Yuri Frota. Cliques, holes and the vertex coloring polytope. Information Processing Letters, 89(4):159–164, February 28, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Collette:2008:IJP

[CCG08] Sébastien Collette, Liliana Cucu, and Joël Goossens. Integrating job parallelism in realtime scheduling theory. Information Processing Letters, 106 (5):180–187, May 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carrabs:2005:LTA

[CCGP05] F. Carrabs, R. Cerulli, M. Gentili, and G. Parlato. A linear time algorithm for the minimum Weighted Feedback Vertex Set on diamonds. Information Processing Letters, 94(1):29–35, April 15, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chung:2002:DCK

[CCKL02] Ilyong Chung, Wankyu Choi, Youngchel Kim, and Mike Lee. The design of conference key distribution system employing a symmetric balanced incomplete block design. Information Processing Letters, 81(6): 313–318, March 31, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/37/
31/abstract.html. See erra-

tum [Chu04].

Cazorla:2001:PAP

[CCVP01] D. Cazorla, F. Cuartero, V. Valero, and F. L. Pelayo. A process algebra for probabilistic and nondeterministic processes. Information Processing Letters, 80(1):15–23, October 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/gej-ng/10/23/
                                        20/84/27/29/abstract.html.

Chan:2007:GOF

[CCY^{+}07] Joseph Wun-Tat Chan, Francis Y. L. Chin, Deshi Ye, Yong Zhang, and Hong Zhu. Greedy online frequency allocation in cellular networks. Information Processing Letters, 102(2–3):55– 61, April 30, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cai:2008:DPS

[CCZT08] Xingjuan Cai, Zhihua Cui, Jianchao Zeng, and Ying Tan. Dispersed particle swarm optimization. Information Processing Letters, 105(6):231–235, March 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carpi:2009:WAN

[CD09] Arturo Carpi and Valerio D’Alonzo. On a word avoiding near repeats. Information Processing Letters, 109(13):679– 682, June 15, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chatterjee:2004:PFO

[CDC04] Krishnendu Chatterjee, Pallab Dasgupta, and P. P. Chakrabarti. The power of first-order quantification over states in branching and linear time temporal logics. Information Processing Letters, 91(5): 201–210, September 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chin:2004:SAC

[CDF^{+}04] Francis Y. L. Chin, Alfredo De Santis, Anna Lisa Ferrara, N. L. Ho, and S. K. Kim. A simple algorithm for the constrained sequence problems. Information Processing Letters, 90(4): 175–179, May 31, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Costagliola:2001:DCP

[CDFG01] G. Costagliola, V. Deufemia, F. Ferrucci, and C. Gravino. Decidability of the consistency problem for regular symbolic picture description languages. Information Processing Letters, 80(3):119–124, November 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/84/35/27/abstract.html.

Cimato:2005:ICV

[CDFM05] Stelvio Cimato, Alfredo De Santis, Anna Lisa Ferrara, and Barbara Masucci. Ideal contrast visual cryptography schemes with reversing. Information Processing Letters, 93(4):199–206, February 28, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chew:2002:WAF

[CDKK02] L. Paul Chew, Haggai David, Matthew J. Katz, and Klara Kedem. Walking around fat obstacles. Information Processing Letters, 83(3):135–140, August 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carpi:2002:CPF

[CdL02] Arturo Carpi and Aldo de Luca. A combinatorial property of the factor poset of a word. Information Processing Letters, 81(1): 35–39, January 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/85/27/
                                        32/abstract.html.

Codenotti:2004:AAH

[CDL^{+}04] Bruno Codenotti, Gianluca De Marcoo, Mauro Leoncini, Manuela Montangero, and Massimo Santini. Approximation algorithms for a hierarchically structured bin packing problem. Information Processing Letters, 89(5):215–221, March 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Costantini:2002:ERA

[CDP02] Stefania Costantini, Ottavio D’Antona, and Alessandro Provetti. On the equivalence and range of applicability of graph-based representations of logic programs. Information Processing Letters, 84(5):241– 249, December 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chandran:2009:CBG

[CDS09] L. Sunil Chandran, Anita Das, and Naveen Sivadasan. On the cubicity of bipartite graphs. Information Processing Letters, 109(9):432–435, April 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cesati:2002:PCC

[Ces02] Marco Cesati. Perfect Code is W [1]-complete. Information Processing Letters, 81(3):163– 168, February 14, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/34/
34/abstract.html.

Cheng:2002:KRG

[CF02] Qi Cheng and Fang Fang. Kolmogorov random graphs only have trivial stable colorings. Information Processing Letters, 81 (3):133–136, February 14, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.

                                        elsevier.com/gej-ng/10/23/
                                        20/85/34/29/abstract.html.

Cichacz:2009:FP

[CF09] Sylwia Cichacz and Dalibor Froncek. Factorization of K_{n,n} into (0, j)-prisms. Information Processing Letters, 109(16):932– 934, July 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2009:ICC

[CFM09] Wenguang Chen, Wenfei Fan, and Shuai Ma. Incorporating cardinality constraints and synonym rules into conditional functional dependencies. Information Processing Letters, 109 (14):783–789, June 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cabello:2009:HOV

[CFS09] S. Cabello, M. Fort, and J. A. Sellarès. Higher-order Voronoi diagrams on triangulated surfaces. Information Processing Letters, 109(9):440–445, April 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2008:RPC

[CFvG08] Taolue Chen, Wan Fokkink, and Rob van Glabbeek. Ready to preorder: The case of weak process semantics. Information Processing Letters, 109(2):104– 111, December 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2001:CHN

[CFW01] Lusheng Chen, Fang-Wei Fu, and Victor K.-W. Wei. On the constructions of highly nonlinear zigzag functions and unbiased functions. Information Processing Letters, 79(3): 135–140, July 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/32/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/80/32/30/article.
pdf.

Castro:2000:PSP

[CG00] Jorge Castro and David Guijarro. PACS, simple-PAC and query learning. Information Processing Letters, 73(1–2):11– 16, January 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/62/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/62/25/27/article.
pdf.

Chockler:2004:LBT

[CG04a] Hana Chockler and Dan Gutfreund. A lower bound for testing juntas. Information Processing Letters, 90(6):301–305, June 30, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crescenzi:2004:MLC

[CG04b] Pilu Crescenzi and Federico Greco. The minimum likely column cover problem. Information Processing Letters, 89 (4):175–179, February 28, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chandran:2005:RMV

[CG05] L. Sunil Chandran and Fabrizio Grandoni. Refined memorization for vertex cover. Information Processing Letters, 93 (3):125–131, February 14, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Choffrut:2006:SRR

[CG06] Christian Choffrut and Serge Grigorieff. Separability of rational relations in A^{∗} × N^{m} by recognizable relations is decidable. Information Processing Letters, 99(1):27–32, July 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Czerwinski:2008:IRF

[CG08] Sebastian Czerwiński and Jarosław Grytczuk. Invisible runners in finite fields. Information Processing Letters, 108(2):64– 67, September 30, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2009:AVD

[CG09] Meirun Chen and Xiaofeng Guo. Adjacent vertex-distinguishing edge and total chromatic numbers of hypercubes. Information Processing Letters, 109(12):599– 602, May 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cabello:2008:PCD

[CGK08] Sergio Cabello, Panos Giannopoulos, and Christian Knauer. On the parameterized complexity of d-dimensional point set pattern matching. Information Processing Letters, 105(2):73–77, January 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cegielski:2006:MSE

[CGM06] Patrick Cégielski, Irène Guessarian, and Yuri Matiyasevich. Multiple serial episodes matching. Information Processing Letters, 98(6):211–218, June 30, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Czerwinski:2009:LLG

[CGŻ09] Sebastian Czerwiński, Jarosław Grytczuk, and Wiktor Żelazny. Lucky labelings of graphs. Information Processing Letters, 109(18):1078–1081, August 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cho:2002:REF

[CH02] Hsun-Jung Cho and Li-Yen Hsu.

Ring embedding in faulty honeycomb rectangular torus. Information Processing Letters, 84 (5):277–284, December 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cho:2003:GHT

[CH03a] Hsun-Jung Cho and Li-Yen Hsu. Generalized honeycomb torus. Information Processing Letters, 86(4):185–190, May 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Choi:2003:TSP

[CH03b] Kwanghoon Choi and Taisook Han. A type system for the push-enter model. Information Processing Letters, 87(4):205– 211, August 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cameron:2006:SCI

[CH06] Kathie Cameron and Chính T. Hoàng. On the structure of certain intersection graphs. Information Processing Letters, 99 (2):59–63, July 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chatterjee:2008:RSP

[CH08] Krishnendu Chatterjee and Thomas A. Henzinger. Reduction of stochastic parity to stochastic mean-payoff games. Information Processing Letters, 106(1):1–7, March 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:2002:MGC

[Cha02a] Byeong-Mo Chang. Managing the granularity of constraintbased analyses by rule transformation. Information Processing Letters, 83(2):79–88, July 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:2002:CPP

[Cha02b] Gerard J. Chang. Corrigendum to “The path-partition problem in block graphs”: [Information Processing Letters 52 (1994) 317–322]. Information Processing Letters, 83(5):293, September 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [YC94].

Chauve:2002:TPM

[Cha02c] Cedric Chauve. Tree pattern matching with a more general notion of occurrence of the pattern. Information Processing Letters, 82(4):197–201, May 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chandran:2003:LBH

[Cha03] L. Sunil Chandran. A lower bound for the hitting set size for combinatorial rectangles and an application. Information Processing Letters, 86(2):75–78, April 30, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chan:2004:NMI

[Cha04a] Timothy M. Chan. A note on maximum independent sets in rectangle intersection graphs. Information Processing Letters, 89(1):19–23, January 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chandran:2004:MCG

[Cha04b] L. Sunil Chandran. Minimum cuts, girth and a spectral threshold. Information Processing Letters, 89(3):105–110, February 14, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chataigner:2005:AMA

[Cha05] Frédéric Chataigner. Approximating the Maximum Agreement Forest on k trees. Information Processing Letters, 93(5): 239–244, March 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chan:2008:WSP

[Cha08]  Timothy M. Chan.  Well-

separated pair decomposition in linear time? Information Processing Letters, 107(5):138–141, August 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chuang:2002:OEF

[CHC02] Yen-Chu Chuang, Lih-Hsing Hsu, and Chung-Haw Chang. Optimal 1-edge fault-tolerant designs for ladders. Information Processing Letters, 84(2):87–92, October 31, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2000:TMW

[Che00a] Tzung-Shi Chen. Task migration in 2D wormhole-routed mesh multicomputers. Information Processing Letters, 73(3– 4):103–110, February 29, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/23/
20/62/27/29/abstract.html;
 http://www.elsevier.nl/gej-
ng/10/23/20/62/27/29/article.
pdf.

Chen:2000:MSS

[Che00b] Weimin Chen. Multi-subsequence searching. Information Processing Letters, 74(5–6):229– 233, June 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/28/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/28/32/article.
pdf.

Chen:2002:SFS

[Che02] Yangjun Chen. Signature files and signature trees. Information Processing Letters, 82(4): 213–221, May 31, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2003:EMM

[Che03] L.-B. Chen. An efficient minimum and maximum global snapshot algorithm. Information Processing Letters, 86(3):155– 159, May 16, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2004:ERF

[Che04a] Yu-Wei Chen. An enhanced recursive frequency splitting broadcasting algorithm for near video-on-demand services. Information Processing Letters, 92 (6):299–302, December 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheung:2004:NCL

[Che04b] K. S. Cheung. New characterization for live and reversible augmented marked graphs. Information Processing Letters, 92 (5):239–243, December 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2006:SAP

[Che06a] Jing-Chao Chen. A simple algorithm for in-place merging. Information Processing Letters, 98 (1):34–40, April 15, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2006:CSS

[Che06b] Yangjun Chen. On the cost of searching signature trees. Information Processing Letters, 99 (1):19–26, July 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2007:CPT

[Che07] Xie-Bin Chen. Cycles passing through prescribed edges in a hypercube with some faulty edges. Information Processing Letters, 104(6):211–215, December 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2009:HPC

[Che09a]  Xie-Bin Chen.  Hamilto-

nian paths and cycles passing through a prescribed path in hypercubes. Information Processing Letters, 110(2):77–82, December 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2009:PBH

[Che09b] Xie-Bin Chen. On path bipancyclicity of hypercubes. Information Processing Letters, 109 (12):594–598, May 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2009:SRT

[Che09c] Xie-Bin Chen. Some results on topological properties of folded hypercubes. Information Processing Letters, 109(8):395–399, March 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Choi:2009:KDK

[CHH^{+}09] Seung Geol Choi, Javier Herranz, Dennis Hofheinz, Jung Yeon Hwang, Eike Kiltz, Dong Hoon Lee, and Moti Yung. The Kurosawa–Desmedt key encapsulation is not chosen-ciphertext secure. Information Processing Letters, 109(16):897–901, July 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chung:2009:IBE

[CHK^{+}09] Byungchun Chung, Junbeom Hur, Heeyoul Kim, Seong-Min Hong, and Hyunsoo Yoon. Improved batch exponentiation. Information Processing Letters, 109(15):832–837, July 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Charbit:2008:NCS

[CHL^{+}08] Pierre Charbit, Michel Habib, Vincent Limouzy, Fabien de Montgolfier, Mathieu Raffinot, and Michaël Rao. A note on computing set overlap classes. Information Processing Letters, 108(4): 186–191, October 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Czumaj:2005:AAO

[CHLN05] Artur Czumaj, Magnús M. Halldórsson, Andrzej Lingas, and Johan Nilsson. Approximation algorithms for optimization problems in graphs with superlogarithmic treewidth. Information Processing Letters, 94 (2):49–53, April 30, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Costa:2007:NSM

[CHLN07] Vaston Costa, Edward Haeusler, Eduardo S. Laber, and Loana Nogueira. A note on the size of minimal covers. Information Processing Letters, 102(2– 3):124–126, April 30, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cho:2002:DRU

[Cho02] Haengrae Cho. Database recovery using incomplete page versions in a multisystem data sharing environment. Information Processing Letters, 83(1):49–55, July 16, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cholvi:2008:SBN

[Cho08] V. Cholvi. Stability bounds in networks with dynamic link capacities. Information Processing Letters, 109(2):151–154, December 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chung:2000:FMW

[Chu00] Kuo-Liang Chung. On finding medians of weighted discrete points. Information Processing Letters, 74(3–4):103– 106, May 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/63/27/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/63/27/27/article.
pdf.

Chung:2004:EDC

[Chu04] Ilyong Chung. Erratum to: “The design of conference key distribution system employing a symmetric balanced incomplete block design”: [Inform. Process. Lett. 81 (2002) 313–318]. Information Processing Letters, 91(6):299–300, September 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [CCKL02].

Chen:2006:OOP

[CHU06] J. Chen, R. M. Hierons, and H. Ural. Overcoming observability problems in distributed test architectures. Information Processing Letters, 98(5):177–182, June 15, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chu:2008:SLT

[Chu08] Frank Pok Man Chu. A simple linear time certifying LBFSbased algorithm for recognizing trivially perfect graphs and their complements. Information Processing Letters, 107(1):7–12, June 30, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cai:2009:ECP

[CHZL09] Jiansheng Cai, Jianfeng Hou, Xia Zhang, and Guizhen Liu. Edge-choosability of planar graphs without non-induced 5cycles. Information Processing Letters, 109(7):343–346, March 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crochemore:2008:CLP

[CI08] Maxime Crochemore and Lucian Ilie. Computing Longest Previous Factor in linear time and applications. Information Processing Letters, 106(2):75–80, April 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cole:2003:SFM

[CIL^{+}03] Richard Cole, Costas Iliopoulos, Thierry Lecroq, Wojciech Plandowski, and Wojciech Rytter. On special families of morphisms related to δ-matching and don’t care symbols. Information Processing Letters, 85 (5):227–233, March 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crochemore:2001:FPB

[CIPR01] Maxime Crochemore, Costas S.

Iliopoulos, Yoan J. Pinzon, and James F. Reid. A fast and practical bit-vector algorithm for the Longest Common Subsequence problem. Information Processing Letters, 80(6):279– 285, December 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/84/37/
28/abstract.html.

Crochemore:2008:OPS

[CIR08] Maxime Crochemore, Costas S. Iliopoulos, and M. Sohel Rahman. Optimal prefix and suffix queries on texts. Information Processing Letters, 108(5):320– 325, November 15, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chung:2002:EII

[CK02] Tae-Sun Chung and Hyoung-Joo Kim. Extracting indexing information from XML DTDs. Information Processing Letters, 81 (2):97–103, January 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/85/33/34/abstract.html.

Chandra:2003:DAD

[CK03] Punit Chandra and Ajay D. Kshemkalyani. Distributed algorithm to detect strong conjunctive predicates. Information Processing Letters, 87(5): 243–249, September 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chatain:2007:WFA

[CK07] Thomas Chatain and Victor Khomenko. On the wellfoundedness of adequate orders used for construction of complete unfolding prefixes. Information Processing Letters, 104 (4):129–136, November 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cohen:2008:GMC

[CK08] Reuven Cohen and Liran Katzir.

The Generalized Maximum Coverage Problem. Information Processing Letters, 108(1):15– 22, September 15, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cho:2002:SSD

[CKCB02] Seung Mo Cho, Hyung Ho Kim, Sung Deok Cha, and Doo Hwan Bae. A semantics of sequence diagrams. Information Processing Letters, 84(3):125–130, November 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Caragiannis:2002:NBS

[CKK02a] Ioannis Caragiannis, Christos Kaklamanis, and Panagiotis Kanellopoulos. New bounds on the size of the minimum feedback vertex set in meshes and butterflies. Information Processing Letters, 83(5):275– 280, September 15, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chowdhury:2002:EDT

[CKK02b] Rezaul Alam Chowdhury, M. Kaykobad, and Irwin King. An efficient decoding technique for Huffman codes. Information Processing Letters, 81(6):305– 308, March 31, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/37/
29/abstract.html.

Caragiannis:2003:LAA

[CKK03a] Ioannis Caragiannis, Christos Kaklamanis, and Panagiotis Kanellopoulos. A logarithmic approximation algorithm for the minimum energy consumption broadcast subgraph problem. Information Processing Letters, 86(3):149–154, May 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chung:2003:EPX

[CKK03b] Yon Dohn Chung, Jong Wook Kim, and Myoung Ho Kim. Efficient preprocessing of XML queries using structured signatures. Information Processing Letters, 87(5):257–264, September 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:2004:ETA

[CKKC04] Ku-Young Chang, Howon Kim, Ju-Sung Kang, and Hyun-Sook Cho. An extension of TYT algorithm for GF((2^{n})^{m}) using precomputation. Information Processing Letters, 92(5):231–234, December 16, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cosmadakis:2001:ODD

[CKL01] Stavros Cosmadakis, Gabriel Kuper, and Leonid Libkin. On the orthographic dimension of definable sets. Information Processing Letters, 79(3):141– 145, July 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/32/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/80/32/31/article.
pdf.

Chun:2003:DLC

[CKL^{+}03] Kilsoo Chun, Seungjoo Kim, Sangjin Lee, Soo Hak Sung, and Seonhee Yoon. Differential and linear cryptanalysis for 2-round SPNs. Information Processing Letters, 87(5):277–282, September 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Calvo:2001:CPS

[CKM^{+}01] Jorge Alberto Calvo, Danny Krizanc, Pat Morin, Michael Soss, and Godfried Toussaint. Convexifying polygons with simple projections. Information Processing Letters, 80(2):81– 86, October 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/80/37/
29/abstract.html.

Cheon:2006:KPC

[CKN06] Jung Hee Cheon, Woo-Hwan Kim, and Hyun Soo Nam. Known-plaintext cryptanalysis of the Domingo–Ferrer algebraic privacy homomorphism scheme. Information Processing Letters, 97(3):118–123, February 14, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Caprara:2000:PMS

[CKP00] Alberto Caprara, Hans Kellerer, and Ulrich Pferschy. A PTAS for the Multiple Subset Sum Problem with different knapsack capacities. Information Processing Letters, 73(3–4):111– 118, February 29, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/62/27/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/62/27/30/article.
                                        pdf.

Cohen:2006:EAG

[CKR06] Reuven Cohen, Liran Katzir, and Danny Raz. An efficient approximation for the Generalized Assignment Problem. Information Processing Letters, 100 (4):162–166, November 30, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cada:2009:DHC

[ČKRR09] Roman Čada, Tomáš Kaiser, Moshe Rosenfeld, and Zdeněk Ryjáček. Disjoint Hamilton cycles in the star graph. Information Processing Letters, 110(1): 30–35, December 1, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cachin:2009:FSC

[CKS09] Christian Cachin, Idit Keidar, and Alexander Shraer. Fork sequential consistency is blocking. Information Processing Letters, 109(7):360–364, March 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Collins:2007:NIS

[CKSY07] Michael J. Collins, David Kempe, Jared Saia, and Maxwell Young. Nonnegative integral subset representations of integer sets. Information Processing Letters, 101(3):129–133, February 14, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cygan:2009:ETA

[CKW09] Marek Cygan, Łukasz Kowalik, and Mateusz Wykurz. Exponential-time approximation of weighted set cover. Information Processing Letters, 109 (16):957–961, July 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chrobak:2006:RGA

[CKY06] Marek Chrobak, Claire Kenyon, and Neal Young. The reverse greedy algorithm for the metric k-median problem. Information Processing Letters, 97(2):68–72, January 31, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheng:2000:DTO

[CL00a] Eddie Cheng and Marc J. Lipman. On the Day-Tripathi orientation of the star graphs: Connectivity. Information Processing Letters, 73(1–2):5–10, January 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/62/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/62/25/26/article.
pdf.

Colbourn:2000:QDC

[CL00b] Charles J. Colbourn and Alan C. H. Ling. Quorums from difference covers. Information Processing Letters, 75(1– 2):9–12, July 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/25/26/article.
pdf.

Cheng:2003:ATM

[CL03] T. C. Edwin Cheng and Zhaohui Liu. 3/2-approximation for two-machine no-wait flowshop scheduling with availability constraints. Information Processing Letters, 88(4):161–165, November 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chunlin:2006:QBR

[CL06] Li Chunlin and Li Layuan. QoS based resource scheduling by computational economy in computational grid. Information Processing Letters, 98(3):119– 126, May 16, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheng:2009:OAA

[CLC09] Chih-Huai Cheng, Hsiao-Fei Liu, and Kun-Mao Chao. Optimal algorithms for the averageconstrained maximum-sum segment problem. Information Processing Letters, 109(3):171–174, January 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cleary:2002:RRD

[Cle02] Sean Cleary. Restricted rotation distance between binary trees. Information Processing Letters, 84(6):333–338, December 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:2004:SLH

[CLHH04] Chung-Haw Chang, Cheng-Kuan Lin, Hua-Min Huang, and Lih-Hsing Hsu. The super laceability of the hypercubes. Information Processing Letters, 92 (1):15–21, October 16, 2004.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:2004:NUB

[CLL04] Maw-Shang Chang, Chin-Hua Lin, and Chuan-Min Lee. New upper bounds on feedback vertex numbers in butterflies. Information Processing Letters, 90 (6):279–285, June 30, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chao:2008:SIE

[CLL08] Yi-Hsiung Chao, Shun-Shii Lin, and Kwei-Jay Lin. Schedulability issues for EDZL scheduling on real-time multiprocessor systems. Information Processing Letters, 107(5):158–164, August 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2009:RAC

[CLM09] Xing Chen, Juan Liu, and Jixiang Meng. The restricted arc connectivity of Cartesian product digraphs. Information Processing Letters, 109(21–22): 1202–1205, October 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chung:2007:EAR

[CLT07] Yun-Sheng Chung, Chin Lung Lu, and Chuan Yi Tang. Efficient algorithms for regular expression constrained sequence alignment. Information Processing Letters, 103(6):240–246, September 15, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:2002:GLE

[CLW02] Ji-Woong Chang, Young-Koo Lee, and Kyu-Young Whang. Global lock escalation in database management systems. Information Processing Letters, 82(4): 179–186, May 31, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2009:LTA

[CLZ09] Lei Chen, Changhong Lu, and Zhenbing Zeng. A linear-time algorithm for paired-domination problem in strongly chordal graphs. Information Processing Letters, 110(1):20–23, December 1, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2007:ILB

[CM07] Wenbin Chen and Jiangtao Meng. An improved lower bound for approximating Shortest Integer Relation in ℓ_{∞} norm (SIR_{∞}). Information Processing Letters, 101(4):174–179, February 28, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chandran:2005:CCG

[CMO05] L. S. Chandran, C. Mannino, and G. Oriolo. On the cubicity of certain graphs. Information Processing Letters, 94(3): 113–118, May 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Calude:2003:TCR

[CMS03] Cristian S. Calude, Solomon Marcus, and Ludwig Staiger. A topological characterization of random sequences. Information Processing Letters, 88(5):245– 250, December 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cai:2000:NNN

[CN00] Jin-Yi Cai and Ajay Nerurkar. A note on the non-NP-hardness of approximate lattice problems under general Cook reductions. Information Processing Letters, 76(1–2):61–66, November 20, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/24/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/24/33/article.
pdf.

Chavez:2003:PPS

[CN03] Edgar Chávez and Gonzalo Navarro. Probabilistic proximity search: Fighting the curse of dimensionality in metric spaces. Information Processing Letters, 85(1):39–46, January 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheng:2006:NAO

[CNK06] T. C. E. Cheng, C. T. Ng, and Vladimir Kotov. A new algorithm for online uniformmachine scheduling to minimize the makespan. Information Processing Letters, 99(3):102–105, August 16, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Coja-Oghlan:2006:IAA

[COK06] Amin Coja-Oghlan and Lars Kuhtz. An improved algorithm for approximating the chromatic number of G_{n,p}. Information Processing Letters, 99(6):234– 238, September 30, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Comellas:2000:DSW

[COP00] Francesc Comellas, Javier Ozón, and Joseph G. Peters. Deterministic small-world communication networks. Information Processing Letters, 76(1–2):83– 90, November 20, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/24/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/24/36/article.
pdf.

Chen:2005:IDA

[COW05] Zhi-Zhong Chen, Yuusuke Okamoto, and Lusheng Wang. Improved deterministic approximation algorithms for Max TSP.

Information Processing Letters, 95(2):333–342, July 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cantu-Paz:2002:OSS

[CP02] Erick Cantú-Paz. Order statistics and selection methods of evolutionary algorithms. Information Processing Letters, 82 (1):15–22, April 15, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/86/27/
                                        29/abstract.html.

Capponi:2005:BFL

[CP05a] Agostino Capponi and Concetta Pilotto. Bounded families for the on-line t-relaxed coloring. Information Processing Letters, 96 (4):141–145, November 30, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Caprara:2005:MSS

[CP05b] Alberto Caprara and Ulrich Pferschy. Modified subset sum heuristics for bin packing. Information Processing Letters, 96 (1):18–23, October 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Choi:2003:ERS

[CPC03] Yong-Jin Choi, Ho-Hyun Park, and Chin-Wan Chung. Estimating the result size of a query to velocity skewed moving objects. Information Processing Letters, 88(6):279–285, December 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2006:RGV

[CPX06] Zhenming Chen, Evanthia Papadopoulou, and Jinhui Xu. Robustness of k-gon Voronoi diagram construction. Information Processing Letters, 97(4):138– 145, February 28, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2009:FTA

[CQ09] Meirun Chen and Jianguo Qian. On f-fault tolerant arcforwarding and optical indices of all-optical folded hypercubes. Information Processing Letters, 109(15):828–831, July 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Caprara:2002:PTB

[CR02] Alberto Caprara and Romeo Rizzi. Packing triangles in bounded degree graphs. Information Processing Letters, 84 (4):175–180, November 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chakaravarthy:2009:AMW

[CR09] Venkatesan T. Chakaravarthy and Sambuddha Roy. Approximating maximum weight Kcolorable subgraphs in chordal graphs. Information Processing Letters, 109(7):365–368, March 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2007:TCP

[CRW07] Min Chen, André Raspaud, and Weifan Wang. Three-coloring planar graphs without short cycles. Information Processing Letters, 101(3):134–138, February 14, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chrobak:2000:SAH

[CS00] Marek Chrobak and Jiří Sgall. A simple analysis of the harmonic algorithm for two servers. Information Processing Letters, 75 (1–2):75–77, July 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/25/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/25/38/article.
pdf.

Chandran:2003:SLB

[CS03a] L. Sunil Chandran and C. R. Subramanian. A spectral lower bound for the treewidth of a graph and its consequences. Information Processing Letters, 87 (4):195–200, August 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crochemore:2003:WTC

[CS03b] M. Crochemore and V. T. Stefanov. Waiting time and complexity for matching patterns with automata. Information Processing Letters, 87(3):119– 125, August 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Codenotti:2005:CCN

[CŠ05] Bruno Codenotti and Daniel Štefankovič. On the computational complexity of Nash equilibria for (0,1) bimatrix games. Information Processing Letters, 94(3):145–150, May 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cleary:2009:RDF

[CS09a] Sean Cleary and Katherine St.

John. Rotation distance is fixedparameter tractable. Information Processing Letters, 109(16): 918–922, July 31, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cojocaru:2009:EDR

[CS09b] Alina Carmen Cojocaru and Igor E. Shparlinski. On the embedding degree of reductions of an elliptic curve. Information Processing Letters, 109(13):652– 654, June 15, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chin:2003:TDC

[CSW03] Francis Y. L. Chin, Hong Shen, and Fu Lee Wang. Transversal of disjoint convex polygons. Information Processing Letters, 85(1):55–60, January 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:2000:ESM

[CT00a] Keh-Ning Chang and Shi-Chun Tsai. Exact solution of a minimal recurrence. Information Processing Letters, 75(1– 2):61–64, July 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/25/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/25/35/article.
pdf.

Chen:2000:NCM

[CT00b] Liang Chen and Naoyuki Tokuda. A note on “category and measure in complexity classes”. Information Processing Letters, 73(5–6):167– 168, March 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/62/28/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/62/28/28/article.
pdf.

Chemillier:2003:CWS

[CT03a] Marc Chemillier and Charlotte Truchet. Computation of words satisfying the “rhythmic oddity property” (after Simha Arom’s works). Information Processing Letters, 86(5):255–261, June 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cleary:2003:BRR

[CT03b] Sean Cleary and Jennifer Taback. Bounding restricted rotation distance. Information Processing Letters, 88(5):251– 256, December 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cai:2004:NQR

[CT04] Jin-Yi Cai and Robert A.

Threlfall. A note on quadratic residuosity and UP. Information Processing Letters, 92(3):127– 131, November 15, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Colannino:2005:ACR

[CT05] Justin Colannino and Godfried Toussaint. An algorithm for computing the restriction scaffold assignment problem in computational biology. Information Processing Letters, 95(4): 466–471, August 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See corrigendum [CT07].

Colannino:2007:CAC

[CT07] Justin Colannino and Godfried Toussaint. Corrigendum to “An algorithm for computing the restriction Scaffold assignment problem in computational biology” [Inform. Process. Lett. 95 (4) (2005) 466–471]. Information Processing Letters, 103 (1):44, June 30, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [CT05].

Connie:2005:PNA

[CTGN05] Tee Connie, Andrew Teoh, Michael Goh, and David Ngo. Palmhashing: a novel approach for cancelable biometrics. Information Processing Letters, 93 (1):1–5, January 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2003:NDL

[CTN03] Liang Chen, Naoyuki Tokuda, and Akira Nagai. A new differential LSI space-based probabilistic document classifier. Information Processing Letters, 88 (5):203–212, December 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Curtis:2004:DHB

[Cur04] S. A. Curtis. Darts and hoopla board design. Information Processing Letters, 92(1):53–56, October 16, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cicalese:2000:IHU

[CV00] Ferdinando Cicalese and Ugo Vaccaro. An improved heuristic for “Ulam–Rényi game”. Information Processing Letters, 73(3–4):119–124, February 29, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/62/27/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/62/27/31/article.
pdf.

Cicalese:2003:BSD

[CV03] Ferdinando Cicalese and Ugo Vaccaro. Binary search with delayed and missing answers. Information Processing Letters, 85 (5):239–247, March 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Canete-Valdeon:2008:CSV

[CVEOV08] José M. Cañete-Valdeón, Fernando Enríquez, Javier Ortega, and Ernesto Veláquez. Clarifying the semantics of value in use cases through Jackson’s Problem Frames. Information Processing Letters, 107(6):221–229, August 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chin:2002:EAT

[CW02] Francis Y. L. Chin and Fu Lee Wang. Efficient algorithm for transversal of disjoint convex polygons. Information Processing Letters, 83(3):141–144, August 16, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [CW03].

Chin:2003:EEA

[CW03] Francis Y. L. Chin and Fu Lee Wang. Erratum to: “Efficient algorithm for transversal of disjoint convex polygons”. Information Processing Letters, 85 (1):53, January 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [CW02].

Cai:2004:RCB

[CW04] Jin-Yi Cai and Osamu Watanabe. Relativized collapsing between BPP and PH under stringent oracle access. Information Processing Letters, 90(3): 147–154, May 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheng:2007:CCN

[CW07] Hai-Lung Cheng and Biing-Feng Wang. On Chen and Chen’s new tree inclusion algorithm. Information Processing Letters, 103 (1):14–18, June 30, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:2009:DGP

[CW09] Jou-Ming Chang and Ro-Yu Wu. On the diameter of geometric path graphs of points in convex position. Information Processing Letters, 109(8):409–413, March 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2000:RTB

[CWC00] Ing-Ray Chen, Ding-Chau Wang, and Chih-Ping Chu. Response time behavior of distributed voting algorithms for managing replicated data. Information Processing Letters, 75(6):247–253, November 30, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/31/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/31/26/article.
pdf.

Chia:2002:PDT

[CWCL02] Tsorng-Lin Chia, Kuang-Bor Wang, Zen Chen, and Der-Chyuan Lou. Parallel distance transforms on a linear array architecture. Information Processing Letters, 82(2): 73–81, April 30, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/86/33/
28/abstract.html.

Chien:2001:CCW

[CWJT01] Hung-Yu Chien, Tzong-Chen Wu, Jinn-Ke Jan, and Yuh-Min Tseng. Cryptanalysis of Chang–Wu’s group-oriented authentication and key exchange protocols. Information Processing Letters, 80(2):113–117, October 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/80/37/
                                        33/abstract.html.

Chiang:2008:IA

[CWT08] M. L. Chiang, S. C. Wang, and L. Y. Tseng. The incremental agreement. Information Processing Letters, 107(5):165–170, August 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chrzastowski-Wachtel:2001:DCC

[CWTHR01] Piotr Chrza̧stowski-Wachtel, Jerzy Tyszkiewicz, Achim Hoffmann, and Arthur Ramer. Definability of connectives in conditional event algebras of Schay-Adams-Calabrese and Goodman-Nguyen-Walker. Information Processing Letters, 79 (4):155–160, August 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

                                        tronic).  URL http://www.
                                        elsevier.nl/gej-ng/10/23/
                                        20/80/33/25/abstract.html;
                                         http://www.elsevier.nl/gej-
                                        ng/10/23/20/80/33/25/article.
                                        pdf.

Chung:2000:NST

[CY00] Kuo-Liang Chung and Wen-Ming Yan. On the number of spanning trees of a multicomplete/star related graph. Information Processing Letters, 76(3):113–119, December 15, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/27/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/27/27/article.
pdf.

Chin:2001:SNP

[CY01a] Kien-Weh Chin and Hsu-Chun Yen. The symmetry number problem for trees. Information Processing Letters, 79(2): 73–79, June 30, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/31/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/80/31/28/article.
pdf.

Chung:2001:EAF

[CY01b] Kuo-Liang Chung and Wen-Ming Yan. An efficient algorithm for the Fourier transform on a compressed image in restricted quadtree and shading format. Information Processing Letters, 79(1):1–5, May 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.elsevier.nl/gej%2Dng/
10/23/20/80/25/25/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/23/20/80/25/
25/article.pdf.

Chung:2002:FDC

[CY02] Kuo-Liang Chung and Wen-Ming Yan. Fast 2D discrete cosine transform on compressed image in restricted quadtree and shading format. Information Processing Letters, 81(1): 13–21, January 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/85/27/
                                        29/abstract.html.

Campeanu:2004:PEP

[CY04] Cezar Câmpeanu and Sheng Yu. Pattern expressions and pattern automata. Information Processing Letters, 92(6):267–274, December 31, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2002:TUB

[CZ02] Zhi-Zhong Chen and Shiqing Zhang. Tight upper bound on the number of edges in a bipartite K_{3,3}-free or K_{5}-free graph with an application. Information Processing Letters, 84(3): 141–145, November 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:2009:DCP

[CZ09] Wěi Chén and Wenhui Zhang. A direct construction of polynomialsize OBDD proof of pigeon hole problem. Information Processing Letters, 109(10):472–477, April 30, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dai:2009:AAM

[Dai09] Liuling Dai. An aggressive algorithm for multiple string matching. Information Processing Letters, 109(11):553–559, May 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dalmau:2009:TNP

[Dal09] Víctor Dalmau. There are no pure relational width 2 constraint satisfaction problems. Information Processing Letters, 109(4):213–218, January 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Damaschke:2002:OMO

[Dam02] Peter Damaschke. Optimizing a mail-order with discount and shipping costs. Information Processing Letters, 82(2): 93–97, April 30, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/86/33/
30/abstract.html.

Damaschke:2006:RSP

[Dam06] Peter Damaschke. A remark on the subsequence problem for arc-annotated sequences with pairwise nested arcs. Information Processing Letters, 100(2): 64–68, October 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Darondeau:2005:ELC

[Dar05] Ph. Darondeau. Equality of languages coincides with isomorphism of reachable state graphs for bounded and persistent Petri nets. Information Processing Letters, 94(6): 241–245, June 30, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Davie:2004:CML

[Dav04] George Davie. Characterising the Martin-Löf random sequences using computably enumerable sets of measure one. Information Processing Letters, 92 (3):157–160, November 15, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dawande:2003:NCP

[Daw03] Milind Dawande. A notion of cross-perfect bipartite graphs. Information Processing Letters, 88(4):143–147, November 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dutra:2006:FRA

[DB06] Renato C. Dutra and Valmir C. Barbosa. Finding routes in anonymous sensor networks. Information Processing Letters, 98 (4):139–144, May 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Diaz-Banez:2006:FWE

[DBLS06] J. M. Díaz-Báñez, M. A. López, and J. A. Sellarès. On finding a widest empty 1-corner corridor. Information Processing Letters, 98(5):199–205, June 15, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deBerg:2008:COS

[dBT08] Mark de Berg and Shripad Thite. Cache-oblivious selection in sorted X + Y matrices. Information Processing Letters, 109 (2):87–92, December 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deBerg:2002:FMS

[dBvdS02] Mark de Berg and A. Frank van der Stappen. On the fatness of Minkowski sums. Information Processing Letters, 81 (5):259–264, March 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/85/36/31/abstract.html.

Deo:2007:PDD

[DC07] Narsingh Deo and Aurel Cami.

Preferential deletion in dynamic models of web-like networks. Information Processing Letters, 102(4):156–162, May 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dermid:2007:HRM

[DCS07] Eric Mc Dermid, Christine Cheng, and Ichiro Suzuki.

Hardness results on the manexchange stable marriage problem with short preference lists. Information Processing Letters, 101(1):13–19, January 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dubrois:2006:EPT

[DD06] Jacques Dubrois and Jean-Guillaume Dumas. Efficient polynomial time algorithms computing industrial-strength primitive roots. Information Processing Letters, 97(2):41–45, January 31, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Deng:2008:CAC

[DD08] Xiaotie Deng and Ye Du. The computation of approximate competitive equilibrium is PPAD-hard. Information Processing Letters, 108(6):369–373, November 30, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DeLoof:2008:RGM

[DDD08] K. De Loof, B. De Baets, and H. De Meyer. On the random generation of monotone data sets. Information Processing Letters, 107(6):216–220, August 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dahlhaus:2004:LTA

[DDR04] Elias Dahlhaus, Peter Dankelmann, and R. Ravi. A lineartime algorithm to compute a MAD tree of an interval graph. Information Processing Letters, 89(5):255–259, March 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Delaune:2006:EID

[Del06] Stéphanie Delaune. Easy intruder deduction problems with homomorphisms. Information Processing Letters, 97(6):213– 218, March 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Deorowicz:2005:CEA

[Deo05] Sebastian Deorowicz. Context exhumation after the Burrows– Wheeler transform. Information Processing Letters, 95(1): 313–320, July 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Deorowicz:2006:STI

[Deo06] Sebastian Deorowicz. Speeding up transposition-invariant string matching. Information Processing Letters, 100(1):14–20, October 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Deorowicz:2009:ASL

[Deo09] Sebastian Deorowicz. An algorithm for solving the longest increasing circular subsequence problem. Information Processing Letters, 109(12):630–634, May 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dolev:2009:IDN

[DEPZ09] Shlomi Dolev, Yuval Elovici, Rami Puzis, and Polina Zilberman. Incremental deployment of network monitors based on Group Betweenness Centrality. Information Processing Letters, 109(20):1172–1176, September 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dereniowski:2009:MVO

[Der09] Dariusz Dereniowski. Maximum vertex occupation time and inert fugitive: Recontamination does help. Information Processing Letters, 109(9):422–426, April 15, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Detti:2009:PAM

[Det09] Paolo Detti. A polynomial algorithm for the multiple knapsack problem with divisible item sizes. Information Processing Letters, 109(11):582–584, May 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Demetrescu:2003:CAF

[DF03] Camil Demetrescu and Irene Finocchi. Combinatorial algorithms for feedback problems in directed graphs. Information Processing Letters, 86(3): 129–136, May 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

daFonseca:2003:KHO

[dFdF03] Guilherme D. da Fonseca and Celina M. H. de Figueiredo. Kinetic heap-ordered trees: Tight analysis and improved algorithms. Information Processing Letters, 85(3):165–169, February 14, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deFigueiredo:2009:EWP

[dFdF09] Celina M. H. de Figueiredo and Guilherme D. da Fonseca. Enclosing weighted points with an almost-unit ball. Information Processing Letters, 109(21–22): 1216–1221, October 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

daFonseca:2004:KH

[dFdFC04] Guilherme D. da Fonseca, Celina M. H. de Figueiredo, and Paulo C. P. Carvalho. Kinetic hanger. Information Processing Letters, 89(3):151–157, February 14, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deFigueiredo:2005:NHS

[dFdS05] Celina M. H. de Figueiredo and Vinícius G. P. de Sá. Note on the Homogeneous Set Sandwich Problem. Information Processing Letters, 93(2):75–81, January 31, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Datar:2003:CAM

[DFG^{+}03] Mayur Datar, Tomás Feder, Aristides Gionis, Rajeev Motwani, and Rina Panigrahy. A combinatorial algorithm for MAX CSP. Information Processing Letters, 85(6):307–315, March 31, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Santis:2004:CKA

[DFM04] Alfredo De Santis, Anna Lisa Ferrara, and Barbara Masucci. Cryptographic key assignment schemes for any access control policy. Information Processing Letters, 92(4):199–205, November 30, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Downey:2008:PAD

[DFMR08] Rodney G. Downey, Michael R. Fellows, Catherine McCartin, and Frances Rosamond. Parameterized approximation of dominating set problems. Information Processing Letters, 109(1): 68–70, December 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Devillers:2000:LLS

[DG00] Raymond Devillers and Joël Goossens. Liu and Layland’s schedulability test revisited. Information Processing Letters, 73(5–6):157–161, March 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/62/28/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/62/28/26/article.
pdf.

Delgrande:2002:UC

[DG02] James P. Delgrande and Arvind Gupta. Updating ≤-chains. Information Processing Letters, 82 (5):261–268, June 15, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Downey:2008:TDR

[DG08] Rod Downey and Noam Greenberg. Turing degrees of reals of positive effective packing dimension. Information Processing Letters, 108(5):298–303, November 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dhamdhere:2003:CEH

[DGG03] Dhananjay M. Dhamdhere, K. Gururaja, and Prajakta G. Ganu. A compact execution history for dynamic slicing. Information Processing Letters, 85 (3):145–152, February 14, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dobson:2007:TOR

[DGHS07] Maria Patricia Dobson, Marisa Gutierrez, Michel Habib, and Jayme L. Szwarcfiter. On transitive orientations with restricted covering graphs. Information Processing Letters, 101(3):119– 125, February 14, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Das:2005:SPE

[DGN05] Sandip Das, Partha P. Goswami, and Subhas C. Nandy. Smallest k-point enclosing rectangle and square of arbitrary orientation. Information Processing Letters, 94(6):259–266, June 30, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dutta:2004:FNB

[DGP04] Partha Dutta, Rachid Guerraoui, and Bastian Pochon. Fast non-blocking atomic commit: an inherent trade-off. Information Processing Letters, 91(4):195– 200, August 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Datta:2005:SMP

[DGR05] Ajoy K. Datta, Maria Gradinariu, and Michel Raynal. Stabilizing mobile philosophers. Information Processing Letters, 95 (1):299–306, July 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deVGarcia:2001:CMP

[dH01] Alex de V. Garcia and Edward Hermann Haeusler. Code migration and program maintainability — A categorical perspective. Information Processing Letters, 79(5):249–254, September 15, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/80/34/
32/abstract.html.

Drake:2003:SAA

[DH03] Doratha E. Drake and Stefan Hougardy. A simple approximation algorithm for the weighted matching problem. Information Processing Letters, 85(4):211– 213, February 28, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Drake:2004:AAT

[DH04] Doratha E. Drake and Stefan Hougardy. On approximation algorithms for the terminal Steiner tree problem. Information Processing Letters, 89(1): 15–18, January 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Deng:2006:CME

[DH06] Xiaotie Deng and Li-Sha Huang.

On the complexity of market equilibria with maximum social welfare. Information Processing Letters, 97(1):4–11, January 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Durand:2008:CCP

[DH08] Arnaud Durand and Miki Hermann. On the counting complexity of propositional circumscription. Information Processing Letters, 106(4):164–170, May 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Derisavi:2003:OSS

[DHS03] Salem Derisavi, Holger Hermanns, and William H. Sanders. Optimal state-space lumping in Markov chains. Information Processing Letters, 87(6):309– 315, September 30, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dietel:2008:NOF

[DHS08] Jana Dietel, Hans-Dietrich Hecker, and Andreas Spillner. A note on optimal floodlight illumination of stages. Information Processing Letters, 105(4):121– 123, February 15, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Daza:2009:FSS

[DHS09] Vanesa Daza, Javier Herranz, and Germán Sáez. Flaws in some self-healing key distribution schemes with revocation. Information Processing Letters, 109(11):523–526, May 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Diaconescu:2004:HTA

[Dia04] Răzvan Diaconescu. Herbrand theorems in arbitrary institutions. Information Processing Letters, 90(1):29–37, April 15, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Diaconescu:2008:CSF

[Dia08] Răzvan Diaconescu. A categorical study on the finiteness of specifications. Information Processing Letters, 108(2):75– 80, September 30, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Diaconescu:2009:EPA

[Dia09] Răzvan Diaconescu. An encoding of partial algebras as total algebras. Information Processing Letters, 109(23–24):1245–1251, November 15, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dijkstra:2001:USL

[Dij01]  Edsger W. Dijkstra.  Un-

der the spell of Leibniz’s dream. Information Processing Letters, 77(2–4):53–61, February 28, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/29/25/article.
pdf.

Diller:2002:EMV

[Dil02] Antoni Diller. Efficient multivariate abstraction using an array representation for combinators. Information Processing Letters, 84(6):311–317, December 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dimauro:2000:RAI

[DIPS00] G. Dimauro, S. Impedovo, G. Pirlo, and A. Salzo. RNS architectures for the implementation of the ‘diagonal function’. Information Processing Letters, 73(5–6):189–198, March 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/62/28/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/62/28/32/article.
                                        pdf.

Dawar:2007:MTF

[DJ07] Anuj Dawar and David Janin. The monadic theory of finite representations of infinite words. Information Processing Letters, 103(3):94–101, July 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dunkelman:2008:TIV

[DK08] Orr Dunkelman and Nathan Keller. Treatment of the initial value in Time-Memory-Data Tradeoff attacks on stream ciphers. Information Processing Letters, 107(5):133–137, August 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Daylight:2009:TBI

[DKV09] Edgar G. Daylight, Wouter M. Koolen, and Paul M. B. Vitányi. Time-bounded incompressibility of compressible strings and sequences. Information Processing Letters, 109(18):1055–1059, August 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Durand-Lose:2000:RUS

[DL00]  Jérôme Durand-Lose.  Ran-

domized uniform self-stabilizing mutual exclusion. Information Processing Letters, 74(5–6): 203–207, June 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/28/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/28/28/article.
pdf.

Durand-Lose:2004:KTS

[DL04] Jérôme Durand-Lose. A Kleene theorem for splitable signals. Information Processing Letters, 89 (5):237–245, March 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dori:2006:CAC

[DL06] Shiri Dori and Gad M. Landau.

Construction of Aho–Corasick automaton in linear time for integer alphabets. Information Processing Letters, 98(2):66–72, April 30, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Delbot:2008:BLH

[DL08] François Delbot and Christian Laforest. A better list heuristic for vertex cover. Information Processing Letters, 107(3– 4):125–127, July 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DiGiacomo:2004:NOD

[DLP04] Emilio Di Giacomo, Giuseppe Liotta, and Maurizio Patrignani. A note on 3D orthogonal drawings with direction constrained edges. Information Processing Letters, 90(2):97–101, April 30, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dourado:2008:IAR

[DLPS08] Mitre C. Dourado, Min Chih Lin, Fábio Protti, and Jayme L. Szwarcfiter. Improved algorithms for recognizing p-Helly and hereditary p-Helly hypergraphs. Information Processing Letters, 108(4):247–250, October 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Diaz:2008:DCN

[DLS08] Josep Díaz, Zvi Lotker, and Maria Serna. The distant-2 chromatic number of random proximity and random geometric graphs. Information Processing Letters, 106(4):144–148, May 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

delaVega:2009:GRG

[dlVT09] W. Fernandez de la Vega and Zsolt Tuza. Groupies in random graphs. Information Processing Letters, 109(7):339–340, March 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Davida:2009:FAU

[DLX09] George Davida, Bruce Litow, and Guangwu Xu. Fast arithmetics using Chinese remaindering. Information Processing Letters, 109(13):660–662, June 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DeSantis:2000:SSS

[DM00a] Alfredo De Santis and Barbara Masucci. On secret set schemes. Information Processing Letters, 74(5–6):243– 251, June 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/28/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/28/34/article.
pdf.

Dimitrakos:2000:GMT

[DM00b] Theodosis Dimitrakos and Tom Maibaum. On a generalized modularization theorem. Information Processing Letters, 74 (1–2):65–71, April 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/63/25/
                                        34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/63/25/34/article.
                                        pdf.

Duarte:2000:RGD

[DM00c] Carlos H. C. Duarte and Tom Maibaum. A rely-guarantee discipline for open distributed systems design. Information Processing Letters, 74(1– 2):55–63, April 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/63/25/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/63/25/33/article.
                                        pdf.

Devroye:2003:CHF

[DM03] Luc Devroye and Pat Morin. Cuckoo hashing: Further analysis. Information Processing Letters, 86(4):215–219, May 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Diaz:2005:OSP

[DM05] R. Durán Díaz and J. Muñoz Masqué. Optimal strong primes. Information Processing Letters, 93(1):47–52, January 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Danvy:2008:EBS

[DM08a] Olivier Danvy and Kevin Millikin. On the equivalence between small-step and big-step abstract machines: a simple application of lightweight fusion. Information Processing Letters, 106(3):100–109, April 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Drysdale:2008:AAF

[DM08b] R. L. Scot Drysdale and Asish Mukhopadhyay. An O(n log n) algorithm for the all-farthest-segments problem for a planar set of points. Information Processing Letters, 105(2): 47–51, January 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Das:2009:SSP

[DM09] Ananya Das and Charles Martel. Stochastic shortest path with unlimited hops. Information Processing Letters, 109 (5):290–295, February 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Das:2009:IAW

[DMN09] Gautam K. Das, Debapriyay Mukhopadhyay, and Subhas C. Nandy. Improved algorithm for the widest empty 1-corner corridor. Information Processing Letters, 109(18):1060–1065, August 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Danvy:2005:CTB

[DN05] Olivier Danvy and Lasse R. Nielsen. CPS transformation of beta-redexes. Information Processing Letters, 94(5):217–224, June 15, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dereniowski:2006:VRC

[DN06] Dariusz Dereniowski and Adam Nadolski. Vertex rankings of chordal graphs and weighted trees. Information Processing Letters, 98(3):96–100, May 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Das:2008:WBL

[DN08] Gautam K. Das and Subhas C. Nandy. Weighted broadcast in linear radio networks. Information Processing Letters, 106(4): 136–143, May 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Doberkat:2004:FSR

[Dob04] Ernst-Erich Doberkat. Factoring stochastic relations. Information Processing Letters, 90(4): 161–166, May 31, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Doerr:2004:GRS

[Doe04] Benjamin Doerr. Global roundings of sequences. Information Processing Letters, 92(3):113– 116, November 15, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Duffy:2008:CAD

[DOS08] K. R. Duffy, N. O’Connell, and A. Sapozhnikov. Complexity analysis of a decentralised graph colouring algorithm. Information Processing Letters, 107(2): 60–63, July 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Doyen:2007:RPR

[Doy07] Laurent Doyen. Robust parametric reachability for timed automata. Information Processing Letters, 102(5):208–213, May 31, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DeMarco:2001:FBU

[DP01] Gianluca De Marco and Andrzej Pelc. Faster broadcasting in unknown radio networks. Information Processing Letters, 79 (2):53–56, June 30, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/31/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/80/31/25/article.
pdf.

Damian:2006:AHD

[DP06] Mirela Damian and Sriram V.

Pemmaraju. APX-hardness of domination problems in circle graphs. Information Processing Letters, 97(6):231–237, March 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dieudonne:2007:CFW

[DP07] Yoann Dieudonné and Franck Petit. Circle formation of weak robots and Lyndon words. Information Processing Letters, 101 (4):156–162, February 28, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dimakopoulos:2005:HCP

[DPP05] Vassilios V. Dimakopoulos, Leonidas Palios, and Athanasios S. Poulakidas. On the Hamiltonicity of the Cartesian product. Information Processing Letters, 96(2):49–53, October 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dovier:2000:NCC

[DPR00] Agostino Dovier, Enrico Pontelli, and Gianfranco Rossi. A necessary condition for Constructive Negation in Constraint Logic Programming. Information Processing Letters, 74(3–4): 147–156, May 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/63/27/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/63/27/33/article.
                                        pdf.

Domaratzki:2002:SFA

[DPS02] Michael Domaratzki, Giovanni Pighizzini, and Jeffrey Shallit. Simulating finite automata with context-free grammars. Information Processing Letters, 84 (6):339–344, December 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dourado:2005:HPS

[DPS05] Mitre C. Dourado, Fábio Protti, and Jayme L. Szwarcfiter. The Helly property on subfamilies of limited size. Information Processing Letters, 93(2):53–56, January 31, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dourado:2006:CAG

[DPS06] Mitre C. Dourado, Fábio Protti, and Jayme L. Szwarcfiter. Complexity aspects of generalized Helly hypergraphs. Information Processing Letters, 99(1):13–18, July 16, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Danvy:2006:OBM

[DR06] Olivier Danvy and Henning Korsholm Rohde. On obtaining the Boyer–Moore stringmatching algorithm by partial evaluation. Information Processing Letters, 99(4):158–162, August 31, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Das:2000:LBA

[DS00] Gautam Das and Michiel Smid.

A lower bound for approximating the geometric minimum weight matching. Information Processing Letters, 74(5–6): 253–255, June 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/28/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/28/35/article.
pdf.

Dumitrescu:2001:STT

[DS01] Adrian Dumitrescu and William Steiger. Space-time trade-offs for some ranking and searching queries. Information Processing Letters, 79(5):237–241, September 15, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/80/34/
30/abstract.html.

Dinur:2004:HAL

[DS04a] Irit Dinur and Shmuel Safra. On the hardness of approximating label-cover. Information Processing Letters, 89(5):247–254, March 16, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dougherty:2004:CCP

[DS04b] Daniel J. Dougherty and Stanley M. Selkow. The complexity of the certification of properties of Stable Marriage. Information Processing Letters, 92 (6):275–277, December 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dolama:2006:OCN

[DS06] Mohammad Hosseini Dolama and Éric Sopena. On the oriented chromatic number of Halin graphs. Information Processing Letters, 98(6):247–252, June 30, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carvalho:2008:EGB

[dSCLP08] Ricardo dos Santos Carvalho, Carlile Lavor, and Fábio Protti. Extending the geometric buildup algorithm for the molecular distance geometry problem. Information Processing Letters, 108(4):234–237, October 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dobrev:2000:EGC

[DSSV00] Stefan Dobrev, Heiko Schröder, Ondrej Sýkora, and Imrich Vrt’o. Evolutionary graph colouring. Information Processing Letters, 76(1–2):91–94, November 20, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/24/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/24/37/article.
pdf.

Droste:2008:WAD

[DSV08] Manfred Droste, Jacques Sakarovitch, and Heiko Vogler. Weighted automata with discounting. Information Processing Letters, 108 (1):23–28, September 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dorn:2009:QQC

[DT09] Sebastian Dörn and Thomas Thierauf. The quantum query complexity of the determinant. Information Processing Letters, 109(6):325–328, February 28, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Du:2004:OPS

[Du04] Donglei Du. Optimal preemptive semi-online scheduling on two uniform processors. Information Processing Letters, 92 (5):219–223, December 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Duquesne:2007:IAE

[Duq07] Sylvain Duquesne. Improving the arithmetic of elliptic curves in the Jacobi model. Information Processing Letters, 104(3): 101–105, October 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Durand:2003:AAO

[Dur03] Marianne Durand. Asymptotic analysis of an optimized quicksort algorithm. Information Processing Letters, 85(2):73–77, January 31, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Durgan:2004:CSS

[Dur04] Bruce K. Durgan. Compact searchable static binary trees. Information Processing Letters, 89(1):49–52, January 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Durand:2004:KLS

[DV04] Bruno Durand and Nikolai Vereshchagin. Kolmogorov– loveland stochasticity for finite strings. Information Processing Letters, 91(6):263–269, September 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Devillers:2006:BUS

[DV06] Raymond Devillers and Laurent Van Begin. Boundedness undecidability for synchronized nets. Information Processing Letters, 99(5):208–214, September 15, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dankelmann:2009:MSG

[DV09] Peter Dankelmann and Lutz Volkmann. Minimum size of a graph or digraph of given radius. Information Processing Letters, 109(16):971–973, July 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dietzfelbinger:2007:CAC

[DW07] Martin Dietzfelbinger and Henning Wunderlich. A characterization of average case communication complexity. Information Processing Letters, 101(6): 245–249, March 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deWerra:2003:USS

[dWH03] D. de Werra and P. Hansen. Using stable sets to bound the chromatic number. Information Processing Letters, 87(3): 127–131, August 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dong:2008:EFD

[DYZ08] Qiang Dong, Xiaofan Yang, and Juan Zhao. Embedding a family of disjoint multi-dimensional meshes into a crossed cube. Information Processing Letters, 108(6):394–397, November 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Eisenstat:2007:VDF

[EA07] David Eisenstat and Dana Angluin. The VC dimension of kfold union. Information Processing Letters, 101(5):181–184, March 16, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fujita:2002:ICC

[eF02] Ken etsu Fujita. An interpretation of λ^{μ}-calculus in λcalculus. Information Processing Letters, 84(5):261–264, December 16, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Efraimidis:2008:CLP

[Efr08] Pavlos S. Efraimidis. The complexity of linear programming in (γ, κ)-form. Information Processing Letters, 105(5):199–201, February 29, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Eisenbrand:2003:DDC

[EG03] Friedrich Eisenbrand and Fabrizio Grandoni. Detecting directed 4-cycles still faster. Information Processing Letters, 87 (1):13–15, July 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Engelfriet:2007:ESC

[EG07] Joost Engelfriet and Tjalling Gelsema. An exercise in structural congruence. Information Processing Letters, 101(1):1–5, January 16, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ehrenfeucht:2000:PSC

[EHHR00] Andrzej Ehrenfeucht, Jurriaan Hage, Tero Harju, and Grzegorz Rozenberg. Pancyclicity in switching classes. Information Processing Letters, 73(5–6): 153–156, March 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/62/28/
                                        25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/62/28/25/article.
                                        pdf.

Ettinger:2004:QQC

[EHK04] Mark Ettinger, Peter Høyer, and Emanuel Knill. The quantum query complexity of the hidden subgroup problem is polynomial. Information Processing Letters, 91(1):43–48, July 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Efrat:2006:GGT

[EHP06] Alon Efrat and Sariel Har-Peled. Guarding galleries and terrains. Information Processing Letters, 100(6):238–245, December 31, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Erdelyi:2009:FCV

[EHRS09] Gábor Erdélyi, Lane A. Hemaspaandra, Jörg Rothe, and Holger Spakowski. Frequency of correctness versus average polynomial time. Information Processing Letters, 109(16):946–949, July 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Eidenbenz:2002:AAT

[Eid02] Stephan Eidenbenz. Approximation algorithms for terrain guarding. Information Processing Letters, 82(2):99– 105, April 30, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/86/33/
31/abstract.html.

Eiter:2002:RDD

[EIM02] Thomas Eiter, Toshihide Ibaraki, and Kazuhisa Makino. Recognition and dualization of disguised bidual Horn functions. Information Processing Letters, 82(6): 283–291, June 30, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Eisenbrand:2001:SVP

[Eis01] Friedrich Eisenbrand. Short vectors of planar lattices via continued fractions. Information Processing Letters, 79(3): 121–126, July 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/32/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/80/32/28/article.
pdf.

Eisenstat:2009:FUL

[Eis09] David Eisenstat. k-Fold unions of low-dimensional concept classes. Information Processing Letters, 109(23–24):1232–1234, November 15, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Efrat:2000:CEB

[EK00] Alon Efrat and Matthew J. Katz. Computing Euclidean bottleneck matchings in higher dimensions. Information Processing Letters, 75(4):169–174, September 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/29/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/29/29/article.
pdf.

Ellis:2000:CCP

[EKF00] John Ellis, Tobias Krahn, and Hongbing Fan. Computing the cycles in the perfect shuffle permutation. Information Processing Letters, 75(5):217– 224, October 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/30/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/30/29/article.
pdf.

Epstein:2008:AFP

[EL08] Leah Epstein and Asaf Levin. Asymptotic fully polynomial approximation schemes for variants of open-end bin packing. Information Processing Letters, 109(1):32–37, December 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Elmasry:2009:CSP

[Elm09] Amr Elmasry. Computing the subset partial order for dense families of sets. Information Processing Letters, 109(18): 1082–1086, August 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Esperet:2008:IUG

[ELO08] Louis Esperet, Arnaud Labourel, and Pascal Ochem. On induceduniversal graphs for the class of bounded-degree graphs. Information Processing Letters, 108 (5):255–260, November 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Elkin:2007:NLB

[ELR07] Michael Elkin, Christian Liebchen, and Romeo Rizzi. New length bounds for cycle bases. Information Processing Letters, 104 (5):186–193, November 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Elbassioni:2006:UBN

[ELS06] Khaled Elbassioni, Zvi Lotker, and Raimund Seidel. Upper bound on the number of vertices of polyhedra with 0, 1-constraint matrices. Information Processing Letters, 100(2):69–71, October 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Engelfriet:2006:EPD

[EM06] Joost Engelfriet and Sebastian Maneth. The equivalence problem for deterministic MSO tree transducers is decidable. Information Processing Letters, 100 (5):206–212, December 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Engelhardt:2008:SBM

[EM08] Kai Engelhardt and Yoram Moses. Single-bit messages are insufficient for data link over duplicating channels. Information Processing Letters, 107(6): 235–239, August 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Escoffier:2006:WCF

[EMP06] Bruno Escoffier, Jérôme Monnot, and Vangelis Th. Paschos. Weighted Coloring: further complexity and approximability results. Information Processing Letters, 97(3):98–103, February 14, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Erk:2007:DCS

[EN07] Katrin Erk and Joachim Niehren. Dominance constraints in stratified context unification. Information Processing Letters, 101 (4):141–147, February 28, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Engebretsen:2004:STA

[Eng04] Lars Engebretsen. Simplified tight analysis of Johnson’s algorithm. Information Processing Letters, 92(4):207–210, November 30, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Engelfriet:2007:KCC

[Eng07] Joost Engelfriet. A Kleene characterization of computability. Information Processing Letters, 101(4):139–140, February 28, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Esperet:2007:OCO

[EO07] Louis Esperet and Pascal Ochem. Oriented colorings of 2-outerplanar graphs. Information Processing Letters, 101(5): 215–219, March 16, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Epstein:2000:NLS

[Eps00]  Leah Epstein.  A note on

on-line scheduling with precedence constraints on identical machines. Information Processing Letters, 76(4–6):149– 153, December 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/28/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/28/24/article.
pdf.

El-Qawasmeh:2004:WPU

[EQ04] Eyas El-Qawasmeh. Word prediction using a clustered optimal binary search tree. Information Processing Letters, 92 (5):257–265, December 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Effler:2003:CAG

[ER03] Scott Effler and Frank Ruskey. A CAT algorithm for generating permutations with a fixed number of inversions. Information Processing Letters, 86(2): 107–112, April 30, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Even:2005:HSW

[ERS05] Guy Even, Dror Rawitz, and Shimon (Moni) Shahar. Hitting sets when the VC-dimension is small. Information Processing Letters, 95(2):358–362, July 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Eiben:2002:EC

[ES02] A. E. Eiben and M. Schoenauer. Evolutionary computing. Information Processing Letters, 82 (1):1–6, April 15, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/86/27/
                                        27/abstract.html.

Efraimidis:2006:WRS

[ES06] Pavlos S. Efraimidis and Paul G.

Spirakis. Weighted random sampling with a reservoir. Information Processing Letters, 97(5): 181–185, March 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Enright:2007:SFG

[ES07] Jessica Enright and Lorna Stewart. Subtree filament graphs are subtree overlap graphs. Information Processing Letters, 104 (6):228–232, December 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ergun:2002:IFR

[ESZ02] Funda Ergun, Rakesh Sinha, and Lisa Zhang. An improved FPTAS for Restricted Shortest Path. Information Processing Letters, 83(5):287–291, September 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Esteban:2003:CCT

[ET03] Juan Luis Esteban and Jacobo Torán. A combinatorial characterization of treelike resolution space. Information Processing Letters, 87(6):295–300, September 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Etessami:2000:NQP

[Ete00] Kousha Etessami. A note on a question of Peled and Wilke regarding stutter-invariant LTL.

Information Processing Letters, 75(6):261–263, November 30, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/31/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/31/28/article.
pdf.

Fukagawa:2005:PAG

[FA05] Daiji Fukagawa and Tatsuya Akutsu. Performance analysis of a greedy algorithm for inferring Boolean functions. Information Processing Letters, 93(1):7–12, January 16, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fan:2002:HCC

[Fan02] Jianxi Fan. Hamilton-connectivity and cycle-embedding of the Möbius cubes. Information Processing Letters, 82(2):113– 117, April 30, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/86/33/
33/abstract.html.

Freschi:2004:LCS

[FB04] Valerio Freschi and Alessandro Bogliolo. Longest common subsequence between runlength-encoded strings: a new algorithm with improved parallelism. Information Processing Letters, 90(4):167–173, May 31, 2004. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Fujito:2004:ANA

[FD04] Toshihiro Fujito and Takashi Doi. A 2-approximation NC algorithm for connected vertex cover and tree cover. Information Processing Letters, 90(2): 59–63, April 30, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fecher:2004:CHT

[Fec04] Harald Fecher. A completed hierarchy of true concurrent equivalences. Information Processing Letters, 89(5):261–265, March 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Feijen:2001:JFM

[Fei01] W. H. J. Feijen. The joy of formula manipulation. Information Processing Letters, 77(2–4): 89–96, February 28, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/29/30/article.
pdf.

Fertin:2000:HBG

[Fer00] Guillaume Fertin. Hierarchical broadcast and gossip networks. Information Processing Letters, 73(3–4):131–136, February 29, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/62/27/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/62/27/33/article.
                                        pdf.

Fernandez:2005:PTR

[Fer05a] Mirtha-Lina Fernández. On proving C_{E}-termination of rewriting by size-change termination. Information Processing Letters, 93(4):155–162, February 28, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fernandez:2005:RMI

[Fer05b] Mirtha-Lina Fernández. Relaxing monotonicity for innermost termination. Information Processing Letters, 93(3):117–123, February 14, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Feofiloff:2007:PDA

[FFFdP07] Paulo Feofiloff, Cristina G. Fernandes, Carlos E. Ferreira, and José Coelho de Pina. Primaldual approximation algorithms for the Prize-Collecting Steiner Tree Problem. Information Processing Letters, 103(5):195–202, August 31, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fredriksson:2006:GCA

[FG06] Kimmo Fredriksson and Szymon Grabowski. A general compression algorithm that supports fast searching. Information Processing Letters, 100(6):226–232, December 31, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fomin:2009:SSE

[FGK^{+}09a] Fedor V. Fomin, Petr A. Golovach, Jan Kratochvíl, Dieter Kratsch, and Mathieu Liedloff. Sort and Search: Exact algorithms for generalized domination. Information Processing Letters, 109(14):795–798, June 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Frati:2009:PPT

[FGK09b] Fabrizio Frati, Markus Geyer, and Michael Kaufmann. Planar packing of trees and spider trees. Information Processing Letters, 109(6):301–307, February 28, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fleischer:2004:FOP

[FGLW04] Rudolf Fleischer, Mordecai Golin, Chin-Tau Lea, and Steven Wong. Finding optimal paths in MREP routing. Information Processing Letters, 89(2):57–63, January 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fertin:2002:MFV

[FGR02] Guillaume Fertin, Emmanuel Godard, and André Raspaud. Minimum feedback vertex set and acyclic coloring. Information Processing Letters, 84(3): 131–139, November 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fertin:2003:ADC

[FGR03] Guillaume Fertin, Emmanuel Godard, and André Raspaud. Acyclic and k-distance coloring of the grid. Information Processing Letters, 87(1):51–58, July 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fomin:2006:PCG

[FH06] Fedor V. Fomin and Kjartan Høie. Pathwidth of cubic graphs and exact algorithms. Information Processing Letters, 97(5): 191–196, March 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fischer:2006:INC

[FHK06] Felix Fischer, Markus Holzer, and Stefan Katzenbeisser. The influence of neighbourhood and choice on the complexity of finding pure Nash equilibria. Information Processing Letters, 99(6):239–245, September 30, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fiedorowicz:2008:AAE

[FHN08] Anna Fiedorowicz, Mariusz Hałuszczak, and Narayanan Narayanan. About acyclic edge colourings of planar graphs. Information Processing Letters, 108(6):412–417, November 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fishburn:2002:SSD

[Fis02] John P. Fishburn. Solving a system of difference constraints with variables restricted to a finite set. Information Processing Letters, 82(3):143–144, May 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/86/34/31/abstract.html.

Fujiwara:2008:OCP

[FIY08] Hiroshi Fujiwara, Kazuo Iwama, and Kouki Yonezawa. Online chasing problems for regular polygons. Information Processing Letters, 108(3):155–159, October 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Faliszewski:2005:PUH

[FJ05] Piotr Faliszewski and Janusz Jarosz. Properties of uniformly hard languages. Information Processing Letters, 95(1):329– 332, July 16, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fan:2008:SCE

[FJ08a] GaoJun Fan and ShiYao Jin. A simple coverage-evaluating approach for wireless sensor networks with arbitrary sensing areas. Information Processing Letters, 106(4):159–161, May 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fitzgerald:2008:CBT

[FJ08b] John S. Fitzgerald and Cliff B. Jones. The connection between two ways of reasoning about partial functions. Information Processing Letters, 107(3–4):128– 132, July 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Flanagan:2001:AIM

[FJL01] Cormac Flanagan, Rajeev Joshi, and K. Rustan M. Leino. Annotation inference for modular checkers. Information Processing Letters, 77(2–4):97–108, February 28, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/29/31/article.
pdf.

Feige:2000:FNB

[FK00a] Uriel Feige and Joe Kilian. Finding OR in a noisy broadcast network. Information Processing Letters, 73(1–2):69– 75, January 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/62/25/
36/abstract.html;   http:

//www.elsevier.nl/gej-ng/
29/28/24/62/25/36/article.
pdf.

Freund:2000:LBI

[FK00b] Ari Freund and Howard Karloff.

A lower bound of 8/(7 + \frac{1}{k−1}) on the integrality ratio of the Călinescu-Karloff-Rabani relaxation for multiway cut. Information Processing Letters, 75(1– 2):43–50, July 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/25/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/25/32/article.
pdf.

Feige:2001:NAM

[FKL01] Uriel Feige, Marek Karpinski, and Michael Langberg. A note on approximating Max-Bisection on regular graphs. Information Processing Letters, 79(4):181–188, August 15, 2001. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/33/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/80/33/29/article.
pdf.

Fayyazi:2006:ALT

[FKM06] Morteza Fayyazi, David Kaeli, and Waleed Meleis. An adjustable linear time parallel algorithm for maximum weight bipartite matching. Information Processing Letters, 97(5): 186–190, March 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fomin:2002:AMC

[FKN02] Fedor V. Fomin, Dieter Kratsch, and Jean-Christophe Novelli. Approximating minimum cocolorings. Information Processing Letters, 84(5):285–290, December 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fulop:2004:BCD

[FKV04] Zoltán Fülöp, Armin Kühnemann, and Heiko Vogler. A bottomup characterization of deterministic top-down tree transducers with regular look-ahead. Information Processing Letters, 91 (2):57–67, July 31, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fink:2009:NLB

[FKW09] Irene Fink, Sven O. Krumke, and Stephan Westphal. New lower bounds for online k-server routing problems. Information Processing Letters, 109(11):563– 567, May 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Furmanczyk:2008:NMT

[FKŻ08] Hanna Furmańczyk, Adrian Kosowski, and Paweł Żyliński. A note on mixed tree coloring. Information Processing Letters, 106(4):133–135, May 16, 2008.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fomin:2002:AAT

[FL02] Fedor V. Fomin and Andrzej Lingas. Approximation algorithms for time-dependent orienteering. Information Processing Letters, 83(2):57–62, July 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fang:2005:EIH

[FL05a] Jywe-Fei Fang and Kuan-Chou Lai. Embedding the incomplete hypercube in books. Information Processing Letters, 96(1): 1–6, October 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Finkel:2005:CHR

[FL05b] Alain Finkel and Jérôme Leroux. The convex hull of a regular set of integer vectors is polyhedral and effectively computable. Information Processing Letters, 96(1):30–35, October 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Feige:2007:IAR

[FL07] Uriel Feige and James R. Lee.

An improved approximation ratio for the minimum linear arrangement problem. Information Processing Letters, 101(1): 26–29, January 16, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Finkel:2009:DPT

[FL09] Olivier Finkel and Dominique Lecomte. Decision problems for Turing machines. Information Processing Letters, 109(23–24): 1223–1226, November 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fan:2005:NPE

[FLJ05] Jianxi Fan, Xiaola Lin, and Xiaohua Jia. Node-pancyclicity and edge-pancyclicity of crossed cubes. Information Processing Letters, 93(3):133–138, February 14, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Focardi:2000:FVS

[FLP00] Riccardo Focardi, Flaminia L. Luccio, and David Peleg. Feedback vertex set in hypercubes. Information Processing Letters, 76(1–2):1–5, November 20, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/67/24/
                                        24/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/67/24/24/article.
                                        pdf.

Fulop:2000:DPA

[FM00] Zoltán Fülöp and Sebastian Maneth. Domains of partial attributed tree transducers. Information Processing Letters, 73(5–6):175–180, March 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/62/28/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/62/28/30/article.
pdf.

Fernau:2003:SRS

[FM03] H. Fernau and A. Meduna. A simultaneous reduction of several measures of descriptional complexity in scattered context grammars. Information Processing Letters, 86(5):235–240, June 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Frandsen:2005:RBC

[FM05] Gudmund Skovbjerg Frandsen and Peter Bro Miltersen. Reviewing bounds on the circuit size of the hardest functions. Information Processing Letters, 95 (2):354–357, July 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fredriksson:2006:EPS

[FM06] Kimmo Fredriksson and Maxim Mozgovoy. Efficient parameterized string matching. Information Processing Letters, 100(3): 91–96, November 15, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Feder:2009:GTP

[FM09] Tomás Feder and Rajeev Motwani. On the graph turnpike problem. Information Processing Letters, 109(14):774– 776, June 30, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fecher:2002:BES

[FMCW02] Harald Fecher, Mila Majster-Cederbaum, and Jinzhao Wu. Bundle event structures: A revised cpo approach. Information Processing Letters, 83(1):7– 12, July 16, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Friedman:2004:WFD

[FMR04] Roy Friedman, Achour Mostefaoui, and Michel Raynal. A weakest failure detector-based asynchronous consensus protocol for f < n. Information Processing Letters, 90(1):39–46, April 15, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Friedman:2005:ABL

[FMR05] Roy Friedman, Achour Mostefaoui, and Michel Raynal. Asynchronous bounded lifetime failure detectors. Information Processing Letters, 94(2):85–91, April 30, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fidytek:2001:ACM

[FMSS01] Robert Fidytek, Andrzej W. Mostowski, Rafał Somla, and Andrzej Szepietowski. Algorithms counting monotone Boolean functions. Information Processing Letters, 79(5): 203–209, September 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/80/34/25/abstract.html.

Fiadeiro:2003:F

[FMT03] José Luiz Fiadeiro, Jan Madey, and Andrzej Tarlecki. Foreword. Information Processing Letters, 88(1–2):1–2, October 17, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Feder:2009:AMC

[FMT09] Tomás Feder, Heikki Mannila, and Evimaria Terzi. Approximating the Minimum Chain Completion problem. Information Processing Letters, 109(17): 980–985, August 16, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fouquet:2002:NOP

[Fou02] Jean-Luc Fouquet. No odd pairs in minimal imperfect NP_{5} graphs. Information Processing Letters, 83(4):201–204, August 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fraczak:2004:CL

[FP04] Wojciech Fraczak and Anna Podolak. A characterization of s-languages. Information Processing Letters, 89(2):65–70, January 31, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fusco:2008:ABA

[FP08] Emanuele G. Fusco and Andrzej Pelc. Acknowledged broadcasting in ad hoc radio networks. Information Processing Letters, 109(2):136–141, December 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fosse:2004:SCP

[FR04] S. Fossé and G. Richomme. Some characterizations of Parikh matrix equivalent binary words. Information Processing Letters, 92(2):77–82, October 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Feige:2006:HAM

[FR06] Uriel Feige and Daniel Reichman. On the hardness of approximating Max-Satisfy. Information Processing Letters, 97(1): 31–35, January 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fertin:2008:ACG

[FR08] Guillaume Fertin and André Raspaud. Acyclic coloring of graphs of maximum degree five: Nine colors are enough. Information Processing Letters, 105 (2):65–72, January 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fredman:2003:NTR

[Fre03a] Michael L. Fredman. The number of tests required to search an unordered table. Information Processing Letters, 87(2):85–88, July 31, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fredriksson:2003:SSM

[Fre03b]  Kimmo Fredriksson.  Shift-

or string matching with superalphabets. Information Processing Letters, 87(4):201–204, August 31, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fredriksson:2005:EDC

[Fre05] Kimmo Fredriksson. Exploiting distance coherence to speed up range queries in metric indexes. Information Processing Letters, 95(1):287–292, July 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Froschle:2005:DBH

[Frö05] S. B. Fröschle. The decidability border of hereditary history preserving bisimilarity. Information Processing Letters, 93(6): 289–293, March 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fertin:2003:OCN

[FRR03] Guillaume Fertin, André Raspaud, and Arup Roychowdhury. On the oriented chromatic number of grids. Information Processing Letters, 85(5):261–266, March 16, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Farr:2008:NGP

[FS08] Graham Farr and Johannes Schmidt. On the number of Go positions on lattice graphs. Information Processing Letters, 105(4):124–130, February 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Feder:2009:NTB

[FS09] Tomás Feder and Carlos Subi. Nearly tight bounds on the number of Hamiltonian circuits of the hypercube and generalizations. Information Processing Letters, 109(5):267–272, February 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fuster-Sabater:2003:BNG

[FSGM03] Amparo Fúster-Sabater and Pedro García-Mochales. On the balancedness of nonlinear generators of binary sequences. Information Processing Letters, 85 (2):111–116, January 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fekete:2003:CEE

[FSW03] Sándor P. Fekete, Martin Skutella, and Gerhard J. Woeginger. The complexity of economic equilibria for house allocation markets. Information Processing Letters, 88(5):219– 223, December 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fu:2008:CFH

[Fu08a] Jung-Sheng Fu. Conditional fault Hamiltonicity of the complete graph. Information Processing Letters, 107(3–4):110– 113, July 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fu:2008:FFC

[Fu08b] Jung-Sheng Fu. Fault-free cycles in folded hypercubes with more faulty elements. Information Processing Letters, 108(5):261– 263, November 15, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fuchs:2003:NTS

[Fuc03] Bernhard Fuchs. A note on the terminal Steiner tree problem. Information Processing Letters, 87(4):219–220, August 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fujito:2001:AICa

[Fuj01a] Toshihiro Fujito. On approximability of the independent/ connected edge dominating set problems. Information Processing Letters, 79(6):261–266, September 30, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.

com/gej-ng/10/23/20/80/37/
28/abstract.html.

Fujito:2001:AICb

[Fuj01b] Toshihiro Fujito. On approximability of the independent/ connected edge dominating set problems. Information Processing Letters, 79(6):261–266, September 30, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/80/37/
28/abstract.html.

Fujita:2003:LGP

[Fuj03] Satoshi Fujita. On-line gridpacking with a single active grid. Information Processing Letters, 85(4):199–204, February 28, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fujiyoshi:2005:LNM

[Fuj05] Akio Fujiyoshi. Linearity and nondeletion on monadic contextfree tree grammars. Information Processing Letters, 93(3):103– 107, February 14, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fung:2008:LBO

[Fun08] Stanley P. Y. Fung. Lower bounds on online deadline scheduling with preemption penalties. Information Processing Letters, 108(4):214–218, October 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Friedman:2003:CCC

[FVC03] Roy Friedman, Roman Vitenberg, and Gregory Chockler. On the composability of consistency conditions. Information Processing Letters, 86(4):169–176, May 31, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Feige:2003:CFB

[FY03] Uriel Feige and Orly Yahalom.

On the complexity of finding balanced oneway cuts. Information Processing Letters, 87(1):1– 5, July 16, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gabow:2000:PBD

[Gab00] Harold N. Gabow. Path-based depth-first search for strong and biconnected components. Information Processing Letters, 74(3– 4):107–114, May 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/63/27/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/63/27/28/article.
pdf.

Gagie:2005:RBS

[Gag05] Travis Gagie. Restructuring binary search trees revisited. Information Processing Letters, 95 (3):418–421, August 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gagie:2006:CPD

[Gag06a]  Travis Gagie.  Compressing

probability distributions. Information Processing Letters, 97 (4):133–137, February 28, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gagie:2006:LAI

[Gag06b] Travis Gagie. Large alphabets and incompressibility. Information Processing Letters, 99(6): 246–251, September 30, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gagie:2007:DSC

[Gag07] Travis Gagie. Dynamic Shannon coding. Information Processing Letters, 102(2–3):113–117, April 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gagie:2008:DAC

[Gag08a] Travis Gagie. Dynamic asymmetric communication. Information Processing Letters, 108 (6):352–355, November 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gagie:2008:SSM

[Gag08b] Travis Gagie. Sorting streamed multisets. Information Processing Letters, 108(6):418–421, November 30, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Galbiati:2003:FCB

[Gal03] Giulia Galbiati. On finding cycle bases and fundamental cycle bases with a shortest maximal cycle. Information Processing Letters, 88(4):155–159, November 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gamess:2000:PTE

[Gam00] Eric Gamess. plapackJava: Towards an efficient Java interface for high performance parallel linear algebra. Information Processing Letters, 75(5): 191–197, October 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/30/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/30/26/article.
pdf.

Gavril:2000:MWI

[Gav00] Fanica Gavril. Maximum weight independent sets and cliques in intersection graphs of filaments. Information Processing Letters, 73(5–6):181– 188, March 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/62/28/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/62/28/31/article.
pdf.

Gavril:2002:AMW

[Gav02] Fanica Gavril. Algorithms for maximum weight induced paths. Information Processing Letters, 81(4):203–208, February 28, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/23/
                                        20/85/35/32/abstract.html.

Gavril:2008:MWF

[Gav08] Fanica Gavril. Minimum weight feedback vertex sets in circle graphs. Information Processing Letters, 107(1):1–6, June 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guan:2000:CCS

[GCC00] D. J. Guan, Chun-Yen Chou, and Chiou-Wei Chen. Computational complexity of similarity retrieval in a pictorial database. Information Processing Letters, 75(3):113–117, August 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/64/28/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/64/28/28/article.
                                        pdf.

Gunther:2000:CPL

[GD00] Wolfgang Günther and Rolf Drechsler. On the computational power of linearly transformed BDDs. Information Processing Letters, 75(3):119– 125, August 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/28/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/28/29/article.
pdf.

Guting:2004:SEI

[GD04] Ralf Hartmut Güting and Zhiming Ding. A simple but effective improvement to the plumbline algorithm. Information Processing Letters, 91(6):251–257, September 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gennaro:2005:SMS

[GD05] Rosario Gennaro and Mario Di Raimondo. Secure multiplication of shared secrets in the exponent. Information Processing Letters, 96(2):71–79, October 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gregor:2008:PPH

[GD08] Petr Gregor and Tomáš Dvořák.

Path partitions of hypercubes. Information Processing Letters, 108(6):402–406, November 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goswami:2007:CDN

[GDN07] Partha P. Goswami, Sandip Das, and Subhas C. Nandy. Chromatic distribution of k-nearest neighbors of a line segment in a planar colored point set. Information Processing Letters, 102 (4):163–168, May 16, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Garcia:2008:ECQ

[GdPL08] Pedro García, Manuel Vázquez de Parga, and Damián López. On the efficient construction of quasi-reversible automata for reversible languages. Information Processing Letters, 107(1): 13–17, June 30, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Georgakopoulos:2008:CST

[Geo08] George F. Georgakopoulos. Chain-splay trees, or, how to achieve and prove log log Ncompetitiveness by splaying. Information Processing Letters, 106(1):37–43, March 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grabowski:2008:BPS

[GF08] Szymon Grabowski and Kimmo Fredriksson. Bit-parallel string matching under Hamming distance in O(n⌈m/w⌉) worst case time. Information Processing Letters, 105(5):182–187, February 29, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Greenberg:2002:AHL

[GG02] Ronald I. Greenberg and Lee Guan. On the area of hypercube layouts. Information Processing Letters, 84(1):41–46, October 16, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gaur:2003:AAS

[GGK03] Daya Ram Gaur, Arvind Gupta, and Ramesh Krishnamurti. A 5/3-approximation algorithm for scheduling vehicles on a path with release and handling times. Information Processing Letters, 86(2):87–91, April 30, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ghosh:2000:FCS

[GH00a] Sukumar Ghosh and Xin He. Fault-containing selfstabilization using priority scheduling. Information Processing Letters, 73(3–4):145– 151, February 29, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/62/27/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/62/27/35/article.
pdf.

Granvilliers:2000:CSP

[GH00b] Laurent Granvilliers and Gaétan Hains. A conservative scheme for parallel interval narrowing. Information Processing Letters, 74(3–4):141–146, May 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/63/27/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/63/27/32/article.
pdf.

Guo:2005:ALT

[GH05] J. Y. Guo and F. K. Hwang. An almost-linear time and linear space algorithm for the longest common subsequence problem. Information Processing Letters, 94(3):131–135, May 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Galindo:2008:SPK

[GH08] David Galindo and Javier Herranz. On the security of public key cryptosystems with a double decryption mechanism. Information Processing Letters, 108 (5):279–283, November 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grauman:2008:HNG

[GHJ^{+}08] Tracy Grauman, Stephen G. Hartke, Adam Jobson, Bill Kinnersley, Douglas B. West, Lesley Wiglesworth, Pratik Worah, and Hehui Wu. The hub number of a graph. Information Processing Letters, 108(4):226–228, October 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gudmundsson:2002:LBA

[GHL02] Joachim Gudmundsson, Thore Husfeldt, and Christos Levcopoulos. Lower bounds for approximate polygon decomposition and minimum gap. Information Processing Letters, 81 (3):137–141, February 14, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/85/34/30/abstract.html.

Guo:2007:FAS

[GHM07] Jiong Guo, Falk Hüffner, and Hannes Moser. Feedback arc set in bipartite tournaments is NP-complete. Information Processing Letters, 102(2–3):62–65, April 30, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gropl:2002:STU

[GHNP02] Clemens Gröpl, Stefan Hougardy, Till Nierhoff, and Hans Jürgen Prömel. Steiner trees in uniformly quasi-bipartite graphs. Information Processing Letters, 83(4):195–200, August 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ghodosi:2007:INP

[Gho07] Hossein Ghodosi. On insecurity of Naor–Pinkas’ distributed oblivious transfer. Information Processing Letters, 104(5):179– 182, November 30, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goh:2002:P

[GHS02] William M. Y. Goh, Paweł Hitczenko, and Ali Shokoufandeh. s-partitions. Information Processing Letters, 82(6): 327–329, June 30, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gu:2002:AMC

[GHZ02] Jun Gu, Xiao-Dong Hu, and Mu-Hong Zhang. Algorithms for multicast connection under multi-path routing model. Information Processing Letters, 84 (1):31–39, October 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Giannella:2009:NIR

[Gia09] Chris Giannella. New instability results for high-dimensional nearest neighbor search. Information Processing Letters, 109(19):1109–1113, September 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gazdag:2009:IUB

[GING09] Zsolt Gazdag, Szabolcs Iván, and Judit Nagy-György. Improved upper bounds on synchronizing nondeterministic automata. Information Processing Letters, 109(17):986–990, August 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gaysinsky:2004:SCA

[GIS04] Alexander Gaysinsky, Alon Itai, and Hadas Shachnai. Strongly competitive algorithms for caching with pipelined prefetching. Information Processing Letters, 91(1):19–27, July 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Giurcaneanu:2004:SPN

[Giu04] Ciprian Doru Giurcăneanu. On some properties of the NML estimator for Bernoulli strings. Information Processing Letters, 89 (4):209–213, February 28, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Giaro:2000:ECS

[GK00] Krzysztof Giaro and Marek Kubale. Edge-chromatic sum of trees and bounded cyclicity graphs. Information Processing Letters, 75(1–2):65– 69, July 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/25/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/25/36/article.
pdf.

Geerts:2004:TFT

[GK04] Floris Geerts and Bart Kuijpers.

Topological formulation of termination properties of iterates of functions. Information Processing Letters, 89(1):31–35, January 16, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gasieniec:2009:FMW

[GKL09] Leszek Gasieniec, Mirosław Kowaluk, and Andrzej Lingas. Faster multi-witnesses for Boolean matrix multiplication. Information Processing Letters, 109(4):242–247, January 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gasieniec:2002:ADG

[GL02] Leszek Gasieniec and Andrzej Lingas. On adaptive deterministic gossiping in ad hoc radio networks. Information Processing Letters, 83(2):89–93, July 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grieser:2004:ILA

[GL04a] Gunter Grieser and Steffen Lange. Incremental learning of approximations from positive data. Information Processing Letters, 89(1):37–42, January 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guillaume:2004:BSA

[GL04b] Jean-Loup Guillaume and Matthieu Latapy. Bipartite structure of all complex networks. Information Processing Letters, 90(5): 215–221, June 15, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goubault-Larrecq:2005:DR

[GL05] Jean Goubault-Larrecq. Deciding H_{∞} by resolution. Information Processing Letters, 95(3): 401–408, August 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gottlob:2007:LAM

[GL07] Georg Gottlob and Stephanie Tien Lee. A logical approach to multicut problems. Information Processing Letters, 103(4):136–141, August 16, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gotthilf:2009:IAS

[GL09a] Zvi Gotthilf and Moshe Lewenstein. Improved algorithms for the k simple shortest paths and the replacement paths problems. Information Processing Letters, 109(7):352–355, March 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gu:2009:PSO

[GL09b] Manzhan Gu and Xiwen Lu.

Preemptive stochastic online scheduling on two uniform machines. Information Processing Letters, 109(7):369–375, March 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Glen:2008:NMC

[GLS08] Amy Glen, Aaron Lauve, and Franco V. Saliola. A note on the Markoff condition and central words. Information Processing Letters, 105(6):241–244, March 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goswami:2002:EMC

[GM02] D. Goswami and R. Mall. An efficient method for computing dynamic program slices. Information Processing Letters, 81 (2):111–117, January 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/85/33/36/abstract.html.

Gronau:2007:OIU

[GM07] Ilan Gronau and Shlomo Moran. Optimal implementations of UPGMA and other common clustering algorithms. Information Processing Letters, 104(6): 205–210, December 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ganguly:2008:DKS

[GM08a] Sumit Ganguly and Anirban Majumder. Deterministic K-set structure. Information Processing Letters, 109(1):27–31, December 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldwasser:2008:SCA

[GM08b] Michael H. Goldwasser and Arundhati Bagchi Misra. A simpler competitive analysis for scheduling equal-length jobs on one machine with restarts. Information Processing Letters, 107(6):240–245, August 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Golic:2009:OCA

[GM09] Jovan Dj. Golić and Guglielmo Morgari. Optimal correlation attack on the multiplexer generator. Information Processing Letters, 109(15):838–841, July 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Galbraith:2002:PKS

[GMLS02] S. Galbraith, J. Malone-Lee, and N. P. Smart. Public key signatures in the multi-user setting. Information Processing Letters, 83(5):263–266, September 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ghodsi:2007:STM

[GMM^{+}07] Mohammad Ghodsi, Hamid Mahini, Kian Mirjalali, Shayan Oveis Gharan, Amin S. Sayedi R., and Morteza Zadimoghaddam. Spanning trees with minimum weighted degrees. Information Processing Letters, 104(3):113– 116, October 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gaspers:2009:CGB

[GMNP09] Serge Gaspers, Margaret-Ellen Messinger, Richard J. Nowakowski, and Paweł Prałat. Clean the graph before you draw it! Information Processing Letters, 109 (10):463–467, April 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gambosi:2000:LSS

[GN00] Giorgio Gambosi and Gaia Nicosia. On-line scheduling with setup costs. Information Processing Letters, 73(1–2):61– 68, January 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/62/25/
                                        35/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/62/25/35/article.
                                        pdf.

Guo:2006:FPT

[GN06] Jiong Guo and Rolf Niedermeier. A fixed-parameter tractability result for multicommodity demand flow in trees. Information Processing Letters, 97 (3):109–114, February 14, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Garcia-Nieto:2009:SSB

[GNAJT09] J. García-Nieto, E. Alba, L. Jourdan, and E. Talbi. Sensitivity and specificity based multiobjective approach for feature selection: Application to cancer diagnosis. Information Processing Letters, 109(16):887– 896, July 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gauwin:2008:STA

[GNR08] Olivier Gauwin, Joachim Niehren, and Yves Roos. Streaming tree automata. Information Processing Letters, 109(1):13–17, December 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guo:2008:TFP

[GNU08] Jiong Guo, Rolf Niedermeier, and Johannes Uhlmann. Two fixed-parameter algorithms for Vertex Covering by Paths on Trees. Information Processing Letters, 106(2):81–86, April 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Griffiths:2005:OBD

[GO05] Evan J. Griffiths and Pekka Orponen. Optimization, block designs and No Free Lunch theorems. Information Processing Letters, 94(2):55–61, April 30, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gobel:2000:RPI

[Göb00] Manfred Göbel. Rings of polynomial invariants of the alternating group have no finite SAGBI bases with respect to any admissible order. Information Processing Letters, 74 (1–2):15–18, April 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/25/27/article.
pdf.

Godard:2002:SSE

[God02] Emmanuel Godard. A selfstabilizing enumeration algorithm. Information Processing Letters, 82(6):299–305, June 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldberg:2000:GLC

[Gol00] Mayer Goldberg. Gödelization in the lambda calculus. Information Processing Letters, 75(1– 2):13–16, July 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/25/27/article.
pdf.

Golic:2001:PCM

[Gol01] Jovan Dj. Golić. A probabilistic cryptanalytic method for a timevariant permutation generator. Information Processing Letters, 80(2):67–73, October 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.

elsevier.com/gej-ng/10/23/
20/80/37/27/abstract.html.

Goldstein:2003:DTD

[Gol03] Darin Goldstein. Determination of the topology of a directed network. Information Processing Letters, 88(3):121–131, November 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldberg:2004:COP

[Gol04] Mayer Goldberg. A construction of one-point bases in extended lambda calculi. Information Processing Letters, 89(6): 281–286, March 31, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goller:2008:RPR

[Göl08] Stefan Göller. Reachability on prefix-recognizable graphs. Information Processing Letters, 108(2):71–74, September 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goodrich:2009:ACM

[Goo09] Michael T. Goodrich. On the algorithmic complexity of the Mastermind game with blackpeg results. Information Processing Letters, 109(13):675– 678, June 15, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gortz:2008:ACM

[Gør08] Inge Li Gørtz. Asymmetric k-center with minimum coverage. Information Processing Letters, 105(4):144–149, February 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gouda:2001:ESC

[Gou01]  Mohamed G. Gouda.  El-

ements of security: Closure, convergence, and protection. Information Processing Letters, 77(2–4):109–114, February 28, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/29/32/article.
pdf.

Green:2001:RSE

[GP01] Frederic Green and Randall Pruim. Relativized separation of EQP from PNP. Information Processing Letters, 80(5): 257–260, December 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/84/36/33/abstract.html.

Gal:2003:ENM

[GP03a] Anna Gál and Pavel Pudlák. Erratum to: “A note on monotone complexity and the rank of matrices”: [Information Processing Letters 87 (2003) 321–326]. Information Processing Letters, 88(5):257, December 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [GP03b].

Gal:2003:NMC

[GP03b] Anna Gál and Pavel Pudlák.

A note on monotone complexity and the rank of matrices. Information Processing Letters, 87(6):321–326, September 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [GP03a].

Gropl:2001:EWC

[GPS01] Clemens Gröpl, Hans Jürgen Prömel, and Anand Srivastav. On the evolution of the worstcase OBDD size. Information Processing Letters, 77(1): 1–7, January 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/24/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/24/24/article.
pdf.

Giannella:2003:NAM

[GR03] Chris Giannella and Edward Robertson. A note on approximation measures for multivalued dependencies in relational databases. Information Processing Letters, 85(3):153– 158, February 14, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gaber:2009:ASC

[GRH09] Mohamed Medhat Gaber, Uwe Roehm, and Karel Herink. An analytical study of central and in-network data processing for wireless sensor networks. Information Processing Letters, 110 (2):62–70, December 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grigni:2001:SLC

[Gri01]  Michelangelo Grigni.  A

Sperner lemma complete for PPA. Information Processing Letters, 77(5–6):255–259, March 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/68/32/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/68/32/29/article.
                                        pdf.

Gronemeier:2006:NDC

[Gro06] André Gronemeier. A note on the decoding complexity of error-correcting codes. Information Processing Letters, 100 (3):116–119, November 15, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldmann:2000:ETR

[GRT00] Mikael Goldmann, Alexander Russell, and Denis Thérien. An ergodic theorem for readonce non-uniform deterministic finite automata. Information Processing Letters, 73(1–2):23– 28, January 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/62/25/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/62/25/29/article.
pdf.

Grosse:2006:CSF

[GRW06] André Große, Jörg Rothe, and Gerd Wechsung. On computing the smallest four-coloring of planar graphs and non-self-reducible sets in P. Information Processing Letters, 99(6): 215–221, September 30, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldwurm:2000:CPU

[GS00] Massimiliano Goldwurm and Massimo Santini. Clique polynomials have a unique root of smallest modulus. Information Processing Letters, 75(3):127– 132, August 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/28/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/28/30/article.
pdf.

Gerbessiotis:2003:RSC

[GS03] Alexandros V. Gerbessiotis and Constantinos J. Siniolakis. Randomized selection in n+C +o(n) comparisons. Information Processing Letters, 88(3):95–100, November 15, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Geiselmann:2004:PAS

[GS04a] Willi Geiselmann and Rainer Steinwandt. Power attacks on a side-channel resistant elliptic curve implementation. Information Processing Letters, 91 (1):29–32, July 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gerbessiotis:2004:PIS

[GS04b] Alexandros V. Gerbessiotis and Constantinos J. Siniolakis. Probabilistic integer sorting. Information Processing Letters, 90 (4):187–193, May 31, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Garcia:2005:CGM

[GS05a] Thierry Garcia and David Semé. A Coarse-Grained Multicomputer algorithm for the detection of repetitions. Information Processing Letters, 93(6): 307–313, March 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gupta:2005:CHD

[GS05b] Kishan Chand Gupta and Palash Sarkar. Construction of high degree resilient S-boxes with improved nonlinearity. Information Processing Letters, 95 (3):413–417, August 16, 2005.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gonen:2009:ASC

[GS09a] Mira Gonen and Yuval Shavitt.

A Θ(log n)-approximation for the set cover problem with set ownership. Information Processing Letters, 109(3):183–186, January 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grassl:2009:CAS

[GS09b] Markus Grassl and Rainer Steinwandt. Cryptanalysis of an authentication scheme using truncated polynomials. Information Processing Letters, 109(15):861– 863, July 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grahne:2009:BRP

[GT09] Gösta Grahne and Alex Thomo.

Bounded regular path queries in view-based data integration. Information Processing Letters, 109(13):739–744, June 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gallier:2003:DFA

[GTM03] Jean H. Gallier, Salvatore La Torre, and Supratik Mukhopadhyay. Deterministic finite automata with recursive calls and DPDAs. Information Processing Letters, 87(4):187–193, August 31, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guerraoui:2001:HFS

[Gue01] Rachid Guerraoui. On the hardness of failure-sensitive agreement problems. Information Processing Letters, 79(2):99– 104, June 30, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/31/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/80/31/32/article.
pdf.

Gupta:2008:FAS

[Gup08] Sushmita Gupta. Feedback arc set problem in bipartite tournaments. Information Processing Letters, 105(4):150–154, February 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gurski:2008:PAP

[Gur08] Frank Gurski. Polynomial algorithms for protein similarity search for restricted mRNA structures. Information Processing Letters, 105(5):170–176, February 29, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gusfield:2002:PDP

[Gus02]  Dan Gusfield.  Partition-

distance: A problem and class of perfect graphs arising in clustering. Information Processing Letters, 82(3):159–164, May 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/86/34/34/abstract.html.

Gutjahr:2002:AAG

[Gut02] Walter J. Gutjahr. ACO algorithms with guaranteed convergence to the optimal solution. Information Processing Letters, 82(3):145–153, May 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/86/34/32/abstract.html.

Gutner:2008:EAA

[Gut08] Shai Gutner. Elementary approximation algorithms for prize collecting Steiner tree problems. Information Processing Letters, 107(1):39–44, June 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Garcia-Villalba:2000:LCS

[GVFS00] L. J. García-Villalba and A. Fúster-Sabater. On the linear complexity of the sequences generated by nonlinear filterings. Information Processing Letters, 76(1–2):67–73, November 20, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/24/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/24/34/article.
pdf.

Gradwohl:2008:WIL

[GY08] Ronen Gradwohl and Amir Yehudayoff. t-Wise independence with local dependencies. Information Processing Letters, 106(5):208–212, May 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guan:2001:TEA

[GYY01] Xudong Guan, Yiling Yang, and Jinyuan You. Typing evolving ambients. Information Processing Letters, 80(5):265– 270, December 15, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/84/36/
                                        35/abstract.html.

Ghodosi:2006:COT

[GZN06] Hossein Ghodosi and Rahim Zaare-Nahandi. Comments on the ‘m out of n oblivious transfer’. Information Processing Letters, 97(4):153–155, February 28, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hage:2003:ESM

[Hag03] Jurriaan Hage. Enumerating submultisets of multisets. Information Processing Letters, 85 (4):221–226, February 28, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hagen:2007:FPT

[Hag07] Matthias Hagen. On the fixedparameter tractability of the equivalence test of monotone normal forms. Information Processing Letters, 103(4):163–167, August 16, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Halman:2003:LTA

[Hal03] Nir Halman. A linear time algorithm for the weighted lexicographic rectilinear 1-center problem in the plane. Information Processing Letters, 86(3): 121–128, May 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hammack:2007:MCB

[Ham07] Richard Hammack. Minimum cycle bases of direct products of complete graphs. Information Processing Letters, 102(5):214– 218, May 31, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Han:2006:IAS

[Han06] Yijie Han. Improved algorithm for the symmetry number problem on trees. Information Processing Letters, 98(4):130–132, May 31, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Han:2008:NTA

[Han08] Yijie Han. A note of an O(n^{3}/ log n) time algorithm for all pairs shortest paths. Information Processing Letters, 105 (3):114–116, January 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Haastad:2000:BOC

[Hås00a] Johan Håstad. On bounded occurrence constraint satisfaction. Information Processing Letters, 74(1–2):1–6, April 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/25/25/article.
pdf.

Hasunuma:2000:EDS

[Has00b] Toru Hasunuma. On edgedisjoint spanning trees with small depths. Information Processing Letters, 75(1–2):71– 74, July 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/25/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/25/37/article.
pdf.

Huang:2001:OFA

[HC01] Shing-Tsaan Huang and Bau-Wen Chen. Optimal 1-fair alternators. Information Processing Letters, 80(3):159–163, November 15, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/84/35/
33/abstract.html.

Holt:2002:MAR

[HC02] John D. Holt and Soon M.

Chung. Mining association rules using inverted hashing and pruning. Information Processing Letters, 83(4):211–220, August 31, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ho:2003:SSS

[HC03a] Ting-Yem Ho and Jou-Ming Chang. Sorting a sequence of strong kings in a tournament. Information Processing Letters, 87(6):317–320, September 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:2003:SBP

[HC03b] Tsung-Chuan Huang and Slo-Li Chu. A statement based parallelizing framework for processorin-memory architectures. Information Processing Letters, 85 (3):159–163, February 14, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

He:2008:PBH

[HC08a] Weiping He and Ing-Ray Chen.

Proxy-based hybrid cache management in Mobile IP systems. Information Processing Letters, 106(1):26–32, March 31, 2008.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hsieh:2008:FMD

[HC08b] Sun-Yuan Hsieh and Chih-Sheng Cheng. Finding a maximum-density path in a tree under the weight and length constraints. Information Processing Letters, 105(5):202–205, February 29, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

He:2002:WSM

[He02] Wei-Hua He. Weaknesses in some multisignature schemes for specified group of verifiers. Information Processing Letters, 83 (2):95–99, July 31, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Heam:2008:NPO

[Héa08] P.-C. Héam. A note on partially ordered tree automata. Information Processing Letters, 108(4): 242–246, October 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hehner:2001:VSC

[Heh01]  Eric C. R. Hehner.  Vari-

ables and scopes considered formally. Information Processing Letters, 79(1):33–38, May 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.elsevier.nl/gej%2Dng/
10/23/20/80/25/30/abstract.
html;   http://www.elsevier.

nl/gej%2Dng/10/23/20/80/25/
30/article.pdf.

Hershberger:2004:KCD

[Her04] John Hershberger. Kinetic collision detection with fast flight plan changes. Information Processing Letters, 92(6):287–291, December 31, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hesselink:2003:SMT

[Hes03] Wim H. Hesselink. Salembier’s Min-tree algorithm turned into breadth first search. Information Processing Letters, 88(5): 225–229, December 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hess:2004:SVE

[Hes04] F. Hess. On the security of the verifiably-encrypted signature scheme of Boneh, Gentry, Lynn and Shacham. Information Processing Letters, 89(3):111– 114, February 14, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hesselink:2007:LTA

[Hes07] Wim H. Hesselink. A lineartime algorithm for Euclidean feature transform sets. Information Processing Letters, 102(5): 181–186, May 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Heun:2008:AMG

[Heu08] Volker Heun. Analysis of a modification of Gusfield’s recursive algorithm for reconstructing ultrametric trees. Information Processing Letters, 108(4): 222–225, October 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hutton:2001:GAL

[HG01] Graham Hutton and Jeremy Gibbons. The generic approximation lemma. Information Processing Letters, 79(4):197– 201, August 15, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/33/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/80/33/31/article.
pdf.

Hajiaghayi:2002:NCO

[HG02] Mohammad Taghi Hajiaghayi and Yashar Ganjali. A note on the Consecutive Ones Submatrix problem. Information Processing Letters, 83(3):163–166, August 16, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hack:2006:ORA

[HG06] Sebastian Hack and Gerhard Goos. Optimal register allocation for SSA-form programs in polynomial time. Information Processing Letters, 98(4): 150–155, May 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Han:2009:NLB

[HGYZ09] Xin Han, He Guo, Dawei Yin, and Yong Zhang. A note on online broadcast scheduling with deadlines. Information Processing Letters, 109(3):204–207, January 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hasunuma:2007:IUB

[HH07] Toru Hasunuma and Misa Hirota. An improved upper bound on the queuenumber of the hypercube. Information Processing Letters, 104(2):41–44, October 16, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See note [PCW08].

Hua:2009:GTL

[HH09] Hongbo Hua and Yaoping Hou.

On graphs with the third largest number of maximal independent sets. Information Processing Letters, 109(4):248–253, January 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Halava:2008:PCP

[HHHK08] Vesa Halava, Tero Harju, Mika Hirvensalo, and Juhani Karhumäki. Post Correspondence Problem for short words. Information Processing Letters, 108(3):115–118, October 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Halava:2008:SFP

[HHK08] Vesa Halava, Tero Harju, and Tomi Kärki. Square-free partial words. Information Processing Letters, 108(5):290–292, November 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Halava:2005:ESP

[HHL05] Vesa Halava, Tero Harju, and Michel Latteux. Equality sets of prefix morphisms and regular star languages. Information Processing Letters, 94(4): 151–154, May 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hung:2003:REF

[HHLH03] Chun-Nan Hung, Hong-Chun Hsu, Kao-Yung Liang, and Lih-Hsing Hsu. Ring embedding in faulty pancake graphs. Information Processing Letters, 86(5): 271–275, June 15, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hartmann:2003:AFD

[HHLS03] Sven Hartmann, Anne Hoffmann, Sebastian Link, and Klaus-Dieter Schewe. Axiomatizing functional dependencies in the Higher-Order Entity-Relationship Model. Information Processing Letters, 87(3): 133–137, August 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huemer:2008:RGA

[HHP08] Clemens Huemer, Ferran Hurtado, and Julian Pfeifle. The rotation graph of k-ary trees is Hamiltonian. Information Processing Letters, 109(2):124–129, December 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:2005:SSC

[HHT05] Shing-Tsaan Huang, Su-Shen Hung, and Chi-Hung Tzeng. Self-stabilizing coloration in anonymous planar networks. Information Processing Letters, 95 (1):307–312, July 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Horiyama:2003:TAC

[HI03] Takashi Horiyama and Toshihide Ibaraki. Translation among CNFs, characteristic models and ordered binary decision diagrams. Information Processing Letters, 85(4):191–198, February 28, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hierons:2004:UMN

[Hie04] R. M. Hierons. Using a minimal number of resets when testing from a finite state machine. Information Processing Letters, 90 (6):287–292, June 30, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hierons:2006:AAT

[Hie06] R. M. Hierons. Applying adaptive test cases to nondeterministic implementations. Information Processing Letters, 98(2): 56–60, April 30, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hamada:2009:IAL

[HIM09] Koki Hamada, Kazuo Iwama, and Shuichi Miyazaki. An improved approximation lower bound for finding almost stable maximum matchings. Information Processing Letters, 109(18): 1036–1040, August 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hitchcock:2003:GSC

[Hit03] John M. Hitchcock. Gales suffice for constructive dimension. Information Processing Letters, 86(1):9–12, April 15, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Herman:2005:SPP

[HJ05] Ted Herman and Colette Johnen. Strategies for peer-to-peer downloading. Information Processing Letters, 94(5):203– 209, June 15, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hoffman:2002:GDG

[HJR02] Alan J. Hoffman, Kate Jenkins, and Tim Roughgarden. On a game in directed graphs. Information Processing Letters, 83 (1):13–16, July 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hedetniemi:2001:MMS

[HJS01] Stephen T. Hedetniemi, David P.

Jacobs, and Pradip K. Srimani. Maximal matching stabilizes in time O(m). Information Processing Letters, 80(5): 221–223, December 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/84/36/27/abstract.html.

Hedetniemi:2003:LTS

[HJS03] Stephen T. Hedetniemi, David P.

Jacobs, and Pradip K. Srimani. Linear time self-stabilizing colorings. Information Processing Letters, 87(5):251–255, September 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Howe:2000:ANC

[HK00] Jacob M. Howe and Andy King.

Abstracting numeric constraints with Boolean functions. Information Processing Letters, 75(1– 2):17–23, July 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/25/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/25/28/article.
pdf.

Hoogeboom:2004:TD

[HK04] Hendrik Jan Hoogeboom and Walter A. Kosters. Tetris and decidability. Information Processing Letters, 89(6):267–272, March 31, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Haddad:2009:SSC

[HK09] Mohammed Haddad and Hamamache Kheddouci. A strict strong coloring of trees. Information Processing Letters, 109(18):1047–1054, August 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hsieh:2009:FRF

[HKC09] Sun-Yuan Hsieh, Che-Nan Kuo, and Hsin-Hung Chou. A further result on fault-free cycles in faulty folded hypercubes. Information Processing Letters, 110 (2):41–43, December 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Horev:2009:PCG

[HKKL09] Elad Horev, Matthew J. Katz, Roi Krakovski, and Maarten Löffler. Polychromatic 4coloring of guillotine subdivisions. Information Processing Letters, 109(13):690–694, June 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Harary:2001:NGC

[HKL01] Frank Harary, Vladik Kreinovich, and Luc Longpré. A new graph characteristic and its application to numerical computability. Information Processing Letters, 77(5–6):277– 282, March 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/32/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/32/32/article.
pdf.

Hon:2004:NSE

[HKL^{+}04] Wing-Kai Hon, Ming-Yang Kao, Tak-Wah Lam, Wing-Kin Sung, and Siu-Ming Yiu. Non-shared edges and nearest neighbor interchanges revisited. Information Processing Letters, 91(3): 129–134, August 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hambrusch:2003:DRS

[HL03] Susanne E. Hambrusch and Chuan-Ming Liu. Data replication in static tree structures. Information Processing Letters, 86 (4):197–202, May 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Honkala:2004:ICH

[HL04] Iiro Honkala and Tero Laihonen. On identifying codes in the hexagonal mesh. Information Processing Letters, 89(1): 9–14, January 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Han:2005:AAI

[HL05] Daewan Han and Moonsik Lee.

An algebraic attack on the improved summation generator with 2-bit memory. Information Processing Letters, 93(1):43–46, January 16, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Honkala:2007:NCI

[HL07] Iiro Honkala and Tero Laihonen. On a new class of identifying codes in graphs. Information Processing Letters, 102(2–3):92– 98, April 30, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Haddadi:2008:CBM

[HL08] Salim Haddadi and Zoubir Layouni. Consecutive block minimization is 1.5-approximable. Information Processing Letters, 108(3):132–135, October 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Habib:2003:NFA

[HLP03] Michel Habib, Emmanuelle Lebhar, and Christophe Paul. A note on finding all homogeneous set sandwiches. Information Processing Letters, 87(3):147– 151, August 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hertz:2003:FAC

[HLS03] A. Hertz, V. Lozin, and D. Schindl. Finding augmenting chains in extensions of clawfree graphs. Information Processing Letters, 86(6):311–316, June 30, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Holmes:2002:LCS

[HLSW02] John H. Holmes, Pier Luca Lanzi, Wolfgang Stolzmann, and Stewart W. Wilson. Learning classifier systems: New models, successful applications. Information Processing Letters, 82 (1):23–30, April 15, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/86/27/
30/abstract.html.

Hsu:2007:GPB

[HLT07] Hong-Chun Hsu, Pao-Lien Lai, and Chang-Hsiung Tsai. Geodesic pancyclicity and balanced pancyclicity of Augmented cubes. Information Processing Letters, 101(6):227–232, March 31, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

He:2009:DAM

[HLY09] Cheng He, Yixun Lin, and Jinjiang Yuan. A DP algorithm for minimizing makespan and total completion time on a seriesbatching machine. Information Processing Letters, 109(12):603– 607, May 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Herman:2001:ASH

[HM01] Ted Herman and Toshimitsu Masuzawa. Available stabilizing heaps. Information Processing Letters, 77(2–4):115– 121, February 28, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/29/33/article.
pdf.

Hart:2002:SIE

[HM02] Andrew G. Hart and Servet Martínez. Sequential iteration of the Erlang fixed-point equations. Information Processing Letters, 81(6):319–325, March 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/37/
32/abstract.html.

Horowitz:2003:ENS

[HM03] Keren Horowitz and Dahlia Malkhi. Estimating network size from local information. Information Processing Letters, 88 (5):237–243, December 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Harvey:2004:DS

[HM04] Nicholas J. A. Harvey and J. Ian Munro. Deterministic SkipNet. Information Processing Letters, 90(4):205–208, May 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Held:2008:TIC

[HM08] Martin Held and Joseph S. B.

Mitchell. Triangulating inputconstrained planar point sets. Information Processing Letters, 109(1):54–56, December 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hocquard:2009:EPG

[HM09] Hervé Hocquard and Mickaël Montassier. Every planar graph without cycles of lengths 4 to 12 is acyclically 3-choosable. Information Processing Letters, 109(21–22):1193–1196, October 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hurink:2008:AMI

[HN08] Johann L. Hurink and Tim Nieberg. Approximating minimum independent dominating sets in wireless networks. Information Processing Letters, 109 (2):155–160, December 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hoare:2001:L

[Hoa01] Tony Hoare. Legacy. Information Processing Letters, 77(2– 4):123–129, February 28, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/23/
20/68/29/34/abstract.html;
http://www.elsevier.nl/gej-

ng/10/23/20/68/29/34/article.
pdf.

Hoepman:2001:COB

[Hoe01]  Jaap-Henk Hoepman.  Can

an operation both update the state and return a meaningful value in the asynchronous PRAM model? Information Processing Letters, 79(4):161– 166, August 15, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/33/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/80/33/26/article.
pdf.

Honkala:2002:EPP

[Hon02] Juha Honkala. The equality problem for Parikh simple algebraic power series. Information Processing Letters, 84(2):57–60, October 31, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Honkala:2005:CHS

[Hon05] Juha Honkala. The class of HDT0L sequences is closed with respect to rational functions. Information Processing Letters, 94 (4):155–158, May 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hong:2007:ANI

[Hon07] Dawei Hong. Analysis of noiseinduced phase synchronization in nervous systems: from algorithmic perspective. Information Processing Letters, 105(1): 35–39, December 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hoogerwoord:2001:FW

[Hoo01] Rob R. Hoogerwoord. Formality works. Information Processing Letters, 77(2–4):137– 142, February 28, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/29/36/article.
pdf.

Horn:2007:DS

[Hor07] Florian Horn. Dicing on the Streett. Information Processing Letters, 104(1):1–9, September 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hoyer:2000:SPF

[Høy00] Peter Høyer. Simplified proof of the Fourier Sampling Theorem. Information Processing Letters, 75(4):139–143, September 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/29/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/29/24/article.
pdf.

Herman:2000:EDC

[HP00a] Ted Herman and Sriram Pemmaraju. Error-detecting codes and fault-containing self-stabilization. Information Processing Letters, 73(1–2):41–46, January 31, 2000. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/62/25/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/62/25/32/article.
pdf.

Hwang:2000:SCC

[HP00b] F. K. Hwang and C. K. Pai. Sequential construction of a circular consecutive-2 system. Information Processing Letters, 75(5):231–235, October 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/30/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/30/31/article.
pdf.

Hitchcock:2005:RBS

[HP05] John M. Hitchcock and A. Pavan. Resource-bounded strong dimension versus resourcebounded category. Information Processing Letters, 95(3):377– 381, August 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hurkens:2008:MDG

[HPW08] Cor A. J. Hurkens, Rudi A. Pendavingh, and Gerhard J. Woeginger. The Magnus–Derek game revisited. Information Processing Letters, 109(1):38–40, December 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hassin:2000:BAM

[HR00] Refael Hassin and Shlomi Rubinstein. Better approximations for max TSP. Information Processing Letters, 75(4):181– 186, September 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/29/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/29/31/article.
pdf.

Hassin:2001:AAM

[HR01] Refael Hassin and Shlomi Rubinstein. Approximation algorithms for maximum linear arrangement. Information Processing Letters, 80(4):171–177, November 30, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/84/34/
27/abstract.html.

Hassin:2002:AAM

[HR02] Refael Hassin and Shlomi Rubinstein. A \frac{7}{8}-approximation algorithm for metric Max TSP. Information Processing Letters, 81(5):247–251, March 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/gej-ng/10/23/
                                        20/85/36/29/abstract.html.

Hernandez:2003:DPD

[HR03] Manuel Hernández and David A. Rosenblueth. Disjunctive partial deduction of a right-to-left string-matching algorithm. Information Processing Letters, 87(5):235–241, September 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Her:2005:EMD

[HR05] Jun-Ho Her and R. S. Ramakrishna. External-memory depthfirst search algorithm for solid grid graphs. Information Processing Letters, 93(4):177–183, February 28, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hajiaghayi:2006:IAA

[HR06a] Mohammad Taghi Hajiaghayi and Harald Räcke. An O(\sqrt{n})- approximation algorithm for directed sparsest cut. Information Processing Letters, 97(4):156– 160, February 28, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hassin:2006:IAA

[HR06b] Refael Hassin and Shlomi Rubinstein. An improved approximation algorithm for the metric maximum clustering problem with given cluster sizes. Information Processing Letters, 98 (3):92–95, May 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hamilton:2008:CHI

[HRC08] Chris H. Hamilton and Andrew Rau-Chaplin. Compact Hilbert indices: Space-filling curves for domains with unequal side lengths. Information Processing Letters, 105(5):155–163, February 29, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hertrampf:2000:NCP

[HRV00] Ulrich Hertrampf, Steffen Reith, and Heribert Vollmer. A note on closure properties of logspace MOD classes. Information Processing Letters, 75(3): 91–93, August 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/28/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/28/24/article.
pdf.

Hellwig:2004:NCL

[HRV04] Angelika Hellwig, Dieter Rautenbach, and Lutz Volkmann. Note on the connectivity of line graphs. Information Processing Letters, 91(1):7–10, July 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hvam:2007:BCS

[HRWZ07] K. Hvam, L. Reinhardt, P. Winter, and M. Zachariasen. Bounding component sizes of twoconnected Steiner networks. Information Processing Letters, 104(5):159–163, November 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hartuv:2000:CAB

[HS00] Erez Hartuv and Ron Shamir. A clustering algorithm based on graph connectivity. Information Processing Letters, 76(4–6):175– 181, December 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/67/28/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/67/28/28/article.
                                        pdf.

Hofstee:2001:DRC

[HS01] H. Peter Hofstee and Jun Sawada. Derivation of a rotator circuit with homogeneous interconnect. Information Processing Letters, 77(2–4):131– 135, February 28, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/68/29/
                                        35/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/68/29/35/article.
                                        pdf.

Hochberg:2003:OOP

[HS03] Robert A. Hochberg and Matthias F.

Stallmann. Optimal one-page tree embeddings in linear time. Information Processing Letters, 87(2):59–66, July 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Heber:2005:CIT

[HS05] Steffen Heber and Carla D. Savage. Common intervals of trees. Information Processing Letters, 93(2):69–74, January 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hsieh:2008:NCE

[Hsi08] Sun-Yuan Hsieh. A note on cycle embedding in folded hypercubes with faulty elements. Information Processing Letters, 108(2): 81, September 30, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hong:2002:PSR

[HSL^{+}02] Seokhie Hong, Jaechul Sung, Sangjin Lee, Jongin Lim, and Jongsu Kim. Provable security for 13 round Skipjack-like structure. Information Processing Letters, 82(5):243–246, June 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ho:2009:CFH

[HSTH09] Tung-Yang Ho, Yuan-Kang Shih, Jimmy J. M. Tan, and Lih-Hsing Hsu. Conditional fault Hamiltonian connectivity of the complete graph. Information Processing Letters, 109(12):585– 588, May 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:2009:TLT

[HSWC09] Jing Huang, Haina Sun, Weifan Wang, and Dong Chen. (2,1)- total labelling of trees with sparse vertices of maximum degree. Information Processing Letters, 109(3):199–203, January 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:2006:CCH

[HSZZ06] Wei Huang, Yaoyun Shi, Shengyu Zhang, and Yufan Zhu. The communication complexity of the Hamming distance problem. Information Processing Letters, 99(4):149–153, August 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Higuchi:2000:FAA

[HT00] Akira Higuchi and Naofumi Takagi. A fast addition algorithm for elliptic curve arithmetic in GF(2n) using projective coordinates. Information Processing Letters, 76(3):101– 103, December 15, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/27/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/

10/23/20/67/27/25/article.
pdf.

Hoffmann:2003:APT

[HT03] Michael Hoffmann and Csaba D.

Tóth. Alternating paths through disjoint line segments. Information Processing Letters, 87(6):287–294, September 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:2003:FAA

[HTT03] Po-Hsueh Huang, Yin Te Tsai, and Chuan Yi Tang. A fast algorithm for the alpha-connected two-center decision problem. Information Processing Letters, 85 (4):205–210, February 28, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hu:2007:LTA

[Hu07] Shiyan Hu. A linear time algorithm for max-min length triangulation of a convex polygon. Information Processing Letters, 101(5):203–208, March 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hellwig:2006:LBV

[HV06a] Angelika Hellwig and Lutz Volkmann. Lower bounds on the vertex-connectivity of digraphs and graphs. Information Processing Letters, 99(2):41–46, July 31, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hougardy:2006:AWM

[HV06b] Stefan Hougardy and Doratha E. Vinkemeier. Approximating weighted matchings in parallel. Information Processing Letters, 99(3):119–123, August 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hofbauer:2006:T

[HW06] Dieter Hofbauer and Johannes Waldmann. Termination of aa → bc, bb → ac, cc → ab. Information Processing Letters, 98 (4):156–158, May 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hwang:2000:CYK

[Hwa00] Min-Shiang Hwang. Cryptanalysis of YCN key assignment scheme in a hierarchy. Information Processing Letters, 73(3–4): 97–101, February 29, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/62/27/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/62/27/28/article.
                                        pdf.

Hsu:2002:IGT

[HWW02] Chien-Lung Hsu, Tzong-Sun Wu, and Tzong-Chen Wu. Improvements of generalization of threshold signature and authenticated encryption for group communications. Information Processing Letters, 81(1):41– 45, January 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/27/
33/abstract.html.

Yeh:2008:STB

[hY08] Jyh haw Yeh. A secure timebound hierarchical key assignment scheme based on RSA public key cryptosystem. Information Processing Letters, 105 (4):117–120, February 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yeh:2009:ENH

[hY09a]  Jyh haw Yeh.  Enforcing

non-hierarchical access policies by hierarchical key assignment schemes. Information Processing Letters, 110(2):46–49, December 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hrubes:2009:MSC

[HY09b] Pavel Hrubeš and Amir Yehudayoff. Monotone separations for constant degree polynomials. Information Processing Letters, 110(1):1–3, December 1, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hu:2007:NPE

[HYCH07] Ken S. Hu, Shyun-Shyun Yeoh, Chiuyuan Chen, and Lih-Hsing Hsu. Node-pancyclicity and edge-pancyclicity of hypercube variants. Information Processing Letters, 102(1):1–7, April 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yeh:2005:MYK

[hYScH05] Jyh haw Yeh, Marion Scheepers, and Wen chen Hu. Modifying YCN key assignment scheme to resist the attack from Hwang. Information Processing Letters, 95(4):435–440, August 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:2007:DPA

[HYT^{+}07] Kuo-Si Huang, Chang-Biau Yang, Kuo-Tsung Tseng, Yung-Hsing Peng, and Hsing-Yen Ann. Dynamic programming algorithms for the mosaic longest common subsequence problem. Information Processing Letters, 102(2–3):99–103, April 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:2008:EAF

[HYT^{+}08] Kuo-Si Huang, Chang-Biau Yang, Kuo-Tsung Tseng, Hsing-Yen Ann, and Yung-Hsing Peng. Efficient algorithms for finding interleaving relationship between sequences. Information Processing Letters, 105(5):188– 193, February 29, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hyyro:2008:IBP

[Hyy08] Heikki Hyyrö. Improving the bit-parallel NFA of Baeza-Yates and Navarro for approximate string matching. Information Processing Letters, 108(5):313– 319, November 15, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hebrard:2003:EAH

[HZ03] Jean-Jacques Hébrard and Bruno Zanuttini. An efficient algorithm for Horn description. Information Processing Letters, 88(4):177–182, November 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hong:2009:VFT

[HZ09] Yanmei Hong and Zhao Zhang.

Vertex fault tolerance of optimalκ graphs and super-κ graphs. Information Processing Letters, 109(20):1151–1155, September 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hericko:2008:AOS

[HŽR08] Marjan Heričko, Aleš Živkovič, and Ivan Rozman. An approach to optimizing software development team size. Information Processing Letters, 108(3):101– 106, October 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Islam:2003:NES

[IAIH03] Rafiqul Islam, Nasim Adnan, Nur Islam, and Shohorab Hossen. A new external sorting algorithm with no additional disk space. Information Processing Letters, 86(5):229–233, June 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ibarra:2009:SAF

[Iba09] Louis Ibarra. A simple algorithm to find Hamiltonian cycles in proper interval graphs. Information Processing Letters, 109(18):1105–1108, August 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Isaak:2009:GRA

[IJN09] Garth Isaak, Robert Jamison, and Darren Narayan. Greedy rankings and arank numbers. Information Processing Letters, 109(15):825–827, July 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Iannella:2005:FIR

[IK05] Nicolangelo Iannella and Lars Kindermann. Finding iterative roots with a spiking neural network. Information Processing Letters, 95(6):545–551, September 30, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Itai:2007:CDC

[IK07] Alon Itai and Irit Katriel.

Canonical density control. Information Processing Letters, 104 (6):200–204, December 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Im:2000:DSF

[IKKY00] Byoung-Mo Im, Myoung Ho Kim, Hyung-Il Kang, and Jae Soo Yoo. Declustering signature files based on a dynamic measure. Information Processing Letters, 74(5–6):235– 241, June 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/28/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/28/33/article.
pdf.

Izumi:2006:WAC

[IM06] Taisuke Izumi and Toshimitsu Masuzawa. A weakly-adaptive condition-based consensus algorithm in asynchronous distributed systems. Information Processing Letters, 100(5):199– 205, December 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ilie:2006:PSL

[IMP06] Lucian Ilie, Solomon Marcus, and Ion Petre. Periodic and Sturmian languages. Information Processing Letters, 98(6):

242–246, June 30, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Iwata:2000:FCS

[IMS00] Satoru Iwata, S. Thomas Mc-Cormick, and Maiko Shigeno. A fast cost scaling algorithm for submodular flow. Information Processing Letters, 74(3–4): 123–128, May 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/63/27/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/63/27/30/article.
                                        pdf.

Nakano:2002:EGP

[iN02] Shin ichi Nakano. Efficient generation of plane trees. Information Processing Letters, 84 (3):167–172, November 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Intrigila:2003:SPE

[IN03] B. Intrigila and M. Nesi. On structural properties of etaexpansions of identity. Information Processing Letters, 87(6): 327–333, September 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Isaak:2004:CTH

[IN04] Garth Isaak and Darren A. Narayan. A classification of tournaments having an acyclic tournament as a minimum feedback arc set. Information Processing Letters, 92(3):107–111, November 15, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ikeda:2006:SNP

[IN06] Satoshi Ikeda and Koji Nakazawa.

Strong normalization proofs by CPS-translations. Information Processing Letters, 99(4):163– 170, August 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nakayama:2007:PTA

[iNM07] Shin ichi Nakayama and Shigeru Masuyama. A polynomial time algorithm for obtaining minimum edge ranking on twoconnected outerplanar graphs. Information Processing Letters, 103(6):216–221, September 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Oum:2009:CRW

[iO09] Sang il Oum. Computing rankwidth exactly. Information Processing Letters, 109(13):745– 748, June 15, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Iyer:2009:LTA

[IP09] K. Viswanathan Iyer and S. Prasanna. A linear time algorithm for finding an optimal degree-bounded subtree of an edge-weighted tree. Information Processing Letters, 109(11):560– 562, May 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Iliopoulos:2008:FIP

[IR08a] Costas S. Iliopoulos and M. Sohel Rahman. Faster index for property matching. Information Processing Letters, 105(6): 218–223, March 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See errata [JLW09].

Iliopoulos:2008:NEA

[IR08b] Costas S. Iliopoulos and M. Sohel Rahman. New efficient algorithms for the LCS and constrained LCS problems. Information Processing Letters, 106 (1):13–18, March 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Imbs:2009:NAB

[IR09] Damien Imbs and Michel Raynal. A note on atomicity: Boosting Test&Set to solve consensus. Information Processing Letters, 109(12):589–591, May 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Irving:2007:CRP

[Irv07] Robert W. Irving. The cycle roommates problem: a hard case of kidney exchange. Information Processing Letters, 103(1): 1–4, June 30, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Israeli:2008:AAS

[IS08] Amos Israeli and Oran Sharon.

An approximation algorithm for sequential rectangle placement. Information Processing Letters, 108(6):407–411, November 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Iwen:2009:NCS

[IS09] M. A. Iwen and C. V. Spencer.

A note on compressed sensing and the complexity of matrix multiplication. Information Processing Letters, 109(10):468– 471, April 30, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Igel:2003:CFW

[IT03] Christian Igel and Marc Toussaint. On classes of functions for which No Free Lunch results hold. Information Processing Letters, 86(6):317–321, June 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Iwama:2004:OCP

[IY04] Kazuo Iwama and Kouki Yonezawa. The orthogonal CNN problem. Information Processing Letters, 90(3):115–120, May 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Iwamoto:2006:SSR

[IY06] Mitsugu Iwamoto and Hirosuke Yamamoto. Strongly secure ramp secret sharing schemes for general access structures. Information Processing Letters, 97 (2):52–57, January 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jackson:2003:WSW

[Jac03a] Michael Jackson. Why software writing is difficult and will remain so. Information Processing Letters, 88(1–2):13–25, October 17, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jacquemard:2003:RCU

[Jac03b] Florent Jacquemard. Reachability and confluence are undecidable for flat term rewriting systems. Information Processing Letters, 87(5):265–270, September 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jimenez:2006:IUF

[JAF06] Ernesto Jiménez, Sergio Arévalo, and Antonio Fernández. Implementing unreliable failure detectors with unknown membership. Information Processing Letters, 100(2):60–63, October 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jain:2009:SOPb

[Jai09a] Sanjay Jain. On some open problems in monotonic and conservative learning. Information Processing Letters, 109(16):923– 926, July 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jain:2009:SOPa

[Jai09b] Sanjay Jain. On some open problems in reflective inductive inference. Information Processing Letters, 109(3):208–211, January 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jung:2004:LPG

[JC04] Joo-Won Jung and Kyung-Yong Chwa. Labeling points with given rectangles. Information Processing Letters, 89(3):115– 121, February 14, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ji:2007:FSJ

[JC07] Min Ji and T. C. E. Cheng.

An FPTAS for scheduling jobs with piecewise linear decreasing processing times to minimize makespan. Information Processing Letters, 102(2–3):41–47, April 30, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ji:2008:FPM

[JC08] Min Ji and T. C. E. Cheng.

An FPTAS for parallel-machine scheduling under a grade of service provision to minimize makespan. Information Processing Letters, 108(4):171–174, October 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jea:2004:EFA

[JCL04] Kuen-Fang Jea, Ming-Yuan Chang, and Ke-Chung Lin. An efficient and flexible algorithm for online mining of large itemsets. Information Processing Letters, 92(6):311–316, December 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jarray:2008:CRH

[JCP08] Fethi Jarray, Marie-Christine Costa, and Christophe Picouleau. Complexity results for the horizontal bar packing problem. Information Processing Letters, 108(6):356–359, November 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ji:2009:MBN

[JCSX09] Zexuan Ji, Qiang Chen, Quan-Sen Sun, and De-Shen Xia. A moment-based nonlocal-means algorithm for image denoising. Information Processing Letters, 109(23–24):1238–1244, November 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jha:2003:PDK

[Jha03] Pranava K. Jha. Perfect rdomination in the Kronecker product of two cycles, with an application to diagonal/toroidal mesh. Information Processing Letters, 87(3):163–168, August 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Juan:2008:SDP

[JHfS08] Justie Su-Tzu Juan, Chun-Ming Huang, and I fan Sun. The strong distance problem on the Cartesian product of graphs. Information Processing Letters, 107(2):45–51, July 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jiang:2005:USP

[Jia05] Minghui Jiang. UPS-k : a set partitioning problem with applications in UPS pickup-delivery system. Information Processing Letters, 93(4):173–175, February 28, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jiang:2006:AMC

[Jia06a] Minghui Jiang. Approximating minimum coloring and maximum independent set in dotted interval graphs. Information Processing Letters, 98(1):29–33, April 15, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jiang:2006:NAA

[Jia06b] Minghui Jiang. A new approximation algorithm for labeling points with circle pairs. Information Processing Letters, 99 (4):125–129, August 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jiao:2008:NRP

[Jia08] Li Jiao. A note on regular Petri nets. Information Processing Letters, 108(3):110–114, October 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jonsson:2002:FCA

[JJ02] Fredrik Jönsson and Thomas Johansson. A fast correlation attack on LILI-128. Information Processing Letters, 81(3):127– 132, February 14, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/85/34/
                                        28/abstract.html.

Jerebic:2005:CPC

[JKŠ05] Janja Jerebic, Sandi Klavžar, and Simon Špacapan. Characterizing r-perfect codes in direct products of two and three cycles. Information Processing Letters, 94(1):1–6, April 15, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jez:2009:TDC

[JŁ09] Artur Jeż and Jakub Łopuszański. On the two-dimensional cow search problem. Information Processing Letters, 109(11):543– 547, May 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jacobsen:2002:ECN

[JLN02] Lars Jacobsen, Kim S. Larsen, and Morten N. Nielsen. On the existence and construction of non-extreme (a, b)-trees. Information Processing Letters, 84 (2):69–73, October 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jeong:2001:PAM

[JLO01] Kil-Hyun Jeong, Jong-Kyu Lee, and Hee-Kuck Oh. Performance analysis of a multicasting protocol with contentionless minislots. Information Processing Letters, 79(3):127– 133, July 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/32/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/80/32/29/article.
pdf.

Juan:2009:EFI

[JLW09] M. T. Juan, J. J. Liu, and Y. L. Wang. Errata for “Faster index for property matching”. Information Processing Letters, 109(18):1027–1029, August 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [IR08a].

Jiang:2007:SCA

[JLY07] M. Jiang, Y. P. Luo, and S. Y.

Yang. Stochastic convergence analysis and parameter selection of the standard particle swarm optimization algorithm. Information Processing Letters, 102 (1):8–16, April 15, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jooyandeh:2009:UVD

[JMM09] Mohammadreza Jooyandeh, Ali Mohades, and Maryam Mirzakhah. Uncertain Voronoi diagram. Information Processing Letters, 109(13):709–712, June 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jamison:2006:ACP

[JMV06] Robert E. Jamison, Gretchen L. Matthews, and John Villalpando. Acyclic colorings of products of trees. Information Processing Letters, 99(1):7–12, July 16, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Janssen:2006:MIB

[JN06] Philippe Janssen and Lhouari Nourine. Minimum implicational basis for Λ-semidistributive lattices. Information Processing Letters, 99(5):199–202, September 15, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jean:2007:GRC

[JN07] Géraldine Jean and Macha Nikolski. Genome rearrangements: a correct algorithm for optimal capping. Information Processing Letters, 104(1):14– 20, September 30, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jana:2002:PPC

[JNK^{+}02] Prasanta K. Jana, B. Damodara Naidu, Shailendra Kumar, Monish Arora, and Bhabani P. Sinha. Parallel prefix computation on extended multi-mesh network. Information Processing Letters, 84(6):295–303, December 31, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Johnson:2004:TFS

[Joh04] L. F. Johnson. Tumble, a fast simple iteration algorithm for Fibonacci. Information Processing Letters, 89(4):187–189, February 28, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jonsson:2002:TSP

[Jon02] Håkan Jonsson. The Traveling Salesman Problem for lines in the plane. Information Processing Letters, 82(3): 137–142, May 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/86/34/
30/abstract.html.

Jones:2003:OSC

[Jon03a] Cliff B. Jones. Operational semantics: Concepts and their expression. Information Processing Letters, 88(1–2):27–32, October 17, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jonsson:2003:ASZ

[Jon03b] Håkan Jonsson. An approximative solution to the Zookeeper’s Problem. Information Processing Letters, 87(6):301–307, September 30, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Joye:2003:CPY

[Joy03] Marc Joye. Cryptanalysis of a pay-as-you-watch system. Information Processing Letters, 88 (3):119–120, November 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jager:2009:NPG

[JP09] Gerold Jäger and Marcin Peczarski. The number of pessimistic guesses in Generalized Mastermind. Information Processing Letters, 109(12):635– 641, May 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jiang:2003:SFA

[JQQ^{+}03] Minghui Jiang, Jianbo Qian, Zhongping Qin, Binhai Zhu, and Robert Cimikowski. A simple factor-3 approximation for labeling points with circles. Information Processing Letters, 87(2): 101–105, July 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jothi:2004:SND

[JR04] Raja Jothi and Balaji Raghavachari.

Survivable network design: the capacitated minimum spanning network problem. Information Processing Letters, 91(4):183– 190, August 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Janiak:2007:LEG

[JR07] Adam Janiak and Radosław Rudek. The learning effect: Getting to the core of the problem. Information Processing Letters, 103(5):183–187, August 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

James:2001:RDO

[JS01] Jerry James and Ambuj K.

Singh. Recovering distributed objects. Information Processing Letters, 77(2–4):143–150, February 28, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/29/37/article.
pdf.

Jancar:2007:NEA

[JS07] Petr Jančar and Zdeněk Sawa.

A note on emptiness for alternating finite automata with a one-letter alphabet. Information Processing Letters, 104(5):164– 167, November 30, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ji:2004:SPM

[JTG04] Mingjun Ji, Huanwen Tang, and Juan Guo. A single-point mutation evolutionary programming. Information Processing Letters, 90(6):293–299, June 30, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jukna:2004:MNN

[Juk04] Stasys Jukna. On the minimum number of negations leading to super-polynomial savings. Information Processing Letters, 89(2):71–74, January 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jukna:2005:PVN

[Juk05] Stasys Jukna. On the P versus NP intersected with co-NP question in communication complexity. Information Processing Letters, 96(6):202–206, December 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jukna:2009:NST

[Juk09] S. Jukna. A nondeterministic space-time tradeoff for linear codes. Information Processing Letters, 109(5):286–289, February 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jung:2005:DFS

[Jun05] Haejae Jung. The d-deap*: a fast and simple cache-aligned dary deap. Information Processing Letters, 93(2):63–67, January 31, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jourdan:2005:MNI

[JUZ05] Guy-Vincent Jourdan, Hasan Ural, and Nejib Zaguia. Minimizing the number of inputs while applying adaptive test cases. Information Processing Letters, 94(4):165–169, May 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Javanian:2006:DNR

[JVA06] Mehri Javanian and Mohammad Q. Vahidi-Asl. Depth of nodes in random recursive kary trees. Information Processing Letters, 98(3):115–118, May 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jiang:2009:PCC

[JX09] Jiuchuan Jiang and Xiaojun Xia. Prominence convergence in the collective synchronization of situated multi-agents. Information Processing Letters, 109 (5):278–285, February 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jun:2004:EGA

[JXL04] Ye Jun, Liu Xiande, and Han Lu. Evolutionary game algorithm for continuous parameter optimization. Information Processing Letters, 91(5):211–219, September 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xue:2009:ORF

[jXyL09] Zhan jun Xue and San yang Liu. An optimal result on faulttolerant cycle-embedding in alternating group graphs. Information Processing Letters, 109(21–22):1197–1201, October 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

etal:2002:OAR

[K^{+}02] N. Kitsios et al. An optimal algorithm for reporting visible rectangles. Information Processing Letters, 81(5):283–288, March 16, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:2002:STE

[KA02a] Yong-Jik Kim and James H. Anderson. A space- and timeefficient local-spin spin lock. Information Processing Letters, 84 (1):47–55, October 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kosch:2002:PMJ

[KA02b] Harald Kosch and Solomon Atnafu. Processing a multimedia join through the method of nearest neighbor search. Information Processing Letters, 82(5): 269–276, June 15, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kikuchi:2006:EBE

[KA06] Yosuke Kikuchi and Toru Araki.

Edge-bipancyclicity and edgefault-tolerant bipancyclicity of bubble-sort graphs. Information Processing Letters, 100(2): 52–59, October 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kaser:2004:CAO

[Kas04a] Owen Kaser. Compressing arrays by ordering attribute values. Information Processing Letters, 92(5):253–256, December 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kaski:2004:PST

[Kas04b] Petteri Kaski. Packing Steiner trees with identical terminal sets. Information Processing Letters, 91(1):1–5, July 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ko:2008:PRF

[KAS08] Youngjoong Ko, Hongkuk An, and Jungyun Seo. Pseudorelevance feedback and statistical query expansion for Web snippet generation. Information Processing Letters, 109(1):

18–22, December 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Katriel:2004:ACS

[Kat04] Irit Katriel. On the algebraic complexity of set equality and inclusion. Information Processing Letters, 92(4):175–178, November 30, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Khachiyan:2007:GPA

[KBEG07] Leonid Khachiyan, Endre Boros, Khaled Elbassioni, and Vladimir Gurvich. A global parallel algorithm for the hypergraph transversal problem. Information Processing Letters, 101(4): 148–155, February 28, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kulkarni:2005:ARW

[KBOR05] Sandeep S. Kulkarni, Chase Bolen, John Oleszkiewicz, and Andrew Robinson. Alternators in read/write atomicity. Information Processing Letters, 93 (5):207–215, March 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kwon:2002:AAG

[KC02] Oh-Heum Kwon and Kyung-Yong Chwa. Approximation algorithms for general parallel task scheduling. Information Processing Letters, 81(3):143– 150, February 14, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/34/
31/abstract.html.

Kim:2003:NCS

[KC03a] Jae-Hoon Kim and Kyung-Yong Chwa. Non-clairvoyant scheduling for weighted flow time. Information Processing Letters, 87 (1):31–37, July 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:2003:ODS

[KC03b] Jae-Hoon Kim and Kyung-Yong Chwa. Online deadline scheduling on faster machines. Information Processing Letters, 85(1): 31–37, January 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ku:2003:ORP

[KCH03] Cheng-Yuan Ku, Din-Yuen Chan, and Lain-Chyr Hwang. Optimal reservation policy for two queues in tandem. Information Processing Letters, 85(1): 27–30, January 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kong:2006:IGA

[KCYL06] Fanyu Kong, Zhun Cai, Jia Yu, and Daxing Li. Improved generalized Atkin algorithm for computing square roots in finite fields. Information Processing Letters, 98(1):1–5, April 15, 2006. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Kadrie:2001:APS

[KDTS01] Ahmad Kadrie, V. R. Dare, D. G. Thomas, and K. G. Subramanian. Algebraic properties of the shuffle over ω-trajectories. Information Processing Letters, 80(3):139–144, November 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/84/35/30/abstract.html.

Kaldewaij:2001:ORT

[KdV01] Anne Kaldewaij and Laurens de Vries. Optimal real-time garbage collection for acyclic pointer structures. Information Processing Letters, 77(2–4):151– 157, February 28, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/29/38/article.
pdf.

Kerenidis:2004:QSP

[KdW04] Iordanis Kerenidis and Ronald de Wolf. Quantum symmetricallyprivate information retrieval. Information Processing Letters, 90 (3):109–114, May 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Keller:2002:HAS

[Kel02] Jörg Keller. A heuristic to accelerate in-situ permutation algorithms. Information Processing Letters, 81(3):119–125, February 14, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/34/
27/abstract.html.

Kawai:2001:FBD

[KFS01] Hiroyuki Kawai, Naohiro Fujikake, and Yukio Shibata. Factorization of de Bruijn digraphs by cycle-rooted trees. Information Processing Letters, 77(5–6): 269–275, March 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/32/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/32/31/article.
pdf.

Khosravi:2004:SPS

[KG04] Ramtin Khosravi and Mohammad Ghodsi. Shortest paths in simple polygons with polygonmeet constraints. Information Processing Letters, 91(4):171– 176, August 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kashyap:2005:IRP

[KG05] Sujatha Kashyap and Vijay K.

Garg. Intractability results in predicate detection. Information Processing Letters, 94(6): 277–282, June 30, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kao:2003:BTF

[KH03] Shin-Shin Kao and Lih-Hsing Hsu. Brother trees: A family of optimal 1_{p}-Hamiltonian and 1-edge Hamiltonian graphs. Information Processing Letters, 86 (5):263–269, June 15, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Koo:2009:SVN

[KHL09] Woo Kwon Koo, Jung Yeon Hwang, and Dong Hoon Lee. Security vulnerability in a noninteractive ID-based proxy reencryption scheme. Information Processing Letters, 109(23–24): 1260–1262, November 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kuo:2009:EAM

[KHLL09] Chi-Jung Kuo, Chiun-Chieh Hsu, Hon-Ren Lin, and Kung-Kuei Lin. An efficient algorithm for minimum feedback vertex sets in rotator graphs. Information Processing Letters, 109(9): 450–453, April 15, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Katayama:2005:ELS

[KHN05] Kengo Katayama, Akihiro Hamamoto, and Hiroyuki Narihisa. An effective local search for the maximum clique problem. Information Processing Letters, 95(5):503–511, September 15, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Khuller:1999:ASC

[Khu99] Samir Khuller. An O(|V |^{2})

algorithm for single connectedness. Information Processing Letters, 72(3–4):105–107, November 26, 1999. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/60/28/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/60/28/29/article.

pdf. See addendum [Khu00].

Khuller:2000:AAS

[Khu00] Samir Khuller. Addendum to “An O(|V |^{2}) algorithm for single connectedness”, [Information Processing Letters 72(3– 4) (1999) 105–107]. Information Processing Letters, 74(5– 6):263, June 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/28/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/28/37/article.

pdf. See [Khu99].

Kim:2003:LTA

[Kim03] Sung Kwon Kim. Linear-time algorithm for finding a maximumdensity segment of a sequence. Information Processing Letters, 86(6):339–342, June 30, 2003.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:2005:FLN

[Kim05] Sung Kwon Kim. Finding a longest nonnegative path in a constant degree tree. Information Processing Letters, 93(6): 275–279, March 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kirsten:2002:FPP

[Kir02] Daniel Kirsten. The finite power problem revisited. Information Processing Letters, 84(6):291– 294, December 31, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kapoor:2004:MVD

[KJ04] Hemangee K. Kapoor and Mark B. Josephs. Modelling and verification of delay-insensitive circuits using CCS and the Concurrency Workbench. Information Processing Letters, 89(6): 293–296, March 31, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Koubkova:2002:ATC

[KK02] A. Koubková and V. Koubek. Algorithms for transitive closure. Information Processing Letters, 81(6):289–296, March 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/85/37/
                                        27/abstract.html.

Kovacs:2004:PTM

[KK04] András Kovács and Tamás Kis.

Partitioning of trees for minimizing height and cardinality. Information Processing Letters, 89(4):181–185, February 28, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Konig:2007:ICC

[KK07] Barbara König and Vitali Kozioura. Incremental construction of coverability graphs. Information Processing Letters, 103(5):203–209, August 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ko:2003:EAE

[KKL03] Jong Ha Ko, Sang Hee Kim, and Jong Kyu Lee. An ENA algorithm to enhance the performance of TCP over satellite links. Information Processing Letters, 85(4):215–219, February 28, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kloks:2000:FCS

[KKM00] Ton Kloks, Dieter Kratsch, and Haiko Müller. Finding and counting small induced subgraphs efficiently. Information Processing Letters, 74(3–4): 115–121, May 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/63/27/

29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/63/27/29/article.
pdf.

Kranakis:2004:AHA

[KKS04] Evangelos Kranakis, Danny Krizanc, and Sunil Shende. Approximate hotlink assignment. Information Processing Letters, 90(3):121–128, May 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kochol:2008:CAE

[KKSŠ08] Martin Kochol, Nad’a Krivoňáková, Silvia Smejová, and Katarína Šranková. Complexity of approximation of 3-edge-coloring of graphs. Information Processing Letters, 108(4):238–241, October 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Knill:2001:QCQ

[KL01] E. Knill and R. Laflamme. Quantum computing and quadratically signed weight enumerators. Information Processing Letters, 79(4):173–179, August 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/33/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/80/33/28/article.
pdf.

Klasing:2004:HRA

[KL04a] Ralf Klasing and Christian Laforest. Hardness results and approximation algorithms of ktuple domination in graphs. Information Processing Letters, 89 (2):75–83, January 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Korsh:2004:CTG

[KL04b] James F. Korsh and Paul S.

LaFollette. Constant time generation of derangements. Information Processing Letters, 90 (4):181–186, May 31, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Klein:2008:SOA

[Kle08] Shmuel T. Klein. Should one always use repeated squaring for modular exponentiation? Information Processing Letters, 106 (6):232–237, June 15, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:2004:ESJ

[KLK04] Jongik Kim, Sang Ho Lee, and Hyoung-Joo Kim. Efficient structural joins with clustered extents. Information Processing Letters, 91(2):69–75, July 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:2006:RFC

[KLK06] Sangwon Kim, Joonwon Lee, and Jinsoo Kim. Runtime feasibility check for non-preemptive real-time periodic tasks. Information Processing Letters, 97 (3):83–87, February 14, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:2003:APB

[KLN^{+}03] Jong Min Kim, Donghee Lee, Sam H. Noh, Sang Lyul Min, Yookun Cho, and Chong Sang Kim. An accurate and practical buffer allocation model for the buffer cache based on marginal gains. Information Processing Letters, 85(2):93–97, January 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:2004:FAG

[KLP04] Sung-Ryul Kim, Inbok Lee, and Kunsoo Park. A fast algorithm for the generalized k-keyword proximity problem given keyword offsets. Information Processing Letters, 91(3):115–120, August 16, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:2000:PSA

[KLY00] Hyun-Sung Kim, Sung-Woo Lee, and Kee-Young Yoo. Partitioned systolic architecture for modular multiplication in GF (2 m ). Information Processing Letters, 76(3):135– 139, December 15, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/27/

30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/27/30/article.
pdf.

Karagiorgos:2002:ADA

[KM02] Gregory Karagiorgos and Nikolaos M. Missirlis. Accelerated diffusion algorithms for dynamic load balancing. Information Processing Letters, 84(2):61–67, October 31, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Koo:2003:EBE

[KM03a] Zakil Koo and Songchun Moon.

Effects of broadcast errors on concurrency control in wireless broadcasting environments. Information Processing Letters, 86 (1):13–21, April 15, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kwek:2003:OSR

[KM03b] Stephen Kwek and Kurt Mehlhorn. Optimal search for rationals. Information Processing Letters, 86(1):23–26, April 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kontkanen:2007:LTA

[KM07] Petri Kontkanen and Petri Myllymäki. A linear-time algorithm for computing the multinomial stochastic complexity. Information Processing Letters, 103(6): 227–233, September 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kang:2004:CMP

[KMCC04] Heum-Geun Kang, Jun-Ki Min, Seok-Ju Chun, and Chin-Wan Chung. A compression method for prefix-sum cubes. Information Processing Letters, 92(2): 99–105, October 31, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kumaraswamy:2004:FDV

[KMF04] Krishna Kumaraswamy, Vasileios Megalooikonomou, and Christos Faloutsos. Fractal dimension and vector quantization. Information Processing Letters, 91 (3):107–113, August 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kneis:2006:PPD

[KMRR06] Joachim Kneis, Daniel Mölle, Stefan Richter, and Peter Rossmanith. Parameterized power domination complexity. Information Processing Letters, 98 (4):145–149, May 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kitsios:2002:OAR

[KMS^{+}02] N. Kitsios, C. Makris, S. Sioutas, A. Tsakalidis, J. Tsaknakis, and B. Vassiliadis. An optimal algorithm for reporting visible rectangles. Information Processing Letters, 81(5):283– 288, March 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/36/
35/abstract.html.

Kawabe:2007:TPA

[KMST07] Yoshinobu Kawabe, Ken Mano, Hideki Sakurada, and Yasuyuki Tsukada. Theorem-proving anonymity of infinite-state systems. Information Processing Letters, 101(1):46–51, January 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kouris:2004:EAD

[KMT04] Ioannis N. Kouris, Christos H.

Makris, and Athanasios K. Tsakalidis. Efficient automatic discovery of ‘hot’ itemsets. Information Processing Letters, 90 (2):65–72, April 30, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kano:2005:PST

[KMU05] M. Kano, C. Merino, and J. Urrutia. On plane spanning trees and cycles of multicolored point sets with few intersections. Information Processing Letters, 93 (6):301–306, March 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kosowski:2006:AAM

[KMŻ06] Adrian Kosowski, Michał Małafiejski, and Paweł Żyliński. An approximation algorithm for maximum P_{3}-packing in subcubic graphs. Information Processing Letters, 99(6):230–233, September 30, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kerenidis:2004:WCF

[KN04a] I. Kerenidis and A. Nayak. Weak coin flipping with small bias. Information Processing Letters, 89 (3):131–135, February 14, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krasikov:2004:FNS

[KN04b] I. Krasikov and S. D. Noble. Finding next-to-shortest paths in a graph. Information Processing Letters, 92(3):117–119, November 15, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Knoblauch:2005:NAM

[Kno05] Andreas Knoblauch. Neural associative memory for brain modeling and information retrieval. Information Processing Letters, 95(6):537–544, September 30, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Koura:2009:NGA

[KOH09] Ippei Koura, Takao Ono, and Tomio Hirata. A note on the Greedy algorithm for finding independent sets of C_{k}-free graphs. Information Processing Letters, 109(10):485–489, April 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Koiran:2002:TTS

[Koi02] Pascal Koiran. Transfer theorems via sign conditions. Information Processing Letters, 81 (2):65–69, January 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/85/33/28/abstract.html.

Koivisto:2006:OCS

[Koi06] Mikko Koivisto. Optimal 2constraint satisfaction via sum– product algorithms. Information Processing Letters, 98(1): 24–28, April 15, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kolman:2003:NGA

[Kol03] Petr Kolman. A note on the greedy algorithm for the unsplittable flow problem. Information Processing Letters, 88(3): 101–105, November 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Koltun:2004:RSG

[Kol04] Vladlen Koltun. Ready, set, go!

The Voronoi diagram of moving points that start from a line. Information Processing Letters, 89(5):233–235, March 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kolliopoulos:2005:MCS

[Kol05] Stavros G. Kolliopoulos. Minimumcost single-source 2-splittable flow. Information Processing Letters, 94(1):15–18, April 15, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Koutis:2005:FPA

[Kou05] Ioannis Koutis. A faster parameterized algorithm for set packing. Information Processing Letters, 94(1):7–9, April 15, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Koutis:2006:PCI

[Kou06] Ioannis Koutis. Parameterized complexity and improved inapproximability for computing the largest j-simplex in a V - polytope. Information Processing Letters, 100(1):8–13, October 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kowalik:2007:AQD

[Kow07] Łukasz Kowalik. Adjacency queries in dynamic sparse graphs. Information Processing Letters, 102(5):191–195, May 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kang:2006:RBS

[KP06] Bo Gyeong Kang and Je Hong Park. On the relationship between squared pairings and plain pairings. Information Processing Letters, 97(6):219–224, March 31, 2006. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:2009:IMQ

[KP09] Eunsang Kim and Kunsoo Park.

Improving multikey Quicksort for sorting strings with many equal elements. Information Processing Letters, 109(9):454– 459, April 15, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:2007:ASA

[KPC07] Seok Woo Kim, Seong-Hun Paeng, and Hee Je Cho. Approximately n-secting an angle. Information Processing Letters, 103(1):19–23, June 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:2004:EPS

[KPL04] Sang-Wook Kim, Dae-Hyun Park, and Heon-Gil Lee. Efficient processing of subsequence matching with the Euclidean metric in time-series databases. Information Processing Letters, 90(5):253–260, June 15, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kovalyov:2002:PAL

[KPS02] Mikhail Y. Kovalyov, Marcus Pattloch, and Günter Schmidt. A polynomial algorithm for lotsize scheduling of two type tasks. Information Processing Letters, 83(4):229–235, August 31, 2002.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Keidar:2003:SPU

[KR03a] Idit Keidar and Sergio Rajsbaum. A simple proof of the uniform consensus synchronous lower bound. Information Processing Letters, 85(1):47–52, January 16, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kralovic:2003:MFV

[KR03b] Rastislav Královič and Peter Ružička. Minimum feedback vertex sets in shuffle-based interconnection networks. Information Processing Letters, 86(4): 191–196, May 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kashem:2004:OPA

[KR04] Md. Abul Kashem and M. Ziaur Rahman. An optimal parallel algorithm for c-vertex-ranking of trees. Information Processing Letters, 92(4):179–184, November 30, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kopparty:2005:FPE

[KR05] Swastik Kopparty and Chinya V. Ravishankar. A framework for pursuit evasion games in R^{n}. Information Processing Letters, 96 (3):114–122, November 15, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krovetz:2006:VUH

[KR06] Ted Krovetz and Phillip Rogaway. Variationally universal hashing. Information Processing Letters, 100(1):36–39, October 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kooshesh:2008:EIA

[KR08] A. Kooshesh and B. Ravikumar.

Efficient implementation of algorithms for approximate exponentiation. Information Processing Letters, 105(4):131–137, February 15, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krajicek:2007:SPT

[Kra07] Jan Krajíček. Substitutions into propositional tautologies. Information Processing Letters, 101 (4):163–167, February 28, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Karmakar:2008:FCS

[KRD08] Arindam Karmakar, Sasanka Roy, and Sandip Das. Fast computation of smallest enclosing circle with center on a query line segment. Information Processing Letters, 108(6):343–346, November 30, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krivelevich:2002:DCE

[Kri02] Michael Krivelevich. Deciding k-colorability in expected polynomial time. Information Processing Letters, 81(1): 1–6, January 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/27/
27/abstract.html.

Konjevod:2001:APM

[KRS01] Goran Konjevod, R. Ravi, and F. Sibel Salman. On approximating planar metrics by tree metrics. Information Processing Letters, 80(4):213–219, November 30, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/84/34/
32/abstract.html.

Kishore:2000:OPA

[KS00] K. V. R. C. N. Kishore and Sanjeev Saxena. An optimal parallel algorithm for general maximal matchings is as easy as for bipartite graphs. Information Processing Letters, 75(4):145– 151, September 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/29/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/29/25/article.
pdf.

Kanda:2002:CVT

[KS02a] Takeshi Kanda and Kokichi Sugihara. Comparison of various trees for nearest-point search with/without the Voronoi diagram. Information Processing Letters, 84(1):17–22, October 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kobayashi:2002:CSG

[KS02b] Zenji Kobayashi and Takeshi Sekiguchi. On a characterization of the standard Gray code by using its edge type on a hypercube. Information Processing Letters, 81(5):231–237, March 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/85/36/27/abstract.html.

Kavvadias:2003:MBD

[KS03a] Dimitris J. Kavvadias and Elias C. Stavropoulos. Monotone Boolean dualization is in co − NP[log^{2} n]. Information Processing Letters, 85(1):1–6, January 16, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kikuchi:2003:DNG

[KS03b] Yosuke Kikuchi and Yukio Shibata. On the domination numbers of generalized de Bruijn digraphs and generalized Kautz digraphs. Information Processing Letters, 86(2):79–85, April 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kaplan:2005:GAS

[KS05] Haim Kaplan and Nira Shafrir.

The greedy algorithm for shortest superstrings. Information Processing Letters, 93(1):13–17, January 16, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kaplan:2006:GAE

[KS06] Haim Kaplan and Nira Shafrir. The greedy algorithm for edit distance with moves. Information Processing Letters, 97(1): 23–27, January 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kowalski:2009:GPP

[KS09] Dariusz R. Kowalski and Michał Strojnowski. Gossiping by processors prone to omission failures. Information Processing Letters, 109(6):308–314, February 28, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kannan:2003:MSI

[KSRI03] Rajgopal Kannan, S. Sarangi, Sibabrata Ray, and S. S. Iyengar. Minimal sensor integrity: Measuring the vulnerability of sensor grids. Information Processing Letters, 86(1):49–55, April 15, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kavvadias:2000:GAM

[KSS00] Dimitris J. Kavvadias, Martha Sideri, and Elias C. Stavropoulos. Generating all maximal models of a Boolean expression. Information Processing Letters, 74(3–4):157– 162, May 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/63/27/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/63/27/34/article.
pdf.

Kim:2000:PTD

[KSY00] Sung Kwon Kim, Chan-Su Shin, and Tae-Cheon Yang. Placing two disks in a convex polygon. Information Processing Letters, 73(1–2):33–39, January 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/62/25/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/62/25/31/article.
pdf.

Kobayashi:2002:NMC

[KTK02] Noritaka Kobayashi, Tatsuhiro Tsuchiya, and Tohru Kikuno. A new method for constructing pair-wise covering designs for software testing. Information Processing Letters, 81(2): 85–91, January 31, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/33/
32/abstract.html.

Kuba:2006:QPS

[Kub06] Markus Kuba. On Quickselect, partial sorting and Multiple Quickselect. Information Processing Letters, 99(5):181– 186, September 15, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kuhn:2008:BSS

[Küh08] Ulrich Kühn. Breaking the Shin–Shin–Rhee remotely keyed encryption schemes. Information Processing Letters, 105(6): 236–240, March 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kujala:2009:AAO

[Kuj09]  Jussi Kujala.  Assembling

approximately optimal binary search trees efficiently using arithmetics. Information Processing Letters, 109(16):962– 966, July 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kundu:2000:OAC

[Kun00] Sukhamay Kundu. An optimal O(N^{2}) algorithm for computing the min-transitive closure of a weighted graph. Information Processing Letters, 74(5–6): 215–220, June 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/63/28/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/63/28/30/article.
                                        pdf.

Kurowski:2003:SEF

[Kur03] Maciej Kurowski. Simple and efficient floor-planning. Information Processing Letters, 86(3): 113–119, May 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kurowski:2004:LBN

[Kur04] Maciej Kurowski. A 1.235 lower bound on the number of points needed to draw all n-vertex planar graphs. Information Processing Letters, 92(2):95–98, October 31, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kurosaki:2008:DDT

[Kur08] Tetsuo Kurosaki. Direct definition of a ternary infinite squarefree sequence. Information Processing Letters, 106(5):175–179, May 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kutzkov:2007:NUB

[Kut07] Konstantin Kutzkov. New upper bound for the #3-sat problem. Information Processing Letters, 105(1):1–5, December 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kristiansen:2003:CCF

[KV03] Lars Kristiansen and Paul J.

Voda. Complexity classes and fragments of C. Information Processing Letters, 88(5):213– 218, December 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kallahalla:2004:ASR

[KV04] Mahesh Kallahalla and Peter J. Varman. Analysis of simple randomized buffer management for parallel I/O. Information Processing Letters, 90(1):47–52, April 15, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Korze:2005:LSP

[KV05] Danilo Korže and Aleksander Vesel. L(2, 1)-labeling of strong products of cycles. Information Processing Letters, 94(4): 183–190, May 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Keil:2008:RNG

[KV08a] J. Mark Keil and Tzvetalin S. Vassilev. The relative neighbourhood graph is a part of every 30^{∘}-triangulation. Information Processing Letters, 109(2): 93–97, December 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kenkre:2008:CPP

[KV08b] Sreyash Kenkre and Sundar Vishwanathan. The common prefix problem on trees. Information Processing Letters, 105 (6):245–248, March 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kravchenko:2001:HAMa

[KW01a] Svetlana A. Kravchenko and Frank Werner. A heuristic algorithm for minimizing mean flow time with unit setups. Information Processing Letters, 79(6): 291–296, September 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/80/37/33/abstract.html.

Kravchenko:2001:HAMb

[KW01b] Svetlana A. Kravchenko and Frank Werner. A heuristic algorithm for minimizing mean flow time with unit setups. Information Processing Letters, 79(6): 291–296, September 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/80/37/33/abstract.html.

Kato:2005:SSR

[KW05] Ryoichi Kato and Osamu Watanabe. Substring search and repeat search using factor oracles. Information Processing Letters, 93(6):269–274, March 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kwon:2002:STC

[Kwo02a] Oh-Heum Kwon. Scheduling time-constrained multicast messages in circuit-switched tree networks. Information Processing Letters, 84(2):109–116, October 31, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kwon:2002:DSA

[Kwo02b] Taekyoung Kwon. Digital signature algorithm for securing digital identities. Information Processing Letters, 82(5):247–252, June 15, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [Kwo03].

Kwon:2003:EDS

[Kwo03] Taekyoung Kwon. Erratum to: “Digital signature algorithm for securing digital identities”: [Inform. Process. Lett. 82 (2002) 247–252]. Information Processing Letters, 88(4): 201–202, November 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Kwo02b].

Kaneko:2000:STR

[KY00] Atsushi Kaneko and Kiyoshi Yoshimoto. On spanning trees with restricted degrees. Information Processing Letters, 73(5–6):163–165, March 31, 2000. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/62/28/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/62/28/27/article.
                                        pdf.

Kawano:2003:WCA

[KY03a] Sinichiro Kawano and Koichi Yamazaki. Worst case analysis of a greedy algorithm for graph thickness. Information Processing Letters, 85(6):333–337, March 31, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:2003:PCM

[KY03b] Ho-Sook Kim and Hwan-Seung Yong. Personalized cache management for mobile computing environments. Information Processing Letters, 87(4):221–228, August 31, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kuo:2006:MMS

[KY06] Wen-Hung Kuo and Dar-Li Yang. Minimizing the makespan in a single machine scheduling problem with a time-based learning effect. Information Processing Letters, 97(2):64–67, January 31, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kuo:2007:SMS

[KY07] Wen-Hung Kuo and Dar-Li Yang. Single machine scheduling with past-sequence-dependent setup times and learning effects. Information Processing Letters, 102(1):22–26, April 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Khuller:2002:GST

[KZ02] Samir Khuller and An Zhu. The General Steiner Tree-Star problem. Information Processing Letters, 84(4):215–220, November 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kasperski:2006:AAI

[KZ06] Adam Kasperski and Paweł Zieliński. An approximation algorithm for interval data minmax regret combinatorial optimization problems. Information Processing Letters, 97(5): 177–180, March 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kasperski:2009:AMR

[KZ09] Adam Kasperski and Paweł Zieliński. On the approximability of minmax (regret) network optimization problems. Information Processing Letters, 109 (5):262–266, February 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leal:2001:SLV

[LA01] William Leal and Anish Arora. State-level and value-level simulations in data refinement. Information Processing Letters, 77(2–4):159–167, February 28, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/

39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/29/39/article.
pdf.

Laing:2007:NIC

[Lai07] Kofi A. Laing. Name-independent compact routing in trees. Information Processing Letters, 103 (2):57–60, July 16, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lai:2009:NPB

[Lai09]  Chia-Jui Lai.  A note on

path bipancyclicity of hypercubes. Information Processing Letters, 109(19):1129–1130, September 15, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lange:2008:PMT

[Lan08] Martin Lange. A purely modeltheoretic proof of the exponential succinctness gap between CTL ^{+} and CTL. Information Processing Letters, 108(5):308– 312, November 15, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lasota:2009:ELB

[Las09] Sławomir Lasota. EXPSPACE lower bounds for the simulation preorder between a communication-free Petri net and a finite-state system. Information Processing Letters, 109 (15):850–855, July 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Latifi:2007:SFT

[Lat07] Shahram Latifi. A study of fault tolerance in star graph. Information Processing Letters, 102 (5):196–200, May 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lavallee:2008:RCC

[Lav08] Sylvain Lavallée. N-rationality of a certain class of formal series. Information Processing Letters, 108(3):124–126, October 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:2001:ECM

[LC01] Seok-Lyong Lee and Chin-Wan Chung. On the effective clustering of multidimensional data sequences. Information Processing Letters, 80(2): 87–95, October 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/80/37/
                                        30/abstract.html.

Liu:2002:SJR

[LC02] Zhaohui Liu and T. C. Edwin Cheng. Scheduling with job release dates, delivery times and preemption penalties. Information Processing Letters, 82(2): 107–111, April 30, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/86/33/
                                        32/abstract.html.

Laih:2003:COP

[LC03a] C. S. Laih and S. Y. Chiou.

Cryptanalysis of an optimized protocol for mobile network authentication and security. Information Processing Letters, 85 (6):339–341, March 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liao:2003:TDG

[LC03b] Chung-Shou Liao and Gerard J.

Chang. k-tuple domination in graphs. Information Processing Letters, 87(1):45–50, July 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lushman:2003:PCR

[LC03c] Brad Lushman and Gordon V.

Cormack. Proof of correctness of Ressel’s adOPTed algorithm. Information Processing Letters, 86(6):303–310, June 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lai:2004:PCT

[LC04] Yung-Ling Lai and Gerard J.

Chang. On the profile of the corona of two graphs. Information Processing Letters, 89(6): 287–292, March 31, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2006:LSS

[LC06] Yuan Li and T. W. Cusick.

Linear structures of symmetric functions over finite fields. Information Processing Letters, 97 (3):124–127, February 14, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:2008:LTA

[LC08] Min-Sheng Lin and Yung-Jui Chen. Linear time algorithms for counting the number of minimal vertex covers with minimum/maximum size in an interval graph. Information Processing Letters, 107(6):257–264, August 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:2009:SMC

[LC09a] Chuan-Min Lee and Maw-Shang Chang. Signed and minus cliquetransversal functions on graphs. Information Processing Letters, 109(8):414–417, March 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:2009:CNV

[LC09b] Min-Sheng Lin and Yung-Jui Chen. Counting the number of vertex covers in a trapezoid graph. Information Processing Letters, 109(21–22):1187–1192, October 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:2006:NSC

[LCC06] Jyh-Shyan Lin, Jen-Chun Chang, and Rong-Jaye Chen. New simple constructions of distanceincreasing mappings from binary vectors to permutations. Information Processing Letters, 100(2):83–89, October 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:2006:BAA

[LCKL06] Jeng Farn Lee, Meng Chang Chen, Ming Tat Ko, and Wanjiun Liao. Bandwidth allocation algorithms for weighted maximum rate constrained link sharing policy. Information Processing Letters, 97(6):238–243, March 31, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:2006:VCP

[LCL06] Guohui Lin, Zhipeng Cai, and Dekang Lin. Vertex covering by paths on trees with its applications in machine translation. Information Processing Letters, 97(2):73–81, January 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lai:2008:FOD

[LCL08] Hsin-Hao Lai, Gerard J. Chang, and Ko-Wei Lih. On fully orientability of 2-degenerate graphs. Information Processing Letters, 105(5):177–181, February 29, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:2009:AAM

[LCLP09] Kangbok Lee, Byung-Cheon Choi, Joseph Y.-T. Leung, and Michael L. Pinedo. Approximation algorithms for multiagent scheduling to minimize total weighted completion time. Information Processing Letters, 109(16):913–917, July 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2009:SRE

[LCM09] Juan Liu, Xing Chen, and Jixiang Meng. Super restricted edge connected Cartesian product graphs. Information Processing Letters, 109(13):655–659, June 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luo:2007:CPG

[LCW07] Xiaofang Luo, Min Chen, and Weifan Wang. On 3-colorable planar graphs without cycles of four lengths. Information Processing Letters, 103(4):150–156, August 16, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Litow:2004:GCZ

[LD04] B. Litow and N. Deo. Graph compression and the zeros of polynomials. Information Processing Letters, 92(1):39–44, October 16, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lai:2002:REN

[LE02] Wei Lai and Peter Eades. Removing edge-node intersections in drawings of graphs. Information Processing Letters, 81 (2):105–110, January 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/85/33/35/abstract.html.

Lechtenborger:2004:CUC

[Lec04] Jens Lechtenbörger. Computing unique canonical covers for simple FDs via transitive reduction. Information Processing Letters, 92(4):169–174, November 30, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lecroq:2007:FES

[Lec07]  Thierry Lecroq.  Fast ex-

act string matching algorithms. Information Processing Letters, 102(6):229–235, June 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:2009:CSM

[Lee09] Chuan-Min Lee. On the complexity of signed and minus total domination in graphs. Information Processing Letters, 109(20): 1177–1181, September 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leino:2001:REN

[Lei01] K. Rustan M. Leino. Real estate of names. Information Processing Letters, 77(2–4):169– 171, February 28, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/
40/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/29/40/article.
pdf.

Leino:2005:EWP

[Lei05] K. Rustan M. Leino. Efficient weakest preconditions. Information Processing Letters, 93(6): 281–288, March 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luccio:2007:RRD

[LEP07] Fabrizio Luccio, Antonio Mesa Enriquez, and Linda Pagli. k-Restricted rotation distance between binary trees. Information Processing Letters, 102(5):175– 180, May 31, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Levin:2006:RTS

[Lev06] Asaf Levin. Real time scheduling with a budget: Parametricsearch is better than binary search. Information Processing Letters, 99(5):187–191, September 15, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Levin:2007:FHI

[Lev07] Asaf Levin. The finite horizon investor problem with a budget constraint. Information Processing Letters, 104(1): 21–28, September 30, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Laber:2002:IBA

[LH02] Eduardo Sany Laber and Leonardo Gomes Holanda. Improved bounds for asymmetric communication protocols. Information Processing Letters, 83 (4):205–209, August 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:2003:MMC

[LH03] SingLing Lee and Hann-Jang Ho. On minimizing the maximum congestion for Weighted Hypergraph Embedding in a Cycle. Information Processing Letters, 87(5):271–275, September 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2004:SAD

[LH04] Ying Li and Yibin Hou. Search audio data with the wavelet pyramidal algorithm. Information Processing Letters, 91(1): 49–55, July 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lorencz:2005:SFA

[LH05] Róbert Lórencz and Josef Hlaváč. Subtraction-free Almost Montgomery Inverse algorithm. Information Processing Letters, 94(1):11–14, April 15, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lai:2008:TED

[LH08] Pao-Lien Lai and Hong-Chun Hsu. The two-equal-disjoint path cover problem of Matching Composition Network. Information Processing Letters, 107 (1):18–23, June 30, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2000:STJ

[LHC^{+}00] Zichen Li, L. C. K. Hui, K. P. Chow, C. F. Chong, W. W. Tsang, and H. W. Chan. Security of Tseng–Jan’s group signature schemes. Information Processing Letters, 75(5):187– 189, October 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/30/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/30/25/article.
pdf.

Li:2001:SWA

[LHC^{+}01] Z. C. Li, L. C. K. Hui, K. P. Chow, C. F. Chong, W. W. Tsang, and H. W. Chan. Security of Wang et al.’s grouporiented (t, n) threshold signature schemes with traceable signers. Information Processing Letters, 80(6):295–298, December 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/84/37/
30/abstract.html.

Lee:2005:IAM

[LHC^{+}05] Chuan-Min Lee, Ling-Ju Hung, Maw-Shang Chang, Chia-Ben Shen, and Chuan-Yi Tang. An improved algorithm for the maximum agreement subtree problem. Information Processing Letters, 94(5):211–216, June 15, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:2002:EAH

[LHK02] Dong-Ho Lee, Shin Heu, and Hyoung-Joo Kim. An efficient algorithm for hyperspherical range query processing in high-dimensional data space. Information Processing Letters, 83 (2):115–123, July 31, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2008:LEL

[LHL08] Bin Liu, Jianfeng Hou, and Guizhen Liu. List edge and list total colorings of planar graphs without short cycles. Information Processing Letters, 108(6): 347–351, November 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:2001:SBA

[LHS01] Minkyu Lee, Dongsoo Han, and Jaeyong Shim. Set-based access conflict analysis of concurrent workflow definition. Information Processing Letters, 80(4): 189–194, November 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/84/34/29/abstract.html.

Lu:2000:GSP

[LHT00] Bor-Kuan Lu, Fang-Rong Hsu, and Chuan Yi Tang. Guarding in a simple polygon. Information Processing Letters, 75(4):153– 158, September 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/29/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/29/26/article.
pdf.

Liu:2007:EDRb

[LHWL07] J. J. Liu, G. S. Huang, Y. L. Wang, and R. C. T. Lee. Edit distance for a run-length-encoded string and an uncompressed string. Information Processing Letters, 105(1):12– 16, December 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2000:ERS

[Li00] Chu Min Li. Equivalency reasoning to solve a class of hard SAT problems. Information Processing Letters, 76(1–2):75– 81, November 20, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/24/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/24/35/article.
pdf.

Li:2006:NSC

[Li06] Rao Li. A new sufficient condition for Hamiltonicity of graphs. Information Processing Letters, 98(4):159–161, May 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2009:UBP

[Li09] Jian Li. An O(log n/ log log n)

upper bound on the price of stability for undirected Shapley network design games. Information Processing Letters, 109(15): 876–878, July 16, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Libkin:2003:CRC

[Lib03] Leonid Libkin. A collapse result for constraint queries over structures of small degree. Information Processing Letters, 86 (5):277–281, June 15, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liberatore:2006:CEC

[Lib06] Paolo Liberatore. On the complexity of extension checking in default logic. Information Processing Letters, 98(2):61–65, April 30, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liedloff:2008:FDS

[Lie08] Mathieu Liedloff. Finding a dominating set on bipartite graphs. Information Processing Letters, 107(5):154–157, August 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lifshits:2003:LBS

[Lif03] Yuri Lifshits. A lower bound on the size of ϵ-free NFA corresponding to a regular expression. Information Processing Letters, 85(6):293–299, March 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lins:2002:EAC

[Lin02] Rafael Dueire Lins. An efficient algorithm for cyclic reference counting. Information Processing Letters, 83(3):145–150, August 16, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:2007:FSA

[Lin07] Min-Sheng Lin. Fast and simple algorithms to count the number of vertex covers in an interval graph. Information Processing Letters, 102(4):143–146, May 16, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:2008:CMN

[Lin08a] Jyhjong Lin. A conceptual model for negotiating in serviceoriented environments. Information Processing Letters, 108(4): 192–203, October 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lins:2008:CRC

[Lin08b] Rafael Dueire Lins. Cyclic reference counting. Information Processing Letters, 109(1):71– 78, December 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Litow:2003:IFB

[Lit03a] B. Litow. Inequality of finite behaviors of rational weight finite automata is in R. Information Processing Letters, 87(3): 139–145, August 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Litow:2003:NCM

[Lit03b] B. Litow. A note on commutative multivariate rational series. Information Processing Letters, 87(6):283–285, September 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2004:NPL

[Liu04a] Ding Liu. A note on point location in arrangements of hyperplanes. Information Processing Letters, 90(2):93–95, April 30, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2004:SLB

[Liu04b] Ding Liu. A strong lower bound for approximate nearest neighbor searching. Information Processing Letters, 92(1):23–29, October 16, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2005:MDT

[Liu05] Pangfeng Liu. Minimum degree triangulation for rectangular domains. Information Processing Letters, 96(4):128–135, November 30, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Livne:2009:NPC

[Liv09] Noam Livne. A note on #pcompleteness of NP-witnessing relations. Information Processing Letters, 109(5):259–261, February 15, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:2003:CHS

[LJ03] Li Lin and Yunfei Jiang. The computation of hitting sets: Review and new algorithms. Information Processing Letters, 86 (4):177–184, May 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2009:CLW

[LJ09] GuanJun Liu and ChangJun Jiang. On conditions for the liveness of weakly persistent nets. Information Processing Letters, 109(16):967–970, July 31, 2009.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lucier:2007:ACA

[LJL07] Brendan Lucier, Tao Jiang, and Ming Li. Average-case analysis of QuickSort and Binary Insertion Tree height using incompressibility. Information Processing Letters, 103(2):45–51, July 16, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:2000:REP

[LK00] Jenn-Wei Lin and Sy-Yen Kuo.

Resolving error propagation in distributed systems. Information Processing Letters, 74(5–6): 257–262, June 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/28/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/28/36/article.
pdf.

Lee:2002:ENN

[LK02] Dong-Ho Lee and Hyoung-Joo Kim. An efficient nearest neighbor search in highdimensional data spaces. Information Processing Letters, 81 (5):239–246, March 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/85/36/28/abstract.html.

Lee:2000:DSS

[LKL^{+}00] Ju-Hong Lee, Deok-Hwan Kim, Seok-Lyong Lee, Chin-Wan Chung, and Guang-Ho Cha. Distributed similarity search algorithm in distributed heterogeneous multimedia databases. Information Processing Letters, 75 (1–2):35–42, July 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/25/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/25/31/article.
pdf.

Lee:2001:PRI

[LKP01] Han Wook Lee, Dae Woong Kim, and Chan-Ik Park. Phased RGSS: An improved disk array scheduling for continuous media retrieval. Information Processing Letters, 80(5):225– 231, December 15, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/84/36/
28/abstract.html.

Lee:2002:DSS

[LKY02] Keon-Jik Lee, Kee-Won Kim, and Kee-Young Yoo. Digitserial-in-serial-out systolic multiplier for Montgomery algorithm. Information Processing Letters, 82(2):65–71, April 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/86/33/27/abstract.html.

Li:2001:EVC

[LL01] Xuandong Li and Johan Lilius. Efficient verification of a class of time Petri nets using linear programming. Information Processing Letters, 77(5–6): 219–224, March 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/32/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/32/24/article.
pdf.

Le:2002:NCS

[LL02a] Hoàng-Oanh Le and Van Bang Le. The NP-completeness of (1, r)-subcolorability of cubic graphs. Information Processing Letters, 81(3):157–162, February 14, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/34/
33/abstract.html.

Lefebvre:2002:CLL

[LL02b] Arnaud Lefebvre and Thierry Lecroq. Compror: On-line lossless data compression with a factor oracle. Information Processing Letters, 83(1):1–6, July 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lingas:2009:EAA

[LL09] Andrzej Lingas and Eva-Marta Lundell. Efficient approximation algorithms for shortest cycles in undirected graphs. Information Processing Letters, 109(10):493– 498, April 30, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lan:2009:IUL

[LLC09] James K. Lan, Victor W. Liu, and Chiuyuan Chen. Improved upper and lower bounds on the optimization of mixed chordal ring networks. Information Processing Letters, 109(13):757– 762, June 15, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lu:2009:IEK

[LLH09] Xianhui Lu, Xuejia Lai, and Dake He. Improved efficiency of Kiltz07-KEM. Information Processing Letters, 109(18):1093– 1096, August 31, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lim:2008:DFA

[LLK08] Sung-Hwa Lim, Byoung-Hoon Lee, and Jai-Hoon Kim. Diversity and fault avoidance for dependable replication systems. Information Processing Letters, 108(1):33–37, September 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:2009:NAA

[LLP09a] Kangbok Lee, Joseph Y.- T. Leung, and Michael L. Pinedo. A note on “An approximation algorithm for the load-balanced semi-matching problem in weighted bipartite graphs”. Information Processing Letters, 109(12):608–610, May 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Low06].

Lee:2009:NGB

[LLP09b] Kangbok Lee, Joseph Y.-T. Leung, and Michael L. Pinedo. A note on graph balancing problems with restrictions. Information Processing Letters, 110(1): 24–29, December 1, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leung:2007:MTW

[LLPZ07] Joseph Y.-T. Leung, Haibing Li, Michael Pinedo, and Jiawei Zhang. Minimizing total weighted completion time when scheduling orders in a flexible environment with uniform machines. Information Processing Letters, 103(3):119–129, July 31, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Latapy:2002:CDL

[LM02] Matthieu Latapy and Clémence Magnien. Coding distributive lattices with Edge Firing Games. Information Processing Letters, 83(3):125–128, August 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lenzi:2004:FAI

[LM04a] Giacomo Lenzi and Erich Monteleone. On Fixpoint Arithmetic and Infinite Time Turing Machines. Information Processing Letters, 91(3):121–128, August 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liskiewicz:2004:NLU

[LM04b] Maciej Liśkiewicz and Bodo Manthey. New lower and upper bounds for the competitive ratio of transmission protocols. Information Processing Letters, 89(6):297–301, March 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lesh:2006:BSH

[LM06] N. Lesh and M. Mitzenmacher. Bubblesearch: A simple heuristic for improving priority-based greedy algorithms. Information Processing Letters, 97(4):161– 169, February 28, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Laue:2008:AHM

[LM08a] Sören Laue and Domagoj Matijević. Approximating k-hop minimum spanning trees in Euclidean metrics. Information Processing Letters, 107(3–4):96– 101, July 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2008:RCG

[LM08b] Rui Li and Jixiang Meng. Reversals Cayley graphs of symmetric groups. Information Processing Letters, 109(2):130–132, December 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2008:SCS

[LM08c] Juan Liu and Jixiang Meng.

Super-connected and super-arc-connected Cartesian product of digraphs. Information Processing Letters, 108(3):90–93, October 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lozin:2009:MIS

[LM09] Vadim Lozin and Raffaele Mosca. Maximum independent sets in subclasses of P _{5}-free graphs. Information Processing Letters, 109(6):319–324, February 28, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2009:CFS

[LMC^{+}09] Hai Liu, Lizhuang Ma, Xuan Cai, Zhihua Chen, and Yang Shen. A closed-form solution to video matting of natural snow. Information Processing Letters, 109(18):1097–1104, August 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leenen:2007:RSC

[LMG07] Louise Leenen, Thomas Meyer, and Aditya Ghose. Relaxations of semiring constraint satisfaction problems. Information Processing Letters, 103(5):177–182, August 31, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lucas:2005:OTC

[LMM05] Salvador Lucas, Claude Marché, and José Meseguer. Operational termination of conditional term rewriting systems. Information Processing Letters, 95(4): 446–453, August 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lesh:2004:EAR

[LMMM04] N. Lesh, J. Marks, A. McMahon, and M. Mitzenmacher. Exhaustive approaches to 2D rectangular perfect packings. Information Processing Letters, 90(1):7– 14, April 15, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lanotte:2004:TCT

[LMST04] Ruggero Lanotte, Andrea Maggiolo-Schettini, and Simone Tini. ϵtransitions in Concurrent Timed Automata. Information Processing Letters, 89(1):1–7, January 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leprevost:2005:GAE

[LMVV05] Franck Leprévost, Jean Monnerat, Sébastien Varrette, and Serge Vaudenay. Generating anomalous elliptic curves. Information Processing Letters, 93 (5):225–230, March 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liazi:2008:CAA

[LMZ08] Maria Liazi, Ioannis Milis, and Vassilis Zissimopoulos. A constant approximation algorithm for the densest k-subgraph problem on chordal graphs. Information Processing Letters, 108(1): 29–32, September 15, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Loding:2001:EMD

[Löd01] Christof Löding. Efficient minimization of deterministic weak omega-automata. Information Processing Letters, 79(3):105– 109, July 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/32/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/80/32/25/article.
pdf.

Long:2003:UBS

[Lon03] Philip M. Long. An upper bound on the sample complexity of PAC-learning halfspaces with respect to the uniform distribution. Information Processing Letters, 87(5):229–234, September 15, 2003. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Long:2008:MMD

[Lon08] Brad Long. Managing module dependencies to facilitate continuous testing. Information Processing Letters, 108(3):127– 131, October 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lopriore:2000:PSA

[Lop00] Lanfranco Lopriore. Protection in a single-address-space environment. Information Processing Letters, 76(1–2):25–32, November 20, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/67/24/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/67/24/28/article.
                                        pdf.

Low:2002:ERS

[Low02] Chor Ping Low. An efficient retrieval selection algorithm for video servers with random duplicated assignment storage technique. Information Processing Letters, 83(6):315–321, September 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Low:2006:AAL

[Low06]  Chor Ping Low.  An ap-

proximation algorithm for the load-balanced semi-matching problem in weighted bipartite graphs. Information Processing Letters, 100(4):154–161, November 30, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See note [LLP09a].

Lozin:2002:MIM

[Loz02] V. V. Lozin. On maximum induced matchings in bipartite graphs. Information Processing Letters, 81(1):7–11, January 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/85/27/28/abstract.html.

Lozin:2005:BC

[Loz05] Vadim V. Lozin. Between 2- and 3-colorability. Information Processing Letters, 94(4):179–182, May 31, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lipsky:2008:AMM

[LP08a] Ohad Lipsky and Ely Porat. Approximate matching in the L_{∞} metric. Information Processing Letters, 105(4):138–140, February 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lipsky:2008:PML

[LP08b] Ohad Lipsky and Ely Porat. L_{1} pattern matching lower bound. Information Processing Letters, 105(4):141–143, February 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luccio:2009:FSB

[LP09] Fabrizio Luccio and Linda Pagli. The Fermat star of binary trees. Information Processing Letters, 109(11):568–571, May 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:2002:SAS

[LPC02] Jae-Ha Lee, Sang-Min Park, and Kyung-Yong Chwa. Simple algorithms for searching a polygon with flashlights. Information Processing Letters, 81(5): 265–270, March 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/85/36/
                                        32/abstract.html.

Li:2005:MCK

[LPM05] Mingyan Li, Radha Poovendran, and David A. McGrew. Minimizing center key storage in hybrid one-way function based group key management with communication constraints. Information Processing Letters, 93 (4):191–198, February 28, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lotker:2008:SRT

[LPSR08] Zvi Lotker, Boaz Patt-Shamir, and Dror Rawitz. Ski rental with two general options. Information Processing Letters, 108 (6):365–368, November 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lamport:2000:WDC

[LPW00] Leslie Lamport, Sharon Perl, and William Weihl. When does a correct mutual exclusion algorithm guarantee mutual exclusion? Information Processing Letters, 76(3):131– 134, December 15, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/27/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/27/29/article.
pdf.

Lingas:2014:CNC

[LPWZ14] Andrzej Lingas, Leonidas Palios, Agnieszka Wasylewicz, and Pawel Zyliński. Corrigendum to “Note on covering monotone orthogonal polygons” [Inf. Process. Lett. 104(6) (2007) 220–227]. Information Processing Letters, 114(11):646– 654, November 2014. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0020019014000593.

See [LWŻ07].

Lehman:2003:SEB

[LR03a] Meir M. Lehman and Juan F.

Ramil. Software evolution— -background, theory, practice. Information Processing Letters, 88(1–2):33–44, October 17, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lozin:2003:SRG

[LR03b] Vadim Lozin and Dieter Rautenbach. Some results on graphs without long induced paths. Information Processing Letters, 88 (4):167–171, November 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liebchen:2005:GAC

[LR05] Christian Liebchen and Romeo Rizzi. A greedy approach to compute a minimum cycle basis of a directed graph. Information Processing Letters, 94(3): 107–112, May 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2006:ICU

[LR06] Hong Liu and Kenneth W. Regan. Improved construction for universality of determinant and permanent. Information Processing Letters, 100(6):233–237, December 31, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lopez:2008:HAC

[LR08] Mario A. Lopez and Shlomo Reisner. Hausdorff approximation of 3D convex polytopes. Information Processing Letters, 107(2):76–82, July 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2002:NBB

[LS02] Xiao Yu Li and Matthias F.

Stallmann. New bounds on the barycenter heuristic for bipartite graph drawing. Information Processing Letters, 82(6): 293–298, June 30, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lange:2006:PDL

[LS06] Martin Lange and Rafał Somla.

Propositional dynamic logic of context-free programs and fixpoint logic with chop. Information Processing Letters, 100(2): 72–75, October 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Laroussinie:2007:SEA

[LS07a] François Laroussinie and Jeremy Sproston. State explosion in almost-sure probabilistic reachability. Information Processing Letters, 102(6):236–241, June 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:2007:FRC

[LS07b] Min Chih Lin and Jayme L.

Szwarcfiter. Faster recognition of clique-Helly and hereditary clique-Helly graphs. Information Processing Letters, 103(1): 40–43, June 30, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2009:GCR

[LS09a] Yue Li and Joe Sawada. Gray codes for reflectable languages.

Information Processing Letters, 109(5):296–300, February 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Linhart:2009:MDC

[LS09b] Chaim Linhart and Ron Shamir. Matching with don’t-cares and a small number of mismatches. Information Processing Letters, 109(5):273–277, February 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2006:IAF

[LSC06] Shisheng Li, Guangzhong Sun, and Guoliang Chen. Improved algorithm for finding next-to-shortest paths. Information Processing Letters, 99(5):192– 194, September 15, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2008:NAC

[LSL08] Banghe Li, Yuefeng Shen, and Bo Li. A new algorithm for computing the minimum Hausdorff distance between two point sets on a line under translation. Information Processing Letters, 106(2):52–58, April 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Long:2007:DLC

[LSS07] Philip M. Long, Rocco A. Servedio, and Hans Ulrich Simon. Discriminative learning can succeed where generative learning fails. Information Processing Letters, 103(4):131–135, August 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luan:2007:AAM

[LSY07] Shuang Luan, Jared Saia, and Maxwell Young. Approximation algorithms for minimizing segments in radiation therapy. Information Processing Letters, 101(6):239–244, March 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Landau:2003:SLC

[LSZU03] Gad M. Landau, Baruch Schieber, and Michal Ziv-Ukelson. Sparse LCS Common Substring Alignment. Information Processing Letters, 88(6): 259–270, December 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lu:2003:NUA

[LT03] Chang-Chun Lu and Shi-Chun Tsai. A note on unscrambling address lines. Information Processing Letters, 85(4):185–189, February 28, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2006:VML

[LT06] (Ben) P. C. Li and M. Toulouse.

Variations of the maximum leaf spanning tree problem for bipartite graphs. Information Processing Letters, 97(4):129–132, February 28, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2007:EDRa

[LT07] ChenGuang Liu and Kazuyuki Tanaka. Eigen-distribution on random assignments for game trees. Information Processing Letters, 104(2):73–77, October 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lai:2008:EFM

[LT08a] Chia-Jui Lai and Chang-Hsiung Tsai. Embedding a family of meshes into twisted cubes. Information Processing Letters, 108(5):326–330, November 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lai:2008:ECF

[LT08b] Chia-Jui Lai and Chang-Hsiung Tsai. On embedding cycles into faulty dual-cubes. Information Processing Letters, 109(2):147– 150, December 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:2005:SES

[LTH05] Kuo-Chu Lee, Chi-Hung Tzeng, and Shing-Tsaan Huang. A space-efficient self-stabilizing algorithm for measuring the size of ring networks. Information Processing Letters, 94(3):125–130, May 16, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:2008:EHP

[LTH08] Chung-Meng Lee, Jimmy J. M.

Tan, and Lih-Hsing Hsu. Embedding Hamiltonian paths in hypercubes with a required vertex in a fixed position. Information Processing Letters, 107(5): 171–176, August 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2001:SCT

[LTHS01] Tseng-Kuei Li, Jimmy J. M.

Tan, Lih-Hsing Hsu, and Ting-Yi Sung. The shuffle-cubes and their generalization. Information Processing Letters, 77(1): 35–41, January 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/24/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/24/30/article.
pdf.

Leu:2000:EES

[LTT00] Fang-Cheng Leu, Yin-Te Tsai, and Chuan Yi Tang. An efficient external sorting algorithm. Information Processing Letters, 75(4):159–163, September 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/29/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/29/27/article.
pdf.

Lin:2006:BMS

[LTT06] Chung-Ming Lin, Yin Te Tsai, and Chuan Yi Tang. Balancing minimum spanning trees and multiple-source minimum routing cost spanning trees on metric graphs. Information Processing Letters, 99(2):64–67, July 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2003:BEF

[LTTH03] Tseng-Kuei Li, Chang-Hsiung Tsai, Jimmy J. M. Tan, and Lih-Hsing Hsu. Bipanconnectivity and edge-fault-tolerant bipancyclicity of hypercubes. Information Processing Letters, 87 (2):107–110, July 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2003:SOR

[LTV03] Ming Li, John Tromp, and Paul Vitányi. Sharpening Occam’s razor. Information Processing Letters, 85(5):267–274, March 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leinders:2004:EPS

[LTV04] Dirk Leinders, Jerzy Tyszkiewicz, and Jan Van den Bussche. On the expressive power of semijoin queries. Information Processing Letters, 91(2):93–98, July 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lucas:2004:DAR

[Luc04a] Joan M. Lucas. A direct algorithm for restricted rotation distance. Information Processing Letters, 90(3):129–134, May 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lucas:2004:SNS

[Luc04b] Salvador Lucas. Strong and NVsequentiality of constructor systems. Information Processing Letters, 89(4):191–201, February 28, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lutz:2004:CVE

[Lut04] Jack H. Lutz. Computability versus exact computability of martingales. Information Processing Letters, 92(5):235–237, December 16, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Laguillaumie:2007:MDV

[LV07] Fabien Laguillaumie and Damien Vergnaud. Multi-designated verifiers signatures: anonymity without encryption. Information Processing Letters, 102(2– 3):127–132, April 30, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lingas:2005:NMI

[LW05] Andrzej Lingas and Martin Wahlen. A note on maximum independent set and related problems on box graphs. Information Processing Letters, 93(4):169– 171, February 28, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:2006:PTA

[LW06] Kang Li and Lusheng Wang. A polynomial time approximation scheme for embedding a directed hypergraph on a ring. Information Processing Letters, 97(5): 203–207, March 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:2008:EGM

[LW08] Hwei-Jen Lin and Hung-Hsuan Wu. Efficient geometric measure of music similarity. Information Processing Letters, 109(2):116– 120, December 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:2009:TAP

[LW09] Wanwei Liu and Ji Wang. A tighter analysis of Piterman’s Büchi determinization. Information Processing Letters, 109(16): 941–945, July 31, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lu:2001:DAT

[LWC01] Erl-Huei Lu, Shao-Wei Wu, and Yi-Chang Cheng. A decoding algorithm for triple-error-correcting binary BCH codes. Information Processing Letters, 80(6):299–303, December 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/84/37/31/abstract.html.

Lingas:2007:NCM

[LWŻ07] Andrzej Lingas, Agnieszka Wasylewicz, and Paweł Żyliński. Note on covering monotone orthogonal polygons with starshaped polygons. Information Processing Letters, 104(6):220– 227, December 16, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See corrigendum [LPWZ14].

Lin:2002:TST

[LX02] Guohui Lin and Guoliang Xue.

On the terminal Steiner tree problem. Information Processing Letters, 84(2):103–107, October 31, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:2000:LSM

[LY00] Keon-Jik Lee and Kee-Young Yoo. Linear systolic multiplier/squarer for fast exponentiation. Information Processing Letters, 76(3):105–111, December 15, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/27/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/27/26/article.
pdf.

Lin:2007:BMC

[LY07] Cho-Chin Lin and Hao-Yun Yin. Bounds on the multi-clients incremental computing for homogeneous decreasing computation sequences. Information Processing Letters, 102(1):35–40, April 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lyon:2002:CTC

[Lyo02] Gordon Lyon. Comparison of two code scalability tests. Information Processing Letters, 81 (3):169–174, February 14, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/gej-ng/10/23/
                                        20/85/34/35/abstract.html.

Lee:2002:PMC

[LYP02] Oukseh Lee, Kwangkeun Yi, and Yunheung Paek. A proof method for the correctness of modularized 0CFA. Information Processing Letters, 81(4):179– 185, February 28, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/85/35/
                                        28/abstract.html.

Lange:2004:FLI

[LZ04] Steffen Lange and Sandra Zilles. Formal language identification: query learning vs. Gold-style learning. Information Processing Letters, 91(6):285–292, September 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liberty:2009:MAN

[LZ09] Edo Liberty and Steven W.

Zucker. The Mailman algorithm: A note on matrix–vector multiplication. Information Processing Letters, 109(3):179–182, January 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Laoutaris:2004:JOP

[LZS04] Nikolaos Laoutaris, Vassilios Zissimopoulos, and Ioannis Stavrakakis. Joint object placement and node dimensioning for Internet content distribution. Information Processing Letters, 89(6):273–279, March 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moffat:2006:BCL

[MA06] Alistair Moffat and Vo Ngoc Anh. Binary codes for locally homogeneous sequences. Information Processing Letters, 99(5):175–180, September 15, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maass:2007:CSL

[Maa07] Moritz G. Maaß. Computing suffix links for suffix trees and arrays. Information Processing Letters, 101(6):250–254, March 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Min:2003:EES

[MAC03] Jun-Ki Min, Jae-Yong Ahn, and Chin-Wan Chung. Efficient extraction of schemas for XML documents. Information Processing Letters, 85(1):7–12, January 16, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Madey:2003:WPP

[Mad03] Jan Madey. WMT—-from a personal perspective. Information Processing Letters, 88(1–2):3–6, October 17, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Madry:2005:DEC

[Mad05] Aleksander Madry. Data exchange: On the complexity of answering queries with inequalities. Information Processing Letters, 94(6):253–257, June 30, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Madsen:2006:AES

[Mad06] Bolette Ammitzbøll Madsen. An algorithm for Exact Satisfiability analysed with the number of clauses as parameter. Information Processing Letters, 97(1): 28–30, January 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maffray:2009:SSC

[Maf09] Frédéric Maffray. Stable sets in k-colorable P_{5}-free graphs. Information Processing Letters, 109(23–24):1235–1237, November 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maitra:2002:HNB

[Mai02] Subhamoy Maitra. Highly nonlinear balanced Boolean functions with good local and global avalanche characteristics. Information Processing Letters, 83(5):281–286, September 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maibaum:2003:WEG

[Mai03] Tom Maibaum. On what exactly goes on when software is developed step-by-step, II: The sequel. Information Processing Letters, 88(1–2):45–51, October 17, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makinen:2000:ILS

[Mäk00] Erkki Mäkinen. On inferring linear single-tree languages. Information Processing Letters, 73(1– 2):1–3, January 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/62/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/62/25/25/article.
pdf.

Minetti:2008:SSR

[MAL08] Gabriela Minetti, Enrique Alba, and Gabriel Luque. Seeding strategies and recombination operators for solving the DNA fragment assembly problem. Information Processing Letters, 108(3):94–100, October 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manthey:2005:NAW

[Man05] Bodo Manthey. Non-approximability of weighted multiple sequence alignment for arbitrary metrics. Information Processing Letters, 95(3):389–395, August 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Martin:2001:TEC

[Mar01]  Alain J. Martin.  Towards

an energy complexity of computation. Information Processing Letters, 77(2–4):181– 187, February 28, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/
42/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/29/42/article.
pdf.

Martel:2002:ECP

[Mar02] Chip Martel. The expected complexity of Prim’s minimum spanning tree algorithm. Information Processing Letters, 81(4):197– 201, February 28, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.

com/gej-ng/10/23/20/85/35/
31/abstract.html.

Marx:2004:LEM

[Mar04] Dániel Marx. List edge multicoloring in graphs with few cycles. Information Processing Letters, 89(2):85–90, January 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Martin:2008:PNR

[Mar08] Álvar Ibeas Martín. On the period of the Naor–Reingold sequence. Information Processing Letters, 108(5):304–307, November 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Masopust:2008:DCM

[Mas08]  Tomáš Masopust.  Descrip-

tional complexity of multiparallel grammars. Information Processing Letters, 108(2):68– 70, September 30, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mayordomo:2002:KCC

[May02] Elvira Mayordomo. A Kolmogorov complexity characterization of constructive Hausdorff dimension. Information Processing Letters, 84(1):1–3, October 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Majumder:2008:DDP

[MB08] Subhashis Majumder and Bhargab B. Bhattacharya. On the density and discrepancy of a 2D point set with applications to thermal analysis of VLSI chips. Information Processing Letters, 107(5):177–182, August 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Malpani:2002:NPC

[MC02a] Navneet Malpani and Jianer Chen. A note on practical construction of maximum bandwidth paths. Information Processing Letters, 83(3):175–180, August 16, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mukherjee:2002:PTO

[MC02b] Maharaj Mukherjee and Kanad Chakraborty. A polynomialtime optimization algorithm for a rectilinear partitioning problem with applications in VLSI design automation. Information Processing Letters, 83(1):41–48, July 16, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Misiolek:2006:TFN

[MC06] Ewa Misiołek and Danny Z. Chen. Two flow network simplification algorithms. Information Processing Letters, 97(5): 197–202, March 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

McCartin:2001:IAJ

[McC01] Catherine McCartin. An improved algorithm for the jump number problem. Information Processing Letters, 79(2): 87–92, June 30, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/31/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/80/31/30/article.
pdf.

McIlroy:2001:MS

[McI01] M. Douglas McIlroy. The music of streams. Information Processing Letters, 77(2–4):189– 195, February 28, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/
43/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/29/43/article.
pdf.

Mukhopadhyay:2006:AFS

[MCL06] Asish Mukhopadhyay, Samidh Chatterjee, and Benjamin Lafreniere. On the All-Farthest-Segments problem for a planar set of points. Information Processing Letters, 100(3):120–123, November 15, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moreno:2009:TDU

[MCS^{+}09] A. Moreno, E. Cesar, J. Sorribes, T. Margalef, and E. Luque. Task distribution using factoring load balancing in Master– Worker applications. Information Processing Letters, 109(16): 902–906, July 31, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

McKellar:2000:RRM

[MD00] Steven McKellar and Robert Davis. Redundancy removal in multicast protocols. Information Processing Letters, 73(5–6): 169–173, March 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/62/28/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/62/28/29/article.
                                        pdf.

Montagne:2004:OSF

[ME04] Eurípides Montagne and Anand Ekambaram. An optimal storage format for sparse matrices. Information Processing Letters, 90 (2):87–92, April 30, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Meer:2007:SAV

[Mee07] Klaus Meer. Simulated Annealing versus Metropolis for a TSP instance. Information Processing Letters, 104(6):216–219, December 16, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Melkonian:2007:FDN

[Mel07a] Vardges Melkonian. Flows in dynamic networks with aggregate arc capacities. Information Processing Letters, 101(1):

30–35, January 16, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Melkonian:2007:LBS

[Mel07b] Vardges Melkonian. LP-based solution methods for the asymmetric TSP. Information Processing Letters, 101(6):233–238, March 31, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mestre:2006:MRC

[Mes06] Julián Mestre. On the multiradius cover problem. Information Processing Letters, 99(5): 195–198, September 15, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Marti-Farre:2007:NSS

[MF07] Jaume Martí-Farré. A note on secret sharing schemes with three homogeneous access structure. Information Processing Letters, 102(4):133–137, May 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Martin:2001:SND

[MG01] Clare Martin and Jeremy Gibbons. On the semantics of nested datatypes. Information Processing Letters, 80(5):233– 238, December 15, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/84/36/
29/abstract.html.

Mosko:2006:FIW

[MGLA06] Marc Mosko and J. J. Garcia-Luna-Aceves. Fraction interpolation walking a Farey tree. Information Processing Letters, 98 (1):19–23, April 15, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ma:2009:SPH

[MH09] Daegun Ma and Jin Hong. Success probability of the Hellman trade-off. Information Processing Letters, 109(7):347–351, March 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mehedy:2007:IDB

[MHK07] Lenin Mehedy, M. Kamrul Hasan, and Mohammad Kaykobad. An improved degree based condition for Hamiltonian cycles. Information Processing Letters, 102(2–3):108–112, April 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Miao:2003:CCR

[Mia03]  Ying Miao.  A combinato-

rial characterization of regular anonymous perfect threshold schemes. Information Processing Letters, 85(3):131–135, February 14, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Miettinen:2008:PNP

[Mie08]  Pauli Miettinen.  On the

Positive–Negative Partial Set Cover problem. Information Processing Letters, 108(4):219– 221, October 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Millen:2003:FD

[Mil03] Jonathan Millen. On the freedom of decryption. Information Processing Letters, 86(6): 329–333, June 30, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Minsker:2005:LTA

[Min05] Steven Minsker. The Little Towers of Antwerpen problem. Information Processing Letters, 94 (5):197–201, June 15, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Minkevicius:2007:IOC

[Min07a] Saulius Minkevičius. The impact of overload conditions on computer network reliability. Information Processing Letters, 104 (1):32–35, September 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Minnameier:2007:LGD

[Min07b] Christoph Minnameier. Local and global deadlock-detection in component-based systems are NP-hard. Information Processing Letters, 103(3):105–111, July 31, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Misra:2001:WSP

[Mis01] Jayadev Misra. A walk over the shortest path: Dijkstra’s Algorithm viewed as fixed-point computation. Information Processing Letters, 77(2–4):197– 200, February 28, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/
44/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/29/44/article.
pdf.

Misra:2003:DPS

[Mis03] Jayadev Misra. Derivation of a parallel string matching algorithm. Information Processing Letters, 85(5):255–260, March 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Miyazaki:2008:ACG

[Miy08] Takunari Miyazaki. On the asymmetric complexity of the group-intersection problem. Information Processing Letters, 107(5):188–193, August 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Meduna:2002:HGR

[MK02] A. Meduna and D. Kolář. Homogeneous grammars with a reduced number of non-context-free productions. Information Processing Letters, 81(5):253– 257, March 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/36/
30/abstract.html.

Michiels:2003:CMM

[MK03] Wil Michiels and Jan Korst.

Complexity of min–max subsequence problems. Information Processing Letters, 87(4): 213–217, August 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Min:2005:PBN

[MK05] Kyung-Sub Min and Hyoung-Joo Kim. A path-based node filtering method for efficient structural joins. Information Processing Letters, 95(5):480–486, September 15, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mukhopadhyay:2008:ISP

[MKGB08] Asish Mukhopadhyay, Chanchal Kumar, Eugene Greene, and Binay Bhattacharya. On intersecting a set of parallel line segments with a convex polygon of minimum area. Information Processing Letters, 105(2):58–64, January 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ma:2009:CEE

[ML09] Meijie Ma and Baodong Liu. Cycles embedding in exchanged hypercubes. Information Processing Letters, 110(2):71–76, December 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ma:2008:SCA

[MLX08] Meijie Ma, Guizhen Liu, and Jun-Ming Xu. The super connectivity of augmented cubes. Information Processing Letters, 106(2):59–63, April 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See note [MTXL09].

Manolios:2001:DMC

[MM01] Panagiotis Manolios and J. Strother Moore. On the desirability of mechanizing calculational proofs. Information Processing Letters, 77(2–4):173–179, February 28, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/
41/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/29/41/article.
pdf.

McIver:2005:EPH

[MM05] Annabelle McIver and Carroll Morgan. An elementary proof that Herman’s Ring is Θ(N^{2}). Information Processing Letters, 94(2):79–84, April 30, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manea:2007:ANP

[MM07a] Florin Manea and Victor Mitrana. All NP-problems can be solved in polynomial time by accepting hybrid networks of evolutionary processors of constant size. Information Processing Letters, 103(3):112–118, July 31, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Masopust:2007:DCS

[MM07b] Tomáš Masopust and Alexander Meduna. Descriptional complexity of semi-conditional grammars. Information Processing Letters, 104(1):29–31, September 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Massart:2008:CPO

[MMV08] Thierry Massart, Cédric Meuter, and Laurent Van Begin. On the complexity of partial order trace model checking. Information Processing Letters, 106(3): 120–126, April 30, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maass:2005:NMA

[MN05] Moritz G. Maaß and Johannes Nowak. A new method for approximate indexing and dictionary lookup with one error. Information Processing Letters, 96 (5):185–191, December 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mahajan:2008:RSE

[MN08] Meena Mahajan and Jayalal Sarma M. N. Rigidity of a simple extended lower triangular matrix. Information Processing Letters, 107(5):149–153, August 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maity:2004:CCF

[MNR04] Soumen Maity, Amiya Nayak, and Bimal K. Roy. Characterization of catastrophic faults in two-dimensional reconfigurable systolic arrays with unidirectional links. Information Processing Letters, 92(4):189–197, November 30, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Meijer:2009:ACS

[MNRR09] Henk Meijer, Yurai Núñez-Rodríguez, and David Rappaport. An algorithm for computing simple k-factors. Information Processing Letters, 109(12): 620–625, May 31, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mosheiov:2006:SMS

[MO06] Gur Mosheiov and Daniel Oron. Single machine scheduling with batch-dependent setup times. Information Processing Letters, 98(2):73–78, April 30, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Monnot:2001:MFD

[Mon01] Jérôme Monnot. The maximum f-depth spanning tree problem. Information Processing Letters, 80(4):179–187, November 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/84/34/28/abstract.html.

Monnot:2002:DAR

[Mon02] Jérôme Monnot. Differential approximation results for the traveling salesman and related problems. Information Processing Letters, 82(5):229–235, June 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Monnot:2005:LPM

[Mon05] Jérôme Monnot. The labeled perfect matching in bipartite graphs. Information Processing Letters, 96(3):81–88, November 15, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Montassier:2006:NCS

[Mon06] Mickaël Montassier. A note on the not 3-choosability of some families of planar graphs. Information Processing Letters, 99 (2):68–71, July 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Montoya:2008:PCP

[Mon08] J. Andrés Montoya. The parameterized complexity of probability amplification. Information Processing Letters, 109(1): 46–53, December 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moreno:2005:BSB

[Mor05] Eduardo Moreno. De Bruijn sequences and De Bruijn graphs for a general language. Information Processing Letters, 96 (6):214–219, December 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Morosan:2006:NBS

[Mor06] Calin D. Morosan. On the number of broadcast schemes in networks. Information Processing Letters, 100(5):188–193, December 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Muller-Olm:2004:CPP

[MOS04] Markus Müller-Olm and Helmut Seidl. Computing polynomial program invariants. Information Processing Letters, 91(5): 233–244, September 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moser:2008:RBM

[Mos08a]  Philippe Moser.  Resource-

bounded measure on probabilistic classes. Information Processing Letters, 106(6):241–245, June 15, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moss:2008:CM

[Mos08b] Lawrence S. Moss. Confusion of memory. Information Processing Letters, 107(3–4):114–119, July 31, 2008. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Mitsuhashi:2006:RRD

[MOY06] Ichiro Mitsuhashi, Michio Oyamaguchi, and Toshiyuki Yamada. The reachability and related decision problems for monadic and semi-constructor TRSs. Information Processing Letters, 98(6):219–224, June 30, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moura:2002:NVA

[MP02] Arnaldo V. Moura and Guilherme A. Pinto. A note on the verification of automata specifications of probabilistic real-time systems. Information Processing Letters, 82(5):223–228, June 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Melancon:2004:GCA

[MP04] Guy Melançon and Fabrice Philippe. Generating connected acyclic digraphs uniformly at random. Information Processing Letters, 90(4):209–213, May 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mathieu:2008:DLB

[MP08a] Claire Mathieu and Charalampos Papamanthou. Distortion lower bounds for line embeddings. Information Processing Letters, 108(4):175–178, October 31, 2008. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Mitchell:2008:MPE

[MP08b] Joseph S. B. Mitchell and Valentin Polishchuk. Minimumperimeter enclosures. Information Processing Letters, 107(3– 4):120–124, July 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Majewski:2009:AET

[MP09] Krzysztof Majewski and Nicholas Pippenger. Attribute estimation and testing quasi-symmetry. Information Processing Letters, 109(4):233–237, January 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mehlhorn:2002:APS

[MPSS02] Kurt Mehlhorn, Volker Priebe, Guido Schäfer, and Naveen Sivadasan. All-pairs shortestpaths computation in the presence of negative cycles. Information Processing Letters, 81 (6):341–343, March 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/gej-ng/10/23/
                                        20/85/37/35/abstract.html.

Myrvold:2001:RUPa

[MR01a] Wendy Myrvold and Frank Ruskey. Ranking and unranking permutations in linear time. Information Processing Letters, 79(6):281–284, September 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/80/37/31/abstract.html.

Myrvold:2001:RUPb

[MR01b] Wendy Myrvold and Frank Ruskey. Ranking and unranking permutations in linear time. Information Processing Letters, 79(6):281–284, September 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/80/37/31/abstract.html.

Montassier:2006:NFC

[MR06] Mickaël Montassier and André Raspaud. A note on 2-facial coloring of plane graphs. Information Processing Letters, 98 (6):235–241, June 30, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mohamed:2008:LEP

[MR08] Adnan Mohamed and R. S. Ramakrishna. Linear election in pancake graphs. Information Processing Letters, 106(3):127– 131, April 30, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Marx:2009:CRF

[MR09a] Dániel Marx and Igor Razgon.

Constant ratio fixed-parameter approximation of the edge multicut problem. Information Processing Letters, 109(20):1161– 1166, September 30, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Massberg:2009:BTC

[MR09b] Jens Maßberg and Dieter Rautenbach. Binary trees with choosable edge lengths. Information Processing Letters, 109(18): 1087–1092, August 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mousavi:2005:SCF

[MRG05] MohammadReza Mousavi, Michel Reniers, and Jan Friso Groote. A syntactic commutativity format for SOS. Information Processing Letters, 93(5):217–223, March 16, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maity:2002:ECF

[MRN02] Soumen Maity, Bimal K. Roy, and Amiya Nayak. On enumeration of catastrophic fault patterns. Information Processing Letters, 81(4):209–212, February 28, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/85/35/
                                        33/abstract.html.

Mantaci:2003:BWT

[MRS03] S. Mantaci, A. Restivo, and M. Sciortino. Burrows–Wheeler transform and Sturmian words. Information Processing Letters, 86(5):241–246, June 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Morig:2009:LBC

[MRST09] Marc Mörig, Dieter Rautenbach, Michiel Smid, and Jan Tusch. An Ω(n log n) lower bound for computing the sum of evenranked elements. Information Processing Letters, 109(16):955– 956, July 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mostefaoui:2000:BCM

[MRT00] Achour Mostefaoui, Michel Raynal, and Frédéric Tronel. From Binary Consensus to Multivalued Consensus in asynchronous message-passing systems. Information Processing Letters, 73(5–6):207–212, March 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/62/28/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/62/28/34/article.
pdf.

Montassier:2008:RHC

[MRWW08] Mickaël Montassier, André Raspaud, Weifan Wang, and Yingqian Wang. A relaxation of Havel’s 3-color problem. Information Processing Letters, 107 (3–4):107–109, July 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Markowitch:2003:CWV

[MS03] Olivier Markowitch and Shahrokh Saeednia. Cryptanalysis of the Wu–Varadhrajan fair exchange protocol. Information Processing Letters, 87(3):169–171, August 16, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Markey:2004:PCM

[MS04] N. Markey and Ph. Schnoebelen. A PTIME-complete matching problem for SLP-compressed words. Information Processing Letters, 90(1):3–6, April 15, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Markey:2006:MCP

[MS06] Nicolas Markey and Philippe Schnoebelen. Mu-calculus path checking. Information Processing Letters, 97(6):225–230, March 31, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Muscholl:2006:CDT

[MSS06] Anca Muscholl, Mathias Samuelides, and Luc Segoufin. Complementing deterministic tree-walking automata. Information Processing Letters, 99(1):33–39, July 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mizuki:2009:ORS

[MSS09] Takaaki Mizuki, Takuya Sato, and Hideaki Sone. A oneround secure message broadcasting protocol through a key sharing tree. Information Processing Letters, 109(15):842–845, July 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Massuthe:2008:CFP

[MSSW08] Peter Massuthe, Alexander Serebrenik, Natalia Sidorova, and Karsten Wolf. Can I find a partner? Undecidability of partner existence for open nets. Information Processing Letters, 108(6):374–378, November 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maggiolo-Schettini:2002:DLT

[MST02] Andrea Maggiolo-Schettini and Simone Tini. On disjunction of literals in triggers of statecharts transitions. Information Processing Letters, 84(6):305–310, December 31, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Metivier:2002:RLE

[MSZ02] Yves Métivier, Nasser Saheb, and Akka Zemmari. Randomized local elections. Information Processing Letters, 82(6): 313–320, June 30, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makino:2002:SMA

[MTF02] Kazuhisa Makino, Takashi Takabatake, and Satoru Fujishige. A simple matching algorithm for regular bipartite graphs. Information Processing Letters, 84 (4):189–193, November 30, 2002.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makris:2003:RMM

[MTT03] Christos Makris, Athanasios Tsakalidis, and Kostas Tsichlas. Reflected min-max heaps. Information Processing Letters, 86 (4):209–214, May 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ma:2009:NSC

[MTXL09] Meijie Ma, Xuegong Tan, Jun-Ming Xu, and Guizhen Liu. A note on “The super connectivity of augmented cubes”. Information Processing Letters, 109(12): 592–593, May 31, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [MLX08].

Monroy:2005:GTT

[MU05] Ruy Fabila Monroy and J. Urrutia. Graham triangulations and triangulations with a center are Hamiltonian. Information Processing Letters, 93(6):295–299, March 31, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mitra:2004:ASS

[MUK04] Pradipta Prometheus Mitra, Muhammad Arshad Ul Abedin, and Md. Abul Kashem. Algorithms for solving the symmetry number problem on trees. Information Processing Letters, 91 (4):163–169, August 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mulzer:2009:NPS

[Mul09] Wolfgang Mulzer. A note on predecessor searching in the pointer machine model. Information Processing Letters, 109(13):726– 729, June 15, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mamut:2008:VVP

[MV08] Aygul Mamut and Elkin Vumar.

Vertex vulnerability parameters of Kronecker products of complete graphs. Information Processing Letters, 106(6):258–262, June 15, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Martin-Vide:2001:SUP

[MVM01] Carlos Martín-Vide and Victor Mitrana. Some undecidable problems for parallel communicating finite automata systems. Information Processing Letters, 77(5–6):239– 245, March 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/32/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/32/27/article.
pdf.

Martin-Vide:2007:RAM

[MVM07] Carlos Martín-Vide and Victor Mitrana. Remarks on arbitrary multiple pattern interpretations.

Information Processing Letters, 101(5):209–214, March 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Meidl:2003:LCP

[MW03] Wilfried Meidl and Arne Winterhof. On the linear complexity profile of explicit nonlinear pseudorandom numbers. Information Processing Letters, 85(1): 13–18, January 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Min:2006:DCN

[MW06] Chen Min and Wang Weifan. The 2-dipath chromatic number of Halin graphs. Information Processing Letters, 99(2):47–53, July 31, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ma:2009:SOD

[MY09] Bin Ma and Hongyi Yao. Seed optimization for i.i.d. similarities is no easier than optimal Golomb ruler design. Information Processing Letters, 109(19): 1120–1124, September 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mandoiu:2000:NMH

[MZ00] Ion I. Măndoiu and Alexander Z. Zelikovsky. A note on the MST heuristic for bounded edge-length Steiner trees with minimum number of Steiner points. Information Processing Letters, 75(4):165– 167, September 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/29/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/29/28/article.
pdf.

Nagamochi:2005:CMC

[Nag05a] Hiroshi Nagamochi. On computing minimum (s, t)-cuts in digraphs. Information Processing Letters, 93(5):231–237, March 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nagy:2005:LES

[Nag05b] Benedek Nagy. On the language equivalence of NE star-patterns. Information Processing Letters, 95(3):396–400, August 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nagoya:2009:NDA

[Nag09] Takayuki Nagoya. New differential approximation algorithm for k-customer vehicle routing problem. Information Processing Letters, 109(8):405–408, March 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Naumann:2001:CSA

[Nau01] David A. Naumann. Calculating sharp adaptation rules.

Information Processing Letters, 77(2–4):201–208, February 28, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/
45/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/29/45/article.
pdf.

Navarro:2005:NBA

[NB05] Gonzalo Navarro and Nieves Brisaboa. New bounds on D-ary optimal codes. Information Processing Letters, 96(5):178–184, December 16, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Narayanaswamy:2007:NHN

[NBCS07] N. S. Narayanaswamy, N. Belkale, L. S. Chandran, and N. Sivadasan. A note on the Hadwiger number of circular arc graphs. Information Processing Letters, 104(1):10–13, September 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Noore:2005:MRL

[NC05] A. Noore and P. L. Cross. Modeling the reliability of large dynamic distributed nonhomogeneous networks. Information Processing Letters, 93 (2):57–61, January 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ng:2002:TSP

[NCBJ02] C. T. Ng, T. C. E. Cheng, A. Bachman, and A. Janiak. Three scheduling problems with deteriorating jobs to minimize the total completion time. Information Processing Letters, 81 (6):327–333, March 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/85/37/33/abstract.html.

Ng:2002:SNH

[NCY02] C. T. Ng, T. C. E. Cheng, and J. J. Yuan. Strong NPhardness of the single machine multi-operation jobs total completion time scheduling problem. Information Processing Letters, 82(4):187–191, May 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ngo:2002:NBH

[NDG02] Hung Quang Ngo, Ding-Zhu Du, and Ronald L. Graham. New bounds on a hypercube coloring problem. Information Processing Letters, 84(5):265–269, December 16, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nebel:2007:LGC

[Neb07] Markus E. Nebel. On the lexicographical generation of compressed codes. Information Processing Letters, 104(3):95–100, October 31, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nock:2003:REP

[NEK03] Richard Nock, Tapio Elomaa, and Matti Kääriäinen. Reduced Error Pruning of branching programs cannot be approximated to within a logarithmic factor. Information Processing Letters, 87(2):73–78, July 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nguyen:2005:CHR

[NG05] Linh Anh Nguyen and Rajeev Goré. Completeness of hyperresolution via the semantics of disjunctive logic programs. Information Processing Letters, 95 (2):363–369, July 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nagy-Gyorgy:2008:OHC

[NGI08] J. Nagy-György and Cs. Imreh. Online hypergraph coloring. Information Processing Letters, 109(1):23–26, December 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nagamochi:2001:MCS

[NII01] Hiroshi Nagamochi, Toshimasa Ishii, and Hiro Ito. Minimum cost source location problem with vertex-connectivity requirements in digraphs. Information Processing Letters, 80 (6):287–293, December 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/84/37/29/abstract.html.

Nikolopoulos:2000:RCT

[Nik00] Stavros D. Nikolopoulos. Recognizing cographs and threshold graphs through a classification of their edges. Information Processing Letters, 74(3–4): 129–139, May 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/63/27/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/63/27/31/article.
pdf.

Nikolik:2006:TSO

[Nik06] Borislav Nikolik. Test suite oscillations. Information Processing Letters, 98(2):47–55, April 30, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nivasch:2004:CDU

[Niv04] Gabriel Nivasch. Cycle detection using a stack. Information Processing Letters, 90(3): 135–140, May 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nagamochi:2007:MCS

[NK07] Hiroshi Nagamochi and Yoko Kamidoi. Minimum cost subpartitions in graphs. Information Processing Letters, 102(2–3):79– 84, April 30, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Na:2009:FLC

[NKS09] Joong Chae Na, Dong Kyue Kim, and Jeong Seop Sim. Finding the longest common nonsuperstring in linear time. Information Processing Letters, 109(18):1066–1070, August 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Na:2009:ILC

[NKSK09] Joong Chae Na, Namhee Kim, Jeong Seop Sim, and Dong Kyue Kim. Improving on-line construction of two-dimensional suffix trees for square matrices. Information Processing Letters, 109(10):504–508, April 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nielsen:2005:FDS

[NN05] Frank Nielsen and Richard Nock. A fast deterministic smallest enclosing disk approximation algorithm. Information Processing Letters, 93(6): 263–268, March 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nielsen:2008:SEI

[NN08] Frank Nielsen and Richard Nock. On the smallest enclosing information disk. Information Processing Letters, 105(3): 93–97, January 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nielson:2007:WFN

[NNP07] Flemming Nielson, Hanne Riis Nielson, and Henrik Pilegaard. What is a free name in a process algebra? Information Processing Letters, 103(5):188–194, August 31, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nikolopoulos:2004:LCN

[NNR04] S. D. Nikolopoulos, C. Nomikos, and P. Rondogiannis. A limit characterization for the number of spanning trees of graphs. Information Processing Letters, 90 (6):307–313, June 30, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nielson:2001:KLE

[NNS01] Flemming Nielson, Hanne Riis Nielson, and Mooly Sagiv. Kleene’s logic with equality. Information Processing Letters, 80 (3):131–137, November 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/84/35/29/abstract.html.

Nagamochi:2007:AMR

[NO07] Hiroshi Nagamochi and Kohei Okada. Approximating the minmax rooted-tree cover in a tree. Information Processing Letters, 104(5):173–178, November 30, 2007. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Novotny:2009:MRR

[NON09] Sarah Novotny, Juan Ortiz, and Darren A. Narayan. Minimal k-rankings and the rank number of P_{n}^{2}. Information Processing Letters, 109(3):193–198, January 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nordstrom:2009:SWP

[Nor09] Jakob Nordström. A simplified way of proving trade-off results for resolution. Information Processing Letters, 109(18):1030– 1035, August 31, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nikolopoulos:2000:PFF

[NP00] Stavros D. Nikolopoulos and Charis Papadopoulos. On the performance of the first-fit coloring algorithm on permutation graphs. Information Processing Letters, 75(6):265–273, November 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/31/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/31/29/article.
pdf.

Namneh:2007:FSF

[NP07] Rami Al Na’mneh and W. David Pan. Five-step FFT algorithm with reduced computational complexity. Information Processing Letters, 101(6):262– 267, March 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nardelli:2001:FCM

[NPW01] Enrico Nardelli, Guido Proietti, and Peter Widmayer. A faster computation of the most vital edge of a shortest path. Information Processing Letters, 79 (2):81–85, June 30, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/31/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/80/31/29/article.
pdf.

Nomikos:2001:RPM

[NPZ01] Christos Nomikos, Aris Pagourtzis, and Stathis Zachos. Routing and path multicoloring. Information Processing Letters, 80 (5):249–256, December 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/84/36/32/abstract.html.

Niedermeier:2000:GMS

[NR00] Rolf Niedermeier and Peter Rossmanith. A general method to speed up fixed-parameter-tractable algorithms. Information Processing Letters, 73(3–4):125–129, February 29, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/62/27/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/62/27/32/article.
                                        pdf.

Nutov:2008:AMS

[NR08] Zeev Nutov and Daniel Reichman. Approximating maximum satisfiable subsystems of linear equations of bounded width. Information Processing Letters, 106(5):203–207, May 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nayak:2000:ITS

[NRS00] A. Nayak, J. Ren, and N. Santoro. An improved testing scheme for catastrophic fault patterns. Information Processing Letters, 73(5–6):199– 206, March 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/62/28/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/62/28/33/article.
                                        pdf.

Nomikos:2009:SEL

[NRW09] Christos Nomikos, Panos Rondogiannis, and William W. Wadge. Strong equivalence of logic programs under the infinite-valued semantics. Information Processing Letters, 109 (11):576–581, May 16, 2009.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Naroska:2002:LSP

[NS02] Edwin Naroska and Uwe Schwiegelshohn.

On an on-line scheduling problem for parallel jobs. Information Processing Letters, 81(6): 297–304, March 31, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/37/
28/abstract.html.

Nagamochi:2004:SRM

[NSI04] Hiroshi Nagamochi, Takahisa Suzuki, and Toshimasa Ishii. A simple recognition of maximal planar graphs. Information Processing Letters, 89(5):223–226, March 16, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Narayanan:2006:PRM

[NSR06] Arvind Narayanan, K. Srinathan, and C. Pandu Rangan. Perfectly Reliable Message Transmission. Information Processing Letters, 100(1): 23–28, October 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Neumann:2002:CWA

[NSW02a] Jakub Neumann, Andrzej Szepietowski, and Igor Walukiewicz. Complexity of weak acceptance conditions in tree automata. Information Processing Letters, 84 (4):181–187, November 30, 2002.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Noga:2002:FLA

[NSW02b] John Noga, Steve Seiden, and Gerhard J. Woeginger. A faster off-line algorithm for the TCP acknowledgement problem. Information Processing Letters, 81 (2):71–73, January 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/85/33/29/abstract.html.

Ng:2000:SDA

[NSZ00] Joseph Kee-Yin Ng, Shibin Song, and Wei Zhao. Statistical delay analysis on an ATM switch with self-similar input traffic. Information Processing Letters, 74(3–4):163– 173, May 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/63/27/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/63/27/35/article.
pdf.

Niehren:2000:RCC

[NTT00] Joachim Niehren, Sophie Tison, and Ralf Treinen. On rewrite constraints and context unification. Information Processing Letters, 74(1– 2):35–40, April 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/25/30/article.
pdf.

Nunes:2004:MRE

[Nun04] I. Nunes. Method redefinition— ensuring alternative behaviors. Information Processing Letters, 92(6):279–285, December 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nutov:2009:NRS

[Nut09] Zeev Nutov. A note on Rooted Survivable Networks. Information Processing Letters, 109(19): 1114–1119, September 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nielsen:2002:LSP

[NWZ02] Benny K. Nielsen, Pawel Winter, and Martin Zachariasen. On the location of Steiner points in uniformly-oriented Steiner trees. Information Processing Letters, 83(5):237–241, September 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nagamochi:2004:CEC

[NY04a] Hiroshi Nagamochi and Nobuyasu Yamada. Counting edge crossings in a 2-layered drawing. Information Processing Letters, 91 (5):221–225, September 2004.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nishimura:2004:PTQ

[NY04b] Harumichi Nishimura and Tomoyuki Yamakami. Polynomial time quantum computation with advice. Information Processing Letters, 90(4):195–204, May 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nutov:2009:WND

[NY09] Zeev Nutov and Ariel Yaroshevitch. Wireless network design via 3-decompositions. Information Processing Letters, 109(19): 1136–1140, September 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Okun:2003:NAA

[OB03] Michael Okun and Amnon Barak. A new approach for approximating node deletion problems. Information Processing Letters, 88(5):231–236, December 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ostrovsky-Berman:2005:TMR

[OB05] Yaron Ostrovsky-Berman. The transportation metric and related problems. Information Processing Letters, 95(4):461– 465, August 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ochem:2004:OCT

[Och04] Pascal Ochem. Oriented colorings of triangle-free planar graphs. Information Processing Letters, 92(2):71–76, October 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ogata:2003:FMK

[OF03] Kazuhiro Ogata and Kokichi Futatsugi. Flaw and modification of the i KP electronic payment protocols. Information Processing Letters, 86(2):57–62, April 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ohlebusch:2000:UFT

[Ohl00]  Enno Ohlebusch.  A uni-

form framework for term and graph rewriting applied to combined systems. Information Processing Letters, 73(1–2):53– 59, January 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/62/25/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/62/25/34/article.
pdf.

Ohlebusch:2002:HTR

[Ohl02] Enno Ohlebusch. Hierarchical termination revisited. Information Processing Letters, 84(4): 207–214, November 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Olaverri:2002:MSV

[OHNT02] Alfredo García Olaverri, Ferran Hurtado, Marc Noy, and Javier Tejel. On the minimum size of visibility graphs. Information Processing Letters, 81 (4):223–230, February 28, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/gej-ng/10/23/
                                        20/85/35/35/abstract.html.

Oh:2004:HCA

[OK04] Seung-Joon Oh and Jae-Yearn Kim. A hierarchical clustering algorithm for categorical sequence data. Information Processing Letters, 91(3):135–140, August 16, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Okhotin:2003:HLC

[Okh03] Alexander Okhotin. The hardest linear conjunctive language. Information Processing Letters, 86 (5):247–253, June 15, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Okubo:2009:NDC

[Oku09] Fumiya Okubo. A note on the descriptional complexity of semi-conditional grammars. Information Processing Letters, 110(1):36–40, December 1, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ochem:2008:OCP

[OP08] Pascal Ochem and Alexandre Pinlou. Oriented colorings of partial 2-trees. Information Processing Letters, 108(2):82– 86, September 30, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Obradovic:2007:EDC

[OPR07] Nenad Obradović, Joseph Peters, and Goran Ružić. Efficient domination in circulant graphs with two chord lengths. Information Processing Letters, 102 (6):253–258, June 15, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

ODwyer:2005:SNN

[OR05] Carl O’Dwyer and Daniel Richardson. Spiking neural nets with symbolic internal state. Information Processing Letters, 95(6):529–536, September 30, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Orlov:2009:ORN

[Orl09] Michael Orlov. Optimized random number generation in an interval. Information Processing Letters, 109(13):722–725, June 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ogryczak:2003:MSL

[OT03] Wlodzimierz Ogryczak and Arie Tamir. Minimizing the sum of the k largest functions in linear time. Information Processing Letters, 85(3):117–122, February 14, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Padhye:2005:PKP

[Pad05] Sahadeo Padhye. Partial known plaintext attack on Koyama scheme. Information Processing Letters, 96(3):96–100, November 15, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Paeng:2003:SCU

[Pae03] Seong-Hun Paeng. On the security of cryptosystem using automorphism groups. Information Processing Letters, 88(6): 293–298, December 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pallo:2000:EUB

[Pal00] Jean Pallo. An efficient upper bound of the rotation distance of binary trees. Information Processing Letters, 73(3–4): 87–92, February 29, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/62/27/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/62/27/26/article.
pdf.

Pallo:2003:GBT

[Pal03a] Jean Pallo. Generating binary trees by Glivenko classes on Tamari lattices. Information Processing Letters, 85(5): 235–238, March 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pallo:2003:RAR

[Pal03b] Jean Marcel Pallo. Right-arm rotation distance between binary trees. Information Processing Letters, 87(4):173–177, August 31, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pallo:2009:WAR

[Pal09] Jean Marcel Pallo. Weak associativity and restricted rotation. Information Processing Letters, 109(10):514–517, April 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pandurangan:2005:SRA

[Pan05] Gopal Pandurangan. On a simple randomized algorithm for finding a 2-factor in sparse graphs. Information Processing Letters, 95(1):321–327, July 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Papakonstantinou:2009:NWP

[Pap09] Periklis A. Papakonstantinou.

A note on width-parameterized SAT: An exact machine-model characterization. Information Processing Letters, 110(1):8–12, December 1, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Parnas:2003:SPM

[Par03] David Lorge Parnas. Structured programming: A minor part of software engineering. Information Processing Letters, 88(1–2): 53–58, October 17, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Parhami:2005:HSO

[Par05] Behrooz Parhami. The Hamiltonicity of swapped (OTIS) networks built of Hamiltonian component networks. Information Processing Letters, 95(4):441– 445, August 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Parhami:2008:ISB

[Par08] Behrooz Parhami. On isomorphisms and similarities between generalized Petersen networks and periodically regular chordal rings. Information Processing Letters, 107(6):246–251, August 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Park:2011:CGR

[Par11] Moonju Park. Comments on “Generalized rate monotonic schedulability bounds using relative period ratios”. Information Processing Letters, 111(7): 334–337, March 1, 2011. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [WLLS08].

Patthak:2002:QCT

[PBD^{+}02] A. C. Patthak, I. Bhattacharya, A. Dasgupta, Pallab Dasgupta, and P. P. Chakrabarti. Quantified Computation Tree Logic. Information Processing Letters, 82(3):123–129, May 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/86/34/28/abstract.html.

Park:2000:CEM

[PC00] Ho-Hyun Park and Chin-Wan Chung. Complexity of estimating multi-way join result sizes for area skewed spatial data. Information Processing Letters, 76(3):121–129, December 15, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/27/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/27/28/article.
pdf.

Park:2004:HPC

[PC04] Chong-Dae Park and Kyung-Yong Chwa. Hamiltonian properties on the class of hypercubelike networks. Information Processing Letters, 91(1):11–17, July 16, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pan:2005:IPN

[PC05a] Jun-Jie Pan and Gerard J.

Chang. Isometric-path numbers of block graphs. Information Processing Letters, 93(2): 99–102, January 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Punnen:2005:BM

[PC05b] Abraham P. Punnen and Olena Chapovska. The bottleneck k-MST. Information Processing Letters, 95(5):512–517, September 15, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Petkovic:2004:MFS

[PĆB04] Tatjana Petković, Miroslav Ćirić, and Stojan Bogdanović. Minimal forbidden subwords. Information Processing Letters, 92(5):211–218, December 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Park:2003:ESA

[PCC03] Yongsu Park, Tae-Sun Chung, and Yookun Cho. An efficient stream authentication scheme using tree chaining. Information Processing Letters, 86(1):1– 8, April 15, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pai:2008:NIU

[PCW08] Kung-Jui Pai, Jou-Ming Chang, and Yue-Li Wang. A Note on “An improved upper bound on the queuenumber of the hypercube”. Information Processing Letters, 108(3):107–109, October 16, 2008. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic). See [HH07].

Pai:2009:UBQ

[PCW09] Kung-Jui Pai, Jou-Ming Chang, and Yue-Li Wang. Upper bounds on the queuenumber of k-ary n-cubes. Information Processing Letters, 110(2):50– 56, December 16, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Panda:2003:LTR

[PD03] B. S. Panda and Sajal K. Das.

A linear time recognition algorithm for proper interval graphs. Information Processing Letters, 87(3):153–161, August 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Panda:2009:PAG

[PD09] B. S. Panda and Sajal K.

Das. A parallel algorithm for generating bicompatible elimination orderings of proper interval graphs. Information Processing Letters, 109(18):1041–1046, August 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Peczarski:2006:ITC

[Pec06] Marcin Peczarski. An improvement of the tree code construction. Information Processing Letters, 99(3):92–95, August 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Peczarski:2007:FJA

[Pec07] Marcin Peczarski. The Ford– Johnson algorithm still unbeaten for less than 47 elements. Information Processing Letters, 101(3):126–128, February 14, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Penna:2002:ATT

[Pen02] Paolo Penna. On the approximability of two tree drawing conventions. Information Processing Letters, 82(5):237–242, June 15, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pestov:2000:GSS

[Pes00] Vladimir Pestov. On the geometry of similarity search: Dimensionality curse and concentration of measure. Information Processing Letters, 73(1–2): 47–51, January 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/62/25/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/62/25/33/article.
pdf.

Pessoa:2008:NCE

[Pes08] Artur Alves Pessoa. A note on the construction of error detecting/correcting prefix codes. Information Processing Letters, 107(1):34–38, June 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Petersen:2002:BED

[Pet02] Holger Petersen. Bounds for the Element Distinctness Problem on one-tape Turing machines. Information Processing Letters, 81(2):75–79, January 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/85/33/30/abstract.html.

Petrovic:2005:SEF

[Pet05] Srdjan Petrovic. Space-efficient FCFS group mutual exclusion. Information Processing Letters, 95(2):343–350, July 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Petersen:2007:SMS

[Pet07] Holger Petersen. String matching with simple devices. Information Processing Letters, 105 (1):32–34, December 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Patil:2006:AGP

[PG06] Shailesh Patil and Vijay K.

Garg. Adaptive general perfectly periodic scheduling. Information Processing Letters, 98 (3):107–114, May 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Petersen:2009:RMR

[PG09] Holger Petersen and Szymon Grabowski. Range mode and range median queries in constant time and sub-quadratic space. Information Processing Letters, 109(4):225–228, January 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Phan:2004:IDC

[Pha04] Raphael C.-W. Phan. Impossible differential cryptanalysis of 7-round Advanced Encryption Standard (AES). Information Processing Letters, 91(1):33–38, July 16, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Puolamaki:2008:ARB

[PHG08] Kai Puolamäki, Sami Hanhijärvi, and Gemma C. Garriga. An approximation ratio for biclustering. Information Processing Letters, 108(2):45– 49, September 30, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Park:2001:RTE

[PHK01] Dong-Joo Park, Shin Heu, and Hyoung-Joo Kim. The RS-tree: An efficient data structure for distance browsing queries. Information Processing Letters, 80 (4):195–203, November 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/84/34/30/abstract.html.

Piotrow:2002:NCB

[Pio02] Marek Piotrów. A note on constructing binary heaps with periodic networks. Information Processing Letters, 83(3): 129–134, August 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Podgorelec:2002:EID

[PK02] Vili Podgorelec and Peter Kokol.

Evolutionary induced decision trees for dangerous software modules prediction. Information Processing Letters, 82(1): 31–38, April 15, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/86/27/
31/abstract.html.

Patkar:2001:NOCa

[PN01a] Sachin B. Patkar and H. Narayanan.

A note on optimal covering augmentation for graphic polymatroids. Information Processing Letters, 79(6):285–290, September 30, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/80/37/
32/abstract.html.

Patkar:2001:NOCb

[PN01b] Sachin B. Patkar and H. Narayanan.

A note on optimal covering augmentation for graphic polymatroids. Information Processing Letters, 79(6):285–290, September 30, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/80/37/
                                        32/abstract.html.

Poland:2003:FSM

[Pol03] Jan Poland. Finding smooth maps is NP-complete. Information Processing Letters, 85(5): 249–253, March 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Poland:2004:CTE

[Pol04] Jan Poland. A coding theorem for Enumerable Output Machines. Information Processing Letters, 91(4):157–161, August 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Poon:2003:VMS

[Poo03] C. K. Poon. Verifying minimum stable circuit values. Information Processing Letters, 86 (1):27–32, April 15, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Parker:2004:ACB

[PP04] Richard Parker and Andrew Plater. Addition chains with a bounded number of registers. Information Processing Letters, 90 (5):247–252, June 15, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pelc:2005:BLB

[PP05a] Andrzej Pelc and David Peleg. Broadcasting with locally bounded Byzantine faults. Information Processing Letters, 93 (3):109–115, February 14, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pietracaprina:2005:OMO

[PP05b] Andrea Pietracaprina and Geppino Pucci. Optimal many-to-one routing on the mesh with constant queues. Information Processing Letters, 96(1):24–29, October 16, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pinchinat:2005:DCP

[PR05] Sophie Pinchinat and Stéphane Riedweg. A decidable class of problems for control under partial observation. Information Processing Letters, 95(4): 454–460, August 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Premkumar:2000:OUA

[Pre00] A. B. Premkumar. On the optimal utilization of all available states in the 2^{n} moduli set. Information Processing Letters, 75 (1–2):51–56, July 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/25/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/25/33/article.
pdf.

Pretzel:2002:CSE

[Pre02] Oliver Pretzel. Characterization of simple edge-firing games. Information Processing Letters, 84 (5):235–236, December 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Procaccia:2008:NQC

[Pro08] Ariel D. Procaccia. A note on the query complexity of the Condorcet winner problem. Information Processing Letters, 108 (6):390–393, November 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Priami:2001:ASN

[PRSS01] Corrado Priami, Aviv Regev, Ehud Shapiro, and William Silverman. Application of a stochastic name-passing calculus to representation and simulation of molecular processes. Information Processing Letters, 80(1):25–31, October 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/gej-ng/10/23/
                                        20/84/27/30/abstract.html.

Padro:2002:LBI

[PS02] Carles Padró and Germán Sáez. Lower bounds on the information rate of secret sharing schemes with homogeneous access structure. Information Processing Letters, 83(6):345– 351, September 30, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pustylnik:2003:MSS

[PS03] Gennady Pustylnik and Micha Sharir. The Minkowski sum of a simple polygon and a segment. Information Processing Letters, 85(4):179–184, February 28, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Perst:2004:MFT

[PS04a] Thomas Perst and Helmut Seidl.

Macro forest transducers. Information Processing Letters, 89 (3):141–149, February 14, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pettie:2004:SLT

[PS04b] Seth Pettie and Peter Sanders.

A simpler linear time 2/3 − ϵ approximation for maximum weight matching. Information Processing Letters, 91(6):271– 276, September 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pinlou:2006:OVA

[PS06] Alexandre Pinlou and Éric Sopena. Oriented vertex and arc colorings of outerplanar graphs. Information Processing Letters, 100(3):97–104, November 15, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Polishchuk:2009:SLA

[PS09] Valentin Polishchuk and Jukka Suomela. A simple local 3approximation algorithm for vertex cover. Information Processing Letters, 109(12):642– 645, May 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Philippou:2001:HRC

[PSL^{+}01] Anna Philippou, Oleg Sokolsky, Insup Lee, Rance Cleaveland, and Scott A. Smolka. Hiding resources that can fail: An axiomatic perspective. Information Processing Letters, 80(1): 3–13, October 15, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/84/27/
28/abstract.html.

Penn:2000:CMO

[PT00] Michal Penn and Moshe Tennenholtz. Constrained multi-object auctions and b-matching. Information Processing Letters, 75 (1–2):29–34, July 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/25/30/article.
pdf.

Peng:2006:TAM

[PT06] Z. S. Peng and H. F. Ting. An O(n log n)-time algorithm for the maximum constrained agreement subtree problem for binary trees. Information Processing Letters, 100(4):137–144, November 30, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ponse:2001:ERS

[PU01] Alban Ponse and Yaroslav S.

Usenko. Equivalence of recursive specifications in process algebra. Information Processing Letters, 80(1):59–65, October 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/84/27/34/abstract.html.

Pudlak:2000:NUD

[Pud00] Pavel Pudlák. A note on the use of determinant for proving lower bounds on the size of linear circuits. Information Processing Letters, 74(5–6): 197–201, June 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/28/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/28/27/article.
pdf.

Pavan:2007:RPC

[PW07] A. Pavan and Fengming Wang.

Robustness of PSPACE-complete sets. Information Processing Letters, 103(3):102–104, July 31, 2007. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Paulus:2009:OOL

[PYZ09] Jacob Jan Paulus, Deshi Ye, and Guochuan Zhang. Optimal online-list batch scheduling. Information Processing Letters, 109(19):1125–1128, September 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Punnen:2009:BFU

[PZ09] Abraham P. Punnen and Ruonan Zhang. Bottleneck flows in unit capacity networks. Information Processing Letters, 109 (6):334–338, February 28, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Qureshi:2001:PAT

[QH01] Kalim Qureshi and Masahiko Hatanaka. A practical approach of task scheduling and load balancing on heterogeneous distributed raytracing systems. Information Processing Letters, 79 (2):65–71, June 30, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/80/31/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/80/31/27/article.
                                        pdf.

Quisquater:2005:SCC

[QPV05] Michaël Quisquater, Bart Preneel, and Joos Vandewalle.

Spectral characterization of cryptographic Boolean functions satisfying the (extended) propagation criterion of degree l and order k. Information Processing Letters, 93(1):25–28, January 16, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Quaglia:2002:RET

[Qua02] Francesco Quaglia. A restriction of the elastic time algorithm. Information Processing Letters, 83(5):243–249, September 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Qian:2006:HMP

[QW06] Jianbo Qian and Cao An Wang.

How much precision is needed to compare two sums of square roots of integers? Information Processing Letters, 100(5):194– 198, December 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Raffinot:2001:MRS

[Raf01] Mathieu Raffinot. On maximal repeats in strings. Information Processing Letters, 80 (3):165–169, November 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/84/35/34/abstract.html.

Rahn:2008:MDI

[Rah08] Mirko Rahn. More decidable instances of Post’s correspondence problem: beyond counting. Information Processing Letters, 106(3):115–119, April 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rampersad:2006:SC

[Ram06] Narad Rampersad. The state complexity of L^{2} and L^{k}. Information Processing Letters, 98 (6):231–234, June 30, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rampersad:2007:CFS

[Ram07] Narad Rampersad. On the context-freeness of the set of words containing overlaps. Information Processing Letters, 102(2–3):74–78, April 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rao:2001:RFM

[Rao01]  Josyula R. Rao.  On the

role of formal methods in security. Information Processing Letters, 77(2–4):209–212, February 28, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/68/29/
                                        46/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/68/29/46/article.
                                        pdf.

Rao:2002:TST

[Rao02] S. Srinivasa Rao. Time-space trade-offs for compressed suffix arrays. Information Processing Letters, 82(6):307–311, June 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rautenbach:2005:LBT

[Rau05]  Dieter Rautenbach.  Lower

bounds on treespan. Information Processing Letters, 96(2): 67–70, October 31, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ravelomanana:2007:APW

[Rav07]  Vlady Ravelomanana.  An-

other proof of Wright’s inequalities. Information Processing Letters, 104(1):36–39, September 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Razgon:2007:APC

[Raz07] Igor Razgon. A 2^{O(k)}poly(n)

algorithm for the parameterized Convex Recoloring problem. Information Processing Letters, 104(2):53–58, October 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rastsvetaev:2002:QCF

[RB02] A. L. Rastsvetaev and L. D.

Beklemishev. On the query complexity of finding a local maximum point. Information Processing Letters, 84(6):327–332, December 31, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Roy:2007:ETT

[RB07] Rahul Roy and Amitava Bagchi. On estimating the time taken to rectify faults in a software package during the system testing phase. Information Processing Letters, 104(5):168–172, November 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Roy:2009:NFH

[RBDN09] Sasanka Roy, Subhasis Bhattacharjee, Sandip Das, and Subhas C. Nandy. A new fast heuristic for labeling points. Information Processing Letters, 109(10): 478–484, April 30, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rijmen:2008:RSA

[RBF08] Vincent Rijmen, Paulo S. L. M. Barreto, and Décio L. Gazzoni Filho. Rotation symmetry in algebraically generated cryptographic substitution tables. Information Processing Letters, 106(6):246–250, June 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ryu:2006:DST

[RC06] Pum-Mo Ryu and Key-Sun Choi. Determining the specificity of terms using inside– outside information: a necessary condition of term hierarchy mining. Information Processing Letters, 100(2):76–82, October 31, 2006. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Ruiz:2007:GNC

[RCA07] D. Ruiz, R. Corchuelo, and J. L. Arjona. Generating nonconspiratorial executions. Information Processing Letters, 103 (4):157–162, August 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Regev:2002:PAM

[Reg02] Oded Regev. Priority algorithms for makespan minimization in the subset model. Information Processing Letters, 84 (3):153–157, November 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Remy:2004:RCS

[Rem04]  Jan Remy.  Resource con-

strained scheduling on multiple machines. Information Processing Letters, 91(4):177–182, August 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reznik:2005:ADA

[Rez05] Yuriy A. Reznik. On the average depth of asymmetric LCtries. Information Processing Letters, 96(3):106–113, November 15, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ren:2009:FCS

[RG09] Yanli Ren and Dawu Gu. Fully CCA2 secure identity based broadcast encryption without random oracles. Information Processing Letters, 109(11):527– 533, May 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Roy:2004:OAS

[RGDN04] Sasanka Roy, Partha P. Goswami, Sandip Das, and Subhas C. Nandy. Optimal algorithm for a special point-labeling problem. Information Processing Letters, 89(2):91–98, January 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rothe:2002:CEP

[RH02] Jörg Rothe and Lane A. Hemaspaandra. On characterizing the existence of partial oneway permutations. Information Processing Letters, 82(3): 165–171, May 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/86/34/
35/abstract.html.

Rick:2000:SFL

[Ric00]  Claus Rick.  Simple and

fast linear space computation of longest common subsequences. Information Processing Letters, 75(6):275–281, November 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/31/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/31/30/article.
pdf.

Richards:2002:NCC

[Ric02] Martin Richards. A note concerning the closest point pair algorithm. Information Processing Letters, 82(4):193–195, May 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Richomme:2003:SNF

[Ric03] G. Richomme. Some non finitely generated monoids of repetitionfree endomorphisms. Information Processing Letters, 85(2): 61–66, January 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rizzi:2003:RVA

[Riz03] Romeo Rizzi. On Rajagopalan and Vazirani’s 3/2-approximation bound for the Iterated 1-Steiner heuristic. Information Processing Letters, 86(6):335–338, June 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reddy:2003:RBT

[RK03] P. Krishna Reddy and Masaru Kitsuregawa. Reducing the blocking in two-phase commit with backup sites. Information Processing Letters, 86(1):39–47, April 15, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rahman:2005:CSI

[RK05a] M. Sohel Rahman and M. Kaykobad. Complexities of some interesting problems on spanning trees. Information Processing Letters, 94 (2):93–97, April 30, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rahman:2005:HCH

[RK05b] M. Sohel Rahman and M. Kaykobad. On Hamiltonian cycles and Hamiltonian paths. Information Processing Letters, 94(1):37–41, April 15, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rapaic:2009:TVP

[RK09] Milan R. Rapaić and Željko Kanović. Time-varying PSO — convergence analysis, convergencerelated parameterization and new parameter adjustment schemes. Information Processing Letters, 109(11):548–552, May 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Raidl:2002:ELS

[RL02] Günther R. Raidl and Ivana Ljubić. Evolutionary local search for the edge-biconnectivity augmentation problem. Information Processing Letters, 82 (1):39–45, April 15, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/86/27/
32/abstract.html.

Rombo:2009:OEM

[Rom09] Simona E. Rombo. Optimal extraction of motif patterns in 2D. Information Processing Letters, 109(17):1015–1020, August 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rosen:2002:NMN

[Ros02] Adi Rosén. A note on models for non-probabilistic analysis of packet switching networks. Information Processing Letters, 84 (5):237–240, December 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rothe:2003:ECE

[Rot03] Jörg Rothe. Exact complexity of Exact-Four-Colorability. Information Processing Letters, 87 (1):7–12, July 16, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ranjan:2003:LAP

[RP03] Desh Ranjan and Enrico Pontelli. The Level-Ancestor problem on Pure Pointer Machines. Information Processing Letters, 85(5):275–283, March 16, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reghizzi:2008:CPP

[RP08] Stefano Crespi Reghizzi and Matteo Pradella. A CKY parser for picture grammars. Information Processing Letters, 105(6): 213–217, March 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Roussel:2000:RTG

[RR00] F. Roussel and I. Rusu. Recognizing i-triangulated graphs in O(mn) time. Information Processing Letters, 76(3):141– 147, December 15, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/27/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/27/31/article.
pdf.

Radhakrishnan:2001:TBS

[RR01] Jaikumar Radhakrishnan and Venkatesh Raman. A tradeoff between search and update in dictionaries. Information Processing Letters, 80(5):243– 247, December 15, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/84/36/
31/abstract.html.

Rautenbach:2009:PSC

[RR09] Dieter Rautenbach and Friedrich Regen. On packing shortest cycles in graphs. Information Processing Letters, 109(14):816– 821, June 30, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Riera:2000:ECV

[RRB00] Andreu Riera, Josep Rifà, and Joan Borrell. Efficient construction of vote-tags to allow open objection to the tally in electronic elections. Information Processing Letters, 75(5):211– 215, October 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/30/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/30/28/article.
pdf.

Riege:2007:IEA

[RRSY07] Tobias Riege, Jörg Rothe, Holger Spakowski, and Masaki Yamamoto. An improved exact algorithm for the domatic number problem. Information Processing Letters, 101(3):101–106, February 14, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rajsbaum:2008:IAF

[RRT08] Sergio Rajsbaum, Michel Raynal, and Corentin Travers. An impossibility about failure detectors in the iterated immediate snapshot model. Information Processing Letters, 108(3): 160–164, October 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ron:2007:HED

[RRV07] Dana Ron, Amir Rosenfeld, and Salil Vadhan. The hardness of the Expected Decision Depth problem. Information Processing Letters, 101(3):112–118, February 14, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rao:1994:USR

[RRW94] Rajesh P. N. Rao, Jörg Rothe, and Osamu Watanabe. Upward separation for FewP and related classes. Information Processing Letters, 52(4):175–180, November 25, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See corrigendum [RRW00].

Rao:2000:CUS

[RRW00] Rajesh P. N. Rao, Jörg Rothe, and Osamu Watanabe. Corrigendum to “Upward separation for FewP and related classes”, Information Processing Letters 52(4) (1994) 175–180. Information Processing Letters, 74 (1–2):89, April 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/63/25/
                                        37/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/63/25/37/article.

pdf. See [RRW94].

Rajasekaran:2005:GPS

[RS05] Sanguthevar Rajasekaran and Sandeep Sen. A generalization of the 0–1 principle for sorting. Information Processing Letters, 94 (1):43–47, April 15, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Raja:2006:CLC

[RS06] N. Raja and R. K. Shyamasundar. A closer look at constraints as processes. Information Processing Letters, 98(5):206–210, June 15, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Raman:2007:IFP

[RS07a] Venkatesh Raman and Saket Saurabh. Improved fixed parameter tractable algorithms for two “edge” problems: MAX-CUT and MAXDAG. Information Processing Letters, 104(2): 65–72, October 16, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Raman:2007:PCI

[RS07b] Venkatesh Raman and Somnath Sikdar. Parameterized complexity of the induced subgraph problem in directed graphs. Information Processing Letters, 104(3):79–85, October 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reyzin:2007:LPA

[RS07c] Lev Reyzin and Nikhil Srivastava. On the longest path algorithm for reconstructing trees from distance matrices. Information Processing Letters, 101 (3):98–100, February 14, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rostamabadi:2008:OPR

[RSGK08] Farshad Rostamabadi, Iman Sadeghi, Mohammad Ghodsi, and Ramtin Khosravi. Optimal point removal in closed-2PM labeling. Information Processing Letters, 105(3):108–113, January 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rapaport:2008:MPI

[RST08] Ivan Rapaport, Karol Suchan, and Ioan Todinca. Minimal proper interval completions. Information Processing Letters, 106(5):195–202, May 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ramos:2007:EQT

[RSV07] J. Guadalupe Ramos, Josep Silva, and Germán Vidal. Ensuring the quasi-termination of needed narrowing computations. Information Processing Letters, 101(5):220–226, March 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ravelomanana:2000:PME

[RT00] Vlady Ravelomanana and Loÿs Thimonier. Patchworks and metablocks enumeration. Information Processing Letters, 73 (3–4):77–86, February 29, 2000.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.nl/gej-ng/10/23/
20/62/27/25/abstract.html;
http://www.elsevier.nl/gej-
ng/10/23/20/62/27/25/article.
pdf.

Ruggieri:2003:CSS

[Rug03] Salvatore Ruggieri. On computing the semi-sum of two integers. Information Processing Letters, 87(2):67–71, July 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Raspaud:2009:GCN

[RW09] Andre Raspaud and Jiaojiao Wu. Game chromatic number of toroidal grids. Information Processing Letters, 109(21–22): 1183–1186, October 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Roh:2008:EII

[RY08] Jong-Won Roh and Byoung-Kee Yi. Efficient indexing of interval time sequences. Information Processing Letters, 109(1):1–12, December 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ramnath:2000:IE

[RZ00] Sarnath Ramnath and Peiyi Zhao. On the isomorphism of expressions. Information Processing Letters, 74(3–4):97– 102, May 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/63/27/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/63/27/26/article.
pdf.

Saeednia:2000:SCG

[Sae00] Shahrokh Saeednia. On the security of a convertible group signature scheme. Information Processing Letters, 73(3–4):93– 96, February 29, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/62/27/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/62/27/27/article.
pdf.

Saeednia:2002:IBS

[Sae02] Shahrokh Saeednia. An identitybased society oriented signature scheme with anonymous signers. Information Processing Letters, 83(6):295–299, September 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See cryptanalysis [Sha03a].

Saeednia:2003:NGS

[Sae03] Shahrokh Saeednia. A note on Girault’s self-certified model. Information Processing Letters, 86(6):323–327, June 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sakai:2006:LSA

[Sak06]  Yoshifumi Sakai.  A linear

space algorithm for computing a longest common increasing subsequence. Information Processing Letters, 99(5):203– 207, September 15, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sanders:2001:SPP

[San01a] Beverly A. Sanders. The shortest path in parallel. Information Processing Letters, 77(2–4):213– 217, February 28, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/29/
47/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/29/47/article.
pdf.

Santhanam:2001:LBC

[San01b]  Rahul Santhanam.  Lower

bounds on the complexity of recognizing SAT by Turing machines. Information Processing Letters, 79(5):243–247, September 15, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/80/34/
31/abstract.html.

Sarbazi-Azad:2001:ACH

[SAOKM01] H. Sarbazi-Azad, M. Ould-Khaoua, and L. M. Mackenzie. Algorithmic construction of Hamiltonians in pyramids. Information Processing Letters, 80(2):75–79, October 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/gej-ng/10/23/
                                        20/80/37/28/abstract.html.

Sarkar:2000:NSC

[Sar00] Palash Sarkar. A note on the spectral characterization of correlation immune Boolean functions. Information Processing Letters, 74(5–6):191– 195, June 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/63/28/
                                        26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/63/28/26/article.
                                        pdf.

Sarkar:2008:GMS

[Sar08] Palash Sarkar. A general mixing strategy for the ECB-Mix-ECB mode of operation. Information Processing Letters, 109(2):121– 123, December 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sasaki:2002:TOD

[Sas02] Atsushi Sasaki. A time-optimal distributed sorting algorithm on a line network. Information Processing Letters, 83(1):21–26, July 16, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sasatte:2008:IFA

[Sas08] Prashant Sasatte. Improved FPT algorithm for feedback vertex set problem in bipartite tournament. Information Processing Letters, 105(3):79–82, January 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Savelli:2004:NDC

[Sav04] A. Savelli. On numerically decipherable codes and their homophonic partitions. Information Processing Letters, 90(3): 103–108, May 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sawitzki:2007:LBO

[Saw07] Daniel Sawitzki. Lower bounds on the OBDD size of two fundamental functions’ graphs. Information Processing Letters, 101 (2):66–71, January 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Saxena:2009:DMS

[Sax09] Sanjeev Saxena. Dominance made simple. Information Processing Letters, 109(9):419–421, April 15, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Srinivasan:2002:DBS

[SB02] Anand Srinivasan and Sanjoy Baruah. Deadline-based scheduling of periodic task systems on multiprocessors. Information Processing Letters, 84 (2):93–98, October 31, 2002.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shin:2004:SWC

[SB04] Yong H. Shin and Hyokyung Bahn. A scalable Web cache sharing scheme. Information Processing Letters, 91(5):227– 232, September 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sun:2006:DMA

[SB06] Zheng Sun and Tian-Ming Bu. On discretization methods for approximating optimal paths in regions with directiondependent costs. Information Processing Letters, 97(4):146– 152, February 28, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schneider:2000:DPS

[Sch00a] Markus Schneider. On distribution properties of sequences with perfect linear complexity profile. Information Processing Letters, 74(3–4):175– 182, May 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/63/27/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/63/27/36/article.
pdf.

Schonhage:2000:VCR

[Sch00b]  Arnold Schönhage.  Varia-

tions on computing reciprocals of power series. Information Processing Letters, 74 (1–2):41–46, April 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/25/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/25/31/article.
pdf.

Schnorr:2001:SGH

[Sch01] C. P. Schnorr. Small generic hardcore subsets for the discrete logarithm: Short secret DL-keys. Information Processing Letters, 79(2):93– 98, June 30, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/31/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/80/31/31/article.
pdf.

Schnoebelen:2002:VLC

[Sch02a] Ph. Schnoebelen. Verifying lossy channel systems has nonprimitive recursive complexity. Information Processing Letters, 83(5):251–261, September 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schwiebert:2002:TNO

[Sch02b] Loren Schwiebert. There is no optimal routing policy for the torus. Information Processing Letters, 83(6):331–336, September 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schoning:2006:SSD

[Sch06] Uwe Schöning. Smaller superconcentrators of density 28. Information Processing Letters, 98 (4):127–129, May 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schwarz:2008:OSS

[Sch08] Ulrich M. Schwarz. Online scheduling on semi-related machines. Information Processing Letters, 108(1):38–40, September 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schweitzer:2009:UIM

[Sch09] Pascal Schweitzer. Using the incompressibility method to obtain local lemma results for Ramsey-type problems. Information Processing Letters, 109 (4):229–232, January 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shih:2008:SMC

[SCHT08] Lun-Min Shih, Chieh-Feng Chiang, Lih-Hsing Hsu, and Jimmy J. M. Tan. Strong Menger connectivity with conditional faults on the class of hypercube-like networks. Information Processing Letters, 106(2):64–69, April 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shan:2007:AGB

[SCK07] Erfang Shan, T. C. E. Cheng, and Liying Kang. Absorbant of generalized de Bruijn digraphs. Information Processing Letters, 105(1):6–11, December 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Saha:2009:CSP

[SD09] Chandan Saha and Sandip Das.

Covering a set of points in a plane using two parallel rectangles. Information Processing Letters, 109(16):907–912, July 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shan:2009:TDN

[SDC09] Erfang Shan, Yanxia Dong, and Yukun Cheng. The twin domination number in generalized de Bruijn digraphs. Information Processing Letters, 109(15):856– 860, July 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Seager:2004:GAD

[Sea04] Suzanne M. Seager. The greedy algorithm for domination in graphs of maximum degree 3. Information Processing Letters, 89(2):53–56, January 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sedjelmaci:2007:JWA

[Sed07] Sidi Mohamed Sedjelmaci.

Jebelean–Weber’s algorithm without spurious factors. Information Processing Letters, 102 (6):247–252, June 15, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Segerlind:2005:ESB

[Seg05] Nathan Segerlind. Exponential separation between Res(k) and Res(k + 1) for k ≤ ϵ log n. Information Processing Letters, 93 (4):185–190, February 28, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Segal:2008:FAM

[Seg08] Michael Segal. Fast algorithm for multicast and data gathering in wireless networks. Information Processing Letters, 107 (1):29–33, June 30, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Seppanen:2005:UBA

[Sep05] Jouni K. Seppänen. Upper bound for the approximation ratio of a class of hypercube segmentation algorithms. Information Processing Letters, 93 (3):139–141, February 14, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Servedio:2001:LETa

[Ser01a] Rocco A. Servedio. On the limits of efficient teachability. Information Processing Letters, 79(6):267–272, September 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/80/37/29/abstract.html.

Servedio:2001:LETb

[Ser01b] Rocco A. Servedio. On the limits of efficient teachability. Information Processing Letters, 79(6):267–272, September 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/80/37/29/abstract.html.

Serre:2003:NWP

[Ser03] Olivier Serre. Note on winning positions on pushdown games with ω-regular conditions. Information Processing Letters, 85 (6):285–291, March 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Seynhaeve:2001:SCT

[Sey01] Franck Seynhaeve. Set constraints and topology. Information Processing Letters, 80(3): 145–150, November 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/84/35/31/abstract.html.

Srivastava:2003:DAF

[SG03] Saurabh Srivastava and R. K.

Ghosh. Distributed algorithms for finding and maintaining a k-tree core in a dynamic network. Information Processing Letters, 88(4):187–194, November 30, 2003. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Sezgin:2005:DSC

[SG05] Ali Sezgin and Ganesh Gopalakrishnan. On the definition of sequential consistency. Information Processing Letters, 96(6): 193–196, December 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shi:2004:ASS

[SGH04] Zhengnan Shi, Wayne Goddard, and Stephen T. Hedetniemi. An anonymous self-stabilizing algorithm for 1-maximal independent set in trees. Information Processing Letters, 91(2):77–83, July 31, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shao:2003:CIB

[Sha03a] Zuhua Shao. Cryptanalysis of “An identity-based society oriented signature scheme with anonymous signers”. Information Processing Letters, 86(6): 295–298, June 30, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Sae02].

Shao:2003:PSS

[Sha03b] Zuhua Shao. Proxy signature schemes based on factoring. Information Processing Letters, 85 (3):137–143, February 14, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shang:2009:CRI

[Sha09a] Yilun Shang. Connectivity in a random interval graph with access points. Information Processing Letters, 109(9):446–449, April 15, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shao:2009:SSC

[Sha09b] Zuhua Shao. Security of selfcertified signatures. Information Processing Letters, 109(20): 1147–1150, September 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sheu:2008:SDR

[SHC08] Jyh-Jian Sheu, Wen-Tzeng Huang, and Chin-Hsing Chen. Strong diagnosability of regular networks under the comparison model. Information Processing Letters, 106(1):19–25, March 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sherstov:2007:PRT

[She07] Alexander A. Sherstov. Powering requires threshold depth 3. Information Processing Letters, 102(2–3):104–107, April 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sung:2007:CIB

[SHH07] Jaechul Sung, Deukjo Hong, and Seokhie Hong. Cryptanalysis of an involutional block cipher using cellular automata. Information Processing Letters, 104 (5):183–185, November 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shi:2000:LBQ

[Shi00] Yaoyun Shi. Lower bounds of quantum black-box complexity and degree of approximating polynomials by influence of Boolean variables. Information Processing Letters, 75(1– 2):79–83, July 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/25/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/25/39/article.
pdf.

Shi:2002:ELB

[Shi02] Yaoyun Shi. Entropy lower bounds for quantum decision tree complexity. Information Processing Letters, 81(1):23– 27, January 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/27/
30/abstract.html.

Shih:2003:FTWa

[Shi03a] Jau-Der Shih. Fault-tolerant wormhole routing for hypercube networks. Information Processing Letters, 86(2):93–100, April 30, 2003. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Shih:2003:FTWb

[Shi03b] Jau-Der Shih. A fault-tolerant wormhole routing scheme for torus networks with nonconvex faults. Information Processing Letters, 88(6):271–278, December 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shim:2008:RKA

[Shi08] Kyung-Ah Shim. Rogue-key attacks on the multi-designated verifiers signature scheme. Information Processing Letters, 107(2):83–86, July 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shparlinski:2000:LCN

[Shp00]  Igor E. Shparlinski.  Lin-

ear complexity of the Naor– Reingold pseudo-random function. Information Processing Letters, 76(3):95–99, December 15, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/27/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/27/24/article.
pdf.

Shparlinski:2002:SMS

[Shp02] Igor E. Shparlinski. Security of most significant bits of g^{x^{2}}. Information Processing Letters, 83 (2):109–113, July 31, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shparlinski:2004:UDD

[Shp04] Igor E. Shparlinski. On the uniformity of distribution of the decryption exponent in fixed encryption exponent RSA. Information Processing Letters, 92 (3):143–147, November 15, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shparlinski:2007:BFC

[Shp07] Igor E. Shparlinski. Bounds on the Fourier coefficients of the weighted sum function. Information Processing Letters, 103 (3):83–87, July 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sibeyn:2004:EMM

[Sib04] Jop F. Sibeyn. External matrix multiplication and all-pairs shortest path. Information Processing Letters, 91(2):99–106, July 31, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sidorowicz:2007:GCN

[Sid07] Elżbieta Sidorowicz. The game chromatic number and the game colouring number of cactuses. Information Processing Letters, 102(4):147–151, May 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Simonsen:2004:CRC

[Sim04] Jakob Grue Simonsen. On confluence and residuals in Cauchy convergent transfinite rewriting. Information Processing Letters, 91(3):141–146, August 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sintzoff:2003:DCO

[Sin03] Michel Sintzoff. On the design of correct and optimal dynamical systems and games. Information Processing Letters, 88(1–2): 59–65, October 17, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shi:2005:ONO

[SJ05] Qingmin Shi and Joseph JaJa.

Optimal and near-optimal algorithms for generalized intersection reporting on pointer machines. Information Processing Letters, 95(3):382–388, August 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shen:2008:LCB

[SJ08] Haibin Shen and Yier Jin.

Low complexity bit parallel multiplier for GF(2^{m}) generated by equally-spaced trinomials. Information Processing Letters, 107(6):211–215, August 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Spriggs:2002:NBM

[SK02] Michael J. Spriggs and J. Mark Keil. A new bound for map labeling with uniform circle pairs. Information Processing Letters, 81(1):47–53, January 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/85/27/34/abstract.html.

Skowronek-Kaziow:2008:CMV

[SK08a] Joanna Skowronek-Kaziów. 1, 2 Conjecture — the multiplicative version. Information Processing Letters, 107(3–4):93–95, July 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Skowronek-Kaziow:2008:SDA

[SK08b] Joanna Skowronek-Kaziów. Some digraphs arising from number theory and remarks on the zero-divisor graph of the ring Z_{n}. Information Processing Letters, 108(3):165–169, October 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shmulevich:2001:AAM

[SKA01] Ilya Shmulevich, Aleksey D. Korshunov, and Jaakko Astola. Almost all monotone Boolean functions are polynomially learnable using membership queries. Information Processing Letters, 79(5):211–213, September 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Seal:2005:OHC

[SKA05] Sudip Seal, Srikanth Komarina, and Srinivas Aluru. An optimal hierarchical clustering algorithm for gene expression data. Information Processing Letters, 93 (3):143–147, February 14, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sung:2002:CPE

[SKP^{+}02] Minyoung Sung, Soyoung Kim, Sangsoo Park, Naehyuck Chang, and Heonshik Shin. Comparative performance evaluation of Java threads for embedded applications: Linux Thread vs. Green Thread. Information Processing Letters, 84(4):221–225, November 30, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Skulrattanakulchai:2002:ECG

[Sku02a] San Skulrattanakulchai. 4-edge-coloring graphs of maximum degree 3 in linear time. Information Processing Letters, 81 (4):191–195, February 28, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/85/35/30/abstract.html.

Skurczynski:2002:CBT

[Sku02b] Jerzy Skurczyński. A characterization of Büchi tree automata. Information Processing Letters, 81(1):29–33, January 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/gej-ng/10/23/
                                        20/85/27/31/abstract.html.

Skulrattanakulchai:2004:ACS

[Sku04] San Skulrattanakulchai. Acyclic colorings of subcubic graphs. Information Processing Letters, 92 (4):161–167, November 30, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Skulrattanakulchai:2006:LVC

[Sku06] San Skulrattanakulchai. Δ-List vertex coloring in linear time. Information Processing Letters, 98(3):101–106, May 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Singh:2001:SRG

[SL01] Gautam B. Singh and Donglin Liu. Sequence retrieval from genomic databases. Information Processing Letters, 79 (1):25–32, May 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

                                        /www.elsevier.nl/gej%2Dng/
                                        10/23/20/80/25/29/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej%2Dng/10/23/20/80/25/
                                        29/article.pdf.

Shiowattana:2006:OLB

[SL06a] Dungjade Shiowattana and Satyanarayana V. Lokam. An optimal lower bound for 2-query locally decodable linear codes.

Information Processing Letters, 97(6):244–250, March 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stump:2006:KBC

[SL06b] Aaron Stump and Bernd Löchner.

Knuth–Bendix completion of theories of commuting group endomorphisms. Information Processing Letters, 98(5):195–198, June 15, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shi:2005:IGN

[SLL^{+}05] X. H. Shi, Y. C. Liang, H. P.

Lee, C. Lu, and L. M. Wang. An improved GA and a novel PSO-GA-based hybrid algorithm. Information Processing Letters, 93 (5):255–261, March 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shi:2007:PSO

[SLL^{+}07] X. H. Shi, Y. C. Liang, H. P. Lee, C. Lu, and Q. X. Wang. Particle swarm optimization-based algorithms for TSP and generalized TSP. Information Processing Letters, 103(5):169–176, August 31, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Su:2008:IAF

[SLT08] Hsin-Hao Su, Chin Lung Lu, and Chuan Yi Tang. An improved algorithm for finding a length-constrained maximumdensity subtree in a tree. Information Processing Letters, 109 (2):161–164, December 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Swain:2002:PIS

[SM02] Anjan Kumar Swain and Alan S. Morris. Performance improvement of self-adaptive evolutionary methods with a dynamic lower bound. Information Processing Letters, 82(1): 55–63, April 15, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/86/27/
34/abstract.html.

Stanica:2003:CCR

[SM03] Pantelimon Stănică and Subhamoy Maitra. A constructive count of rotation symmetric functions. Information Processing Letters, 88(6):299–304, December 31, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Smart:2001:NCP

[Sma01] N. P. Smart. A note on the xcoordinate of points on an elliptic curve in characteristic two. Information Processing Letters, 80(5):261–263, December 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/84/36/34/abstract.html.

Smorodinsky:2008:NOF

[Smo08] Shakhar Smorodinsky. A note on the online First-Fit algorithm for coloring k-inductive graphs. Information Processing Letters, 109(1):44–45, December 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Solomonoff:2008:PUN

[Sol08] Ray J. Solomonoff. The probability of “undefined” (nonconverging) output in generating the universal probability distribution. Information Processing Letters, 106(6):238–240, June 15, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sopena:2002:TEO

[Sop02] Éric Sopena. There exist oriented planar graphs with oriented chromatic number at least sixteen. Information Processing Letters, 81(6):309– 312, March 31, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/37/
30/abstract.html.

Song:2007:BNN

[SPKL07] Moon Bae Song, Kwang Jin Park, Ki-Sik Kong, and Sang Keun Lee. Bottom-up nearest neighbor search for R-trees. Information Processing Letters, 101(2): 78–85, January 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Srinivasan:2008:NDN

[Sri08] Aravind Srinivasan. A note on the distribution of the number of prime factors of the integers. Information Processing Letters, 109(2):133–135, December 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Serra-Sagrista:2000:ELP

[SS00]  Joan Serra-Sagristà.  Enu-

meration of lattice points in l_{1} norm. Information Processing Letters, 76(1–2):39–44, November 20, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/67/24/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/67/24/30/article.
                                        pdf.

Schwarzer:2002:COS

[SS02] Fabian Schwarzer and Achim Schweikard. On the complexity of one-shot translational separability. Information Processing Letters, 83(4):187–194, August 31, 2002. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sanders:2003:BLT

[SS03a] Peter Sanders and Jop F. Sibeyn. A bandwidth latency tradeoff for broadcast and reduction. Information Processing Letters, 86(1):33–38, April 15, 2003. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Sawada:2003:SEO

[SS03b] J. Sawada and J. P. Spinrad.

From a simple elimination ordering to a strong elimination ordering in linear time. Information Processing Letters, 86(6): 299–302, June 30, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schobbens:2002:TLT

[SSSS02] P. Y. Schobbens, G. Saake, A. Sernadas, and C. Sernadas. A two-level temporal logic for evolving specifications. Information Processing Letters, 83(3): 167–172, August 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shim:2004:ECC

[SSYM04] Jeong Min Shim, Seok Il Song, Jae Soo Yoo, and Young Soo Min. An efficient cache conscious multi-dimensional index structure. Information Processing Letters, 92(3):133–142, November 15, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Saks:2000:LDS

[SSZZ00] Michael Saks, Aravind Srinivasan, Shiyu Zhou, and David Zuckerman. Low discrepancy sets yield approximate minwise independent permutation families. Information Processing Letters, 73(1–2):29–32, January 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/62/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/62/25/30/article.
pdf.

Schonlau:2000:DMI

[ST00] Matthias Schonlau and Martin Theus. Detecting masquerades in intrusion detection based on unpopular commands. Information Processing Letters, 76(1–2): 33–38, November 20, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/24/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/24/29/article.
pdf.

Sung:2002:EGR

[ST02] Shao-Chin Sung and Keisuke Tanaka. An exponential gap with the removal of one negation gate. Information Processing Letters, 82(3):155–157, May 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/86/34/33/abstract.html.

Sung:2004:LNB

[ST04a] Shao Chin Sung and Keisuke Tanaka. Limiting negations in bounded-depth circuits: An extension of Markov’s theorem. Information Processing Letters, 90 (1):15–20, April 15, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szepietowski:2004:NOC

[ST04b] Andrzej Szepietowski and Monika Targan. A note on the oriented chromatic number of grids. Information Processing Letters, 92 (2):65–70, October 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Srivastava:2005:TBP

[ST05] Nikhil Srivastava and Alan D.

Taylor. Tight bounds on plurality. Information Processing Letters, 96(3):93–95, November 15, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shioura:2006:EPE

[ST06] Akiyoshi Shioura and Takeshi Tokuyama. Efficiently pricing European-Asian optionsultimate implementation and analysis of the AMO algorithm. Information Processing Letters, 100(6):213–219, December 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Saule:2009:AST

[ST09] Erik Saule and Denis Trystram. Analyzing scheduling with transient failures. Information Processing Letters, 109(11):539– 542, May 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Staiger:2005:CDE

[Sta05] Ludwig Staiger. Constructive dimension equals Kolmogorov complexity. Information Processing Letters, 93(3):149–153, February 14, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stefankovic:2000:AOD

[Šte00] Daniel Štefankovič. Acyclic orientations do not lead to optimal deadlock-free packet routing algorithms. Information Processing Letters, 73(5–6):221– 225, March 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/28/24/62/28/
                                        36/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/28/24/62/28/36/article.
                                        pdf.

Stewart:2008:FPT

[Ste08] Iain A. Stewart. On the fixedparameter tractability of parameterized model-checking problems. Information Processing Letters, 106(1):33–36, March 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shih:2007:EBC

[STH07] Lun-Min Shih, Jimmy J. M. Tan, and Lih-Hsing Hsu. Edgebipancyclicity of conditional faulty hypercubes. Information Processing Letters, 105(1):20– 25, December 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stinson:2004:ACS

[Sti04] D. R. Stinson. Attack on a concast signature scheme. Information Processing Letters, 91 (1):39–41, July 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Studer:2009:CKD

[Stu09]  Thomas Studer.  Common

knowledge does not have the Beth property. Information Processing Letters, 109(12):611– 614, May 31, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Subramanian:2006:AHA

[Sub06] C. R. Subramanian. Analysis of a heuristic for acyclic edge colouring. Information Processing Letters, 99(6):227– 229, September 30, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sung:2001:INC

[Sun01] Pantelimon Stănicăand Soo Hak Sung. Improving the nonlinearity of certain balanced Boolean functions with good local and global avalanche characteristics. Information Processing Letters, 79(4):167–172, August 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/33/
27/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/80/33/27/article.
                                        pdf.

Suomela:2007:AIC

[Suo07] Jukka Suomela. Approximability of identifying codes and locating–dominating codes. Information Processing Letters, 103(1):28–33, June 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Slivnik:2005:PLP

[SV05] Boštjan Slivnik and Boštjan Vilfan. Producing the left parse during bottom-up parsing. Information Processing Letters, 96 (6):220–224, December 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Steinwandt:2008:OTS

[SV08] Rainer Steinwandt and Viktória I. Villányi. A one-time signature using run-length encoding. Information Processing Letters, 108(4):179–185, October 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Smith:2009:EPS

[SV09] Rebecca Smith and Vincent Vatter. The enumeration of permutations sortable by pop stacks in parallel. Information Processing Letters, 109(12):626–629, May 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Servedio:2005:CSP

[SW05a] Rocco A. Servedio and Andrew Wan. Computing sparse permanents faster. Information Processing Letters, 96(3):89–92, November 15, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shparlinski:2005:LCB

[SW05b] Igor E. Shparlinski and Arne Winterhof. On the linear complexity of bounded integer sequences over different moduli. Information Processing Letters, 96(5):175–177, December 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shen:2007:SCP

[SW07a] Liang Shen and Yingqian Wang.

A sufficient condition for a planar graph to be 3-choosable. Information Processing Letters, 104(4):146–151, November 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shparlinski:2007:QPR

[SW07b] Igor E. Shparlinski and Arne Winterhof. Quantum period reconstruction of approximate sequences. Information Processing Letters, 103(6):211–215, September 15, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Salamon:2008:FST

[SW08] Gábor Salamon and Gábor Wiener. On finding spanning trees with few leaves. Information Processing Letters, 105 (5):164–169, February 29, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Spoerhase:2009:ASM

[SW09] J. Spoerhase and H.-C. Wirth. An O(n(log n)^{2}/ log log n) algorithm for the single maximum coverage location or the (1, X_{p})- medianoid problem on trees. Information Processing Letters, 109(8):391–394, March 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sui:2004:CBA

[SWCC04] Hongfei Sui, Jianxin Wang, Jianer Chen, and Songqiao Chen. The cost of becoming anonymous: on the participant payload in Crowds. Information Processing Letters, 90(2):81–86, April 30, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Saia:2008:RCC

[SY08] Jared Saia and Maxwell Young. Reducing communication costs in robust peer-to-peer networks. Information Processing Letters, 106(4):152–158, May 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Su:2004:APL

[SYL04] Yih-Ching Su, Chu-Sing Yang, and Chen-Wei Lee. The analysis of packet loss prediction for Gilbert-model with loss rate uplink. Information Processing Letters, 90(3):155–159, May 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

parl:2004:LCA

[ŠŽ04] Petra Šparl and Janez Žerovnik.

2-local 5/4-competitive algorithm for multicoloring trianglefree hexagonal graphs. Information Processing Letters, 90(5): 239–246, June 15, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szabo:2003:SPN

[Sza03] Tibor Szabó. On the spectrum of projective norm-graphs. Information Processing Letters, 86 (2):71–74, April 30, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szepietowski:2006:NAO

[Sze06] Andrzej Szepietowski. A note on alternating one-pebble Turing machines with sublogarithmic space. Information Processing Letters, 98(5):174–176, June 15, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szepietowski:2008:FTM

[Sze08] Andrzej Szepietowski. Fooling Turing machines with sublogarithmic space: a note on “For completeness, sublogarithmic space is no space” by M.

Agrawal. Information Processing Letters, 106(4):162–163, May 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Agr02].

Shachnai:2009:NGR

[SZM09] Hadas Shachnai, Lisa Zhang, and Tomomi Matsui. A note on generalized rank aggregation. Information Processing Letters, 109(13):647–651, June 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sun:2008:MHA

[SZQ08] Li Sun, Fuji Zhang, and Jianguo Qian. Multi-hop all-to-all optical routings in Cartesian product networks. Information Processing Letters, 107(6):252–256, August 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Takenaga:2008:PCE

[TA08] Yasuhiko Takenaga and Shigeru Arai. PSPACE-completeness of an escape problem. Information Processing Letters, 108(4): 229–233, October 31, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Takamura:2004:SVA

[TAI04] Masataka Takamura, Tom Altman, and Yoshihide Igarashi. Speedup of Vidyasankar’s algorithm for the group k-exclusion problem. Information Processing Letters, 91(2):85–91, July 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Takahashi:2000:FAC

[Tak00] Daisuke Takahashi. A fast algorithm for computing large Fibonacci numbers. Information Processing Letters, 75(6):243– 246, November 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/31/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/31/25/article.
pdf.

Takaoka:2005:TAA

[Tak05] Tadao Takaoka. An O(n^{3} log log n/ log n)

time algorithm for the all-pairs shortest path problem. Information Processing Letters, 96 (5):155–161, December 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Talbot:2000:FFO

[Tal00] Jean-Marc Talbot. The ∃∀^{2} fragment of the first-order theory of atomic set constraints is Π_{1}^{0}-hard. Information Processing Letters, 74(1–2):27– 33, April 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/25/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/

10/23/20/63/25/29/article.
pdf.

Tambouratzis:2000:SHC

[Tam00]  T. Tambouratzis.  Solving

the Hamiltonian cycle problem via an artificial neural network. Information Processing Letters, 75(6):237–242, November 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/31/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/31/24/article.
pdf.

Tan:2000:OBB

[Tan00]  Xuehou Tan.  On optimal

bridges between two convex regions. Information Processing Letters, 76(4–6):163–168, December 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/28/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/28/26/article.
pdf.

Tan:2001:FCS

[Tan01a] Xuehou Tan. Fast computation of shortest watchman routes in simple polygons. Information Processing Letters, 77(1): 27–33, January 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.

nl/gej-ng/10/23/20/68/24/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/24/29/article.
pdf.

Tan:2001:OCV

[Tan01b] Xuehou Tan. Optimal computation of the Voronoi diagram of disjoint clusters. Information Processing Letters, 79(3): 115–119, July 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/32/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/80/32/27/article.
pdf.

Tan:2001:SZR

[Tan01c] Xuehou Tan. Shortest zookeeper’s routes in simple polygons. Information Processing Letters, 77(1):23–26, January 31, 2001. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/24/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/24/28/article.
pdf.

Tan:2006:AIS

[Tan06a] Chik How Tan. Analysis of improved signcryption scheme with key privacy. Information Processing Letters, 99(4):135–138, August 31, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tan:2006:AAZ

[Tan06b] Xuehou Tan. A 2-approximation algorithm for the Zookeeper’s Problem. Information Processing Letters, 100(5):183–187, December 16, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tan:2007:SSP

[Tan07] Xuehou Tan. Sweeping simple polygons with the minimum number of chain guards. Information Processing Letters, 102 (2–3):66–71, April 30, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tan:2008:NIC

[Tan08] Jinsong Tan. A note on the inapproximability of correlation clustering. Information Processing Letters, 108(5):331–335, November 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tao:2006:IPC

[Tao06] Yunfeng Tao. Infinity problems and countability problems for ω-automata. Information Processing Letters, 100(4):151–153, November 30, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tarlecki:2004:E

[Tar04] Andrzej Tarlecki. Editorial. Information Processing Letters, 90 (1):1–2, April 15, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tatti:2006:CCQ

[Tat06] Nikolaj Tatti. Computational complexity of queries based on itemsets. Information Processing Letters, 98(5):183–187, June 15, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

terBeek:2005:MTA

[tBK05] Maurice H. ter Beek and Jetty Kleijn. Modularity for teams of I/O automata. Information Processing Letters, 95(5):487– 495, September 15, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tongchim:2002:PGA

[TC02] Shisanu Tongchim and Prabhas Chongstitvatana. Parallel genetic algorithm with parameter adaptation. Information Processing Letters, 82(1): 47–54, April 15, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/86/27/
33/abstract.html.

Toroslu:2004:DPS

[TC04] Ismail H. Toroslu and Ahmet Cosar. Dynamic programming solution for multiple query optimization problem. Information Processing Letters, 92(3): 149–155, November 15, 2004.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

tenCate:2009:NEP

[tC09] Balder ten Cate. A note on the expressibility problem for modal logics and star-free regular expressions. Information Processing Letters, 109(10):509– 513, April 30, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tian:2009:OSU

[TCNY09] Ji Tian, T. C. E. Cheng, C. T. Ng, and Jinjiang Yuan. Online scheduling on unbounded parallel-batch machines to minimize the makespan. Information Processing Letters, 109(21–22): 1211–1215, October 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tian:2009:NSI

[TD09] Cong Tian and Zhenhua Duan. A note on stutter-invariant PLTL. Information Processing Letters, 109(13):663–667, June 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tan:2002:OOA

[TH02] Zhiyi Tan and Yong He. Optimal online algorithm for scheduling on two identical machines with machine availability constraints. Information Processing Letters, 83(6):323–329, September 30, 2002. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Tan:2003:FSS

[TH03] Xuehou Tan and Tomio Hirata.

Finding shortest safari routes in simple polygons. Information Processing Letters, 87(4): 179–186, August 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Turau:2009:SSA

[TH09] Volker Turau and Bernd Hauck.

A self-stabilizing algorithm for constructing weakly connected minimal dominating sets. Information Processing Letters, 109 (14):763–767, June 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

ETAC:2002:CLP

[The02] The Eindhoven Tuesday Afternoon Club. On computing a longest path in a tree. Information Processing Letters, 81(2): 93–96, January 31, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/33/
33/abstract.html.

Thomasian:2005:RVR

[Tho05] Alexander Thomasian. Reconstruct versus read-modify writes in RAID. Information Processing Letters, 93(4):163–168, February 28, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsai:2007:PBH

[TJ07] Chang-Hsiung Tsai and Shu-Yun Jiang. Path bipancyclicity of hypercubes. Information Processing Letters, 101(3): 93–97, February 14, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Teoh:2006:RBM

[TJC^{+}06] Andrew Teoh, Beng Jin, Tee Connie, David Ngo, and Chek Ling. Remarks on BioHash and its mathematical foundation. Information Processing Letters, 100(4):145–150, November 30, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tzeng:2007:SSE

[TJH07] Chi-Hung Tzeng, Jehn-Ruey Jiang, and Shing-Tsaan Huang. A self-stabilizing (Δ + 4)-edge-coloring algorithm for planar graphs in anonymous uniform systems. Information Processing Letters, 101(4):168–173, February 28, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tzeng:2009:SSA

[TJH09] Chi-Hung Tzeng, Jehn-Ruey Jiang, and Shing-Tsaan Huang. A self-stabilizing algorithm for the maximum planarization problem in complete bipartite networks. Information Processing Letters, 109(10):518–522, April 30, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsuchiya:2002:BQS

[TK02] Tatsuhiro Tsuchiya and Tohru Kikuno. Byzantine quorum systems with maximum availability. Information Processing Letters, 83(2):71–77, July 31, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tomita:2009:RVG

[TK09] Kohji Tomita and Haruhisa Kurokawa. On the reachability of a version of graph-rewriting system. Information Processing Letters, 109(14):777–782, June 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tanaka:2007:IFK

[TKS07] Yuuki Tanaka, Hiroyuki Kawai, and Yukio Shibata. Isomorphic factorization, the Kronecker product and the line digraph. Information Processing Letters, 101(2):72–77, January 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsai:2002:LFF

[TLH02] Yin-Te Tsai, Yaw-Ling Lin, and F. R. Hsu. The on-line firstfit algorithm for radio frequency assignment problems. Information Processing Letters, 84(4): 195–199, November 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Thomasian:2005:ENQ

[TLZ05] Alexander Thomasian, Yue Li, and Lijuan Zhang. Exact k-NN queries on clustered SVD datasets. Information Processing Letters, 94(6):247–252, June 30, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tiplea:2001:NSE

[ŢM01] Ferucio Laurenţiu Ţiplea and Erkki Mäkinen. A note on synchronized extension systems. Information Processing Letters, 79(1):7–9, May 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.elsevier.nl/gej%2Dng/
10/23/20/80/25/26/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/23/20/80/25/
26/article.pdf.

Tiplea:2005:SSW

[ŢM05] Ferucio Laurenţiu Ţiplea and Dan Cristian Marinescu. Structural soundness of workflow nets is decidable. Information Processing Letters, 96(2):54–58, October 31, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tian:2009:OHV

[TM09] Yingzhi Tian and Jixiang Meng. λ_{c}-Optimally half vertex transitive graphs with regularity k. Information Processing Letters, 109(13):683–686, June 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tavares:2008:HRT

[TMSO08] Eduardo Tavares, Paulo Maciel, Bruno Silva, and Meuse N. Oliveira, Jr. Hard real-time tasks’ scheduling considering voltage scaling, precedence and exclusion relations. Information Processing Letters, 108(2): 50–59, September 30, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

To:2009:UFA

[To09] Anthony Widjaja To. Unary finite automata vs. arithmetic progressions. Information Processing Letters, 109(17):1010– 1014, August 16, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tonien:2006:TTS

[Ton06] Dongvu Tonien. On a traitor tracing scheme from ACISP 2003. Information Processing Letters, 100(1):21–22, October 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Touzet:2003:TED

[Tou03] Hélène Touzet. Tree edit distance with gaps. Information Processing Letters, 85(3):123– 129, February 14, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Trelea:2003:PSO

[Tre03] Ioan Cristian Trelea. The particle swarm optimization algorithm: convergence analysis and parameter selection. Information Processing Letters, 85(6): 317–325, March 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tripakis:2004:UPD

[Tri04] Stavros Tripakis. Undecidable problems of decentralized observation and control on regular languages. Information Processing Letters, 90(1):21–28, April 15, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tripakis:2006:FTD

[Tri06] Stavros Tripakis. Folk theorems on the determinization and minimization of timed automata. Information Processing Letters, 99(6):222–226, September 30, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ta-Shma:2002:SIE

[TS02] Amnon Ta-Shma. Storing information with extractors. Information Processing Letters, 83(5):267–274, September 15, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsai:2003:CLC

[Tsa03] Yin-Te Tsai. The constrained longest common subsequence problem. Information Processing Letters, 88(4):173–176, November 30, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsaban:2006:FGD

[Tsa06] Boaz Tsaban. Fast generators for the Diffie–Hellman key agreement protocol and malicious standards. Information Processing Letters, 99(4):145– 148, August 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsai:2007:CEH

[Tsa07] Chang-Hsiung Tsai. Cycles embedding in hypercubes with node failures. Information Processing Letters, 102(6):242–246, June 15, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tse:2005:SNL

[Tse05] Savio S. H. Tse. A short note on the lower bound of dilation for O(log n)-label interval routing. Information Processing Letters, 95(2):351–353, July 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsin:2002:SRD

[Tsi02] Y. H. Tsin. Some remarks on distributed depth-first search. Information Processing Letters, 82(4):173–178, May 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsin:2004:FED

[Tsi04] Y. H. Tsin. On finding an ear decomposition of an undirected graph distributively. Information Processing Letters, 91(3): 147–153, August 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsin:2007:ISS

[Tsi07]  Yung H. Tsin.  An im-

proved self-stabilizing algorithm for biconnectivity and bridgeconnectivity. Information Processing Letters, 102(1):27–34, April 15, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsur:2008:FAG

[Tsu08] Dekel Tsur. Faster algorithms for guided tree edit distance. Information Processing Letters, 108(4):251–254, October 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsai:2004:SCP

[TTH04] Chang-Hsiung Tsai, Jimmy J. M. Tan, and Lih-Hsing Hsu. The super-connected property of recursive circulant graphs. Information Processing Letters, 91 (6):293–298, September 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsai:2002:FTH

[TTLH02] Chang-Hsiung Tsai, Jimmy J. M. Tan, Tyne Liang, and Lih-Hsing Hsu. Fault-tolerant Hamiltonian laceability of hypercubes. Information Processing Letters, 83(6):301–306, September 30, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Turski:2003:IWF

[Tur03] Władysław M. Turski. It was fun. Information Processing Letters, 88(1–2):7–12, October 17, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Turau:2007:LSS

[Tur07]  Volker Turau.  Linear self-

stabilizing algorithms for the independent and dominating set problems using an unfair distributed scheduler. Information Processing Letters, 103(3): 88–93, July 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Trahan:2002:SMA

[TV02] Jerry L. Trahan and Ramachandran Vaidyanathan. Scaling multiple addition and prefix sums on the reconfigurable mesh. Information Processing Letters, 82(6):277–282, June 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Takenaga:2006:TNC

[TW06] Yasuhiko Takenaga and Toby Walsh. Tetravex is NPcomplete. Information Processing Letters, 99(5):171–174, September 15, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tan:2008:NAS

[TYP08] Xuegong Tan, Shun-Zheng Yu, and Jin Han Park. A note about some properties of BC graphs. Information Processing Letters, 108(6):398–401, November 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tang:2001:EAS

[TYW01] Shyue-Ming Tang, Fu-Long Yeh, and Yue-Li Wang. An efficient algorithm for solving the homogeneous set sandwich problem. Information Processing Letters, 77(1):17– 22, January 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/24/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/24/27/article.
pdf.

Telelis:2005:AEA

[TZ05] Orestis A. Telelis and Vassilis Zissimopoulos. Absolute o(log m) error in approximating random set covering: an average case analysis. Information Processing Letters, 94(4):171–177, May 31, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Telelis:2008:DBO

[TZ08] Orestis A. Telelis and Vassilis Zissimopoulos. Dynamic bottleneck optimization for k-edge and 2-vertex connectivity. Information Processing Letters, 106(6): 251–257, June 15, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Uehara:2000:PAA

[UC00] Ryuhei Uehara and Zhi-Zhong Chen. Parallel approximation algorithms for maximum weighted matching in general graphs. Information Processing Letters, 76(1–2):13–17, November 20, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/24/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/24/26/article.
pdf.

Ultes-Nitsche:2007:PSC

[UN07] Ulrich Ultes-Nitsche. A powerset construction for reducing Büchi automata to nondeterminism degree two. Information Processing Letters, 101 (3):107–111, February 14, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ukkonen:2009:RAA

[UPGM09] Antti Ukkonen, Kai Puolamäki, Aristides Gionis, and Heikki Mannila. A randomized approximation algorithm for computing bucket orders. Information Processing Letters, 109(7):356–359, March 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ural:2003:DTE

[UW03] Hasan Ural and David Whittier. Distributed testing without encountering controllability and observability problems. Information Processing Letters, 88 (3):133–141, November 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vagvolgyi:2006:DRT

[Vág06] Sándor Vágvölgyi. Descendants of a recognizable tree language for sets of linear monadic term rewrite rules. Information Processing Letters, 99(3):111–118, August 16, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vagvolgyi:2008:MTR

[Vág08] Sándor Vágvölgyi. Murg term rewrite systems. Information Processing Letters, 108(5):264– 272, November 15, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vahrenhold:2007:PAK

[Vah07] Jan Vahrenhold. An in-place algorithm for Klee’s measure problem in two dimensions. Information Processing Letters, 102(4):169–174, May 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vajnovszki:2008:MRG

[Vaj08] Vincent Vajnovszki. More restrictive Gray codes for necklaces and Lyndon words. Information Processing Letters, 106 (3):96–99, April 30, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Valiant:2004:LRC

[Val04a] Paul Valiant. The Log-Rank Conjecture and low degree polynomials. Information Processing Letters, 89(2):99–103, January 31, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Valiente:2004:TUS

[Val04b] Gabriel Valiente. Trading uninitialized space for time. Information Processing Letters, 92(1): 9–13, October 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

VandenBussche:2001:RQU

[Van01] Jan Van den Bussche. Rewriting queries using views over monadic database schemas. Information Processing Letters, 79 (3):111–114, July 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/80/32/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/80/32/26/article.
pdf.

Vanherpe:2003:SOP

[Van03] Jean-Marie Vanherpe. Some optimization problems on weakbisplit graphs. Information Processing Letters, 88(6):305–310, December 31, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

VanRoy:2007:M

[Van07] Benjamin Van Roy. MIN. Information Processing Letters, 102 (2–3):72–73, April 30, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vassilevska:2009:EAC

[Vas09] Virginia Vassilevska. Efficient algorithms for clique problems. Information Processing Letters, 109(4):254–257, January 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vieira:2008:CRF

[VB08] Gustavo M. D. Vieira and Luiz E. Buzato. On the coordinator’s rule for Fast Paxos. Information Processing Letters, 107(5):183–187, August 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vandeBult:2008:PMC

[vdBW08] Fokko J. van de Bult and Gerhard J. Woeginger. The problem of the moody chess players. Information Processing Letters, 108(5):336–337, November 15, 2008. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

vanDijk:2006:ICS

[vDKST06] Marten van Dijk, Tom Kevenaar, Geert-Jan Schrijen, and Pim Tuyls. Improved constructions of secret sharing schemes by applying (λ, ω)- decompositions. Information Processing Letters, 99(4):154– 157, August 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanDeursen:2009:NFP

[vDR09] Ton van Deursen and Saša Radomirović. On a new formal proof model for RFID location privacy. Information Processing Letters, 110(2):57–61, December 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanderZwaag:2001:CFP

[vdZ01] Mark B. van der Zwaag. The cones and foci proof technique for timed transition systems. Information Processing Letters, 80(1):33–40, October 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/84/27/31/abstract.html.

Veanes:2000:FTR

[Vea00]  Margus Veanes.  Farmer’s

Theorem revisited. Information Processing Letters, 74(1– 2):47–53, April 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/63/25/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/63/25/32/article.
                                        pdf.

Verma:2002:ARR

[Ver02] Rakesh Verma. Algorithms and reductions for rewriting problems. II. Information Processing Letters, 84(4):227–233, November 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Verhoeven:2006:EAL

[Ver06] Yves F. Verhoeven. Enhanced algorithms for Local Search. Information Processing Letters, 97 (5):171–176, March 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vereshchagin:2007:KCE

[Ver07] Nikolai K. Vereshchagin. Kolmogorov complexity of enumerating finite sets. Information Processing Letters, 103(1):34– 39, June 30, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vestergaard:2006:CAS

[Ves06] René Vestergaard. A constructive approach to sequential Nash equilibria. Information Processing Letters, 97(2):46–51, January 31, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Veloso:2002:LMI

[VFV02] Paulo A. S. Veloso, José L. Fiadeiro, and Sheila R. M. Veloso. On local modularity and interpolation in entailment systems. Information Processing Letters, 82(4):203–211, May 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vigneron:2000:AFM

[VGG^{+}00] Antoine Vigneron, Lixin Gao, Mordecai J. Golin, Giuseppe F. Italiano, and Bo Li. An algorithm for finding a k-median in a directed tree. Information Processing Letters, 74(1– 2):81–88, April 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/25/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/25/36/article.
pdf.

Viana:2008:QEP

[Via08] Raquel Viaña. Quick encoding of plane graphs in log_{2} 14 bits per edge. Information Processing Letters, 108(3):150–154, October 16, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vidyasankar:2003:SGM

[Vid03]  K. Vidyasankar.  A simple

group mutual l-exclusion algorithm. Information Processing Letters, 85(2):79–85, January 31, 2003. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vigneron:2003:RIA

[Vig03] Antoine Vigneron. Reporting intersections among thick objects. Information Processing Letters, 85(2):87–92, January 31, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vinodchandran:2004:AMN

[Vin04] N. V. Vinodchandran. AMexp ⊈ (NP ∩ coNP)/poly. Information Processing Letters, 89(1):43–47, January 16, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vincent:2004:IUS

[VL04] Millist W. Vincent and Jixue Liu. Irrelevant updates and selfmaintainability in transitive closure database views. Information Processing Letters, 89(1): 25–29, January 16, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Voorhies:2006:FSM

[VLK06] Seth Voorhies, Hyunyoung Lee, and Andreas Klappenecker. Fair service for mice in the presence of elephants. Information Processing Letters, 99(3): 96–101, August 16, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Voorhoeve:2001:IFD

[VM01] Marc Voorhoeve and Sjouke Mauw. Impossible futures and determinism. Information Processing Letters, 80(1): 51–58, October 15, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/84/27/
33/abstract.html.

Vogler:2008:ASP

[Vog08] Walter Vogler. Another short proof of optimality for the MIN cache replacement algorithm. Information Processing Letters, 106(5):219–220, May 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Volzer:2004:CPF

[Völ04] Hagen Völzer. A constructive proof for FLP. Information Processing Letters, 92(2):83–87, October 31, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Volkmann:2007:RAC

[Vol07] Lutz Volkmann. Restricted arcconnectivity of digraphs. Information Processing Letters, 103(6):234–239, September 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vasco:2001:CPK

[VS01] María Isabel González Vasco and Rainer Steinwandt. Clouds over a public key cryptosystem based on Lyndon words. Information Processing Letters, 80 (5):239–242, December 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/84/36/30/abstract.html.

Verstraeten:2005:IWR

[VSSV05] D. Verstraeten, B. Schrauwen, D. Stroobandt, and J. Van Campenhout. Isolated word recognition with the Liquid State Machine : a case study. Information Processing Letters, 95(6):521–528, September 30, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vu:2005:CSP

[Vu05] Thuy Duong Vu. The compression structure of a process. Information Processing Letters, 96 (6):225–229, December 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Veloso:2001:LMV

[VV01] Paulo A. S. Veloso and Sheila R. M. Veloso. On local modularity variants and Πinstitutions. Information Processing Letters, 77(5–6):247– 253, March 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/32/
28/abstract.html;   http:

//www.elsevier.nl/gej-ng/
10/23/20/68/32/28/article.
pdf.

Vesel:2002:ILB

[VŽ02] Aleksander Vesel and Janez Žerovnik. Improved lower bound on the Shannon capacity of C_{7}. Information Processing Letters, 81(5):277–282, March 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/85/36/34/abstract.html.

Wongngamnit:2001:RLG

[WA01] Chinda Wongngamnit and Dana Angluin. Robot localization in a grid. Information Processing Letters, 77(5–6):261– 267, March 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/32/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/32/30/article.
pdf.

Wang:2001:OAC

[Wan01] D. P. Wang. An optimal algorithm for constructing an optimal bridge between two simple rectilinear polygons. Information Processing Letters, 79(5): 229–236, September 15, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/80/34/29/abstract.html.

Wang:2004:UFG

[Wan04] Guilin Wang. Universal forgery on a group signature scheme using self-certified public keys. Information Processing Letters, 89(5):227–231, March 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:2007:BEI

[Wan07] Junhu Wang. Binary equality implication constraints, normal forms and data redundancy. Information Processing Letters, 101(1):20–25, January 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:2008:ANN

[Wan08] Yusu Wang. Approximating nearest neighbor among triangles in convex position. Information Processing Letters, 108 (6):379–385, November 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:2001:HTM

[WC01] Guoqing Wang and T. C. Edwin Cheng. Heuristics for two-machine no-wait flowshop scheduling with an availability constraint. Information Processing Letters, 80(6):305– 309, December 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/84/37/
                                        32/abstract.html.

Walczak:2003:MVW

[WC03] Krzysztof Walczak and Wojciech Cellary. Modeling virtual worlds in databases. Information Processing Letters, 88(1–2): 67–72, October 17, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:2008:FFH

[WC08] Wen-Qing Wang and Xie-Bin Chen. A fault-free Hamiltonian cycle passing through prescribed edges in a hypercube with faulty edges. Information Processing Letters, 107(6):205–210, August 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:2009:TNT

[WC09] Weifan Wang and Dong Chen.

(2, 1)-total number of trees with maximum degree three. Information Processing Letters, 109 (14):805–810, June 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wu:2008:FCH

[WCFC08] Ruei-Yu Wu, Gen-Huey Chen, Jung-Sheng Fu, and Gerard J. Chang. Finding cycles in hierarchical hypercube networks. Information Processing Letters, 109(2):112–115, December 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:2008:ILB

[WCH08] Jianxin Wang, Jianer Chen, and Min Huang. An improved lower bound on approximation algorithms for the Closest Substring problem. Information Processing Letters, 107(1):24–28, June 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Weisman:2004:VDM

[WCK04] Amit Weisman, L. Paul Chew, and Klara Kedem. Voronoi diagrams of moving points in the plane and of lines in space: tight bounds for simple configurations. Information Processing Letters, 92(5):245–251, December 16, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wan:2002:NML

[WCX02] Yingyu Wan, Guoliang Chen, and Yinlong Xu. A note on the minimum label spanning tree. Information Processing Letters, 84(2):99–101, October 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:2000:TMW

[WCY00] Cao An Wang, Francis Y. Chin, and Boting Yang. Triangulations without minimum-weight drawing. Information Processing Letters, 74(5–6):183– 189, June 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/63/28/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/63/28/25/article.
pdf.

Wegener:2000:WCE

[Weg00] Ingo Wegener. Worst case examples for operations on OBDDs. Information Processing Letters, 74(3–4):91– 96, May 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/28/24/63/27/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/28/24/63/27/25/article.
pdf.

Wegener:2002:SCP

[Weg02] Ingo Wegener. A simplified correctness proof for a well-known algorithm computing strongly connected components. Information Processing Letters, 83 (1):17–19, July 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Weston:2004:FPT

[Wes04]  Mark Weston.  A fixed-

parameter tractable algorithm for matrix domination. Information Processing Letters, 90(5): 267–272, June 15, 2004. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Westphal:2008:NCT

[Wes08] Stephan Westphal. A note on the k-Canadian Traveller Problem. Information Processing Letters, 106(3):87–89, April 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wu:2003:IMT

[WHLH03] Tzong-Sun Wu, Chien-Lung Hsu, Han-Yu Lin, and Po-Sheng Huang. Improvement of the Miyazaki–Takaragi threshold digital signature scheme. Information Processing Letters, 88 (4):183–186, November 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wehrman:2009:GMS

[WHO09] Ian Wehrman, C. A. R. Hoare, and Peter W. O’Hearn. Graphical models of separation logic. Information Processing Letters, 109(17):1001–1004, August 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wu:2004:EAM

[WHZ04] Bang Ye Wu, Zheng-Nan Huang, and Fu-Jie Zhan. Exact algorithms for the minimum latency problem. Information Processing Letters, 92(6):303–309, December 31, 2004. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Williams:2009:FPL

[Wil09] Ryan Williams. Finding paths of length k in O^{∗}(2^{k}) time. Information Processing Letters, 109 (6):315–318, February 28, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Winkowski:2003:ACI

[Win03] Józef Winkowski. An algebraic characterization of independence of Petri net processes. Information Processing Letters, 88(1–2):73–81, October 17, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Windsor:2004:SPF

[Win04] Aaron Windsor. A simple proof that finding a maximal independent set in a graph is in NC. Information Processing Letters, 92 (4):185–187, November 30, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wirth:2003:HSC

[Wir03]  Niklaus Wirth.  Hardware/

software co-design then and now. Information Processing Letters, 88(1–2):83–87, October 17, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:2002:AAB

[WL02] Lusheng Wang and Zimao Li.

An approximation algorithm for a bottleneck k-Steiner tree problem in the Euclidean plane. Information Processing Letters, 81 (3):151–156, February 14, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/85/34/32/abstract.html.

Wu:2003:SLD

[WL03] Chin-Chia Wu and Wen-Chiung Lee. Scheduling linear deteriorating jobs to minimize makespan with an availability constraint on a single machine. Information Processing Letters, 87(2):89–93, July 31, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:2008:OD

[WL08] Shiying Wang and Shangwei Lin. λ′-optimal digraphs. Information Processing Letters, 108 (6):386–389, November 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:2008:NCP

[WLC08] Yingqian Wang, Huajing Lu, and Ming Chen. A note on 3choosability of planar graphs. Information Processing Letters, 105(5):206–211, February 29, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wei:2008:GRM

[WLLS08] Hsin-Wen Wei, Kwei-Jay Lin, Wan-Chen Lu, and Wei-Kuan Shih. Generalized rate monotonic schedulability bounds using relative period ratios. Information Processing Letters, 107 (5):142–148, August 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Par11].

Wan:2007:SIC

[WLP07] Guohua Wan, Joseph Y.-T. Leung, and Michael L. Pinedo. Scheduling imprecise computation tasks on uniform processors. Information Processing Letters, 104(2):45–52, October 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wagh:2000:CCW

[WMG00] Meghanad D. Wagh, Prakash Math, and Osman Guzide. Cyclic-cubes and wrap-around butterflies. Information Processing Letters, 75(1–2):25– 27, July 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/25/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/25/29/article.
pdf.

Wang:2007:FTP

[WMX07] Wei-Wei Wang, Mei-Jie Ma, and Jun-Ming Xu. Faulttolerant pancyclicity of augmented cubes. Information Processing Letters, 103(2):52– 56, July 16, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Williams:2000:BBC

[WNR00] Poul Frederick Williams, Macha Nikolskaïa, and Antoine Rauzy. Bypassing BDD construction for reliability analysis. Information Processing Letters, 75(1– 2):85–89, July 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/64/25/
                                        40/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/64/25/40/article.
                                        pdf.

Woeginger:2001:RPT

[Woe01] Gerhard J. Woeginger. The reconstruction of polyominoes from their orthogonal projections. Information Processing Letters, 77(5–6):225–229, March 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/68/32/
                                        25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/68/32/25/article.
                                        pdf.

Woeginger:2009:CPM

[Woe09]  Gerhard J. Woeginger.  A

comment on parallel-machine scheduling under a grade of service provision to minimize makespan. Information Processing Letters, 109(7):341–342, March 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wolfovitz:2006:CDC

[Wol06] Guy Wolfovitz. The complexity of depth-3 circuits computing symmetric Boolean functions. Information Processing Letters, 100(2):41–46, October 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wotawa:2001:VRH

[Wot01]  Franz Wotawa.  A variant

of Reiter’s hitting-set algorithm. Information Processing Letters, 79(1):45–51, May 31, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.elsevier.nl/gej%2Dng/
10/23/20/80/25/32/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/23/20/80/25/
32/article.pdf.

Wu:2001:ESAa

[WS01a] Jie Wu and Li Sheng. An efficient sorting algorithm for a sequence of kings in a tournament. Information Processing Letters, 79(6):297–299, September 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

elsevier.com/gej-ng/10/23/
20/80/37/34/abstract.html.

Wu:2001:ESAb

[WS01b] Jie Wu and Li Sheng. An efficient sorting algorithm for a sequence of kings in a tournament.

Information Processing Letters, 79(6):297–299, September 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/23/
                                        20/80/37/34/abstract.html.

Wu:2006:LCD

[WS06] Jigang Wu and Thambipillai Srikanthan. Low-complex dynamic programming algorithm for hardware/software partitioning. Information Processing Letters, 98(2):41–46, April 30, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Worah:2007:LTD

[WS07] Pratik Worah and Sandeep Sen. A linear time deterministic algorithm to find a small subset that approximates the centroid. Information Processing Letters, 105(1):17–19, December 31, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:2005:BPL

[WTC05] Xianbing Wang, Yong Meng Teo, and Jiannong Cao. A bivalency proof of the lower bound for uniform consensus. Information Processing Letters, 96 (5):167–174, December 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wu:2000:PTA

[Wu00] Bang Ye Wu. Polynomial time algorithms for some minimum latency problems. Information Processing Letters, 75(5):225– 229, October 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/64/30/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/64/30/30/article.
pdf.

Wu:2007:NC

[Wu07] Zhilin Wu. A note on the characterization of TL[EF]. Information Processing Letters, 102(2– 3):48–54, April 30, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wu:2008:EIM

[Wu08] Xiaodong Wu. Efficient intensity map splitting algorithms for intensity-modulated radiation therapy. Information Processing Letters, 106(5):188–194, May 31, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wu:2009:OAM

[Wu09a] Bang Ye Wu. An optimal algorithm for the maximum-density path in a tree. Information Processing Letters, 109(17):975– 979, August 16, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wu:2009:AUB

[Wu09b] Yufeng Wu. An analytical upper bound on the minimum number of recombinations in the history of SNP sequences in populations. Information Processing Letters, 109(9):427–431, April 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wu:2003:OAI

[WW03] Jer-Shyan Wu and Yu-Kuo Wang. An optimal algorithm to implement the Hanoi towers with parallel moves. Information Processing Letters, 86(6): 289–293, June 30, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wu:2009:CDS

[WW09] J. Wu and R. Wei. Comments on “Distributed symmetric key management for mobile ad hoc networks”. Information Processing Letters, 109(14):822– 824, June 30, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:2004:FVS

[WWC04] Fu-Hsing Wang, Yue-Li Wang, and Jou-Ming Chang. Feedback vertex sets in star graphs. Information Processing Letters, 89 (4):203–208, February 28, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wong:2005:FEF

[WWGF05] Angeline Wong, Leejay Wu, Phillip B. Gibbons, and Christos Faloutsos. Fast estimation of fractal dimension and correlation integral on stream data. Information Processing Letters, 93(2):91–97, January 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:2008:SPG

[WWW08] De-Qiang Wang, Yu-Peng Wen, and Ke-Lun Wang. A smaller planar graph without 4-, 5cycles and intersecting triangles that is not 3-choosable. Information Processing Letters, 108 (3):87–89, October 16, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:2005:SJU

[WX05] Ji-Bo Wang and Zun-Quan Xia.

Scheduling jobs under decreasing linear deterioration. Information Processing Letters, 94 (2):63–69, April 30, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Weifan:2006:LMF

[WY06] Wang Weifan and Wang Yiqiao.

L(p, q)-labelling of K_{4}-minor free graphs. Information Processing Letters, 98(5):169–173, June 15, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wan:2001:SMSa

[WYL01a] Guohua Wan, Benjamin P.-C.

Yen, and Chung-Lun Li. Single machine scheduling to minimize total compression plus weighted flow cost is NP-hard. Information Processing Letters, 79(6): 273–280, September 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/80/37/30/abstract.html.

Wan:2001:SMSb

[WYL01b] Guohua Wan, Benjamin P.-C. Yen, and Chung-Lun Li. Single machine scheduling to minimize total compression plus weighted flow cost is NP-hard. Information Processing Letters, 79(6): 273–280, September 30, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/80/37/30/abstract.html.

Wu:2006:PTA

[WYL06] Gang Wu, Jia-Huai You, and Guohui Lin. A polynomial time algorithm for the minimum quartet inconsistency problem with O(n) quartet errors. Information Processing Letters, 100 (4):167–171, November 30, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xu:2005:EFT

[XDX05] Jun-Ming Xu, Zheng-Zhong Du, and Min Xu. Edge-fault-tolerant edge-bipancyclicity of hypercubes. Information Processing Letters, 96(4):146–150, November 30, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xu:2005:LBE

[XDZ05] Yinfeng Xu, Wenqiang Dai, and Binhai Zhu. A lower bound on the edge l_{∞} radius of Saitou and Nei’s method for phylogenetic reconstruction. Information Processing Letters, 94(5): 225–230, June 15, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xiao:2005:SPA

[XH05] Lu Xiao and Howard M. Heys.

A simple power analysis attack against the key schedule of the Camellia block cipher. Information Processing Letters, 95(3): 409–412, August 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xiaofang:2007:CTG

[Xia07] Luo Xiaofang. The 4-choosability of toroidal graphs without intersecting triangles. Information Processing Letters, 102(2–3):85– 91, April 30, 2007. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xirotiri:2006:SSS

[Xir06] Olga Xirotiri. Simulation of simultaneous safe recursion over an arbitrary structure. Information Processing Letters, 99 (2):72–81, July 31, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xu:2006:ADB

[XKS06] Guangjun Xu, Liying Kang, and Erfang Shan. Acyclic domination on bipartite permutation graphs. Information Processing Letters, 99(4):139–144, August 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xue:2000:NFB

[XL00] Jue Xue and Ju Liu. A networkflow-based lower bound for the minimum weighted integer coloring problem. Information Processing Letters, 76(1–2):45– 50, November 20, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/24/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/24/31/article.
pdf.

Xu:2005:SCL

[XLMH05] Jun-Ming Xu, Min Lü, Meijie Ma, and Angelika Hellwig. Super connectivity of line graphs. Information Processing Letters, 94(4):191–195, May 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xu:2006:PMC

[XML06] Jun-Ming Xu, Meijie Ma, and Min Lü. Paths in Möbius cubes and crossed cubes. Information Processing Letters, 97(3):94–97, February 14, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xiao:2006:CGM

[XP06] Wenjun Xiao and Behrooz Parhami. Cayley graphs as models of deterministic small-world networks. Information Processing Letters, 97(3):115–117, February 14, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xiang:2000:ELG

[XUT00] Limin Xiang, Kazuo Ushijima, and Changjie Tang. Efficient loopless generation of Gray codes for k-ary trees. Information Processing Letters, 76(4–6):169–174, December 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/67/28/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/67/28/27/article.
pdf.

Xiang:2001:GAT

[XUT01] Limin Xiang, Kazuo Ushijima, and Changjie Tang. On generating k-ary trees in computer representation. Information Processing Letters, 77(5–6):231– 238, March 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/23/20/68/32/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/23/20/68/32/26/article.
pdf.

Xu:2005:LRA

[XX05a] Guang Xu and Jinhui Xu. An LP rounding algorithm for approximating uncapacitated facility location problem with penalties. Information Processing Letters, 94(3):119–123, May 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xu:2005:EPM

[XX05b] Min Xu and Jun-Ming Xu. Edge-pancyclicity of Möbius cubes. Information Processing Letters, 96(4):136–140, November 30, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xu:2007:FIA

[XX07] Min Xu and Jun-Ming Xu. The forwarding indices of augmented cubes. Information Processing Letters, 101(5):185–189, March 16, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xu:2005:FDC

[XXH05] Min Xu, Jun-Ming Xu, and Xin-Min Hou. Fault diameter of Cartesian product graphs. Information Processing Letters, 93 (5):245–248, March 16, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xu:2005:SCS

[XXZ05] Jun-Ming Xu, Min Xu, and Qiang Zhu. The super connectivity of shuffle-cubes. Information Processing Letters, 96 (4):123–127, November 30, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xu:2007:FDP

[XY07] Jun-Ming Xu and Chao Yang.

Fault diameter of product graphs. Information Processing Letters, 102(6):226–228, June 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Xu:2007:FTA

[XZX07] Jun-Ming Xu, Qiang Zhu, and Min Xu. Fault-tolerant analysis of a class of networks. Information Processing Letters, 103(6): 222–226, September 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yamaguchi:2004:FMC

[YAM04] Atsuko Yamaguchi, Kiyoko F.

Aoki, and Hiroshi Mamitsuka. Finding the maximum common subgraph of a partial k-tree and a graph with a polynomially bounded number of spanning trees. Information Processing Letters, 92(2):57–63, October 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:2000:CBQ

[Yan00] Jae Dong Yang. A conceptbased query evaluation with indefinite fuzzy triples. Information Processing Letters, 74(5–6): 209–214, June 30, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/63/28/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/63/28/29/article.
                                        pdf.

Yanovsky:2008:AAC

[Yan08] Vladimir Yanovsky. Approximation algorithm for coloring of dotted interval graphs. Information Processing Letters, 108(1): 41–44, September 15, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:2009:EFT

[Yan09] Ming-Chien Yang. Edge-fault-tolerant node-pancyclicity of twisted cubes. Information Processing Letters, 109(21–22): 1206–1210, October 31, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yan:1994:PPP

[YC94] Jing Ho Yan and Gerard J. Chang. The path-partition problem in block graphs. Information Processing Letters, 52(6):317–322, December 23, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See corrigendum [Cha02b].

Yilmaz:2009:ETB

[YC09] Okan Yilmaz and Ing-Ray Chen.

Elastic threshold-based admission control for QoS satisfaction with reward optimization for servicing multiple priority classes in wireless networks. Information Processing Letters, 109(15):868–875, July 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ying:2007:CQW

[YCFD07] Mingsheng Ying, Jianxin Chen, Yuan Feng, and Runyao Duan. Commutativity of quantum weakest preconditions. Information Processing Letters, 104 (4):152–158, November 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yi:2007:ESC

[YCKK07] Kwangkeun Yi, Hosik Choi, Jaehwang Kim, and Yongdai Kim. An empirical study on classification methods for alarms from a bug-finding static C analyzer. Information Processing Letters, 102(2–3):118–123, April 30, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:2006:MNG

[YCML06] Xiaofan Yang, Jianqiu Cao, Graham M. Megson, and Jun Luo. Minimum neighborhood in a generalized cube. Information Processing Letters, 97(3):88–93, February 14, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:2004:GHT

[YELM04] Xiaofan Yang, David J. Evans, Hongjian Lai, and Graham M. Megson. Generalized honeycomb torus is Hamiltonian. Information Processing Letters, 92 (1):31–37, October 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:2005:MIS

[YEM05] Xiaofan Yang, David J. Evans, and Graham M. Megson. Maximum induced subgraph of a recursive circulant. Information Processing Letters, 95(1): 293–298, July 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hsieh:2002:VRS

[yH02] Sun yuan Hsieh. On vertex ranking of a starlike graph. Information Processing Letters, 82(3): 131–135, May 16, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/86/34/
                                        29/abstract.html.

Yang:2005:FAC

[YHC05] I-Hsuan Yang, Chien-Pin Huang, and Kun-Mao Chao. A fast algorithm for computing a longest common increasing subsequence. Information Processing Letters, 93(5):249–253, March 16, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yokoyama:2004:DSO

[YHT04] Tetsuo Yokoyama, Zhenjiang Hu, and Masato Takeichi. Deterministic second-order patterns. Information Processing Letters, 89(6):309–314, March 31, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

YijieHan:2004:IAA

[Yij04] YijieHan. Improved algorithm for all pairs shortest paths. Information Processing Letters, 91 (5):245–250, September 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yang:2003:FTC

[YLTH03] Ming-Chien Yang, Tseng-Kuei Li, Jimmy J. M. Tan, and Lih-Hsing Hsu. Fault-tolerant cycleembedding of crossed cubes. Information Processing Letters, 88 (4):149–154, November 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yoshinaka:2009:EPG

[Yos09] Ryo Yoshinaka. An elementary proof of a generalization of double Greibach normal form. Information Processing Letters, 109(10):490–492, April 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Youn:2008:WRB

[YPKL08] Taek-Young Youn, Young-Ho Park, Changhan Kim, and Jongin Lim. Weakness in a RSAbased password authenticated key exchange protocol. Information Processing Letters, 108 (6):339–342, November 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yoshida:2003:NMC

[YSK03] Kiyoaki Yoshida, Yasumasa Sujaku, and Tohru Kohda. Number of mutual connections in neighborhoods and its application to self-diagnosable systems. Information Processing Letters, 88(3):89–94, November 15, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yamashita:2004:SPR

[YSK04] Masafumi Yamashita, Ichiro Suzuki, and Tiko Kameda. Searching a polygonal region by a group of stationary ksearchers. Information Processing Letters, 92(1):1–8, October 16, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yunes:2008:SAS

[Yun08] Jean-Baptiste Yunès. A 4states algebraic solution to linear cellular automata synchronization. Information Processing Letters, 107(2):71–75, July 16, 2008. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Yuster:2006:FCC

[Yus06] Raphael Yuster. Finding and counting cliques and independent sets in r-uniform hypergraphs. Information Processing Letters, 99(4):130–134, August 31, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yan:2009:AME

[YWW09] Guofeng Yan, Jianxin Wang, and Weiping Wang. An analytical model for end-to-end communication channel over PLCN based on QBDs. Information Processing Letters, 109(23–24): 1252–1259, November 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ye:2003:LSE

[YZ03] Deshi Ye and Guochuan Zhang.

On-line scheduling with extendable working time on a small number of machines. Information Processing Letters, 85 (4):171–177, February 28, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ye:2007:MTP

[YZ07] Deshi Ye and Guochuan Zhang.

Maximizing the throughput of parallel jobs on hypercubes. Information Processing Letters, 102(6):259–263, June 15, 2007.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yan:2008:EMG

[YZ08] Jun Yan and Jian Zhang. An efficient method to generate feasible paths for basis path testing. Information Processing Letters, 107(3–4):87–92, July 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhou:2009:LDG

[yZyLZ09] Qi yan Zhou, San yang Liu, and Qiang Zhu. Local diagnosability of generic star-pyramid graph. Information Processing Letters, 109(13):695–699, June 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zuhily:2007:OMP

[ZB07] Areej Zuhily and Alan Burns. Optimal (D − J)-monotonic priority assignment. Information Processing Letters, 103(6):247– 250, September 15, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2004:AIB

[ZC04] Fangguo Zhang and Xiaofeng Chen. Attack on an IDbased authenticated group key agreement scheme from PKC 2004. Information Processing Letters, 91(4):191–193, August 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2009:CII

[ZC09a] Fangguo Zhang and Xiaofeng Chen. Cryptanalysis and improvement of an ID-based adhoc anonymous identification scheme at CT-RSA 05. Information Processing Letters, 109(15): 846–849, July 16, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2009:BCA

[ZC09b] Jialin Zhang and Wei Chen.

Bounded cost algorithms for multivalued consensus using binary consensus instances. Information Processing Letters, 109(17):1005–1009, August 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2009:IUR

[ZC09c] Jialin Zhang and Wei Chen.

Implementing uniform reliable broadcast with binary consensus in systems with fair-lossy links. Information Processing Letters, 110(1):13–19, December 1, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zheng:2006:TLB

[ZCF^{+}06] Feifeng Zheng, Francis Y. L.

Chin, Stanley P. Y. Fung, Chung Keung Poon, and Yinfeng Xu. A tight lower bound for job scheduling with cancellation. Information Processing Letters, 97(1):1–3, January 16, 2006. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Zeadally:2000:PEJ

[Zea00] S. Zeadally. Performance evaluation of a Java-based networking Application Programming Interface (API). Information Processing Letters, 75(5):199– 209, October 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/23/20/64/30/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/23/20/64/30/27/article.
                                        pdf.

Zelke:2007:OPE

[Zel07] Mariano Zelke. Optimal peredge processing times in the semi-streaming model. Information Processing Letters, 104(3): 106–112, October 31, 2007. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zemmari:2008:HRG

[Zem08] Akka Zemmari. On handshakes in random graphs. Information Processing Letters, 108(3):119– 123, October 16, 2008. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zanuttini:2002:UFS

[ZH02] Bruno Zanuttini and Jean-Jacques Hébrard. A unified framework for structure identification. Information Processing Letters, 81(6):335–339, March 31, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/85/37/
34/abstract.html.

Zhang:2005:VRP

[ZH05] Huaming Zhang and Xin He.

Visibility representation of plane graphs via canonical ordering tree. Information Processing Letters, 96(2):41–48, October 31, 2005. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2006:SSL

[ZH06] Huaming Zhang and Xin He. On simultaneous straight-line grid embedding of a planar graph and its dual. Information Processing Letters, 99(1):1–6, July 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhou:2007:CAS

[ZH07] Yuren Zhou and Jun He. Convergence analysis of a selfadaptive multi-objective evolutionary algorithm based on grids. Information Processing Letters, 104(4):117–122, November 15, 2007. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhao:2006:NDN

[Zha06] Yunlei Zhao. A note on the Dwork–Naor timed deniable authentication. Information Processing Letters, 100(1):1–7, October 16, 2006. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhou:2008:EFG

[Zho08]  Xiang Zhou.  Ehrenfeucht–

fraïssé games in finite set theory. Information Processing Letters, 108(1):3–9, September 15, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zobel:2001:MHT

[ZHW01] Justin Zobel, Steffen Heinz, and Hugh E. Williams. In-memory hash tables for accumulating text vocabularies. Information Processing Letters, 80(6):271– 277, December 31, 2001. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

                                        URL http://www.elsevier.
                                        com/gej-ng/10/23/20/84/37/
                                        27/abstract.html.

Ziegler:2008:AAR

[Zie08] Valentin Ziegler. Approximation algorithms for restricted Bayesian network structures. Information Processing Letters, 108(2):60–63, September 30, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ziegler:2009:AOB

[Zie09] Valentin Ziegler. Approximating optimum branchings in linear time. Information Processing Letters, 109(3):175–178, January 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zivny:2009:SPO

[Živ09] Stanislav Živný. Structural properties of oracle classes. Information Processing Letters, 109(19):1131–1135, September 15, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhou:2005:HEI

[ZJ05] Sheng Zhou and Christopher B.

Jones. HCPO: an efficient insertion order for incremental Delaunay triangulation. Information Processing Letters, 93(1): 37–42, January 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2006:FET

[ZLJ06] Xianchao Zhang, Weifa Liang, and He Jiang. Flow equivalent trees in undirected nodeedge-capacitated planar graphs. Information Processing Letters, 100(3):110–115, November 15, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2009:BNC

[ZLM09] Xindong Zhang, Juan Liu, and Jixiang Meng. The bondage number in complete t-partite digraphs. Information Processing Letters, 109(17):997–1000, August 16, 2009. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zheng:2005:DSG

[ZLR^{+}05] Jun Zheng, Shahram Latifi, Emma Regentova, Kai Luo, and Xiaolong Wu. Diagnosability of star graphs under the comparison diagnosis model. Information Processing Letters, 93(1): 29–36, January 16, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zimmermann:2002:SSG

[ZLT02] W. Zimmermann, W. Löwe, and D. Trystram. On scheduling send-graphs and receive-graphs under the LogP-model. Information Processing Letters, 82 (2):83–92, April 30, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.
com/gej-ng/10/23/20/86/33/
29/abstract.html.

Zhao:2005:QTD

[ZLZ05] Jianhua Zhao, Xuandong Li, and Guoliang Zheng. A quadratic-time DBM-based successor algorithm for checking timed automata. Information Processing Letters, 96(3):101– 105, November 15, 2005. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhao:2009:SCP

[ZLZ09] Kewen Zhao, Yue Lin, and Ping Zhang. A sufficient condition for pancyclic graphs. Information Processing Letters, 109(17): 991–996, August 16, 2009. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhou:2003:DFD

[ZM03] Jianjun Zhou and Martin Müller. Depth-first discovery algorithm for incremental topological sorting of directed acyclic graphs. Information Processing Letters, 88(4):195–200, November 30, 2003. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhao:2003:LTA

[ZNI03] Liang Zhao, Hiroshi Nagamochi, and Toshihide Ibaraki. A linear time 5/3-approximation for the minimum strongly-connected spanning subgraph problem. Information Processing Letters, 86 (2):63–70, April 30, 2003. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhang:2008:CPG

[ZS08] Haihui Zhang and Zhiren Sun.

On 3-choosability of planar graphs without certain cycles. Information Processing Letters, 107(3–4):102–106, July 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhou:2009:FCG

[ZS09] Sizhong Zhou and Qiqing Shen.

On fractional (f, n)-critical graphs. Information Processing Letters, 109(14):811–815, June 30, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhou:2002:EMS

[ZSN02] Hai Zhou, Narendra Shenoy, and William Nicholls. Efficient minimum spanning tree construction without Delaunay triangulation. Information Processing Letters, 81(5):271–276, March 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic).  URL http://www.
elsevier.com/gej-ng/10/23/
20/85/36/33/abstract.html.

Zhang:2008:IPS

[ZSZY08] Changsheng Zhang, Jigui Sun, Xingjun Zhu, and Qingyun Yang. An improved particle swarm optimization algorithm for flowshop scheduling problem. Information Processing Letters, 108(4):204–209, October 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhou:2005:NWS

[ZXC05] Fengfeng Zhou, Yinlong Xu, and Guoliang Chen. No-wait scheduling in single-hop multichannel LANs. Information Processing Letters, 93(1):19–24, January 16, 2005. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zheng:2008:HMC

[ZXZ08] Feifeng Zheng, Yinfeng Xu, and E. Zhang. How much can lookahead help in online single machine scheduling. Information Processing Letters, 106(2):70– 74, April 15, 2008. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zarrabi-Zadeh:2008:FPT

[ZZ08] Hamid Zarrabi-Zadeh. Flying over a polyhedral terrain. Information Processing Letters, 105 (3):103–107, January 31, 2008. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zanuttini:2009:NSC

[ZŽ09] Bruno Zanuttini and Stanislav Živný. A note on some collapse results of valued constraints. Information Processing Letters, 109(11):534–538, May 16, 2009. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).